### v1

# Asymptotic condition numbers for linear ODEs

Abstract.
We are interested in the (relative) conditioning of the linear problem $y_{0}\mapsto\mathrm{e}^{tA}y_{0}$, i.e. the conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector. The present paper is a qualitative study of the long-time behavior of this conditioning. In other words, we are interested to study the propagation to the solution $y(t)$ of perturbations of the initial value for a linear ODE $y^{\prime}(t)=Ay(t)$, by measuring these perturbations with relative errors instead of absolute errors. We introduce three condition numbers: the first considers a specific initial value and a specific direction of perturbation; the second considers a specific initial value and the worst case by varying the direction of perturbation; and the third considers the worst case by varying both the initial value and the direction of perturbation. The long-time behaviors of these three condition numbers are studied.
Keywords: linear ordinary differential equations, matrix exponential, relative error, asymptotic behavior, condition numbers.
MSC2020 classification: 15A12, 15A16, 15A18, 15A21, 34A30, 34D05.
1. Introduction
We are interested to understand how a perturbation of the initial value $y_{0}$ of the linear $n-$dimensional Ordinary Differential Equation (ODE)
$\left\{\begin{array}[]{l}y^{\prime}\left(t\right)=Ay\left(t\right),\ t\in\mathbb{R},\\ y\left(0\right)=y_{0},\end{array}\right.$ (1)
where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$, is propagated to the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time $t$. This perturbation, propagating along the solution, is measured by a relative error rather than an absolute error. In other words, we study the (relative) conditioning of the linear problem
$y_{0}\mapsto\mathrm{e}^{tA}y_{0}$ (2)
for large time $t$.
The conditioning of the problem (2) has received little attention in literature despite of its significance in the context of ODEs. In contrast, the study of the conditioning of the matrix exponential $\mathrm{e}^{tA}$ is quite extensive; see, for example, [10], [15], [9], [11], [14], [2], [16], [3], [5], and [1]. However, none of these papers examined how the conditioning depends on $t$, a crucial aspect for ODEs. This type of analysis, for $A$ normal, was addressed in [12] for the problem (2) and in [7] for the problem $A\mapsto\mathrm{e}^{tA}y_{0}$. In the present paper, for the problem (2), we consider a general matrix $A$.
The present paper complements the work presented in [13]. The paper [13] offers a non-technical exposition without proofs of the main results found in the present paper. It includes motivations, practical examples, and numerical tests. Consequently, these aspects are not addressed in the present paper, which focuses on providing the proofs of all results and the necessary theoretical development for obtaining them.
1.1. The condition numbers
[…]
$\varepsilon=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$
of $\widetilde{y}_{0}$ and the normwise relative error
$\delta\left(t\right)=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$
of $\widetilde{y}(t)$. By writing
$\widetilde{y}_{0}=y_{0}+\varepsilon\|y_{0}\|\widehat{z}_{0},$
where $\widehat{z}_{0}\in\mathbb{C}^{n}$ with $\|\widehat{z}_{0}\|=1$ is the direction of perturbation, we obtain
$\delta\left(t\right)=K\left(t,y_{0},\widehat{z}_{0}\right)\cdot\varepsilon,$
where
$K\left(t,y_{0},\widehat{z}_{0}\right):=\frac{\left\|\mathrm{e}^{tA}\widehat{z}_{0}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$ (3)
with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$, is called the directional condition number of the problem (2).
Along with the condition number (3), we also introduce other two condition numbers: the condition number of the problem (2) (see [4] for the definition of condition number of a general problem) given by
$K\left(t,y_{0}\right):=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}K\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|\mathrm{e}^{tA}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$ (4)
[…]
$K\left(t\right):=\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}K\left(t,y_{0}\right)=\left\|\mathrm{e}^{tA}\right\|\cdot\left\|\mathrm{e}^{-tA}\right\|=\kappa\left(\mathrm{e}^{tA}\right),$ (5)
which equals the standard condition number $\kappa\left(\mathrm{e}^{tA}\right)$ of the matrix $\mathrm{e}^{tA}$.
Observe that $K\left(t,y_{0}\right)$ is the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$ and $K(t)$ is the worst condition number $K(t,y_{0})$ by varying $y_{0}$, i.e. the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying both $y_{0}$ and $\widehat{z}_{0}$.
The paper [12] studied the condition numbers (3), (4) and (5) in case of $A$ normal and the Euclidean norm as vector norm on $\mathbb{R}^{n}$. The present paper study the general case where $A$ (complex) and the vector norm are arbitrary.
Remark 1.
We also consider the case of a real ODE (1). In this case, we replace $\mathbb{C}$ by $\mathbb{R}$ and we have $A\in\mathbb{R}^{n\times n}$ and $y_{0},y(t),\widetilde{y}_{0},\widetilde{y}(t)\in\mathbb{R}^{n}$. Moreover, the condition numbers (4) and (5) are defined by maxima over $\widehat{z}_{0}\in\mathbb{R}^{n}$ with $\|\widehat{z}_{0}\|=1$ and $y_{0}\in\mathbb{R}^{n}$ with $y_{0}\neq 0$, respectively.
1.2. Plan of the paper
The present paper is quite long and technical. For this reason, it should be read after or alongside the previously cited expository work [13], which presents the main results given here without proofs and extensive technical details.
Here is the plan of the paper. Section 2 introduces some necessary technical linear algebra results concerning the Jordan Canonical Form (JCF) of $A$. Section 3 developes notions and notations for understanding the asymptotic (long-time) behavior considered in the next sections. Section 4 analyzes the asymptotic behaviors of the three conditions numbers $K(t,y_{0},$ $\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$, by introducing the three asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,y_{0})$ and $K_{\infty}(t)$. Section 5 summarizes the most important result of Section 4. From Section 6 to the last Section 10, the results of the asymptotic behavior analysis are applied in a generic situation for a real matrix $A$.
2. Linear Algebra Results
We introduce linear algebra notations and results necessary for the analysis developed in this paper. Although the content pertains to the well-known topic of the JCF of a matrix, it addresses very specific aspects that are either not known or not sufficiently detailed for the purposes of this paper.
This section can be skipped and referred back to as needed while reading the subsequent sections.
The main goal of the section is to derive a formula for the matrix exponential $\mathrm{e}^{tA}$, where it is explicitly identified how $\mathrm{e}^{tA}$ depends on $t$. Since it is based on the JCF of $A$, first we revise such a form.
2.1. The JCF of the matrix $A$
Let $A\in\mathbb{C}^{n\times n}$ and let $\lambda_{1},\ldots,\lambda_{p}$ be the distinct eigenvalues of $A$. The matrix $A$ is similar to a matrix $J\in\mathbb{C}^{n\times n}$, called a Jordan Canonical Form (JCF) of $A$, with the following structure.
  • The matrix $J$ is block-diagonal with $p$ blocks $J^{\left(1\right)},\ldots,J^{\left(p\right)}$ called Jordan blocks:
    $J=\mathrm{diag}\left(J^{(1)},\ldots,J^{(p)}\right)\in\mathbb{C}^{n\times n}.$
  • For any $i\in\{1,\ldots,p\}$, the Jordan block $J^{\left(i\right)}$ has dimension $\nu_{i}$, where $\nu_{i}$ is the algebraic multiplicity of $\lambda_{i}$, and it is block-diagonal with $d_{i}$ blocks $J^{\left(i,1\right)},\ldots,J^{\left(i,d_{i}\right)}$ called Jordan mini-blocks, where $d_{i}$ is the geometric multiplicity of $\lambda_{i}$:
    $J^{(i)}=\mathrm{diag}\left(J^{\left(i,1\right)},\ldots,J^{\left(i,d_{i}\right)}\right)\in\mathbb{C}^{\nu_{i}\times\nu_{i}}.$
  • For any $i\in\{1,\ldots,p\}$ and for any $j\in\{1,\ldots,d_{i}\}$, the Jordan mini-block $J^{\left(i,j\right)}$ is upper bi-diagonal:
    $J^{\left(i,j\right)}=\left[\begin{array}[]{ccccc}\lambda_{i}&1&&&\\ &.&1&&\\ &&.&1&\\ &&&.&1\\ &&&&\lambda_{i}\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}.$
    The dimension of $J^{\left(i,j\right)}$ is denoted by $m_{ij}$.
There is a unique JCF of $A$, except for permutations of the blocks and permutations of the mini-blocks within the blocks.
For any $i\in\{1,\ldots,p\}$, we have
$\sum\limits_{j=1}^{d_{i}}m_{ij}=\nu_{i}$
and we call
$m_{i}:=\max_{j\in\{1,\ldots,d_{i}\}}m_{ij}$ (6)
the ascent (or index) of $\lambda_{i}$.
An eigenvalue $\lambda_{i}$, $i\in\{1,\ldots,p\}$, is called defective if $d_{i}<\nu_{i}$ and non-defective if $d_{i}=\nu_{i}$. Clearly, $\lambda_{i}$ is non-defective if and only if $m_{i}=1$, i.e. $J^{(i)}$ is diagonal. The matrix $A$ is diagonalizable if and only if all the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ are non-defective, i.e. $J$ is diagonal.
An eigenvalue $\lambda_{i}$, $i\in\{1,\ldots,p\}$, is called simple if $d_{i}=\nu_{i}=1$.
2.1.1. Three-index notation
In the following, we denote an index $l\in\{1,\ldots,n\}$ by three indices $\left(i,j,k\right)$, where (see Figure 1):
  • $i\in\{1,\ldots,p\}$ is the index of the block $J^{(i)}$ traversed by the $l$-th column (or row) of $J$.;
  • $j\in\left\{1,\ldots,d_{i}\right\}$ is the index of the mini-block $J^{(i,j)}$ of $J^{(i)}$ traversed by the $l$-th column (or row) of $J$;
  • $k\in\left\{1,\ldots,m_{ij}\right\}$ is the index of the column (or row) of the mini-block $J^{(i,j)}$ included in the $l$-th column (row) of $J$.
Observe that the triples $(i,j,k)$ appear lexicographically ordered when the index $l$ moves from $1$ to $n$.
Figure 1. The index $l$ and the three indices $(i,j,k)$.
2.1.2. The Jordan basis
Since $A$ is similar to $J$, there exists $V\in\mathbb{C}^{n\times n}$ non-singular such that
$J=V^{-1}AV.$ (7)
Let
$V=\left[v^{(1)}\ \cdots\ v^{(n)}\right].$
The $n$ columns $v^{(1)},\ldots,v^{(n)}$ of $V$ constitute a Jordan basis of the space $\mathbb{C}^{n}$. In the three-index notation, the $n$ columns $v^{(l)}$, $l\in\{1,\ldots,n\}$, are denoted by
$v^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\}.$ (8)
For $i\in\{1,\ldots,p\}$, the vectors
$v^{\left(i,j,1\right)},\ j\in\{1,\ldots,d_{i}\},$
are eigenvectors corresponding to the eigenvalue $\lambda_{i}$: they constitute a basis for the eigenspace corresponding to the eigenvalue $\lambda_{i}$. The vectors
$v^{\left(i,j,k\right)},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$
are generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$: they constitute a basis for the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$.
2.1.3. Jordan chains
The Jordan basis (8) is partitioned in the Jordan chains
$\left(v^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}}=\left(v^{\left(i,j,1\right)},v^{\left(i,j,2\right)},\ldots,v^{\left(i,j,m_{ij}\right)}\right)$
$i\in\{1,\ldots,p\}\text{\ and\ }j\in\{1,\ldots,d_{i}\}.$
The elements of a Jordan chain satisfy
$\left(A-\lambda_{i}I\right)v^{\left(i,j,1\right)}=0$
$\left(A-\lambda_{i}I\right)v^{\left(i,j,k+1\right)}=v^{\left(i,j,k\right)},\ k\in\{1,\ldots,m_{ij}-1\}.$
The chain stops with $v^{\left(i,j,m_{ij}\right)}$ since the system
$\left(A-\lambda_{i}I\right)x=v^{\left(i,j,m_{ij}\right)}$
has not a solution.
Observe that (7) is equivalent to have (2.1.3), for $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, and that there is a correspondence one-to-one between mini-blocks and Jordan chains.
2.1.4. The matrix $zA$
The next proposition describes the JCF and a Jordan basis of the matrix $zA$, where $z\in\mathbb{C}\setminus\{0\}$, in terms of the JCF and a Jordan basis of $A$.
Proposition 2.
Let $z\in\mathbb{C}\setminus\{0\}$. The distinct eigenvalues of $zA$ are $z\lambda_{i}$$i\in\{1,\ldots,p\}$. For any $i\in\{1,\ldots,p\}$, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ in the JCF of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ in the JCF of $A$. A Jordan basis of $zA$ is
$z^{-(k-1)}v^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\}.$ (10)
Proof.
Let $i\in\{1,\ldots,p\}$. Given a Jordan chain
$\left(v^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}}$
of $A$ corresponding to the eigenvalue $\lambda_{i}$, we have that
$\left(z^{-(k-1)}v^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}}$ (11)
is a Jordan chain of $zA$ corresponding to the eigenvalue $z\lambda_{i}$.
In fact, we have (see (2.1.3))
$\left(zA-z\lambda_{i}I\right)v^{\left(i,j,1\right)}=0$
$\left(zA-z\lambda_{i}I\right)z^{-k}v^{\left(i,j,k+1\right)}=z^{-(k-1)}v^{\left(i,j,k\right)},\ k\in\{1,\ldots,m_{ij}-1\}.$
Moreover, the system
$\left(zA-z\lambda_{i}I\right)x=z^{-(m_{ij}-1)}v^{\left(i,j,m_{ij}\right)}$
has no solution; otherwise, the system
$\left(A-\lambda_{i}I\right)y=v^{\left(i,j,m_{ij}\right)}$
would have the solution $y=z^{m_{ij}}x$.
By exchanging the role of $A$ and $zA$, i.e. we consider $zA$ and $z^{-1}(zA)$, we also see that given a Jordan chain
$\left(u^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}(zA)}$
of $zA$ corresponding to the eigenvalue $z\lambda_{i}$,
$\left(z^{k-1}u^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}(zA)}$
is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i}$ (here, $m_{ij}(zA)$ denotes the length of a Jordan chain of $zA$). This shows that the Jordan chains of $zA$ are all of type (11). Consequently, the number and the lenghts of the Jordan chains corresponding to the eigenvalue $z\lambda_{i}$ of $\alpha A$ are equal to the number and the lengths of the Jordan chains corresponding to the eigenvalue $\lambda_{i}$ of $A$. Therefore, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ of $A$.
A Jordan basis for $zA$ is given by collecting all the Jordan chains (11). Thus, we obtain the Jordan basis (10) for $zA$. ∎
2.1.5. The real case
When $A$ is a real matrix, the distinct complex eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are divided in real eigenvalues and complex conjugate pairs of eigenvalues. The next proposition describes the JCF and a Jordan basis of $A$, when $A$ is real, in terms of these real eigenvalues and complex conjugate pairs of eigenvalues.
Here and in the following, for a vector or matrix $Z$, $\overline{Z}$ denotes the vector or matrix given by the complex conjugates of the elements of $Z$.
Proposition 3.
Assume $A\in\mathbb{R}^{n\times n}$. For $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e. $\lambda_{i_{1}}$ and $\lambda_{i_{2}}$ form a complex conjugate pair of eigenvalues, we have
$\nu_{i_{2}}=\nu_{i_{1}},\ d_{i_{2}}=d_{i_{1}}\text{\ and\ }m_{i_{2}j}=m_{i_{1}j},\ j\in\{1,\ldots,d_{i_{2}}\}.$ (12)
Moreover, there exists a Jordan basis of $A$ such that:
  • for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$, we have
    $v^{\left(i,j,k\right)}\in\mathbb{R}^{n},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\};$
  • for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, we have
    $v^{\left(i_{2},j,k\right)}=\overline{v^{\left(i_{1},j,k\right)}},\ j\in\{1,\ldots,d_{i_{2}}\}\text{\ and\ }k\in\{1,\ldots,m_{i_{2}j}\}.$ (13)
    Observe that $d_{i_{2}}=d_{i_{1}}$ and $m_{i_{2}j}=m_{i_{1}j}$ hold in (13).
Proof.
Consider a complex conjugate pair given by $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$. Given a Jordan chain
$\left(v^{\left(i_{1},j,k\right)}\right)_{k=1,\ldots,m_{i_{1}j}}$
of $A$ corresponding to the eigenvalue $\lambda_{i_{1}}$, we have that
$\left(\overline{v^{\left(i_{1},j,k\right)}}\right)_{k=1,\ldots,m_{i_{1}j}}$ (14)
is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
In fact, by conjugating both sides in (2.1.3), we have
$\left(A-\overline{\lambda_{i_{1}}}I\right)\overline{v^{\left(i_{1},j,1\right)}}=0$
$\left(A-\overline{\lambda_{i_{1}}}I\right)\overline{v^{\left(i_{1},j,k+1\right)}}=\overline{v^{\left(i_{1},j,k\right)}},\ k\in\{1,\ldots,m_{i_{1}j}-1\}.$
Moreover, the system
$\left(A-\overline{\lambda_{i_{1}}}I\right)x=\overline{v^{\left(i_{1},j,m_{i_{1}j}\right)}}$
has no solution; otherwise, the system
$\left(A-\lambda_{i_{1}}I\right)y=v^{\left(i_{1},j,m_{i_{1}j}\right)}$
would have the solution $y=\overline{x}$.
By exchanging the role of $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e. we consider $\lambda_{i_{2}}$ and $\overline{\lambda_{i_{2}}}$, we also see that given a Jordan chain
$\left(u^{\left(i_{2},j,k\right)}\right)_{k=1,\ldots,m_{i_{2}j}}$
corresponding to the eigenvalue $\lambda_{i_{2}}$,
$\left(\overline{u^{\left(i_{2},j,k\right)}}\right)_{k=1,\ldots,m_{i_{2}j}}$
is a Jordan chain corresponding to the eigenvalue $\lambda_{i_{1}}=\overline{\lambda_{i_{2}}}$. This shows that the Jordan chains corresponding to $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ are all of type (14). This implies (12). By collecting all the Jordan chains (14) corresponding to $\lambda_{i_{1}}$, we obtain a Jordan basis satisfying (13).
The result about a real eigenvalue $\lambda_{i}$ follows from the next two facts:
  • the eigenspace of $\lambda_{i}$ in $\mathbb{C}^{n}$ has a basis of real eigenvalues;
  • if $v^{(i,j,k)}$ is real and the linear system
    $(A-\lambda_{i}I)x=v^{(i,j,k)}$
    has a solution in $\mathbb{C}^{n}$, then it also has a solution in $\mathbb{R}^{n}$.
Thus, we can have Jordan chains corresponding to $\lambda_{i}$ constituted by real generalized eigenvectors. ∎
In the following, in case of a real matrix $A$, we assume to have a Jordan basis as that described in the previous proposition.
2.2. The matrices $V^{(i)}$, $W^{(i)}$, $N^{(i,l)}$ and the vector $\alpha^{(i)}(u)$
The formula of our interest for the matrix exponential $\mathrm{e}^{tA}$ is constructed by using the matrices $V^{(i)}$, $W^{(i)}$ and $N^{(i,l)}$ now introduced. Here, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ (remind that $m_{i}$ is the ascent of $\lambda_{i}$ defined in (6)). We also introduce the vector $\alpha^{(i)}(u)$ of components of $u$ in the Jordan basis.
Remind that $V$ is the matrix whose columns (8) constitute a Jordan basis.
  • For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
    $V^{(i,j)}:=\left[v^{(i,j,1)}\ \cdots\ v^{(i,j,m_{ij})}\right]\in\mathbb{C}^{n\times m_{ij}}$
    and let
    $V^{(i)}:=\left[V^{(i,1)}\ \cdots\ V^{(i,d_{i})}\right]\in\mathbb{C}^{n\times\nu_{i}}.$ (15)
    Observe that
    $V=\left[V^{(1)}\ \cdots\ V^{(p)}\right].$
  • Let $W=V^{-1}$ and let
    $w^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$ (16)
    be the rows of $W$ in the three-index notation. The $n$ row vectors in (16) are called left generalized eigenvectors of $A$, whereas, as we have already seen, the $n$ columns vectors in (8) are called (right) generalized eigenvectors of $A$. For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
    $W^{(i,j)}:=\left[\begin{array}[]{c}w^{(i,j,1)}\\ \vdots\\ w^{(i,j,m_{ij})}\end{array}\right]\in\mathbb{C}^{m_{ij}\times n}$
    and let
    $W^{(i)}:=\left[\begin{array}[]{c}W^{(i,1)}\\ \vdots\\ W^{(i,d_{i})}\end{array}\right]\in\mathbb{C}^{\nu_{i}\times n}.$ (17)
    Observe that
    $W=\left[\begin{array}[]{c}W^{(1)}\\ \vdots\\ W^{(p)}\end{array}\right].$
  • For $u\in\mathbb{C}^{n}$, let
    $\alpha(u):=Wu.$
    Observe that $\alpha(u)$ is the vector of the components of $u$ in the Jordan basis. In the three-index notation, the components of $\alpha(u)$ are
    $\alpha_{ijk}(u),\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\}.$
    $\alpha_{ijk}(u)$ is the component of $u$ along $v^{(i,j,k)}$. For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
    $\alpha^{(i,j)}(u):=\left[\begin{array}[]{c}\alpha_{ij1}(u)\\ \vdots\\ \alpha_{ijm_{ij}}(u)\end{array}\right]\in\mathbb{C}^{m_{ij}}.$
    and let
    $\alpha^{(i)}(u):=\left[\begin{array}[]{c}\alpha^{(i,1)}(u)\\ \vdots\\ \alpha^{(i,d_{i})}(u)\end{array}\right]\in\mathbb{C}^{\nu_{i}}.$
    Observe that
    $\alpha(u):=\left[\begin{array}[]{c}\alpha^{(1)}(u)\\ \vdots\\ \alpha^{(p)}(u)\end{array}\right].$
  • For $i\in\{1,\ldots,p\}$, $l\in\{0,\ldots,m_{i}-1\}$ and $j\in\{1,\ldots,d_{i}\}$, let
    $N^{(i,l,j)}:=\left\{\begin{array}[]{l}\left[\begin{array}[]{ccccccc}0&.&0&1&0&.&0\\ &.&.&.&.&.&.\\ &&.&.&.&.&0\\ &&&.&.&.&1\\ &&&&.&.&0\\ &&&&&.&.\\ &&&&&&0\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }l\leq m_{ij}-1\\ \\ 0\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }m_{ij}\leq l\leq m_{i}-1,\end{array}\right.$ (18)
    where the upper diagonal of elements equal to $1$ is the $l$-th upper diagonal. For $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$, let
    $N^{(i,l)}:=\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\in\mathbb{C}^{\nu_{i}\times\nu_{i}}.$ (19)
    Observe that the matrix $N^{(i,l)}$ has the same dimensions $\nu_{i}\times\nu_{i}$ of the Jordan block $J^{(i)}=\mathrm{diag}\left(J^{(i,1)},\ldots,J^{(i,d_{i})}\right)$ and the diagonal blocks $N^{(i,l,1)},$ $\ldots$,… , $N^{(i,l,d_{i})}$ have the same dimensions of the Jordan mini-blocks $J^{(i,1)},\ldots,$ $J^{(i,d_{i})}$, respectively. Thus, as it is illustrated in Figure 2, the matrix $N^{(i,l)}$ has $1$ in the intersection of the $l$-th upper diagonal with the (frame of the) Jordan mini-blocks and 00 in all the other places.
Figure 2. The matrix $N^{(i,l)}$.
2.2.1. The matrix $zA$
When we replace $A$ by $zA$, $z\in\mathbb{C}\setminus\{0\}$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2). Moreover, a Jordan basis for $zA$ is given in (10). Therefore, we know how it is transformed the matrix $V$ by replacing $A$ by $zA$. The next proposition says this and, in addition, how it is transformed the matrix $W=V^{-1}$.
Proposition 4.
Let $z\in\mathbb{C}\setminus\{0\}$. The matrix $V(zA)$ corresponding to $zA$ has columns
$z^{-(k-1)}v^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$ (20)
and the matrix $W(zA)$ corresponding to $zA$ has rows
$z^{k-1}w^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\}.$ (21)
Proof.
The columns of the matrix $V(zA)$ are given in (10). For $i_{1},i_{2}\in\{1,\ldots,p\}$, $j_{1}\in\{1,\ldots,d_{i_{1}}\}$, $j_{2}\in\{1,\ldots,d_{i_{2}}\}$, $k_{1}\in\{1,\ldots,m_{i_{1}j_{1}}\}$ and $k_{2}\in\{1,\ldots,m_{i_{2}j_{2}}\}$, we have
$z^{k_{1}-1}w^{\left(i_{1},j_{1},k_{1}\right)}z^{-(k_{2}-1)}v^{\left(i_{2},j_{2},k_{2}\right)}=\left\{\begin{array}[]{l}1\ \ \text{if}\ \ (i_{1},j_{1},k_{1})=(i_{2},j_{2},k_{2})\\ 0\ \ \text{otherwise}\end{array}\right.$
since $W$, whose rows are (16), is the inverse of $V$, whose columns are (8). This shows that the matrix of rows (21) is the inverse of the matrix of columns (20). ∎
2.2.2. The real case
The next proposition describes, when the matrix $A$ and the vector $u$ are real, the matrices $V^{(i)}$ and $W^{(i)}$ and the vector $\alpha^{(i)}(u)$ in terms of the real eigenvalues and the complex conjugate pairs of eigenvalues of $A$.
Proposition 5.
Assume $A\in\mathbb{R}^{n\times n}$ and $u\in\mathbb{R}^{n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
For $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$, we have
$V^{(i)}\in\mathbb{R}^{n\times\nu_{i}},\ W^{(i)}\in\mathbb{R}^{\nu_{i}\times n}\text{\ \ and\ \ }\alpha^{(i)}(u)\in\mathbb{R}^{\nu_{i}}.$
For $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, we have
$V^{(i_{2})}=\overline{V^{(i_{1})}},\ W^{(i_{2})}=\overline{W^{(i_{1})}}\text{\ \ and\ \ }\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}.$
Proof.
Since we are assuming to have a Jordan basis as that described in Proposition 3, we have $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
Now, we prove that for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ we have $\alpha^{(i)}(u)\in\mathbb{R}$, and for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$ we have $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$. By conjugating both sides of
$u=V\alpha(u)=\sum\limits_{i=1}^{p}V^{(i)}\alpha^{(i)}(u),$
we obtain
$u=\sum\limits_{i=1}^{p}\overline{V^{(i)}}\ \overline{\alpha^{(i)}(u)}.$
So, by separating real eigenvalues and complex conjugate pairs of eigenvalues, we have
$u=\sum\limits_{i=1}^{p}V^{(i)}\alpha^{(i)}(u)=\sum\limits_{\lambda_{i}\in\mathbb{R}}V^{(i)}\alpha^{(i)}(u)+\sum\limits_{\begin{subarray}{c}\lambda_{i_{1}}\in\mathbb{C}\\ \mathrm{Im}\left(\lambda_{i_{1}}\right)>0\end{subarray}}\left(V^{(i_{1})}\alpha^{(i_{1})}(u)+V^{(i_{2})}\alpha^{(i_{2})}(u)\right),$
as well as
$u=\sum\limits_{i=1}^{p}\overline{V^{(i)}}\ \ \overline{\alpha^{(i)}(u)}=\sum\limits_{\lambda_{i}\in\mathbb{R}}V^{(i)}\ \overline{\alpha^{(i)}(u)}+\sum\limits_{\begin{subarray}{c}\lambda_{i_{1}}\in\mathbb{C}\\ \mathrm{Im}\left(\lambda_{i_{1}}\right)>0\end{subarray}}\left(V^{(i_{2})}\overline{\alpha^{(i_{1})}(u)}+V^{(i_{1})}\overline{\alpha^{(i_{2})}(u)}\right)$
by reminding that $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$. Since $u$ can be expressed as a linear combination of the Jordan basis in a unique manner, we obtain $\alpha^{(i)}(u)=\overline{\alpha^{(i)}(u)}$, i.e. $\alpha^{(i)}(u)\in\mathbb{R}$, for $\lambda_{i}\in\mathbb{R}$ and $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i_{1}}\right)>0$.
Finally, we show that, for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, we have $W^{(i_{2})}=\overline{W^{(i_{1})}}$. This includes the case $i_{1}=i_{2}$ and $\lambda_{i_{1}}\in\mathbb{R}$, for which we can conclude that $W^{(i_{1})}$ is a real matrix. We have, with $e^{\left(1\right)},\ldots,e^{\left(n\right)}$ the real vectors of the canonical basis of $\mathbb{C}^{n}$,
$W^{(i_{2})}$ $=$ $W^{(i_{2})}\left[e^{\left(1\right)}\ \cdots\ e^{\left(n\right)}\right]=\left[\alpha^{\left(i_{2}\right)}\left(e^{\left(1\right)}\right)\ \cdots\ \alpha^{\left(i_{2}\right)}\left(e^{(n)}\right)\right]$
$=$ $\left[\overline{\alpha^{\left(i_{1}\right)}\left(e^{\left(1\right)}\right)}\ \cdots\ \overline{\alpha^{\left(i_{1}\right)}\left(e^{(n)}\right)}\right]=\overline{\left[\alpha^{\left(i_{1}\right)}\left(e^{\left(1\right)}\right)\ \cdots\ \alpha^{\left(i_{1}\right)}\left(e^{(n)}\right)\right]}$
$=$ $\overline{W^{(i_{1})}\left[e^{\left(1\right)}\ \cdots\ e^{\left(n\right)}\right]}=\overline{W^{(i_{1})}}.$
2.3. The formula for $\mathrm{e}^{tA}$
Next proposition provides the announced formula for the matrix exponential $\mathrm{e}^{tA}$. For taste of generality, we consider the general situation of a function of matrix $f(A)$, where $f(z)$, $z\in\mathcal{D}\subseteq\mathbb{C}$, is an analytic complex function of $z$. The domain $\mathcal{D}$ of $f$ is an open subset of $\mathbb{C}$ and we assume that the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are contained in $\mathcal{D}$.
Proposition 6.
We have
$f(A)=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{m_{i}-1}\frac{f^{(l)}(\lambda_{i})}{l!}P_{il},$ (22)
where, for $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$,
$P_{il}:=V^{\left(i\right)}N^{(i,l)}W^{\left(i\right)}\in\mathbb{C}^{n\times n}$ (23)
with $V^{\left(i\right)}$, $W^{\left(i\right)}$ and $N^{(i,l)}$ defined in (15), (17) and (19), respectively.
Proof.
By recalling the definition of function of matrix by the JCF (see, e.g., [8]), we have
$f(A)=Vf(J)W=\sum\limits_{i=1}^{p}V^{\left(i\right)}f(J^{(i)})W^{\left(i\right)},$ (24)
where
$f\left(J\right)=\mathrm{diag}\left(f\left(J^{(1)}\right),\ldots,f\left(J^{(p)}\right)\right)\in\mathbb{C}^{n\times n}$
and
$f\left(J^{(i)}\right)=\mathrm{diag}\left(f\left(J^{(i,1)}\right),\ldots,f\left(J^{(i,d_{i})}\right)\right)\in\mathbb{C}^{\nu_{i}\times\nu_{i}},\ i\in\{1,\ldots,p\},$
with
$f\left(J^{(i,j)}\right):=\left[\begin{array}[]{ccccc}f(\lambda_{i})&f^{\prime}(\lambda_{i})&\frac{f^{\prime\prime}(\lambda_{i})}{2}&\cdots&\frac{f^{(m_{ij}-1)}(\lambda_{i})}{\left(m_{ij}-1\right)!}\\ &.&.&.&\vdots\\ &&.&.&\frac{f^{\prime\prime}(\lambda_{i})}{2}\\ &&&.&f^{\prime}(\lambda_{i})\\ &&&&f(\lambda_{i})\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}$
$j\in\{1,\ldots,d_{i}\}.$
We can write
$f\left(J^{(i,j)}\right)=\sum\limits_{l=0}^{m_{ij}-1}\frac{f^{(l)}(\lambda_{i})}{l!}N^{(i,l,j)}=\sum\limits_{l=0}^{m_{i}-1}\frac{f^{(l)}(\lambda_{i})}{l!}N^{(i,l,j)},$
where the matrices $N^{(i,l,j)}$, $l\in\{0,\ldots,m_{i}-1\}$, are defined in (18).
We conclude that, for $i\in\{1,\ldots,p\}$, we have
$f\left(J^{(i)}\right)$ $=$ $\mathrm{diag}\left(\sum\limits_{l=0}^{m_{i}-1}\frac{f^{(l)}(\lambda_{i})}{l!}N^{(i,l,1)},\ldots,\sum\limits_{l=0}^{m_{i}-1}\frac{f^{(l)}(\lambda_{i})}{l!}N^{(i,l,d_{i})}\right)$
$=$ $\sum\limits_{l=0}^{m_{i}-1}\frac{f^{(l)}(\lambda_{i})}{l!}N^{(i,l)}.$
Now (22) follows by (24). ∎
For the case $f(A)=\mathrm{e}^{tA}$ of our interest, formula (22) becomes
$\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il}.$ (26)
2.3.1. The matrices $P_{i0}$
For $i\in\{1,\ldots,p\}$, since $N^{(i,0)}=I_{n}$, we have
$P_{i0}=V^{\left(i\right)}W^{\left(i\right)}=\sum\limits_{j=1}^{d_{i}}\sum\limits_{k=1}^{m_{ij}}v^{(i,j,k)}w^{(i,j,k)}.$ (27)
The matrix $P_{i0}$ is the projection onto the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e. the subspace spanned by the generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$. In fact, we have, for $u\in\mathbb{C}^{n}$,
$P_{i0}u=V^{\left(i\right)}W^{\left(i\right)}u=V^{\left(i\right)}\alpha^{\left(i\right)}(u)=\sum\limits_{j=1}^{d_{i}}\sum\limits_{k=1}^{m_{ij}}\alpha_{ijk}(u)v^{(i,j,k)}.$
2.3.2. The case $A$ diagonalizable
When $A$ is diagonalizable, the formula (22) simplifies to the well-known formula
$f(A)=\sum\limits_{i=1}^{p}f(\lambda_{i})P_{i0}.$ (28)
where, for $i\in\{1,\ldots,p\}$,
$P_{i0}=\sum\limits_{j=1}^{d_{i}}v^{(i,j,1)}w^{(i,j,1)}.$
In this case of $A$ diagonalizable, the matrix $P_{i0}$ is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e. the subspace spanned by the eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
For the case $f(A)=\mathrm{e}^{tA}$ of our interest, formula (28) becomes
$\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}P_{i0}.$
2.4. The matrices $P_{il}$
In this subsection, we see some properties of the matrices $P_{il}$ defined in (23).
2.4.1. An alternative expression for $P_{il}$
Next proposition gives an alternative expression for $P_{il}$, more informative than the definition (23).
Proposition 7.
For $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$, we have
$P_{il}=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}v^{\left(i,j,k\right)}w^{\left(i,j,l+k\right)}.$ (29)
Moreover, for $u\in\mathbb{C}^{n}$, we have
$P_{il}u=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}\alpha_{ij\,l+k}(u)v^{\left(i,j,k\right)}.$ (30)
The indeces $j$ in the outer sum
$\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}$
in (29) and (30) are the indeces $j$ of the mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal. The indeces $k$ in the inner sum
$\sum\limits_{k=1}^{m_{ij}-l}$
are the row indeces of the elements of the mini-block $J^{(i,j)}$ on this $l$-th upper diagonal (see Figure 3). The column indeces $l+k$ of these elements appear as third indeces in $w^{(i,j,l+k)}$ in (29) and $\alpha_{ij\,l+k}(u)$ in (30). In other words, with reference to the previous Figure 2, in the double sum in (29) and (30) we are summing over all elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with the mini-blocks: the three-index notation $(i,j,k)$ of the row of these elements appears in $v^{(i,j,k)}$ and the three-index notation $(i,j,l+k)$ of the columns appears in $w^{(i,j,l+k)}$ for (29) and in $\alpha_{ijl+k}(u)$ for (30).
Proof.
As for (29), by the definition (23) we have
$P_{il}$ $=$ $\sum\limits_{j=1}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
$=$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
$=$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}v^{\left(i,j,k\right)}w^{\left(i,j,l+k\right)}$
by reminding the form (18) of $N^{(i,l,j)}$.
As for (30), we have
$P_{il}u=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}v^{\left(i,j,k\right)}\underset{=\alpha_{ij\,l+k}(u)}{\underbrace{w^{\left(i,j,l+k\right)}u}}.$
Figure 3. Indeces $j$ and $k$ in (29) and (30): $j$ is an index of mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal and $k$ is an index of row of elements in the $l$-th upper diagonal.
Remark 8.
  • 1.
    For $l=0$, formula (29) reduces to formula (27).
  • 2.
    If there exists an unique $j\in\{1,\ldots,d_{i}\}$ such that $m_{i}=m_{ij}$, then
    $P_{i\,m_{i}-1}=v^{\left(i,j,1\right)}w^{\left(i,j,m_{i}\right)}$
    and, for $u\in\mathbb{C}^{n}$,
    $P_{i\,m_{i}-1}u=\alpha_{ijm_{i}}(u)v^{\left(i,j,1\right)}.$
  • 3.
    As we can see in (30), $P_{il}u$ is not the projection of $u$ onto the subspace spanned by the vectors
    $v^{(i,j,k)},\ j\in\{1,\ldots,d_{i}\}\text{\ with \ }m_{ij}\geq l+1\text{\ and\ }k\in\{1,\ldots,m_{ij}-l\},$
    since these vectors are multiplied by coefficients different from the corresponding components $\alpha_{ijk}(u)$. The coefficients are the components $\alpha_{ijl+k}(u)$ shifted by $l$ in the third index.
2.4.2. The matrix $zA$
The next proposition shows how the matrices $P_{il}$ are transformed when the matrix $A$ is replaced by $zA$, $z\in\mathbb{C}\setminus\{0\}$. Observe that, when $A$ is replaced by $zA$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2). Thus, the indeces $i$ and $l$ for the matrices $P_{il}(zA)$ corresponding to $zA$ are the same indeces $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ for the matrices $P_{il}$ corresponding to $A$.
Proposition 9.
Let $z\in\mathbb{C}\setminus\{0\}$. We have
$P_{il}(zA)=z^{l}P_{il},\ i\in\{1,\ldots,p\}\text{\ and\ }l\in\{0,\ldots,m_{i}-1\}.$
Proof.
The columns of the matrix $V(zA)$ corresponding to $zA$ are given in (20) and the rows of the matrix $W(zA)$ corresponding to $zA$ are given in (21). Now, use the formula (29). ∎
2.4.3. The real case
The next proposition describes the matrices $P_{il}$, when $A$ is real, in terms of real eigenvalues and complex conjugate pairs of eigenvalues of $A$.
Proposition 10.
Assume $A\in\mathbb{R}^{n\times n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
For $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$, we have
$P_{il}\in\mathbb{R}^{n\times n},\ l\in\{0,\ldots,m_{i}-1\}.$
For $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, we have
$P_{i_{2}l}=\overline{P_{i_{1}l}},\ l\in\{0,\ldots,m_{i_{2}}-1\}.$
Observe that $m_{i_{2}}=m_{i_{1}}$ holds by Proposition 3.
Proof.
We prove that, for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, we have
$P_{i_{2}l}=\overline{P_{i_{1}l}},\ l\in\{0,\ldots,m_{i_{2}}-1\}.$
This includes the case $i_{1}=i_{2}$ and $\lambda_{i_{1}}\in\mathbb{R}$, for which we can conclude that $P_{i_{1}l}$, $l\in\{0,\ldots,m_{i_{1}}-1\}$, is a real matrix.
Since (see Proposition 3)
$d_{i_{2}}=d_{i_{1}}\text{\ and\ }m_{i_{2}j}=m_{i_{1}j},\ j\in\{1,\ldots,d_{i_{1}}\},\text{\ and\ }m_{i_{2}}=m_{i_{1}}$
we obtain, for $l\in\{0,\ldots,m_{i_{2}}-1\}$,
$N^{(i_{2},l)}=N^{(i_{1},l)}$
and then (see Proposition 5)
$P_{i_{2}l}=V^{\left(i_{2}\right)}N^{(i_{2},l)}W^{\left(i_{2}\right)}=\overline{V^{\left(i_{1}\right)}}N^{(i_{1},l)}\overline{W^{\left(i_{1}\right)}}=\overline{V^{\left(i_{1}\right)}N^{(i_{1},l)}W^{\left(i_{1}\right)}}=\overline{P_{i_{1}l}}.$
2.4.4. The kernel of the matrices $P_{il}$
Next three propositions concern the kernel of the matrices $P_{il}$. The first proposition describes this kernel.
Proposition 11.
For $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$, we have
$\mathrm{ker}\left(P_{il}\right)=\left\{u\in\mathbb{C}^{n}:\alpha_{ijk}(u)=0\ \text{for all }j\in\{1,\ldots,d_{i}\}\ \text{with $m_{ij}\geq l+1$}\right.$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.\ \text{and for all }k\in\{l+1,\ldots,m_{ij}\}\right\}.$
In other words, with reference to the previous Figure 2, the proposition says that the kernel of $P_{il}$ is constituted by the vectors $u\in\mathbb{C}^{n}$ with zero component along all the generalized eigenvectors $v^{(i,j,k)}$ whose third index $k$ (in the three-index notation) is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
Proof.
For $u\in\mathbb{C}^{n}$, we have
$P_{il}u=0$$\Leftrightarrow$ $V^{\left(i\right)}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
$\Leftrightarrow$ $N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
by the linear independence of the vectors $v^{(i,j,k)}$ in (8)
$\Leftrightarrow$ $N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0\ \text{for\ all\ }j=1,\ldots,d_{i},$
$\Leftrightarrow$ $\alpha_{ijk}\left(u\right)=0\ \text{for\ all }j\in\{1,\ldots,d_{i}\}\ \text{with $m_{ij}\geq l+1$}$
 and for all $\text{\ and\ for all\ }k\in\{l+1,\ldots,m_{ij}\}.$
Next two propositions are immediate consequences of the previous proposition.
Proposition 12.
For $i\in\{1,\ldots,p\}$ and $l_{1},l_{2}\in\{0,\ldots,m_{i}-1\}$, with $l_{1}<l_{2}$, we have
$\mathrm{ker}\left(P_{il_{1}}\right)\subseteq\mathrm{ker}\left(P_{il_{2}}\right).$
Proposition 13.
For $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$, we have
$u\in\mathrm{ker}\left(P_{i0}\right)\ \ \Leftrightarrow\ \ \alpha^{\left(i\right)}\left(u\right)=0.$
The Proposition 13 confirms our previous observation in Subsection 2.4 that $P_{i0}$ is the projection on the generalized eigenspace corresponding to $\lambda_{i}$.
2.5. The index $l_{i}(u)$
In view of the Propositions 12 and 13, for $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{(i)}(u)\neq 0$, we define the index
$l_{i}\left(u\right):=\max\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}$
(see Figure 4).
Indeed, if $\alpha^{(i)}(u)=0$, then
$\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}=\emptyset$
by Propositions 12 and 13. In this case, the index $l_{i}(u)$ cannot be defined. On the other hand, if $\alpha^{(i)}(u)\neq 0$, then
$\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}\neq\emptyset,$
since $u\notin\mathrm{ker}(P_{i0})$ by Proposition 13. In this case, the index $l_{i}(u)$ can be defined. By Proposition 12, we have
$u\notin\mathrm{ker}(P_{il}),\ l\in\{0,\ldots,l_{i}(u)\},$
and
$u\in\mathrm{ker}(P_{il}),\ l\in\{l_{i}(u)+1,\ldots,m_{i}\}.$
Observe that if $\lambda_{i}$ is a non-defective eigenvalue, i.e. $m_{i}=0$, then $l_{i}\left(u\right)=0$.
Figure 4. The index $l_{i}(u)$. $\mathrm{ker}(P_{il})^{c}$ denotes the complementary set of $\mathrm{ker}(P_{il})$. We have $\mathrm{ker}\left(P_{i0}\right)^{c}\supseteq\mathrm{ker}\left(P_{i1}\right)^{c}\supseteq\mathrm{ker}\left(P_{i2}\right)^{c}\supseteq\mathrm{ker}\left(P_{i3}\right)^{c}\supseteq\cdots$. In this case $l_{i}(u)=2$, since $u\in\mathrm{ker}\left(P_{i2}\right)^{c}$ but $u\notin\mathrm{ker}\left(P_{i3}\right)^{c}$.
The following proposition gives $l_{i}\left(u\right)$ in terms of the zero components of $\alpha^{\left(i\right)}\left(u\right)$.
Proposition 14.
For $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{\left(i\right)}\left(u\right)\neq 0$, we have
$l_{i}\left(u\right)+1$
$=\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\}.$
In other words, the proposition says that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
$(v^{(i,j,k)})_{k=1,\ldots,m_{ij}},\ j\in\{1,\ldots,d_{i}\},$
corresponding to the eigenvalue $\lambda_{i}$, such that $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$. Alternatively, with reference to Figure 2, the proposition says that $l_{i}(u)$ is the maximum index $l$ such that $u$ has a non-zero component along some generalized eigenvector $v^{(i,j,k)}$ whose third index $k$ is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
Proof.
The proof is straightforward by recalling the interpretation of the statement of Proposition 14 from a few lines above, referencing Figure 2, and the observation about $\mathrm{ker}(P_{il})$, also referencing Figure 2, immediately following Proposition 11. However, a more formal proof is now presented below.
Suppose $l_{i}\left(u\right)=m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{im_{i}-1}\right)$. By Proposition 11, there exists $\overline{j}\in\{1,\ldots,d_{i}\}$ such that $m_{i\overline{j}}=m_{i}$ and $\alpha_{i\overline{j}m_{i}}(u)\neq 0$. Thus, we have
$m_{i}$
$=\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\},$
i.e. (31) holds.
Suppose $l_{i}\left(u\right)<m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$ and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$. By Proposition 11 and $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$, there exist $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}(u)+1$ and $\overline{k}\in\{l_{i}\left(u\right)+1,\ldots,m_{i\overline{j}}\}$ such that $\alpha_{i\overline{j}\ \overline{k}}\neq 0$. Then
$l_{i}(u)+1$
$\leq\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\}.$
On the other hand, by Proposition 11 and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$, we have $\alpha_{ijk}\left(u\right)=0$for all $j\in\{1,\ldots,d_{i}\}$ with $m_{ij}\geq l_{i}(u)+2$ and for all $k\in\{l_{i}\left(u\right)+2,\ldots,m_{ij}\}$. Then
$l_{i}(u)+1$
$\geq\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\}.$
Thus, (31) holds. ∎
2.6. Linear independence
This subsection deals with the linear independence of the vectors $P_{il}u$ and the matrices $P_{il}$.
Proposition 15.
For $u\in\mathbb{C}^{n}$, the vectors
$P_{il}u,\ i\in\{1,\ldots,p\}\text{\ such that $\alpha^{(i)}(u)\neq 0$ and\ }l\in\{0,\ldots,l_{i}\left(u\right)\},$ (32)
are linearly independent in the vector space $\mathbb{C}^{n}$.
Proof.
Consider a zero linear combination of the vectors (32):
$0=\sum\limits_{\begin{subarray}{c}i=1\\ \alpha^{(i)}(u)\neq 0\end{subarray}}^{p}\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}P_{il}u=\sum\limits_{\begin{subarray}{c}i=1\\ \alpha^{(i)}(u)\neq 0\end{subarray}}^{p}V^{\left(i\right)}\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right),$ (33)
where the second inequality follows by (23). Since $V$ has linearly independent columns, we obtain, for $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$,
00 0 $=$ $\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\alpha^{\left(i\right)}\left(u\right)$
$=$ $\left(\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,1)}\alpha^{\left(i,1\right)}\left(u\right),\ldots,\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,d_{i})}\alpha^{\left(i,d_{i}\right)}\left(u\right)\right)$
and then, for $j\in\{1,\ldots,d_{i}\}$,
$\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0.$
Now, fix $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$. By Proposition 14, there exists $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}\left(u\right)+1$ such that $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$ and $\alpha_{i\overline{j}k}(u)=0$ for $k\in\{l_{i}(u)+2,\ldots,m_{i\overline{j}}\}$. Thus
$0=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,\overline{j})}\alpha^{\left(i,\overline{j}\right)}\left(u\right)$
$=\left[\begin{array}[]{ccccccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}&|&0&\cdot&0\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&|&c_{il_{i}\left(u\right)}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{il_{i}\left(u\right)-1}&\cdot&0\\ \cdot&\cdot&\cdot&c_{i0}&c_{i1}&|&\cdot&\cdot&c_{il_{i}\left(u\right)}\\ \cdot&\cdot&\cdot&\cdot&c_{i0}&|&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ -&-&-&-&-&&-&-&-\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{i0}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&\cdot&\cdot&c_{i1}\\ 0&\cdot&\cdot&\cdot&\cdot&|&\cdot&0&c_{i0}\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}l_{i}\left(u\right)+1}\left(u\right)\\ -\\ 0\\ \cdot\\ 0\end{array}\right]$
and then
$0=\left[\begin{array}[]{ccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 0&0&\cdot&c_{i0}&c_{i1}\\ 0&0&\cdot&0&c_{i0}\\ &&&&\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\\ \end{array}\right].$
Now, since $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$, we obtain $c_{il}=0$ successively for $l=0,1,\ldots,l_{i}\left(u\right)$.
We have proved that the coefficients of the zero linear combination (33) are all zero. ∎
Proposition 16.
The matrices
$P_{il},\ i\in\{1,\ldots,p\}\text{\ and \ }l\in\{0,\ldots,m_{i}-1\},$ (36)
are linearly independent in the vector space $\mathbb{C}^{n\times n}$.
Proof.
Consider a zero linear combination of the matrices (36):
$0=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{m_{i}-1}c_{il}P_{il}.$
Consider $u\in\mathbb{C}^{n}$ such that $\alpha_{ijk}(u)\neq 0$ for $i\in\{1,\ldots,p\}$, $j\in\{1,\ldots,d_{i}\}$ and $k\in\{1,\ldots,m_{ij}\}$. For example, one can take $u=Ve$, where $e=(1,\ldots,1)$, and then $\alpha(u)=e$. We have $l_{i}(u)=m_{i}-1$ for $i\in\{1,\ldots,p\}$. Since
$0=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{m_{i}-1}c_{il}P_{il}u=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{l_{i}(u)}c_{il}P_{il}u$
and $P_{il}u$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,l_{i}(u)\}$, are linearly independent (see previous Proposition 15), we have $c_{il}=0$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$. ∎
2.6.1. Linear combinations depending on $t$
In the present paper, we deal with linear combinations of some of the vectors $P_{il}u$ in (32) or some of the matrices $P_{il}$ in (36), whose coefficients depend on time $t$ and they are not all zero. For a fixed $t$, such a linear combination is not zero since the vectors and the matrices are linearly independent (recall the two previous Propositions 15 and 16).
However, in the analysis of the asymptotic behavior of the condition numbers, we require that the norm of the linear combination is away from zero, uniformly with respect to $t$. This is the content of the next lemma.
Lemma 17.
[…]
$\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}$
of $a_{1},\ldots,a_{K}$.
If $a_{1},\ldots,a_{K}$ are linearly independent and
$\inf_{t\in I}\|f\left(t\right)\|_{\infty}>0,$ (37)
then
$\inf_{t\in I}\left\|\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}\right\|>0.$
Proof.
Suppose $a_{1},\ldots,a_{K}$ linearly independent and (37). Let
$C=\left\{z\in\mathbb{C}^{K}:\left\|z\right\|_{\infty}=1\right\}.$
Consider the function $g:\mathbb{C}^{K}\rightarrow\mathbb{R}$ given by
$g\left(z\right)=\left\|\sum\limits_{k=1}^{K}z_{k}a_{k}\right\|,\ z\in\mathbb{C}^{K}.$
Since $C$ is a compact subset of $\mathbb{C}^{K}$ and $g$ is a continuous function, the extreme value theorem says that
$m=\inf_{z\in C}g\left(z\right)=g\left(\overline{z}\right)$
for some $\overline{z}\in C$. Since $a_{1},\ldots,a_{K}$ are linearly independent, we have $m>0$.
Now, let $t\in I$. We have
$\frac{f(t)}{\|f(t)\|_{\infty}}=\left(\frac{f_{1}\left(t\right)}{\|f(t)\|_{\infty}},\ldots,\frac{f_{K}\left(t\right)}{\|f(t)\|_{\infty}}\right)\in C,$
which implies
$g\left(\frac{f(t)}{\|f(t)\|_{\infty}}\right)=\left\|\sum\limits_{k=1}^{K}\frac{f_{k}\left(t\right)}{\|f(t)\|_{\infty}}a_{k}\right\|\geq m$
and then
$\left\|\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}\right\|\geq m\|f(t)\|_{\infty}.$
We conclude that
$\inf_{t\in I}\left\|\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}\right\|\geq m\inf_{t\in I}\|f(t)\|_{\infty}>0.$
In the paper, we apply this lemma to the case where $a_{1},\ldots,a_{K}$ are some of the linearly independent vectors $P_{il}u$ in (32), or some of the linearly independent matrices $P_{il}$ in (36), and $f_{1}(t),\ldots,f_{K}(t)$ are of the form $\mathrm{e}^{\sqrt{-1}\omega t}$, with $\sqrt{-1}$ the imaginary unit and $\omega\in\mathbb{R}$, and then $\|f(t)\|_{\infty}=1$.
3. Asymptotic forms
In the next Section 4, we analyze the asymptotic (long-time) behavior of the three condition numbers (3), (4) and (5), i.e. what happens to (3), (4) and (5) as $t$ approaches infinity (becomes large).
This analysis determines the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (3), of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (4) and of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ in (5). Then, the asymptotic behavior of the condition numbers is described by substituting in (3), (4) and (5) these asymptotic forms and by defining asymptotic condition numbers the new expressions obtained by these substitutions.
All of this might seem straightforward, but from a mathematical perspective it is not a simple task. Specifically:
  • It is necessary to precisely define what we mean by asymptotic form and how to determine it. This is particularly complicated in the non-diagonalizable case, where generalized eigenvectors are involved.
  • We also want to quantify how dominant the asymptotic form is at finite times.
  • The core of our analysis is the qualitative study of the asymptotic condition numbers. Understanding them requires significant mathematical effort in case of complex eigenvalues.
  • If we are interested in defining asymptotic condition numbers for the problem (2), the asymptotic condition numbers derived from (4) and (5) as $t\rightarrow+\infty$ may not be appropriate, since they represent the asymptotic worst cases of (3) and (4), respectively. Instead, we are concerned with the worst asymptotic cases of (3) and (4) by varying $\widehat{z}_{0}$ and $y_{0}$. In other words: do “asymptotic” and “worst” commute?
In this section, we determine the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, where $u$ is a vector. In order to define them some preliminary notions and results need to be introduced.
3.1. Notations $\sim$ and $\approx$
In this subsection, we precise what we mean for asymptotic form.
Let $f(t)$ and $g(t)$ be scalar, vector or matrix functions of $t\in\mathbb{R}$. For $g$ such that $g(t)\neq 0$ for $t$ in a neighborhood of +$\infty$+ ∞, we write
$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$ (38)
when
$\lim\limits_{t\rightarrow+\infty}\frac{\|f(t)-g(t)\|}{\|g(t)\|}=0.$
[…]
We interpret (38) as indicating that $g$ is an asymptotic form of $f$.
Observe that (38) means that the relative error of $f$ with respect to its asymptotic form $g$ asymptotically vanishes. At a finite time, one may ask how dominant the asymptotic form is, i.e., how large the relative error of $f$ with respect to $g$ is. Therefore, we introduce the following notation. For $t\in\mathbb{R}$ such that $g(t)\neq 0$ and $\epsilon\geq 0$, we write
$f\left(t\right)\approx g\left(t\right)\textit{\ with precision\ }\epsilon$ (39)
when
$\frac{\|f\left(t\right)-g\left(t\right)\|}{\|g\left(t\right)\|}\leq\epsilon.$
Observe that (39) means that the relative error of $f$ with respect to $g$ at the time $t$ is not larger than $\epsilon$. The notation (39) serves to quantify how much dominant the asymptotic form $g$ of $f$ is at the finite time $t$.
Remark 18.
Note that, for vector or matrix functions $f(t)$ and $g(t)$,
$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$
implies
$\|f\left(t\right)\|\sim\|g\left(t\right)\|,\ t\rightarrow+\infty,$
and, for $\epsilon\geq 0$,
$f\left(t\right)\approx g\left(t\right)\text{\ with precision\ }\epsilon$
implies
$\|f\left(t\right)\|\approx\|g\left(t\right)\|\text{\ with precision\ }\epsilon.$
This follows by
$\left|\left\|f(t)\right\|-\left\|g(t)\right\|\right|\leq\left\|f(t)-g(t)\right\|.$
3.2. Partition of the spectrum and formula for the matrix exponential
In this subsection, we introduce the tool for determining the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, $u\in\mathbb{C}^{n}$.
The spectrum $\Lambda=\left\{\lambda_{1},\ldots,\lambda_{p}\right\}$ of $A$, where $\lambda_{1},\ldots,\lambda_{p}$ are the distinct eigenvalues of $A$, is partitioned by decreasing real parts (see Figure 5) in the subsets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, given by
$\Lambda_{j}:=\{\lambda_{i_{j-1}+1},\lambda_{i_{j-1}+2},\ldots,\lambda_{i_{j}}\}$
$\mathrm{Re}\left(\lambda_{i_{j-1}+1}\right)=\mathrm{Re}\left(\lambda_{i_{j-1}+2}\right)=\cdots=\mathrm{Re}\left(\lambda_{i_{j}}\right)=r_{j},$
where the $q$ distinct real parts $r_{j}$, $j\in\{1,\ldots,q\}$, of the eigenvalues of $A$ satisfy $r_{1}>r_{2}>\cdots>r_{q}$.
Figure 5. Partition of the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ by decreasing real part. The eigenvalues are marked by “x”.
Now, we recall the formula (26) for the matrix exponential $\mathrm{e}^{tA}$:
$\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$
where $m_{i}$ is the ascent of the eigenvalue $\lambda_{i}$ (maximum dimension of the mini-blocks corresponding to $\lambda_{i}$ in the JCF of $A$) and the matrices $P_{il}$ are defined and studied in Section 2. By using the partition of the spectrum given above, we have
$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$ (40)
where $\sqrt{-1}$ denotes the imaginary unit and $\omega_{i}$ denotes the imaginary part of the eigenvalue $\lambda_{i}$.
Now, by exchanging the two inner sums we can rewrite (40) as
$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t),$ (41)
where
$L_{j}:=\max_{\lambda_{i}\in\Lambda_{j}}m_{i}-1,\ j\in\{1,\ldots,q\},$
and
$Q_{jl}(t):=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il},\ j\in\{1,\ldots,q\}\text{\ and\ }l\in\{0,\ldots,L_{j}\}.$ (42)
The formula (41) is used to determine the asymptotic forms in (3), (4) and (5).
3.3. Notations for the asymptotic form of $\mathrm{e}^{tA}u$
The following notations are crucial for determining the asymptotic form of $\mathrm{e}^{tA}u$.
For $j\in\{1,\ldots,q\}$ and $u\in\mathbb{C}^{n}$:
  • Let
    $\Lambda_{j}(u):=\{\lambda_{i}\in\Lambda_{j}:\alpha^{(i)}(u)\neq 0\},$
    where $\alpha^{(i)}(u)$, defined in Section 2, is the vector of the Jordan basis components of $u$ along the generalized eigenvectors corresponding to $\lambda_{i}$. In other words, $\Lambda_{j}(u)$ is obtained from $\Lambda_{j}$ by including only the eigenvalues for which $u$ has a non-zero projection onto the corresponding generalized eigenspace.
  • In the case where $\Lambda_{j}(u)\neq\emptyset$, let
    $L_{j}(u):=\max\{l_{i}(u):\lambda_{i}\in\Lambda_{j}(u)\},$
    where
    $l_{i}(u)=\max\{l\in\{0,\ldots,m_{i}-1\}:P_{il}u\neq 0\}$
    is defined and studied in Section 2, and it is such that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
    $(v^{(i,j,k)})_{k=1,\ldots,m{ij}},\ j\in\{1,\ldots,d_{j}\},$
    corresponding to the eigenvalue $\lambda_{i}$, for which $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$. In other words, $L_{j}(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j}$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain.
3.3.1. Indeces of dominance
For $u\in\mathbb{C}^{n}\setminus\{0\}$:
  • Let
    $j(u):=\min\{j\in\left\{1,\ldots,q\right\}:\Lambda_{j}(u)\neq\emptyset\}.$
    In other words, $r_{j(u)}$ is the rightmost real part of an eigenvalue of $A$ such that $u$ has a non-zero projection on the corresponding generalized eigenspace. We call $j(u)$ the primary index of dominance of $u$.
  • Let
    $L(u):=L_{j(u)}(u).$
    In other words, $L(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j(u)}(u)$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain. We call $L(u)$ the secondary index of dominance of $u$.
Similarly to Section 2, the next two subsections, 3.4 and 3.5, can be skipped and referred back to when necessary.
3.4. Properties of the matrices $Q_{jl}(t)$
In our study of asymptotic forms and asymptotic condition numbers, the following properties of the matrices $Q_{jl}(t)$ in (42) are fundamental. We collect these properties in four propositions.
The first proposition relates the indices $j(u)$ and $L_{j}(u)$ to the condition $Q_{jl}(t)u=0$.
Proposition 19.
Let $u\in\mathbb{C}^{n}$. We have:
  • 1)
    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)=\emptyset$ and $l\in\{0,\ldots,L_{j}\}$;
  • 2)
    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,j(u)-1\}$ and $l\in\{0,\ldots,L_{j}\}$;
  • 3)
    $Q_{jl}(t)u=0$ for $j\in\{j(u),\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
In 2) and 3), we assume $u\neq 0$ in order to have $j(u)$ defined.
Proof.
For $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{j}\}$, we have
$Q_{jl}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u$
by recalling (42).
Proof of point 1). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)=\emptyset$ and $l\in\{0,\ldots,L_{j}\}$. $\Lambda_{j}(u)=\emptyset$ implies, for any $\lambda_{i}\in\Lambda_{j}$, $\alpha^{(i)}(u)=0$ and then, by Propositions 12 and 13, for any $\lambda_{i}\in\Lambda_{j}$ we have $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$. In particular, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, we have $P_{il}u=0$.
Point 2) follows by point 1).
Proof of point 3). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$. For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$. For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)\neq 0$, we also have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{l_{i}(u)+1,\ldots,m_{i}-1\}$ by the definition of $l_{i}(u)$, and $l\geq L_{j}(u)+1\geq l_{i}(u)+1$. ∎
The second proposition says that the matrices $Q_{jl}(t)$ and their actions on vectors remain bounded and away from zero, by varying $t$.
Proposition 20.
Let $j\in\{1,\ldots,q\}$, $l\in\{0,\ldots,L_{j}\}$ and $u\in\mathbb{C}^{n}$. We have:
$\sup\limits_{t\in\mathbb{R}}\left\|Q_{jl}(t)\right\|<+\infty$;
$\sup\limits_{t\in\mathbb{R}}\left\|Q_{jl}(t)u\right\|<+\infty$;
$\inf\limits_{t\in\mathbb{R}}\left\|Q_{jl}(t)\right\|>0$;
$\inf\limits_{t\in\mathbb{R}}\left\|Q_{jl}(t)u\right\|>0$ if $\Lambda_{j}(u)\neq\emptyset$ and $l\leq L_{j}(u)$.
Proof.
Since (see (42))
$\left\|Q_{jl}(t)\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}\right\|\text{\ \ and\ \ }\left\|Q_{jl}(t)u\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}u\right\|$
hold, the points 1) and 2) follow.
Proposition 16 and Lemma 17, as applied to the linear combination
$Q_{jl}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il},$
imply 3).
Finally, suppose $\Lambda_{j}(u)\neq\emptyset$ and $l\leq L_{j}(u)$. We obtain
$Q_{jl}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}(u)\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}(u)\\ m_{i}\geq l+1\\ l_{i}(u)\geq l\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u,$
where the second equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$ we have $P_{il}u=0$; and the third equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, $\alpha^{(i)}(u)\neq 0$ and $l_{i}(u)<l$ we have $P_{il}u=0$ by the definition of $l_{i}(u)$. Now, Proposition 15 and Lemma 17 imply 4). ∎
The third proposition says how the matrices $Q_{jl}(t)$ are transformed when we replace the matrix $A$ by $-A$. Observe that the matrix $-A$ has opposite eigenvalues, i.e. eigenvalues with opposite imaginary and real parts, with respect to the matrix $A$ and the dimensions of blocks and mini-blocks in the JCF of $-A$ are the same as in the JCF of $A$ (see Proposition 2 with $z=-1$). Therefore, we have the same number $q$ of different real parts for the eigenvalues of $-A$ and $A$. Moreover, the set $\Lambda_{j}(-A)$ and the numbers $r_{j}(-A)$ and $L_{j}(-A)$, $j\in\{1,\ldots,q\}$, corresponding to $-A$ are
$\Lambda_{j}(-A)=-\Lambda_{q+1-j},\ r_{j}(-A)=-r_{q+1-j}\text{\ \ and\ \ }L_{j}(-A)=L_{q+1-j},$ (43)
where $\Lambda_{q+1-j}$, $r_{q+1-j}$ and $L_{q+1-j}$ correspond to $A$. The indeces $j$ and $l$ for the matrices $Q_{jl}(t,-A)$ corresponding to $-A$ range over $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, respectively.
Proposition 21.
We have
$Q_{jl}(t,-A)=(-1)^{l}Q_{q+1-j\ l}(-t),\ j\in\{1,\ldots,q\}\text{\ and\ }l\in\{0,\ldots,L_{q+1-j}\}.$
Proof.
The matrix $Q_{jl}(t,-A)$, $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, is given by
$Q_{jl}(t,-A)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{q+1-j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ (-\omega_{i})t}P_{il}(-A).$
Now, use Proposition 9 with $z=-1$. ∎
In the case of a real matrix $A$, the fourth proposition explains how to rewrite the expression (42) that defines $Q_{jl}(t)$ in terms of the real eigenvalues and complex conjugate pairs of eigenvalues of $A$. For a matrix $Z$, we denote by $\mathrm{Re}(Z)$ the matrix given by the real parts of the elements of $Z$.
Proposition 22.
Assume $A\in\mathbb{R}^{n\times n}$. For $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{j}\}$, we have
$Q_{jl}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ \lambda_{i}\mathrm{\ is\ real}\\ m_{i}\geq l+1\end{subarray}}P_{il}+2\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ \omega_{i}>0\\ m_{i}\geq l+1\end{subarray}}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}\right).$ (44)
Proof.
For $\lambda_{i}\in\Lambda_{j}$ such that $\lambda_{i}$ is real, we have $\omega_{i}=0$. For a complex conjugate pairs $\lambda_{i_{1}},\lambda_{i_{2}}\in\Lambda_{j}$ with $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ and $\omega_{i_{1}}>0$, by Proposition 10 we have
$\mathrm{e}^{\sqrt{-1}\omega_{i_{2}}t}P_{i_{2}l}$ $=$ $\mathrm{e}^{-\sqrt{-1}\ \omega_{i_{1}}t}\overline{P_{i_{1}l}}=\overline{\mathrm{e}^{\sqrt{-1}\ \omega_{i_{1}}t}P_{i_{1}l}}$
and then
$\mathrm{e}^{\sqrt{-1}\omega_{i_{1}}t}P_{i_{1}l}+\mathrm{e}^{\sqrt{-1}\ \omega_{i_{2}}t}P_{i_{2}l}=2\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{i_{1}}t}P_{i_{1}l}\right).$
Remark 23.
In (44) the sum
$\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ \lambda_{i}\text{ is real}\\ m_{i}\geq l+1\end{subarray}}P_{il}$
has zero or one term $P_{il}$, which is real (see Proposition 10). Moreover, in the other sum, each term
$\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{i}t}P_{il}\right)=\cos\omega_{i}t\cdot\mathrm{Re}(P_{il})-\sin\omega_{i}t\cdot\mathrm{Im}(P_{il})$
is a periodic function of $t$ of period $\frac{2\pi}{\omega_{i}}$.
3.5. The matrices $Q^{e}_{j}(t)$
The contents of this section are used for defining in Subsection 4.6 the global asymptotic condition number of the problem (2).
For $j\in\{1,\ldots,q\}$, we introduce
$Q^{e}_{j}(t):=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{i}^{e},$
where
$P_{i}^{e}:=V^{(i)}_{e}W^{(i)}_{e},$
is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, with
$V^{(i)}_{e}:=\left[v^{(i,1,1)}\ \cdots\ v^{(i,d_{i},1)}\right]\in\mathbb{C}^{n\times d_{i}}\text{\ \ and\ \ }W^{(i)}_{e}:=\left[\begin{array}[]{c}w^{(i,1,1)}\\ \vdots\\ w^{(i,d_{i},1)}\end{array}\right]\in\mathbb{C}^{d_{i}\times n}$
the matrices of the (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$ (see Section 2). The superscript or subscript $e$ in the notation denotes “eigen”.
Let $U^{e}_{j}$ be the sum of the eigenspaces corresponding to eigenvalues in $\Lambda_{j}$. Observe that the linear operator $Q^{e}_{j}(t)$ maps $\mathbb{C}^{n}$ in the subspace $U^{e}_{j}$. Therefore, $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$. The next proposition states that the restriction $Q^{e}_{j}(t)|_{U^{e}_{j}}$ is invertible.
Proposition 24.
Let $j\in\{1,\ldots,q\}$. The linear operator $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$ is invertible and the inverse is
$\left(Q^{e}_{j}(t)|_{U^{e}_{j}}\right)^{-1}=Q^{e}_{j}(-t)|_{U^{e}_{j}}.$
Proof.
We have
$Q^{e}_{j}(-t)Q^{e}_{j}(t)$ $=$ $\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{-\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)$
$=$ $\sum\limits_{\lambda_{i}\in\Lambda_{j}}P^{e}_{i}$
and then
$Q_{j}^{e}(-t)|_{U_{j}^{e}}Q_{j}^{e}(t)|_{U_{j}^{e}}=I_{U_{j}^{e}}.$
Remark 25.
  • 1.
    As a consequence of the previous proposition, we have
    $\min\limits_{\begin{subarray}{c}\widehat{u}\in U^{e}_{j}\\ \left\|\widehat{u}\right\|=1\end{subarray}}\left\|Q^{e}_{j}(t)\widehat{u}\right\|=\frac{1}{\left\|Q^{e}_{j}(-t)|_{U^{e}_{j}}\right\|}.$
  • 2.
    In the case where $\Lambda_{j}$ consists of a real eigenvalue $\lambda_{i}$, we have
    $Q^{e}_{j}(t)|_{U^{e}_{j}}=P_{i}^{e}|_{U^{e}_{j}}=I_{U^{e}_{j}}.$
Next proposition considers the linear operators $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$, $\lambda_{i}\in\Lambda_{j}$.
Proposition 26.
Let $j\in\{1,\ldots,q\}$. The linear operators
$P_{i}^{e}|_{U^{e}_{j}},\ \lambda_{i}\in\Lambda_{j},$ (45)
are linearly independent in the vector space of the linear operators $U^{e}_{j}\rightarrow U^{e}_{j}$.
The proof of this proposition is similar to the proof of Proposition 16.
Proof.
Consider a zero linear combination of the linear operators (45):
$0=\sum\limits_{i=1}^{p}c_{i}P_{i}^{e}|_{U^{e}_{j}}.$
Consider $u\in U^{e}_{j}$ such that $\alpha_{ij^{\prime}1}(u)\neq 0$ for $\lambda_{i}\in\Lambda_{j}$ and $j^{\prime}\in\{1,\ldots,d_{i}\}$. Since $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$. Thus $l_{i}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, by Proposition 14. Since (recall Subsection 2.3.1)
$P_{i}^{e}u=P_{i0}u,\ \lambda_{i}\in\Lambda_{j},$
we have the zero linear combination
$0=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i}^{e}u=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i0}u=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\sum\limits_{l=0}^{l_{i}(u)=0}c_{il}P_{il}u,$
where we set $c_{i0}=c_{i}$, $i\in\{1,\ldots,p\}$. By using Proposition 15, we obtain $c_{i}=c_{i0}=0$, $\lambda_{i}\in\Lambda_{j}$. ∎
Remark 27.
As a consequence of the previous proposition and Lemma 17, we obtain
$\inf\limits_{t\in\mathbb{R}}\left\|Q^{e}_{j}(t)|_{U^{e}_{j}}\right\|>0.$
3.6. The asymptotic form of $\mathrm{e}^{tA}$
By looking at the formula (41), we immediately identify
$\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$ (46)
as the asymptotic form of $\mathrm{e}^{tA}$, since it is the dominant term as $t\rightarrow+\infty$: it is the term with the highest order of infinite. Next proposition precises that (46) is indeed an asymptotic form of $\mathrm{e}^{tA}$ in our definition (38) of asymptotic form and quantifies its dominance at a finite time.
Proposition 28.
We have
$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
with precision
$\epsilon(t):=\sum\limits_{l=0}^{L_{1}-1}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{1l}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}+\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\sum\limits_{l=0}^{L_{j}}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{jl}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}.$
Moreover, we have
$\epsilon(t)\rightarrow 0,\ t\rightarrow+\infty,$ (47)
and then
$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t),\ t\rightarrow+\infty.$
Proof.
By (41), we can write
$\mathrm{e}^{tA}$ $=$ $\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)$
$=$ $\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)+\mathrm{e}^{r_{1}t}\sum\limits_{l=0}^{L_{1}-1}\frac{t^{l}}{l!}Q_{1l}(t)+\sum\limits_{j=2}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t).$
The first part of the proposition regarding $\approx$ follows. By 1) and 3) in Proposition 20, we obtain (47) and then the second part regarding $\sim$ follows. ∎
Remark 29.
If $A$ is diagonalizable, then we have
$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}Q_{10}(t)$
with precision
$\epsilon(t)=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{\left\|Q_{j0}(t)\right\|}{\left\|Q_{10}(t)\right\|}$
and
$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}Q_{10}(t),\ t\rightarrow+\infty,$
where
$Q_{j0}(t)=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{i0},\ j\in\{1,\ldots,q\},$
with $P_{i0}$ the projection onto the eigenspace corresponding to $\lambda_{i}$.
3.7. The asymptotic form of $\mathrm{e}^{tA}u$
Let $u\in\mathbb{C}^{n}\setminus\{0\}$. When we use the formula (41) for $\mathrm{e}^{tA}$ in $\mathrm{e}^{tA}u$, we obtain
$\mathrm{e}^{tA}u=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u.$
By 2) and 3) in Proposition 19 we identify
$\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$ (48)
as the asymptotic form of $\mathrm{e}^{tA}u$, since it is the dominant term as $t\rightarrow+\infty$. Next proposition precises that (48) is indeed an asymptotic form in our definition and quantifies its dominance.
Proposition 30.
For $u\in\mathbb{C}^{n}\setminus\{0\}$, we have
$\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$
with precision
$\epsilon(t,u)$ ::=:= $\sum\limits_{l=0}^{L\left(u\right)-1}\frac{L\left(u\right)!}{l!}t^{l-L\left(u\right)}\frac{\left\|Q_{j(u)l}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}$
$+\sum\limits_{\begin{subarray}{c}j=j\left(u\right)+1\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\sum\limits_{l=0}^{L_{j}(u)}\frac{L(u)!}{l!}t^{l-L(u)}\frac{\left\|Q_{jl}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}.$
Moreover, we have
$\epsilon(t,u)\rightarrow 0,\ t\rightarrow+\infty,$ (49)
and then
$\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u,\ t\rightarrow+\infty.$
Proof.
By (41) and Proposition 19, we can write
$\mathrm{e}^{tA}u$ $=$ $\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u=\sum\limits_{j=j(u)}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u$
$=$ $\sum\limits_{\begin{subarray}{c}j=j(u)\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}(u)}\frac{t^{l}}{l!}Q_{jl}(t)u$
$=$ $\mathrm{e}^{r_{j(u)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u+\mathrm{e}^{r_{j(u)}t}\sum\limits_{l=0}^{L(u)-1}\frac{t^{l}}{l!}Q_{j(u)l}(t)u$
$+\sum\limits_{\begin{subarray}{c}j=j(u)+1\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}(u)}\frac{t^{l}}{l!}Q_{jl}(t)u.$
The first part of the proposition regarding $\approx$ follows. By 2) and 4) in Proposition 20, we obtain (49) and then the second part regarding $\sim$ follows. ∎
Remark 31.
  • 1.
    The asymptotic form (48) of $\mathrm{e}^{tA}u$ is determined by the primary and secondary indices of dominance of $u$. In particular, the smaller the primary index of dominance $j(u)$, the higher the order of infinity of (48), and for a fixed primary index, the larger the secondary index of dominance $L(u)$, the higher the order of infinity of (48).
  • 2.
    The contents of Propositions 28 and 30 are deeper than what might be apparent at first glance. We are not simply stating the obvious facts that (46) and (48) are the dominant terms in $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, respectively. Rather, we are asserting that they represent the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ according to our definition (38). In other words, the relative errors of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ with respect to (46) and (48), respectively, approach zero asymptotically. Proving these deeper conclusions requires Propositions 15 and 16 and Lemma 17, which were developed through the work in Section 2.
  • 3.
    If $A$ is diagonalizable, then we have
    $\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u$
    with precision
    $\epsilon(t,u)=\sum\limits_{j=j\left(u\right)+1}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\frac{\left\|Q_{j0}(t)u\right\|}{\left\|Q_{j(u)0}(t)u\right\|}$
    and
    $\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u,\ t\rightarrow+\infty.$
4. The asymptotic behavior of the condition numbers
This section forms the core of our analysis. Building on the preparatory work of the previous two sections, we can now easily describe the asymptotic behavior of the condition numbers.
4.1. The asymptotic directional condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$
We set
$j^{\ast}:=j\left(\widehat{y}_{0}\right)=j\left(y_{0}\right)\text{\ \ and\ \ }L^{\ast}:=L\left(\widehat{y}_{0}\right)=L\left(y_{0}\right)$
as well as
$j^{\ast\ast}:=j\left(\widehat{z}_{0}\right)\text{\ \ and\ \ }L^{\ast\ast}:=L\left(\widehat{z}_{0}\right),$
i.e. $j^{\ast}$ and $j^{\ast\ast}$ are the primary indeces of dominance, and $L^{\ast}$ and $L^{\ast\ast}$ are the secondary indeces of dominance, of $y_{0}$ and $\widehat{z}_{0}$, respectively.
The next theorem describes the long-time behavior of the directional condition number $K\left(t,y_{0},\widehat{z}_{0}\right)$.
Theorem 32.
We have
$K\left(t,y_{0},\widehat{z}_{0}\right)\approx K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),$
where
$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right):=\frac{L^{\ast}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{j^{\ast}}\right)t}t^{L^{\ast\ast}-L^{\ast}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
with precision
$\frac{\epsilon(t,\widehat{z}_{0})+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})},$
whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t,u)$, $u=\widehat{z}_{0},\widehat{y}_{0}$, defined in Proposition 30). Moreover, we have
$K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\rightarrow+\infty.$
Proof.
Proposition 30 says what are the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (3) and how much they are dominant at a finite time $t$: we have
$\mathrm{e}^{tA}\widehat{z}_{0}\approx\mathrm{e}^{r_{j^{\ast\ast}}t}\frac{t^{L^{\ast\ast}}}{L^{\ast\ast}!}Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}$
with precision $\epsilon(t,\widehat{z}_{0})$ and
$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
with precision $\epsilon(t,\widehat{y}_{0})$. The first part of the theorem regarding $\approx$ follows by using Remark 18. The second part regarding $\sim$ follows since $\epsilon(t,\widehat{y}_{0})\rightarrow 0$ and $\epsilon(t,\widehat{z}_{0})\rightarrow 0$, $t\rightarrow+\infty$. ∎
We define the function
$t\rightarrow K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\in\mathbb{R},$
as the asymptotic directional condition number of the problem (2).
Remark 33.
  • 1.
    The asymptotic directional condition number of the problem (2):
    • is bounded and away from zero as $t$ varies (remind 2) and 4) in Proposition 20) if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}=L^{\ast\ast}$;
    • decays polynomially to zero, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}>L^{\ast\ast}$;
    • diverges polynomially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}<L^{\ast\ast}$;
    • decays exponentially to zero, as $t\rightarrow+\infty$, if $j^{\ast}<j^{\ast\ast}$;
    • diverges exponentially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}>j^{\ast\ast}$.
    Hence, whether the asymptotic condition number decreases to zero, diverges to infinity, or exhibits different behavior depends on which between $y_{0}$ and $\widehat{z}_{0}$ is more dominant. This is determined by which of $y_{0}$ and $\widehat{z}_{0}$ possesses the smaller primary index of dominance, or, in the case of equal primary indices, the larger secondary index of dominance.
  • 2.
    The generic case for $y_{0}$ and $\widehat{z}_{0}$ is $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$. In this generic case, we have
    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
4.2. The asymptotic condition number $K\left(t,y_{0}\right)$
The next theorem describes the long-time behavior of the condition number $K\left(t,y_{0}\right)$, worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$.
Theorem 34.
We have
$K\left(t,y_{0}\right)\approx K_{\infty}\left(t,y_{0}\right),$
where
$K_{\infty}\left(t,y_{0}\right):=\frac{L^{\ast}!}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{j^{\ast}}\right)t}t^{L_{1}-L^{\ast}}\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
with precision
$\frac{\epsilon(t)+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t)$ defined in Proposition 28 and $\epsilon(t,\widehat{y}_{0})$ defined in Proposition 30). Moreover, we have
$K\left(t,y_{0}\right)\sim K_{\infty}\left(t,y_{0}\right),\ t\rightarrow+\infty.$
Proof.
Proposition 28 says what is the asymptotic form of $\mathrm{e}^{tA}$ and how much it is dominant: we have
$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
with precision $\epsilon(t)$. Proposition 30 says what is the asymptotic form of $\mathrm{e}^{tA}\widehat{y}_{0}$ and how much it is dominant: we have
$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
with precision $\epsilon(t,\widehat{y}_{0})$. The first part of the theorem regarding $\approx$ follows by using Remark 18. The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,\widehat{y}_{0})\rightarrow 0$, $t\rightarrow+\infty$. ∎
We define the function
$t\rightarrow K_{\infty}(t,y_{0}),\ t\in\mathbb{R},$
as the asymptotic condition number of the problem (2).
Remark 35.
  • 1.
    The asymptotic condition number of the problem (2):
    • is bounded and away from zero as $t$ varies if $j^{\ast}=1$ and $L^{\ast}=L_{1}$;
    • diverges polynomially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}=1$ and $L^{\ast}<L_{1}$;
    • diverges exponentially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}>1$.
    Therefore, the asymptotic condition number does not diverge to infinity if and only if $y_{0}$ is as dominant as possible, meaning that $y_{0}$ has the smallest possible primary index of dominance (i.e. $j^{\ast}=1$) and simultaneously the largest possible secondary index of dominance (i.e. $L^{\ast}=1$).
  • 2.
    The generic case for $y_{0}$ is $j^{\ast}=1$ and $L^{\ast}=L_{1}$. In this generic case, we have
    $K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
4.3. Is $K_{\infty}\left(t,y_{0}\right)$ the worst $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$?
By definition, the condition number $K\left(t,y_{0}\right)$ is the worst condition number $K\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$. Hence, an interesting question is the following. Does this fact hold asymptotically as $t\rightarrow+\infty$? Specifically, is the asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$ the worst asymptotic condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$? In other words, do
[…]
commute? The answer is YES and it is given by the next theorem, which considers the ratio
$\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=\frac{L_{1}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{1}\right)t}t^{L^{\ast\ast}-L_{1}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\right\|}$ (50)
independent of $y_{0}$.
Theorem 36.
We have
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=1.$ (51)
In particular:
  • a)
    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}<1$ or $j^{\ast\ast}=1$ and $L^{\ast\ast}<L_{1}$, we have
    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=0.$
  • b)
    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, we have
    $\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}\leq 1.$
  • c)
    There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a sequence $\{t_{m}\}_{m}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m},y_{0}\right)}=1.$
Proof.
Points a), b) and c) imply (51). Points a) and b) immediately follows by (50). Now, we prove c).
Consider a sequence $\{t_{k}\}_{k}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{z}_{0k}\}_{k}$ such that $\widehat{z}_{0k}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0k}\right\|=1$ and
$\left\|Q_{1L_{1}}(t_{k})\right\|=\left\|Q_{1L_{1}}(t_{k})\widehat{z}_{0k}\right\|.$
By the compactness of the unit sphere in $\mathbb{C}^{n}$, there exists a subsequence $\{\widehat{z}_{0k_{m}}\}_{m}$ of $\{\widehat{z}_{0k}\}_{k}$ converging to some $\widehat{z}_{0\infty}\in\mathbb{C}^{n}$ with $\left\|\widehat{z}_{0\infty}\right\|=1$.
By Proposition 19 (points 2) and 3)) and Proposition 20 (points 1) and 3)), we obtain $j\left(\widehat{z}_{0\infty}\right)=1$ and $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$. In fact, for any index $m$, we have
$\left|\ \left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|-\left\|Q_{1L_{1}}(t_{k_{m}})\right\|\ \right|\leq\sup\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|,$
where the right-hand side goes to zero as $m\rightarrow\infty$. Therefore, there exists an index $m$ such that
$\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|\geq\frac{1}{2}\inf\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|,$
where the right-hand side is positive. Since $Q_{1L_{1}}(t_{k_{m}})$ $\widehat{z}_{0\infty}\neq 0,$ we cannot have $j\left(\widehat{z}_{0\infty}\right)>1$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$. Hence, $j\left(\widehat{z}_{0\infty}\right)=1$. Once again since $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\neq 0,$ we cannot have $L_{1}\left(\widehat{z}_{0\infty}\right)<L_{1}$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$. Hence, $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$.
By using as direction of perturbation $\widehat{z}_{0\infty}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$. So
$\left|\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}-1\right|$ $=$ $\left|\frac{\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|}{\left\|Q_{1L_{1}}(t_{k_{m}})\right\|}-1\right|$
$\leq$ $\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|.$
We conclude that
$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}=1.$
Remark 37.
In the case where $\Lambda_{1}$ consists of a real eigenvalue, the point c) is modified to:
  • $c^{\prime}$)
    There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, such that
    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right).$ (52)
In fact, in this case $Q_{1L_{1}}(t)=Q_{1L_{1}}$ is independent of $t$ and so there exists $\widehat{z}_{0}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0}\right\|=1$, independent of $t$ such that
$\left\|Q_{1L_{1}}\right\|=\left\|Q_{1L_{1}}\widehat{z}_{0}\right\|.$
By using as direction of perturbation $\widehat{z}_{0}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ and then we obtain (52) by (50).
4.4. The asymptotic global condition number $K_{\infty}^{+}\left(t,\right)$
Next theorem describes the long-time behavior of the global condition number $K\left(t\right)$, worst condition number $K\left(t,y_{0}\right)$ by varying $y_{0}$.
Theorem 38.
We have
$K\left(t\right)\approx K_{\infty}^{+}(t),$
where
$K_{\infty}^{+}\left(t\right):=\frac{1}{L_{1}!L_{q}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}+L_{q}}\left\|Q_{1L_{1}}(t)\right\|\ \cdot\left\|Q_{qL_{q}}(-t)\right\|$ (53)
with precision
$\epsilon(t)+\epsilon(t,-A)+\epsilon(t)\epsilon(t,-A),$
where $\epsilon(t,-A)$ is $\epsilon(t)$ for the matrix $-A$. Moreover, we have
$K\left(t\right)\sim K^{+}_{\infty}(t),\ t\rightarrow+\infty.$
Proof.
Propositions 28 says what are the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ and how much they are dominant: we have
$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
with precision $\epsilon(t)$ and
$\mathrm{e}^{-tA}=\mathrm{e}^{t(-A)}\approx\mathrm{e}^{-r_{q}t}\frac{t^{L_{q}}}{L_{q}!}(-1)^{L_{q}}Q_{qL_{q}}(-t)$
with precision $\epsilon(t,-A)$. For the latter, remind (43) and Proposition 21. The first part of the theorem regarding $\approx$ follows by using Remark 18. The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,-A)\rightarrow 0$, $t\rightarrow+\infty$. ∎
We define the function
$t\rightarrow K^{+}_{\infty}(t),\ t\in\mathbb{R},$
as the asymptotic global condition number of the problem (2).
4.5. Is $K^{+}_{\infty}\left(t\right)$ the worst $K_{\infty}\left(t,y_{0}\right)$?
By definition, the condition number $K\left(t\right)$ is the worst condition number $K\left(t,y_{0}\right)$, by varying $y_{0}$. Hence, as in Subsection 4.3, an interesting question is the following. Does this fact hold asymptotically as $t\rightarrow+\infty$? Specifically, is the asymptotic condition number $K^{+}_{\infty}\left(t\right)$ the worst asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$? In other words, do
[…]
commute? Unlike Subsection 4.3, here the answer is NO and it is given by the next theorem, which considers
$K_{\infty}\left(t\right):=\frac{1}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}}\left\|Q_{1L_{1}}(t)\right\|\cdot\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|,$ (54)
where $Q^{e}_{q}(-t)|_{U^{e}_{q}}$ is defined in Subsection 3.5, and the ratio
$\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=L^{\ast}!\mathrm{e}^{\left(r_{q}-r_{j^{\ast}}\right)t}t^{-L^{\ast}}\frac{1}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}.$ (55)
Theorem 39.
We have
$\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=1.$ (56)
In particular:
  • a)
    For any initial value $y_{0}$ such that $j^{\ast}<q$ or $j^{\ast}=q$ and $L^{\ast}>0$, we have
    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=0.$
  • b)
    For any initial value $y_{0}$ such that $j^{\ast}=q$ and $L^{\ast}=0$, we have
    $\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}\leq 1.$
  • c)
    There exist an initial value $y_{0}$, with $j^{\ast}=q$ and $L^{\ast}=0$, and a sequence $\{t_{m}\}_{m}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0}\right)}{K_{\infty}\left(t_{m}\right)}=1.$
  • d)
    For the initial value $y_{0}$ at point c), there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}_{s}$ of the sequence $\{t_{m}\}_{m}$ at point c), such that
    $\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=1.$
Proof.
Points a), b) and c) imply (56). Point a) follows by (55) (remind point 4 in Proposition 20 and Remark 27).
Point b) follows by point 1 in Remark 25 and by observing that, for such $y_{0}$, we have $y_{0}\in U^{e}_{q}$ and then
$Q_{q0}(t)\widehat{y}_{0}=Q^{e}_{q}(t)\widehat{y}_{0}.$
For the point c), we proceed as in the proof of Theorem 36. Consider a sequence $\{t_{k}\}_{k}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{y}_{0k}\}_{k}$, where $\widehat{y}_{0k}\in U^{e}_{q}$ and $\left\|\widehat{y}_{0k}\right\|=1$, such that
$\left\|Q_{q0}(t_{k})\widehat{y}_{0k}\right\|=\left\|Q^{e}_{q}(t_{k})\widehat{y}_{0k}\right\|=\frac{1}{\left\|Q^{e}_{q}(-t_{k})|_{U^{e}_{q}}\right\|}$ (57)
(recall point 1 in Remark 25). There exists a subsequence $\{\widehat{y}_{0k_{m}}\}_{m}$ of $\{\widehat{y}_{0k}\}_{k}$ converging to some $\widehat{y}_{0\infty}\in U^{e}_{q}$ with $\left\|\widehat{y}_{0\infty}\right\|=1$. By using $\widehat{y}_{0\infty}$ as initial value, we have $j^{\ast}=q$ and $L^{\ast}=0$ (we have $l_{i}(\widehat{y}_{0\infty})=0$ for any $\lambda_{i}\in\Lambda_{q}$, see Proposition 14). Therefore,
$\left|\frac{K_{\infty}\left(t_{k_{m}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}}\right)}-1\right|$ $=$ $\left|\frac{1}{\left\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\right\|\left\|Q^{e}_{q}(-t_{k_{m}})|_{U^{e}_{q}}\right\|}-1\right|$
$=$ $\left|\frac{\|Q_{q0}(t_{k_{m}})\widehat{y}_{0k_{m}}\|}{\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\|}-1\right|$
[…]
We conclude that
$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{m}\right)}=1.$
For the point d), repeat the proof of point c) in Theorem 36 with the sequence $\{t_{k}\}_{k}$ replaced by the sequence $\{t_{m}\}_{m}$ above in point c) of this proof. In this way, we show that there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}_{s}$ of the sequence $\{t_{m}\}_{m}$, such that
$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1$
and then
$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{m_{s}}\right)}\cdot\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1.$
Remark 40.
In the case where $\Lambda_{q}$ consists of real eigenvalue, by recalling (55) and point 2 in Remark 25, the point c) is modified to:
  • $c^{\prime}$)
    For any initial value $y_{0}\in U^{e}_{q}$, we have
    $K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
In addition, in the case where $\Lambda_{1}$ consists of a real eigenvalue, the point d) is modified to:
  • $d^{\prime}$)
    There exist a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ such that, for any initial value $y_{0}\in U^{e}_{q}$, we have
    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
See Remark 37.
4.6. The global asymptotic condition number $K_{\infty}\left(t\right)$
Theorem 39 says that $K_{\infty}\left(t\right)$ in (54), not $K^{+}_{\infty}\left(t\right)$ in (53), is the worst asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$, namely the worst asymptotic condition $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying both $y_{0}$ and $\widehat{z}_{0}$.
We define the function
$t\rightarrow K_{\infty}(t),\ t\in\mathbb{R},$
as the global asymptotic condition number of the problem (2).
Remark 41.
  • 1.
    Observe that $K_{\infty}(t)$ is the global asymptotic condition number, i.e. the worst-case asymptotic condition number $K_{\infty}(t,y_{0})$ as $y_{0}$ varies. In contrast, $K_{\infty}^{+}(t)$ is the asymptotic global condition number, i.e. the asymptotic form of the worst-case condition number $K(t,y_{0})$ as $y_{0}$ varies.
  • 2.
    Both the global asymptotic condition number and the asymptotic global condition number of the problem (2):
    • are bounded and away from zero as $t$ varies if $q=1$ and $L_{1}=0$;
    • diverge polynomially to infinite, as $t\rightarrow+\infty$, if $q=1$ and $L_{1}>0$;
    • diverge exponentially to infinite, as $t\rightarrow+\infty$, if $q>1$.
    Therefore, the they do not diverges to infinite if and only if all the eigenvalues of $A$ have the same real part, i.e. they lie in a vertical line of the complex plane, and are simple.
Observe that $K^{+}_{\infty}(t)$ can be significantly larger than $K_{\infty}(t)$: we have
$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=L_{q}!t^{-L_{q}}\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{qL_{q}}(-t)\right\|}$
and then
$\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=0$
for $L_{q}>0$ and
$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}=\frac{\left\|Q_{q0}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}\leq 1$
for $L_{q}=0$.
In the next example, we illustrate the difference between $K^{+}_{\infty}(A)$ and $K_{\infty}(A)$.
Example 42.
Consider the one Jordan block matrix
$A_{1}=V\left[\begin{array}[]{rr}1&1\\ 0&1\end{array}\right]W$
and the diagonalizable matrix
$A_{2}=V\left[\begin{array}[]{rr}1&0\\ 0&-1\end{array}\right]W,$
where
$V=\left[\begin{array}[]{rr}1&\frac{1}{2}\\ 1&1\end{array}\right]\text{\ \ and\ \ }W=V^{-1}=2\left[\begin{array}[]{rr}1&-\frac{1}{2}\\ -1&1\end{array}\right].$
For the matrix $A_{1}$, we have
$\Lambda=\{1\},\ q=1\text{\ \ and\ \ }L_{1}=L_{q}=1$
and then
$K_{\infty}(t)=t\left\|Q_{11}(t)\right\|\left\|Q^{e}_{1}(-t)|_{U^{e}_{1}}\right\|$
$K^{+}_{\infty}(t)=t^{2}\left\|Q_{11}(t)\right\|\left\|Q_{11}(-t)\right\|$
where
$Q_{11}(t)=Q_{11}(-t)=P_{11}=v^{(1,1,1)}w^{(1,1,2)}=\left[\begin{array}[]{rr}-2&2\\ -2&2\end{array}\right]$
(remind point 2 in Remark 8)
$Q^{e}_{1}(-t)|_{U^{e}_{1}}=I|_{U^{e}_{1}}$
(remind point 2 in Remark 25).
Thus, for the Euclidean norm as vector norm, we have, since $\|P_{11}\|_{2}=4$,
$K_{\infty}(t)=4t$
$K^{+}_{\infty}(t)=16t^{2}.$
For the matrix $A_{2}$, we have
$\Lambda=\{1,-1\},\ q=2\text{\ \ and\ \ }L_{1}=L_{q}=0$
and then
$K_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q^{e}_{2}(-t)|_{U^{e}_{2}}\right\|$
$K^{+}_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q_{20}(-t)\right\|,$
where
$Q_{10}(t)=P_{10}=v^{(1,1,1)}w^{(1,1,1)}=\left[\begin{array}[]{rr}2&-1\\ 2&-1\end{array}\right]$
$Q_{20}(-t)=P_{20}=v^{(2,1,1)}w^{(2,1,1)}=\left[\begin{array}[]{rr}-1&1\\ -2&2\end{array}\right]$
$Q^{e}_{2}(-t)|_{U^{e}_{2}}=I|_{U^{e}_{2}}.$
Thus, for the Euclidean norm as vector norm, we have, since $\|P_{10}\|_{2}=\|P_{20}\|_{2}=\sqrt{10}$,
$K_{\infty}(t)=\sqrt{10}\mathrm{e}^{2t}$
$K^{+}_{\infty}(t)=10\mathrm{e}^{2t}.$
In Figure 6 for the matrix $A_{1}$, we see in logarithmic scale the condition numbers $K(t)$ (blu dashed line) , $K^{+}_{\infty}(t)$ (red dashed line) and $K_{\infty}(t)$ (red dashed line), $t\in[0,100]$. From the beginning, $K(t)$ is not distinguishable from its asymptotic version $K^{+}_{\infty}(t)$.
Figure 6. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{1}$.
The asymptotic global condition number $K^{+}_{\infty}(t)$ is unrelated to the worst asymptotic behavior of the condition numbers $K(t,y_{0},\widehat{z}_{0})$. In fact, in Figure 6 we also see, for each $\tau\in\{10,20,30,40\}$, the condition number $K(t,y_{0\tau},\widehat{z}_{0\tau})$ (black solid lines) for an initial value $y_{0\tau}$ and a direction of perturbation $\widehat{z}_{0\tau}$ such that $K(\tau,y_{0\tau},\widehat{z}_{0\tau})=K(\tau)$. To obtain this, we take the initial value $y_{0\tau}$ such that
$\|\mathrm{e}^{-\tau A}\|=\frac{1}{\|\mathrm{e}^{\tau A}\widehat{y}_{0\tau}\|}$
(we take $y_{0\tau}=\mathrm{e}^{-\tau A}x_{0\tau}$, where $x_{0\tau}$ is such that
$\|\mathrm{e}^{-\tau A}\|=\frac{\|\mathrm{e}^{-\tau A}x_{0\tau}\|}{\|x_{0\tau}\|})$
and the direction of perturbation $\widehat{z}_{0\tau}$ such that
$\|\mathrm{e}^{\tau A}\|=\|\mathrm{e}^{\tau A}\widehat{z}_{0\tau}\|.$
The condition number $K(t,y_{0\tau},\widehat{z}_{0\tau})$ is the worst condition number $K(t,y_{0},\widehat{z}_{0})$, by varying the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$, only at $t=\tau$. But, asymptotically it goes below $K_{\infty}(t)$.
On the other hand, the global asymptotic condition number $K_{\infty}(t)$ is the worst asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$, by varying the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$. Indeed, we have $K_{\infty}(t)=K_{\infty}(t,y_{0},\widehat{z}_{0})$ for $y_{0}\in U^{e}_{2}$ and for $\widehat{z}_{0}$ such that
$\left\|Q_{11}(t)\right\|=\left\|P_{11}\right\|=\left\|P_{11}\widehat{z}_{0}\right\|$
(see point d’) in Remark 40 and Remark 37).
In Figure 7, we see the same for the matrix $A_{2}$. Here $\tau\in\{1,2,3,4\}$.
Figure 7. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{2}$.
5. Some considerations about the asymptotic condition numbers
Suppose that the matrix $A$ in the ODE (1) does not have all eigenvalues with the same real part, meaning $q>1$ (recall that $q$ is the number of distinct real parts). The global asymptotic condition number $K_{\infty}\left(t\right)$ grows exponentially with $t$. Hence, the relative error $\varepsilon$ in the initial value is exponentially magnified in the relative error $\delta(t)$ in the solution, in the worst case for the initial value and the perturbation of the initial value.
However, this conclusion is too pessimistic, since this exponential magnification of $\varepsilon$ in $\delta(t)$ appears only in a non-generic situation for the initial value $y_{0}$. In fact, as illustrated in point 2 of Remark 33, in the generic situation $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$ for $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$, we have
$\frac{\delta(t)}{\varepsilon}=K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|},\ t\rightarrow+\infty,$ (61)
where $Q_{1}(t):=Q_{1L_{1}}(t)$ to semplify the notation. In this generic situation, the magnification of the initial error remains bounded as well as away from zero by varying $t$ (see point 1 of Remark 33).
We have introduced the three asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,$ $y_{0})$ and $K_{\infty}(t)$. The most important asymptotic condition number is $K_{\infty}(t,y_{0})$. In fact, $K_{\infty}(t,y_{0},\widehat{z}_{0})$ is given in terms of the (in general) unknown direction of perturbation $\widehat{z}_{0}$ and $K_{\infty}(t)$ is a worst asymptotic magnification factor which applies in a non-generic situation for $y_{0}$.
As illustrated in point 2 of Remark 35, in the generic situation $j^{\ast}=1$ and $L^{\ast}=L_{1}$ for $y_{0}$, this most important asymptotic condition number $K_{\infty}(t,y_{0})$, representing the asympotic magnification factor $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$ in (61) for the worst perturbation of $y_{0}$, is given by
$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1}(t)\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|}.$ (62)
Going into more detail, in (61) and (62), the vectors $Q_{1}(t)u$, with $u=\widehat{y}_{0}$ or $u=\widehat{z}_{0}$, and the matrix $Q_{1}(t)$ are given by (see (42), (29) and (30))
$Q_{1}(t)u$ $=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}$
$=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}$
and
$Q_{1}(t)$ $=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)}$
$=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)},$
where
$M_{1}:=L_{1}+1=\max\limits_{\lambda_{i}\in\Lambda_{1}}m_{i}=\max\limits_{\lambda_{i}\in\Lambda_{1}}\max\limits_{j\in\{1,\ldots,d_{i}\}}m_{ij}$
to semplify the notation. This shows that the asymptotic behavior of the magnification in the solution of the initial relative error is determined by
  • the imaginary parts $\omega_{i}$ of the rightmost eigenvalues of $A$ with the longest Jordan chains (chains of length $M_{1}$); in other words, with the largest mini-blocks in the JCF (mini-blocks of order $M_{1}$);
  • the eigenvectors $v^{\left(i,j,1\right)}$ of these longest Jordan chains;
  • the components $\alpha_{ijM_{1}}(u)$ of $u=\widehat{y}_{0},\widehat{z}_{0}$ along the last generalized eigenvectors of these longest Jordan chains;
  • for the worst case of $\widehat{z}_{0}$, the last left generalized eigenvectors $w^{(i,j,M_{1})}$ corresponnding to these longest Jordan chains.
For $u\in\mathbb{C}^{n}$, we say that $u$ satisfies the Rightmost Last Generalized Eigenvector (RLGE) condition if $j(u)=1$ and $L(u)=L_{1}$, i.e. there is a non-zero component $\alpha_{ijM_{1}}(u)$ along the last generalized eigenvctors of the longest Jordan chain of the rightmost eigenvalues of $A$. Therefore, we can say that (61) holds if $y_{0}$ (or $\widehat{y}_{0}$) and $\widehat{z}_{0}$ satisfy the RLGE condition and (62) holds if $y_{0}$ (or $\widehat{y}_{0}$) satisfies the RLGE condition. This terminology is utilized in [13].
6. The generic real case. Part I: introduction
In this section, Part I, and the following four sections, Parts II, III, IV, and V, we tailor the results from the previous Section 4 to a real ODE (1) in a generic situation for the matrix $A$, the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$.
Since, as noted in the previous Section 5, the worst-case scenario of relative error magnification described by the asymptotic condition number $K_{\infty}(t)$ occurs in a non-generic situation for the initial value $y_{0}$, we focus our analysis on the other two more significant asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$.
Parts I and II introduce concepts and present results essential for the developments in Parts III and IV, where asymptotic condition numbers are addressed. Parts I and III consider an arbitrary vector norm, while Parts II and IV explore the concepts introduced in Parts I and III, respectively, in the case of the Euclidean vector norm and a complex conjugate pairs of eigenvalues.
The concluding section, Part V, investigates the onset of asymptotic behavior, specifically studying when a condition number begins to be close to its asymptotic form and identifying the factors that influence its rapid convergence towards this form.
6.1. Real and complex induced norms
[…]
$\|B\|=\max_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \left\|u\right\|=1\end{subarray}}\left\|Bu\right\|,$ (63)
instead of the complex induced norm
$\max_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \left\|u\right\|=1\end{subarray}}\left\|Bu\right\|.$
For a complex row vectors $b\in\mathbb{C}^{1\times n}$, we utilize both the real induced norm
$\left\|b\right\|_{\mathbb{R}}:=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left|bu\right|$
and the complex induced norm
$\left\|b\right\|=\max_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \left\|u\right\|=1\end{subarray}}\left|bu\right|.$
Observe that we have
$\left\|b\right\|_{\mathbb{R}}\leq\left\|b\right\|.$ (64)
Remark 43.
[…]
$\left\|b\right\|=\left\|b^{T}\right\|_{q},$
where $\|\ \cdot\ \|_{q}$ is the vector $q$-norm with $\frac{1}{p}+\frac{1}{q}=1$.
6.2. The generic $\Lambda_{j}$
Assume $A\in\mathbb{R}^{n\times n}$ in (1). Let $j\in\{1,\ldots,q\}$ be such that the set $\Lambda_{j}$ consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues. We say that $\Lambda_{j}$ is real in the first case and complex in the second case.
We consider a set $\Lambda_{j}$ of this type, since it is a generic situation for the real matrix $A$ to have all sets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, of such a form.
When $\Lambda_{j}$ is real, we denote by $\lambda_{j}$ the real eigenvalue in $\Lambda_{j}$. When $\Lambda_{j}$ is complex, we denote by $\lambda_{j}$ and $\overline{\lambda_{j}}$ the complex conjugate pair in $\Lambda_{j}$, where $\lambda_{j}$ is the eigenvalue with positive imaginary part. This positive imaginary part is denoted by $\omega_{j}$.
Let $V$ be the matrix whose columns form a Jordan basis for the matrix $A$ and let $W=V^{-1}$ (see Section 2). Let $v^{(j)}\in\mathbb{C}^{n}$ be the eigenvector of $A$ corresponding to the simple eigenvalue $\lambda_{j}$ (real eigenvalue or complex eigenvalue with positive imaginary part in the complex conjugate pair) appearing as a column in the matrix $V$. Let $w^{(j)}\in\mathbb{C}^{1\times n}$ be the corresponding row in the matrix $W$, i.e. the index of $v^{(j)}$ in $V$ matches the index of $w^{(j)}$ in $W$. Observe that $w^{(j)}$ is a left eigenvector of the matrix $A$ corresponding to the eigenvalue $\lambda_{j}$.
If $\Lambda_{j}$ is real, then $v^{(j)}$ and $w^{(j)}$ are real vectors (see Proposition 5).
The projection onto the eigenspace of the simple (real or complex) eigenvalue $\lambda_{j}$ is $v^{(j)}w^{(j)}$ (see (27)). Thus, by recalling Proposition 22, we have
$Q_{j0}(t)=v^{(j)}w^{(j)}\in\mathbb{R}^{n\times n},$ (65)
when $\Lambda_{j}$ is real and
$Q_{j0}(t)=2\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}\right)\in\mathbb{R}^{n\times n},$ (66)
when $\Lambda_{j}$ is complex.
Remark 44.
Recall Propostion 22. When $\Lambda_{j}$ is real, the projection on the eigenspace corresponding to $\lambda_{j}$ is $Q_{j0}(t)$, independent of $t$. When $\Lambda_{j}$ is complex, the projection onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$ is
$Q_{j0}(0)=v^{(j)}w^{(j)}+\overline{v^{(j)}}\ \overline{w^{(j)}}=2\mathrm{Re}\left(v^{(j)}w^{(j)}\right).$
6.3. The unit vectors $\widehat{v}^{\left(j\right)}$ and $\widehat{w}^{\left(j\right)}$ and the number $f_{j}$
Let
$\widehat{v}^{\left(j\right)}:=\frac{v^{\left(j\right)}}{\left\|v^{\left(j\right)}\right\|},\ \widehat{w}^{\left(j\right)}:=\frac{w^{\left(j\right)}}{\left\|w^{\left(j\right)}\right\|}\ \text{and}\ f_{j}:=\left\|w^{\left(j\right)}\right\|\cdot\left\|v^{\left(j\right)}\right\|.$ (67)
In (67), when $\Lambda_{j}$ is real, $\left\|w^{\left(j\right)}\right\|$ denotes the real induced norm
$\left\|w^{\left(j\right)}\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
of the real row vector $w^{(j)}$. On the other hand, when $\Lambda_{j}$ is complex, $\left\|w^{\left(j\right)}\right\|$ denotes the complex induced norm
$\left\|w^{\left(j\right)}\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \|u\|=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
of the complex row vector $\widehat{w}^{(j)}$.
6.4. The numbers $g_{j}(t,u)$ and $g_{j}(t)$, the vector $\widehat{\Theta}_{j}(t,u)$ and the matrix $\widehat{\Theta}_{j}(t)$.
In the following, in both cases for $\Lambda_{j}$, real or complex, we write the norm of the real vector $Q_{j0}(t)u$, where $u\in\mathbb{R}^{n}$, as
$\|Q_{j0}(t)u\|=f_{j}\left|\widehat{w}^{(j)}u\right|g_{j}(t,u)$ (68)
for a suitable number $g_{j}(t,u)$. Analogously, we write the real induced norm (63) of the real matrix $Q_{j0}(t)$ as
$\|Q_{j0}(t)\|=f_{j}g_{j}(t)$ (69)
for a suitable number $g_{j}(t)$.
Next two lemmas say what $g_{j}(t,u)$ and $g_{j}(t)$ are.
Lemma 45.
Let $j\in\{1,\ldots,q\}$. If $\Lambda_{j}$ is real, then
$g_{j}(t,u)=1,\ u\in\mathbb{R}^{n},$
$g_{j}(t)=1.$
Proof.
For $u\in\mathbb{R}^{n}$, by (65) we have
$Q_{j0}(t)u=v^{(j)}w^{(j)}u=\left(w^{(j)}u\right)v^{(j)}=f_{j}\left(\widehat{w}^{(j)}u\right)\widehat{v}^{(j)}$
and then
$\left\|Q_{j0}(t)u\right\|=\left|w^{(j)}u\right|\left\|v^{(j)}\right\|=f_{j}\left|\widehat{w}^{(j)}u\right|.$ (70)
By the second equality in (70), we obtain $g_{j}(t,u)=1$. By the first equality, we obtain
$\left\|Q_{j0}(t)\right\|=\left\|w^{(j)}\right\|\left\|v^{(j)}\right\|=f_{j}$
and then $g_{j}(t)=1$. ∎
Lemma 46.
Let $j\in\{1,\ldots,q\}$. If $\Lambda_{j}$ is complex, then
$g_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|,\ u\in\mathbb{R}^{n},$
$g_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|,$
where
$\widehat{\Theta}_{j}\left(t,u\right)=\left(\left|\widehat{v}_{k}^{\left(j\right)}\right|\cos\left(\omega_{j}t+\alpha_{jk}+\gamma_{j}\left(u\right)\right)\right)_{k=1,\ldots,n}\in\mathbb{R}^{n}$
$\widehat{\Theta}_{j}\left(t\right)=\left(\left|\widehat{v}_{k}^{\left(j\right)}\right|\left|\widehat{w}_{l}^{\left(j\right)}\right|\cos\left(\omega_{j}t+\alpha_{jk}+\beta_{jl}\right)\right)_{k,l=1,\ldots,n}\in\mathbb{R}^{n\times n}$
with $\alpha_{jk}$, $\beta_{jl}$ and $\gamma_{j}(u)$ defined by
$\widehat{v}_{k}^{\left(j\right)}=\left|\widehat{v}_{k}^{\left(j\right)}\right|\mathrm{e}^{\sqrt{-1}\alpha_{jk}},\ k\in\{1,\ldots,n\},$
$\widehat{w}_{l}^{\left(j\right)}=\left|\widehat{w}_{l}^{\left(j\right)}\right|\mathrm{e}^{\sqrt{-1}\beta_{jl}},\ l\in\{1,\ldots,n\},$
$\widehat{w}^{\left(j\right)}u=\left|\widehat{w}^{\left(j\right)}u\right|\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)},$
i.e. they are, respectively, the angles in the polar forms of the components of the complex vector $\widehat{v}^{(j)}$, the angles in the polar forms of the components of the complex vector $\widehat{w}^{(j)}$ and the angle in the polar form of the complex number $\widehat{w}^{\left(j\right)}u$.
Moreover, we have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left(\left|\widehat{w}^{\left(j\right)}u\right|\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|\right).$ (71)
In this lemma, $\left\|\widehat{\Theta}_{j}(t)\right\|$ is the real induced norm (63) of the real matrix $\widehat{\Theta}_{j}(t)$.
Proof.
By (66), we have
$Q_{j0}(t)=2\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}\right)=2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right).$ (72)
Thus, for $u\in\mathbb{R}^{n}$, we have
$Q_{j0}(t)u$ $=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right)u$
$=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}u\right)$
$=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\left(\widehat{w}^{\left(j\right)}u\right)\widehat{v}^{(j)}\right)$
$=$ $2f_{j}\left|\widehat{w}^{\left(j\right)}u\right|\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)}\widehat{v}^{(j)}\right)$
and
$\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)}\widehat{v}^{(j)}\right)=\widehat{\Theta}_{j}\left(t,u\right).$
We have obtained
$Q_{j0}(t)u=2f_{j}\left|\widehat{w}^{\left(j\right)}u\right|\widehat{\Theta}_{j}\left(t,u\right).$ (73)
and then $g_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|$.
Moreover, in (72) we have
$\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right)=\widehat{\Theta}_{j}(t).$
Therefore, we obtain
$Q_{j0}(t)=2f_{j}\widehat{\Theta}_{j}\left(t\right)$ (74)
and then $g_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|$.
Finally, by (74) and (73), we have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|=\frac{\left\|Q_{j0}(t)\right\|}{2f_{j}}=\frac{1}{2f_{j}}\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left\|Q_{j0}(t)u\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left(\left|\widehat{w}^{\left(j\right)}u\right|\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|\right).$
Remark 47.
  • 1.
    The functions $t\mapsto\widehat{\Theta}_{j}(t,u)$ and $t\mapsto\widehat{\Theta}_{j}(t)$ are periodic functions of period $\frac{2\pi}{\omega_{j}}$ and the functions $t\mapsto\|\widehat{\Theta}_{j}(t,u)\|$ and $t\mapsto\|\widehat{\Theta}_{j}(t)\|$ are periodic functions of period $\frac{\pi}{\omega_{j}}$.
  • 2.
    When the vector norm is a $p$-norm, we have
    $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{p}=\left(\sum\limits_{k=1}^{n}\left|\widehat{v}_{k}^{\left(j\right)}\right|^{p}\left|\cos\left(\omega_{j}t+\alpha_{jk}+\gamma_{j}\left(u\right)\right)\right|^{p}\right)^{\frac{1}{p}}\leq\left\|\widehat{v}^{\left(j\right)}\right\|_{p}=1.$
  • 3.
    When the vector norm is a $p$-norm, we have $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{p}\leq 1$. This follows by (71),
[…]
    and the previous point 2.
7. The generic real case. Part II: introduction (euclidean norm and complex conjugate eigenvalues)
[…]
In the following, for a complex row or column vector $z$, $z^{T}$ denotes the transpose of $z$, representing only transposition without conjugation. The conjugate transpose of $z$ is denoted by $z^{H}$.
7.1. The set $\mathcal{W}_{j}$ and the ellipse $\mathcal{E}_{j}$.
For $u\in\mathbb{R}^{n}$, we have (recall (68) and Lemma 46)
$\left\|Q_{j0}(t)u\right\|_{2}=2f_{j}\left|\widehat{w}^{(j)}u\right|\left\|\widehat{\Theta}_{j}(t,u)\right\|_{2}.$ (75)
The complex number $\widehat{w}^{(j)}u$ plays a crucial role in this formula: its modulus is part of it, and the vector $\widehat{\Theta}_{j}(t,u)$ is defined in Lemma 46 by means of the angle $\gamma_{j}(u)$ in the polar form of $\widehat{w}^{(j)}u$.
For this reason, now we study the set
$\mathcal{W}_{j}:=\{\widehat{w}^{(j)}u:u\in\mathbb{R}^{n}\text{\ and\ }\left\|u\right\|_{2}=1\}$
of the complex numbers $\widehat{w}^{(j)}u$, by varying the unit real vector $u$.
Next proposition describes the set $\mathcal{W}_{j}$ and states the real induced norm
$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
of the complex row vector $\widehat{w}^{(j)}$, whose complex induced norm is equal to $1$.
Proposition 48.
Assume that $A$ is a real matrix. Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex. Consider, in the complex plane, the ellipse $\mathcal{E}_{j}$ centered at the origin with semi-axes the vectors $\sigma_{j}\alpha^{(j)}$ and $\mu_{j}\beta^{(j)}$, where $\sigma_{j}\geq\mu_{j}\geq 0$ are the singular values of the $2\times n$ real matrix
$R_{j}:=\left[\begin{array}[]{c}\mathrm{Re}(\widehat{w}^{(j)})\\ \mathrm{Im}(\widehat{w}^{(j)})\end{array}\right]$
and $\alpha^{(j)},\beta^{(j)}\in\mathbb{R}^{2}$ are the corresponding left singular vectors. We have
$\sigma_{j}=\sqrt{\frac{1+W_{j}}{2}}\text{\ \ and\ \ }\mu_{j}=\sqrt{\frac{1-W_{j}}{2}},$ (76)
where $W_{j}$ is the modulus of the complex number $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$.
If $n=2$, the set $\mathcal{W}_{j}$ is the ellipse $\mathcal{E}_{j}$. If $n>2$, the set $\mathcal{W}_{j}$ is the union of the ellipse $\mathcal{E}_{j}$ and its interior. Moreover, we have
$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\sigma_{j}=\sqrt{\frac{1+W_{j}}{2}}.$
Proof.
We begin by proving (76). The singular values $\sigma_{j}$ and $\mu_{j}$ of the matrix $R_{j}$ are the square roots of the largest eigenvalue and the smallest eigenvalue, respectively, of the matrix
$R_{j}R_{j}^{T}=\left[\begin{array}[]{c}\mathrm{Re}(w^{(j)})\\ \mathrm{Im}(w^{(j)})\end{array}\right]\left[\mathrm{Re}(w^{(j)})^{T}\ \mathrm{Im}(w^{(j)})^{T}\right]=\left[\begin{array}[]{cc}a_{j}^{2}&p_{j}\\ p_{j}&b_{j}^{2}\end{array}\right],$
where
$a_{j}=\left\|\mathrm{Re}\left(\widehat{w}^{(j)}\right)^{T}\right\|_{2},\ b_{j}=\left\|\mathrm{Im}\left(\widehat{w}^{(j)}\right)^{T}\right\|_{2}\text{\ \ and\ \ }p_{j}=\mathrm{Re}\left(\widehat{w}^{(j)}\right)\mathrm{Im}\left(\widehat{w}^{(j)}\right)^{T}.$
Therefore, we obtain
$\sigma_{j}^{2}=\frac{1+\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}}{2}\text{\ and\ }\mu_{j}^{2}=\frac{1-\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}}{2}.$
by observing that $a_{j}^{2}+b_{j}^{2}=\left\|\widehat{w}^{(j)}\right\|_{2}^{2}=1$. Since
$\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}=a_{j}^{2}-b_{j}^{2}+\sqrt{-1}\cdot 2p_{j},$
we get
$W_{j}=\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}$
and (76) follows.
Now, we pass to describe the set $\mathcal{W}_{j}$ in terms of the ellipse $\mathcal{E}_{j}$. By looking to a complex number as the pair constituted by the real and imaginary parts, the complex number $\widehat{w}^{(j)}u$, $u\in\mathbb{R}^{n}$, is $R_{j}u$. By introducing the SVD of $R_{j}$, we see that
$R_{j}u=\sigma_{j}c_{j}(u)\alpha^{(j)}+\mu_{j}d_{j}(u)\beta^{(j)}$
where $c_{j}(u)$ and $d_{j}(u)$ are the components of $u$ along the first and second right singular vectors of $R_{j}$, i.e. the right singular vectors corresponding to $\sigma_{j}$ and $\mu_{j}$, respectively. Thus,
$\mathcal{W}_{j}=\{\sigma_{j}c\alpha^{(j)}+\mu_{j}d\beta^{(j)}:(c,d)\in\mathbb{R}^{2}\text{\ and\ }\left\|(c,d)\right\|_{2}=1\}$ (77)
when $n=2$ and
$\mathcal{W}_{j}=\{\sigma_{j}c\alpha^{(j)}+\mu_{j}d\beta^{(j)}:(c,d)\in\mathbb{R}^{2}\text{\ and\ }\left\|(c,d)\right\|_{2}\leq 1\}$ (78)
when $n>2$.
In the orthonormal basis of $\mathbb{R}^{2}$ given by the vector $\alpha^{(j)}$ and $\beta^{(j)}$, the coordinates $x=\sigma_{j}c$ and $y=\mu_{j}d$ of the elements of the set (77) satisfies the equation
$\frac{x^{2}}{\sigma_{j}^{2}}+\frac{y^{2}}{\mu_{j}^{2}}=1,$
which is the equation of an ellipse centered at the origin with lengths of the semi-axes $\sigma_{j}$ and $\mu_{j}$. Thus, we have $\mathcal{W}_{j}=\mathcal{E}_{j}$. On the other hand, the coordinates $x=\sigma_{j}c$ and $y=\mu_{j}d$ of the elements of the set (78) satisfies the inequality
$\frac{x^{2}}{\sigma_{j}^{2}}+\frac{y^{2}}{\mu_{j}^{2}}\leq 1.$
Thus, $\mathcal{W}_{j}$ is the union of the ellipse $\mathcal{E}_{j}$ and its interior.
Finally, we have
$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left|\widehat{w}^{\left(j\right)}u\right|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left\|R_{j}u\right\|_{2}=\left\|R_{j}\right\|_{2}=\sigma_{j}.$
Remark 49.
  • 1.
    By the Cauchy-Schwarz inequality and Remark 43, or only by Remark 43, we have
    $W_{j}=\left|\widehat{w}^{\left(j\right)}\left(\widehat{w}^{\left(j\right)}\right)^{T}\right|=\left|\left(\left(\overline{\widehat{w}^{\left(j\right)}}\right)^{T}\right)^{H}\left(\widehat{w}^{\left(j\right)}\right)^{T}\right|\leq 1.$
    Moreover, $W_{j}<1$ holds, otherwise $\widehat{w}^{\left(j\right)}$ and $\overline{\widehat{w}^{\left(j\right)}}$ were linearly dependent, which is impossible since $w^{\left(j\right)}$ and $\overline{w^{\left(j\right)}}$ are the rows of the non-singular matrix $W$ corresponding to the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively.
  • 2.
    Remind the inequality (64). By the previous point 1, we have
    $\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\sqrt{\frac{1+W_{j}}{2}}<1$
    for the real induced norm of the row vector $\widehat{w}^{(j)}$. On the other hand,
    $\left\|\widehat{w}^{(j)}\right\|_{2}=1.$
    for the complex induced norm of $\widehat{w}^{(j)}$.
7.2. Projection onto the eigenspaces of $\lambda_{j}$ and $\overline{\lambda}_{j}$
The projection of $u\in\mathbb{R}^{n}$ onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda}_{j}$ is (see Remark 44)
$f_{j}\left(\widehat{w}^{(j)}u\right)\widehat{v}^{(j)}+f_{j}\overline{\left(\widehat{w}^{(j)}u\right)}\ \overline{\widehat{v}^{(j)}},$
where the first and the second terms are the projections onto the eigenspace corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively. Since $\widehat{v}^{(j)}$ and $\overline{\widehat{v}^{(j)}}$ are linearly independent, the projection onto the sum of the eigenspaces is zero if and only if the projection onto the eigenspace corresponding to $\lambda_{j}$ is zero, i.e. $\widehat{w}^{(j)}u=0$.
The Euclidean norm of the projection onto the eigenspace corresponding to $\lambda_{j}$ is $f_{j}\left|\widehat{w}^{(j)}u\right|$, equal to the norm of the projection onto the eigenspace corresponding to $\overline{\lambda_{j}}$. The Euclidean norm of the projection on the sum of the eigenspaces is given in Remark 53 below.
7.3. The right singular vectors of the matrix $R_{j}$
Vectors in $\mathbb{R}^{n}$ can be expressed by real components in the orthonormal basis of $\mathbb{R}^{n}$ of the right singular vectors of the matrix $R_{j}$ of Proposition 48.
As we saw in the proof of Proposition 48, for a vector $u\in\mathbb{R}^{n}$ we have
$\widehat{w}^{(j)}u=a+\sqrt{-1}b$ (79)
with
$(a,b)=R_{j}u=\sigma_{j}c_{j}(u)\alpha^{(j)}+\mu_{j}d_{j}(u)\beta^{(j)},$
where: $\sigma_{j}$ and $\mu_{j}$ are the two singular values of $R_{j}$; $c_{j}(u)$ and $d_{j}(u)$ are the components of $u$ along the first and second, respectively, right singular vectors of $R_{j}$; and $\alpha^{(j)}$ and $\beta^{(j)}$ are the left singular vectors of $R_{j}$. Therefore, $u$ has projection zero onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda}_{j}$, i.e. $\widehat{w}^{(j)}u=0$ by reminding the previous Subsection 7.2, if and only if $u$ is spanned by the last $n-2$ right singular vectors of $R_{j}$.
By (79), we have
$\left|\widehat{w}^{(j)}u\right|=\left\|R_{j}u\right\|_{2}=\sqrt{\sigma_{j}^{2}c_{j}(u)^{2}+\mu_{j}^{2}d_{j}(u)^{2}}.$ (80)
Remark 50.
Suppose $n>2$. By considering the components $c_{j}(u)$ and $d_{j}(u)$ of $u\in\mathbb{R}^{n}$ with $\left\|u\right\|_{2}=1$ along the first and the second right singular vectors of $R_{j}$, the set $\mathcal{W}_{j}$ can be partitioned by level curves of
$\sigma_{j}^{2}c_{j}(u)^{2}+\mu_{j}^{2}d_{j}(u)^{2}=\left|\widehat{w}^{(j)}u\right|^{2},$
whose set of possible values by varying $u$ is $[0,\sigma_{j}^{2}]$, or by level curves of
$c_{j}(u)^{2}+d_{j}(u)^{2},$
whose set of possible values is $[0,1]$. In the first case, the level curves are concentric circles with center the origin. In the second case, the level curves are ellipses concentric to the ellipse $\mathcal{E}_{j}$.
7.4. Notations
Before continuing with our analysis, we recall two important notations and introduce a few others.
Here are the two previously introduced notations.
  • The modulus $W_{j}$ of the complex number $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$, introduced in Proposition 48;
  • The angle $\gamma_{j}(u)$ in the polar form of the complex number $\widehat{w}^{(j)}u$, where $u\in\mathbb{R}^{n}$, introduced In Lemma 46.
Here are the new notations.
  • The modulus $V_{j}$ of the complex number $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$.
  • The angle $\delta_{j}$ in the polar form of the complex number $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$.
  • The angle $\theta_{j}$ in the complex plane between the real axis and the major axis $\sigma_{j}\alpha^{(j)}$ of the ellipse $\mathcal{E}_{j}$ of Proposition 48.
  • $x_{j}(t):=2(\omega_{j}t+\theta_{j})+\delta_{j}.$ (81)
  • For $u\in\mathbb{R}^{n}$,
    $\Delta_{j}(u):=2(\gamma_{j}(u)-\theta_{j}).$ (82)
Remark 51.
  • 1.
    Observe that $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{H}=1$ and $\left(\widehat{v}^{\left(j\right)}\right)^{H}\widehat{v}^{\left(j\right)}=1$. On the other hand, $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$ and $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$ are complex numbers and $W_{j}$ and $V_{j}$ are their modules, respectively.
  • 2.
    At point 1 of Remark 49, we have seen that $W_{j}<1$, The same holds for $V_{j}$. In fact, by the Cauchy-Schwarz inequality we have
    $V_{j}=\left|\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}\right|=\left|\left(\overline{\widehat{v}^{\left(j\right)}}\right)^{H}\widehat{v}^{\left(j\right)}\right|\leq 1.$
    Moreover, $V_{j}<1$ holds, otherwise $\widehat{v}^{\left(j\right)}$ and $\overline{\widehat{v}^{\left(j\right)}}$ were linearly dependent, which is impossible since $v^{\left(j\right)}$ and $\overline{v^{\left(j\right)}}$ are the columns of the non-singular matrix $V$ corresponding to the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively. Therefore, we have $V_{j},W_{j}\in[0,1)$.
7.5. The norm $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$
Next proposition provides an expression for the norm $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$.
Proposition 52.
Assume that $A$ is a real matrix. Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex. For $u\in\mathbb{R}^{n}$, we have
$\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\Delta_{j}(u)\right)}{2}},$ (83)
where $x_{j}(t)$ and $\Delta_{j}(u)$ are defined in (81) and (82).
Proof.
Let
$\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}=V_{j}\mathrm{e}^{\sqrt{-1}\delta_{j}}$
be the polar form of $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$ (remind Subsection 7.4). We have (remind (66))
$\left\|Q_{j0}(t)u\right\|_{2}^{2}=\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u+\overline{\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u}\right\|_{2}^{2}$
$=2\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right\|_{2}^{2}+2\mathrm{Re}\left(\left(\overline{\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u}\right)^{H}\mathrm{e}^{\sqrt{-1}\omega_{j}}v^{(j)}w^{(j)}u\right)$
$=2\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right\|_{2}^{2}+2\mathrm{Re}\left(\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right)^{T}\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right)$
$=2\left(\left\|v^{(j)}\right\|_{2}^{2}\left|w^{(j)}u\right|^{2}+\mathrm{Re}\left(\mathrm{e}^{2\sqrt{-1}\omega_{j}t}\left(\left(v^{(j)}\right)^{T}v^{(j)}\right)\left(w^{(j)}u\right)^{2}\right)\right)$
$=2\left(f_{j}^{2}|\widehat{w}^{(j)}u|^{2}+\mathrm{Re}\left(f_{j}^{2}|\widehat{w}^{(j)}u|^{2}\mathrm{e}^{2\sqrt{-1}\omega_{j}t}V_{j}\mathrm{e}^{\sqrt{-1}\delta_{j}}\mathrm{e}^{2\sqrt{-1}\gamma_{j}(u)}\right)\right)$
$=2f_{j}^{2}\left|\widehat{w}^{(j)}u\right|^{2}\left(1+V_{j}\cos\left(2\omega_{j}t+\delta_{j}+2\gamma_{j}(u)\right)\right)$ (84)
$=2f_{j}^{2}\left|\widehat{w}^{(j)}u\right|^{2}\left(1+V_{j}\cos\left(x_{j}(t)+\Delta_{j}(u)\right)\right).$
Now, (83) follows by (75).
When $V_{j}=0$, we have
$\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1}{2}}$
On the other hand, when $V_{j}\neq 0$, $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$, considered as $\frac{\pi}{\omega_{j}}$-periodic function of $t$, oscillates monotonically between the values
$\max\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1+V_{j}}{2}}$ (85)
and
$\min\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1-V_{j}}{2}}.$ (86)
The maximum value is achieved if and only if $x_{j}(t)+\Delta_{j}(u)$ is an even multiple di $\pi$ and the minimum value is achieved if and only if $x_{j}(t)+\Delta_{j}(u)$ is an odd multiple di $\pi$.
Remark 53.
By using $t=0$ in (84), we obtain
$\left\|Q_{j0}(0)u\right\|_{2}=f_{j}\left|\widehat{w}^{(j)}u\right|\sqrt{2\left(1+V_{j}\cos\left(\delta_{j}+2\gamma_{j}(u)\right)\right)}.$
and this is norm of the projection of $u\in\mathbb{R}^{n}$ onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$ (remind Remark 44). The norm of the projection on each eigenspace is $f_{j}\left|\widehat{w}^{(j)}u\right|$ (see Subsection 7.2).
7.6. The functions $f_{VW}$, $f_{VW}^{\max}$ and $f_{VW}^{\min}$
For the analysis of the norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$, as well as for the analysis of the oscillating terms in Section 9, we introduce the functions $f_{VW}$, $f_{VW}^{\max}$ and $f_{VW}^{\min}$, where $V$ and $W$ are parameters in the interval $[0,1)$. (In this subsection, $V$ is NOT the matrix $V$, with columns the Jordan basis, and W is NOT the inverse of the matrix $V$). We use the symbols $V$ and $W$ because we have in mind the values $V_{j},W_{j}\in[0,1)$.
The function $f_{VW}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is given by
$f_{VW}\left(\alpha,x\right):=\frac{1+V\cos\left(x+\alpha\right)}{1-W\cos\alpha},\ \left(\alpha,x\right)\in\mathbb{R}^{2}.$ (87)
For $x\in\mathbb{R}$, let
$U_{VW}(x):=V\mathrm{e}^{\sqrt{-1}x}+W.$
If $U_{VW}(x)=0$, i.e. $V=W=0$ or $V=W\neq 0$ and $x$ is an odd multiple of $\pi$, then
$f_{VW}\left(\alpha,x\right)=1,\ \alpha\in\mathbb{R}.$
The next lemma states the form of the $2\pi$-periodic function $f_{VW}(\cdot,x)$, for a given $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$.
Lemma 54.
Let $V,W\in[0,1)$ and let $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$. The function $f_{VW}(\cdot,x)$ is not constant and it oscillates monotonically between its maximum value $f_{VW}(\alpha^{\max}_{VW}(x),x)$ and minimum value $f_{VW}(\alpha^{\min}_{VW}(x),x)$, where $\alpha^{\max}_{VW}(x)$ and $\alpha^{\min}_{VW}(x)$ are given by
$\alpha^{\max}_{VW}(x)=\arcsin\frac{VW\sin x}{|U_{VW}(x)|}-\theta_{VW}(x)$
and
$\alpha^{\min}_{VW}(x)=\pi-\arcsin\frac{VW\sin x}{|U_{VW}(x)|}-\theta_{VW}(x)$
with $\theta_{VW}(x)$ the angle in the polar form of $U_{VW}(x)$.
Proof.
In this proof, to simplify the notation, we write $U(x)$ for $U_{VW}(x)$ and $\theta(x)$ for $\theta_{VW}(x)$.
We determine the stationary points of $f_{VW}(\cdot,x)$. By deriving $f_{VW}$ with respect to the argument $\alpha$, we obtain
$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)-W\sin\alpha+VW\sin x}{\left(1-W\cos\alpha\right)^{2}},\ \alpha\in\mathbb{R}.$
Thus, $\alpha\in\mathbb{R}$ is a stationary point of $f_{VW}(\cdot,x)$ if and only if
$V\sin(x+\alpha)+W\sin\alpha=VW\sin x,$ (88)
which can equivalently rewritten as
$\mathrm{Im}\left(U(x)\mathrm{e}^{\sqrt{-1}\alpha}\right)=VW\sin x$
since
$V\sin(x+\alpha)+W\sin\alpha=\mathrm{Im}\left(U(x)\mathrm{e}^{\sqrt{-1}\alpha}\right)$ (89)
By writing $U(x)=|U(x)|\mathrm{e}^{\sqrt{-1}\theta(x)}$ in polar form, we obtain
$\sin(\alpha+\theta(x))=\frac{VW\sin x}{|U(x)|}$
and then the function $f_{VW}(\cdot,x)$ has the stationary points
$\alpha=\arcsin\frac{VW\sin x}{|U(x)|}-\theta(x)+2\pi k$ (90)
and
$\alpha=\pi-\arcsin\frac{VW\sin x}{|U(x)|}-\theta(x)+2\pi k,$ (91)
where $k\in\mathbb{Z}$. In Appendix A, we show that
$\left|\frac{VW\sin x}{|U(x)|}\right|=\frac{|VW\sin x|}{|U(x)|}<1.$ (92)
The maximum and the minimum points of $f_{VW}(\cdot,x)$ are these stationary points. Since (recall (89))
$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=|U(x)|\frac{-\sin(\alpha+\theta(x))+\frac{VW\sin x}{|U(x)|}}{\left(1-W\cos\alpha\right)^{2}},\ \alpha\in\mathbb{R},$
the second derivative of $f_{VW}$ with respect to the argument $\alpha$ is
$\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}\left(\alpha,x\right)=|U(x)|\frac{-\cos(\alpha+\theta(x))}{\left(1-W\cos\alpha\right)^{2}},$
when $\alpha$ is stationary point. Thus (90) are maximum points and (91) are minimum points. ∎
Remark 55.
Observe that the functions $\alpha_{VW}^{\max}$ and $\alpha_{VW}^{\min}$ are differentiable at every point where they are defined, i.e. at every point $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$. This is a consequence of the fact that the argument of $\mathrm{arcsin}$ in the definitions of $\alpha_{VW}^{\max}$ and $\alpha_{VW}^{\min}$ in Lemma 54 is neither $-1$ nor $1$, as can be seen in (92).
The functions $f_{VW}^{\max},f_{VW}^{\min}:\mathbb{R}\rightarrow\mathbb{R}$ are given by
$f_{VW}^{\max}(x)=\max\limits_{\alpha\in\mathbb{R}}f_{VW}\left(\alpha,x\right),\ x\in\mathbb{R},$ (93)
and
$f_{VW}^{\min}(x)=\min\limits_{\alpha\in\mathbb{R}}f_{VW}\left(\alpha,x\right),\ x\in\mathbb{R}.$ (94)
If $V=0$ or $W=0$, then
$f_{VW}^{\max}(x)=\frac{1+V}{1-W},\ x\in\mathbb{R},$ (95)
and
$f_{VW}^{\min}(x)=\frac{1-V}{1+W},\ x\in\mathbb{R},$
are constant functions. Thus, assume $V\neq 0$ and $W\neq 0$. The forms of the $2\pi$-periodic functions $f_{VW}^{\max}$ and $f_{VW}^{\min}$ are stated in the next Lemma.
Lemma 56.
Let $V,W\in[0,1)$ such that $V\neq 0$ and $W\neq 0$.
The function $f_{VW}^{\max}$ oscillates monotonically between the values
$\max\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1-W}$
and
$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\left\{\begin{array}[]{l}\frac{1-V}{1-W}\text{\ if\ }V\leq W\\ \\ \frac{1+V}{1+W}\text{\ if\ }V\geq W.\end{array}\right.$
The maximum value is achieved if and only if $x$ is an even multiple of $\pi$ and the minimum value is achieved if and only if $x$ is an odd multiple of $\pi$.
The function $f_{VW}^{\min}$ oscillates monotonically between the values
$\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\left\{\begin{array}[]{l}\frac{1+V}{1+W}\text{\ if\ }V\leq W\\ \\ \frac{1-V}{1-W}\text{\ if\ }V\geq W\end{array}\right.$
and
$\min\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1+W}.$
The maximum value is achieved if and only if $x$ is an odd multiple of $\pi$ and the minimum value is achieved if and only if $x$ is an even multiple of $\pi$.
Proof.
We determine the stationary points of two-variable function $f_{VW}$. We have, for $(\alpha,x)\in\mathbb{R}^{2}$,
$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)-W\sin\alpha+VW\sin x}{\left(1-W\cos\alpha\right)^{2}}$
$\frac{\partial f_{VW}}{\partial x}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)}{1-W\cos\alpha}$ (98)
Hence, $(\alpha,x)\in\mathbb{R}^{2}$ is a stationary point of $f_{VW}$ if and only if
$-V\sin(x+\alpha)-W\sin\alpha+VW\sin x=0$
$\sin(x+\alpha)=0,$
which is equivalent to $\alpha$ and $x$ being multiples of $\pi$.
In the following table, we see the values of
$f_{VW}(\alpha,x)=\frac{1+V\cos(x+\alpha)}{1-W\cos\alpha}$
at the stationary points $(\alpha,x)$.
$x$ even multiple of $\pi$ $x$ odd multiple of $\pi$
$\alpha$ even multiple of $\pi$ $\frac{1+V}{1-W}$ $\frac{1-V}{1-W}$
$\alpha$ odd multiple of $\pi$ $\frac{1-V}{1+W}$ $\frac{1+V}{1+W}$
Observe that
$\max\limits_{x\in\mathbb{R}}f^{\max}_{VW}(x)=\max\limits_{(\alpha,x)\in\mathbb{R}^{2}}f_{VW}(\alpha,x)\text{\ \ and\ \ }\min\limits_{x\in\mathbb{R}}f^{\min}_{VW}(x)=\min\limits_{(\alpha,x)\in\mathbb{R}^{2}}f_{VW}(\alpha,x)$
are values of stationary points of $f_{VW}$. Since
$\frac{1+V}{1-W}\text{\ \ and\ \ }\frac{1-V}{1+W}$
are, respectively, the maximum and the minimum among the values in (99), we conclude that
$\max\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1-W}\text{\ \ \ and\ \ \ }\min\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1+W}$ (100)
with both the maximum and minimum achieved for $x$ being an even multiple of $\pi$.
Observe that
$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)\text{\ and \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)$ (101)
also are values of stationary points of $f_{VW}$. In fact, for $x\in\mathbb{R}$, we have (remind Lemma 54 and Remark 55)
$\left(f_{VW}^{\max}\right)^{\prime}\left(x\right)$ $=$ $\frac{d}{dx}f_{VW}(\alpha^{\max}_{VW}(x),x)$
$=$ $\frac{\partial f_{VW}}{\partial\alpha}(\alpha^{\max}_{VW}(x),x)(\alpha^{\max}_{VW})^{\prime}(x)+\frac{\partial f_{VW}}{\partial x}(\alpha^{\max}_{VW}(x),x)$
$=$ $\frac{\partial f_{VW}}{\partial x}(\alpha^{\max}_{VW}(x),x).$
and, analogously,
$\left(f_{VW}^{\min}\right)^{\prime}\left(x\right)=\frac{\partial f_{VW}}{\partial x}(\alpha^{\min}_{VW}(x),x).$
Therefore, the values (101) are those in (99) different from (100), i.e. those in the second column.
Indeed, when $V=W$, $x$ cannot be an odd multiple of $\pi$ in the previous argument, since $\alpha^{\max}_{VW}(x)$ and $\alpha^{\min}_{VW}(x)$ are not defined in this case. However, when $V=W$, we can still conclude that the values in (101) are stationary points of $f_{VW}$. In fact, for $x\in\mathbb{R}$ that is not an odd multiple of $\pi$, the previous argument works and then $\left(f_{VW}^{\max}\right)^{\prime}(x)$ and $\left(f_{VW}^{\min}\right)^{\prime}(x)$ are zero only when $x$ is a multiple of $\pi$, indeed an even multiple of $\pi$. Thus, the maximum and minimum values of $\left(f_{VW}^{\max}\right)(x)$ and $\left(f_{VW}^{\min}\right)(x)$ as $x$ varies can be achieved only for $x$ being a multiple of $\pi$. Now, we immediately conclude that the values in (101) are stationary points of $f_{VW}$.
When $V=W$, the values in the second column of (100) are both equal to $1$ and then the values (101) are both equal to $1$ and both achieved for $x$ being an odd multiple of $\pi$.
Suppose $V\neq W$. Since
$\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}\left(\alpha,x\right)=\frac{-V\cos(x+\alpha)-W\cos\alpha}{(1-W\cos\alpha)^{2}}$ (102)
at a stationary point $(\alpha,x)$, we obtain the following table of the values of $\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}(\alpha,x)$ at stationary points.
$x$ even multiple of $\pi$ $x$ odd multiple of $\pi$
$\alpha$ even multiple of $\pi$ $-\frac{V+W}{(1-W)^{2}}$ $\frac{V-W}{(1-W)^{2}}$
$\alpha$ odd multiple of $\pi$ $\frac{V+W}{(1+W)^{2}}$ $\frac{-V+W}{(1+W)^{2}}$
where a positive value indicates that $\alpha$ is a minimum point of the function $f_{VW}(\cdot,x)$ and then $f_{VW}(\alpha,x)$ is a value of $f^{\min}_{VW}$, while a negative value indicates that $\alpha$ is a maximum point and then $f_{VW}(\alpha,x)$ is a value of $f^{\max}_{VW}$. This confirms (100) and shows that if $V<W$, then
$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1-V}{1-W}\text{\ \ and\ \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1+V}{1+W},$
and if $V>W$, then
$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1+W}\text{\ \ and\ \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1-W}.$
with both the maximum and minimum achieved for $x$ being an odd multiple of $\pi$. ∎
Remark 57.
When $x$ is an even multiple of $\pi$, $f_{VW}^{\max}(x)$ and $f_{VW}^{\min}(x)$ are at their maximum value and minimum value, respectively. On the other hand, when $x$ is an odd multiple of $\pi$, $f_{VW}^{\max}(x)$ and $f_{VW}^{\min}(x)$ are at their minimum value and maximum value, respectively.
7.7. The norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$
Next proposition provides an expression for the norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$.
Proposition 58.
Assume that $A$ is a real matrix. Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex. We have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1-W_{j}^{2}}{4}f_{V_{j}W_{j}}^{\max}\left(x_{j}(t)\right)},$ (103)
where $f_{V_{j}W_{j}}^{\max}$ is defined in the previous Subsection 7.6 and $x_{j}(t)$ is defined in (81).
Proof.
By (71), Proposition 48 and Proposition 52, we have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}},$
where $\mathcal{W}_{j}$ is the ellipse $\mathcal{E}_{j}$ of Proposition 48 when $n=2$ and $\mathcal{W}_{j}$ is the union of $\mathcal{E}_{j}$ and its interior when $n>2$. We now show that (103) holds.
Suppose $n>2$. By using, in the complex plane, the cartesian coordinate with axes the axes of the ellipse $\mathcal{E}_{j}$, the set $\mathcal{W}_{j}$ is the set of point $(x,y)$ such that
$\left(\frac{x}{\sigma_{j}}\right)^{2}+\left(\frac{y}{\mu_{j}}\right)^{2}\leq 1,$
where $\sigma_{j}$ and $\mu_{j}$ are the lenghts of the axes of the ellipse. Thus
$\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}$$\Leftrightarrow$ $\left(\frac{\rho\cos(\theta-\theta_{j})}{\sigma_{j}}\right)^{2}+\left(\frac{\rho\sin(\theta-\theta_{j})}{\mu_{j}}\right)^{2}\leq 1$
$\Leftrightarrow$ $\rho^{2}\leq R_{j}\left(\theta\right)^{2},$
where
$R_{j}\left(\theta\right):=\sqrt{\frac{\sigma_{j}^{2}\mu_{j}^{2}}{\mu^{2}_{j}\cos^{2}(\theta-\theta_{j})+\sigma^{2}_{j}\sin^{2}(\theta-\theta_{j})}}.$ (104)
Thus
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
$=\max\limits_{\theta\in[0,2\pi]}\max\limits_{\rho\in\left[0,R_{j}\left(\theta\right)\right]}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
$=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}.$
In the case $n=2$, we have
$\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}\Leftrightarrow\rho^{2}=R_{j}\left(\theta\right)^{2}$
and then
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
in this case as well.
Recall (76). In (104), we have
$\sigma_{j}^{2}\mu_{j}^{2}=\frac{1-W^{2}_{j}}{4}$
and, by introducing $\alpha=2(\theta-\theta_{j})$,
$\mu^{2}_{j}\cos^{2}(\theta-\theta_{j})+\sigma^{2}_{j}\sin^{2}(\theta-\theta_{j})$
$=\mu^{2}_{j}\cos^{2}\frac{\alpha}{2}+\sigma^{2}_{j}\sin^{2}\frac{\alpha}{2}$
$=\mu^{2}_{j}\frac{1+\cos\alpha}{2}+\sigma_{j}^{2}\frac{1-\cos\alpha}{2}$
$=\frac{\mu^{2}_{j}+\sigma^{2}_{j}+\left(\mu^{2}_{j}-\sigma^{2}_{j}\right)\cos\alpha}{2}$
$=\frac{1-W_{j}\cos\alpha}{2}.$
Now, we can write
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
$=\max\limits_{\alpha\in[-2\theta_{j},2(2\pi-\theta_{j})]}\sqrt{\frac{1-W_{j}^{2}}{2\left(1-W_{j}\cos\alpha\right)}}\cdot\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\alpha\right)}{2}}$
$=\max\limits_{\alpha\in[0,2\pi]}\sqrt{\frac{1-W_{j}^{2}}{2\left(1-W_{j}\cos\alpha\right)}}\cdot\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\alpha\right)}{2}},$
i.e. (103) by recalling the definition (93)-(87) of $f^{\max}_{V_{j}W_{j}}$. ∎
Recall Lemma 56 (and the observation just before it) and the previous Proposition 58. When $V_{j}=0$, we have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1+W_{j}}{4}}.$
When $W_{j}=0$, we have
$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1+V_{j}}{4}}.$
On the other hand, when $V_{j}\neq 0$ and $W_{j}\neq 0$, $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$, considered as a $\frac{\pi}{\omega_{j}}$-periodic function of $t$, oscillates monotonically between the values
$\max\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{(1+W_{j})(1+V_{j})}{4}}$ (105)
and
$\min\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{a(V_{j},W_{j})}{4}},$ (106)
where
$a(V_{j},W_{j}):=\left\{\begin{array}[]{l}(1+W_{j})(1-V_{j})\text{\ if\ }V_{j}\leq W_{j}\\ \\ (1-W_{j})(1+V_{j})\text{\ if\ }V_{j}\geq W_{j}.\end{array}\right.$
The maximum value is achieved if and only if $x_{j}(t)$ is an even multiple di $\pi$ and the minimum value is achieved if and only if $x_{j}(t)$ is an odd multiple di $\pi$.
8. The generic real case. Part III: asymptotic condition numbers
In this section, we describe the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$ for a real matrix $A$ and in a generic situation for $A$, $y_{0}$ and $\widehat{z}_{0}$. This description is given in the next theorem.
Theorem 59.
Assume that $A$ is a real matrix, $\Lambda_{1}$ consists of either a real simple eigenvalue(i.e. $\Lambda_{1}$ is real), or a single pair of complex conjugate smple eigenvalues (i.e. $\Lambda_{1}$ is complex), and $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$. We have (recall Subsections 6.3 and 6.4 )
$K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{g_{1}\left(t,\widehat{z}_{0}\right)}{g_{1}\left(t,\widehat{y}_{0}\right)}$ (107)
and
$K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{g_{1}\left(t\right)}{g_{1}\left(t,\widehat{y}_{0}\right)}.$ (108)
Proof.
By reminding that $v^{(1)}w^{(1)}$ is the projection onto the eigenspace of $\lambda_{1}$, the projections of $y_{0}$ and $\widehat{z}_{0}$ onto the eigenspace of $\lambda_{1}$ are
$v^{(1)}w^{(1)}y_{0}=\left(w^{(1)}y_{0}\right)v^{(1)}\neq 0$
and
$v^{(1)}w^{(1)}\widehat{z}_{0}=\left(w^{(1)}\widehat{z}_{0}\right)v^{(1)}\neq 0,$
respectively, and then $j^{\ast}=j^{\ast\ast}=1$. By points 2 of Remarks 33 and 35, we have
$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{10}(t)\widehat{z}_{0}\right\|}{\left\|Q_{10}(t)\widehat{y}_{0}\right\|}.$
and
$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{10}(t)\right\|}{\left\|Q_{10}(t)\widehat{y}_{0}\right\|}.$
Now, (107) and (108) follows by (68) and (69). ∎
Remark 60.
The previous theorem considers a generic situation for the real matrix $A$, the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$. In fact, as previously noted, it is a generic situation for $A$ real to have, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ consisting of either a real simple eigenvalue, or a single pair of complex conjugate simple eigenvalues. Moreover it is a generic situation for $y_{0}$ to satisfy $w^{(1)}y_{0}\neq 0$ and it is a generic situation for $\widehat{z}_{0}$ to satisfy $w^{(1)}\widehat{z}_{0}\neq 0$.
Now, we describe the result of the previous theorem by separating the two cases of $\Lambda_{1}$real and $\Lambda_{1}$ complex.
  • If $\Lambda_{1}$ is real, then
    $K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
    $K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
    by Lemma 45.
  • If $\Lambda_{1}$ is complex, then
    $K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}$ (109)
    $K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}$ (110)
    by Lemma 46.
8.1. The oscillation scale factors and the oscillating terms
Consider the case where $\Lambda_{1}$ is complex. We can rewrite (109) and (110) as
$K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z_{0}})\cdot\mathrm{OT}(t,y_{0},\widehat{z_{0}})$
$K_{\infty}(t,y_{0})=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
where the constants
$\mathrm{OSF}(y_{0},\widehat{z_{0}}):=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$ (111)
$\mathrm{OSF}(y_{0}):=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$ (112)
are called the oscillation scale factors and the functions
$t\mapsto\mathrm{OT}(t,y_{0},\widehat{z_{0}}):=\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|},\ t\in\mathbb{R},$ (113)
$t\mapsto\mathrm{OT}(t,y_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|},\ t\in\mathbb{R},$ (114)
which are periodic functions of period $\frac{\pi}{\omega_{1}}$, are called the oscillating terms.
Remark 61.
The oscillation scale factor $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ and the oscillating term $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ are two completely unrelated quantities, separating moduli and angles in the polar forms of $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$. Specifically, $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ depends on $y_{0}$ and $\widehat{z}_{0}$ through the moduli of the complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$. On the other hand, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ depends on $y_{0}$ and $\widehat{z}_{0}$ through the angles $\gamma_{1}(\widehat{y}_{0})$ and $\gamma_{1}(\widehat{z}_{0})$ in these polar forms (refer to the definition of $\widehat{\Theta}(t,u)$ in Lemma 46).
Similarly, $\mathrm{OSF}(y_{0})$ depends on $y_{0}$ through the modulus of the complex number $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\mathrm{OT}\left(t,y_{0}\right)$ depends on $y_{0}$ through the angle $\gamma_{1}(\widehat{y}_{0})$ in this polar form.
9. The generic real case. Part IV: oscillation scale factors and oscillating terms in case of the euclidean norm
In this section, we study the oscillating scale factors and the oscillating terms in case of the Euclidean norm as vector norm.
9.1. The oscillation scale factor $\mathrm{OSF}(y_{0},\widehat{z}_{0})$
The oscillating scale factor
$\mathrm{OSF}(y_{0},\widehat{z}_{0})$ in (111) is the ratio of the modules of the non-zero complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$. In case of the Euclidean norm as vector norm, such complex numbers can be arbitrary non-zero points in the set $\mathcal{W}_{1}$ of Proposition 48.
Thus, as $y_{0}$ and $\widehat{z}_{0}$ vary, $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ can be any number in $(0,+\infty)$ when $n>2$ and any number in
$\left[\frac{\mu_{1}}{\sigma_{1}},\frac{\sigma_{1}}{\mu_{1}}\right]=\left[\sqrt{\frac{1-W_{1}}{1+W_{1}}},\sqrt{\frac{1+W_{1}}{1-W_{1}}}\right],$
where $\sigma_{1}$ and $\mu_{1}$ are the lengths of the semi-axes of the ellipse $\mathcal{E}_{1}$ of Proposition 48, when $n=2$.
Remark 62.
By (80), we can express $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ in terms of the components $c_{1}(u)$ and $d_{1}(u)$ of $u=\widehat{y}_{0},\widehat{z}_{0}$ along the first two right singular vectors of the matrix $R_{1}$ of Proposition 48. We have
$\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{(1+W_{1})c_{1}(\widehat{z}_{0})^{2}+(1-W_{1})d_{1}(\widehat{z}_{0})^{2}}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$ (115)
9.2. The oscillating term $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$
In case of the Euclidean norm as vector norm, by reminding Proposition 52, we can write the oscillating term (113) as
$\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}\left(\alpha,x\right)},$
where $f_{V_{1}V_{1}}$ is the function defined in (87),
$\alpha=x_{1}(t)+\Delta_{1}(\widehat{y}_{0})+\pi,$
with $x_{1}(t)$ and $\Delta_{1}(\widehat{y}_{0})$ defined in (81) and (82), and
$x=\Delta_{1}(\widehat{z}_{0})-\Delta_{1}(\widehat{y}_{0})-\pi=2\left(\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})\right)-\pi.$
In the case $V_{1}=0$, we have
$\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1.$
Therefore, we assume $V_{1}\neq 0$. We are interested to the maximum and minimum values
$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}^{\max}\left(x\right)}$ (116)
and
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}^{\min}\left(x\right)}$ (117)
of the oscillating term, where $f_{V_{1}V_{1}}^{\max}$ and $f_{V_{1}V_{1}}^{\min}$ are defined in (93) and (94). Observe that (116) and (117) hold since
$\alpha=x_{1}(t)+\Delta_{1}(\widehat{y}_{0})+\pi=2(\omega_{1}t+\gamma_{1}(\widehat{y}_{0}))+\delta_{1}+\pi$
varies in $\mathbb{R}$, as $t$ varies in $\mathbb{R}$.
Given $y_{0}$, we do not determine (116) and (117) for any $\widehat{z}_{0}$ but we look only at their maximum and minimum values, by varying $\widehat{z}_{0}$. The next theorem gives such maximum and minumum values. They are independent of $y_{0}$.
Theorem 63.
Assume that $A$ is a real matrix, $\Lambda_{1}$ is complex and $w^{(1)}y_{0}\neq 0$. Assume $V_{1}\neq 0$.
We have
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1+V_{1}}{1-V_{1}}}$
and
$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1.$
The maximum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$ and the minimum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$.
Moreover, we have
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$
and
$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1-V_{1}}{1+V_{1}}},$
The maximum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is multiple of $\pi$ and the minimum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$.
Proof.
As $\widehat{z}_{0}$ varies in $\mathbb{R}^{n}$ with $\|\widehat{z}_{0}\|_{2}=1$ and $w^{(1)}\widehat{z}_{0}\neq 0$, $\widehat{w}^{(1)}\widehat{z}_{0}$ can be an arbitrary non-zero point in the set $\mathcal{W}_{1}$ of Proposition 48. Then, $\gamma_{1}\left(\widehat{z}_{0}\right)$ varies in $[0,2\pi]$ and then
$x=2\left(\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})\right)-\pi$
in (116)-(117) varies in $[-2\gamma_{1}\left(\widehat{y}_{0}\right)-\pi,-2\gamma_{1}\left(\widehat{y}_{0}\right)+3\pi]$, over two periods of the functions $f_{V_{1}V_{1}}^{\max}$ and $f_{V_{1}V_{1}}^{\min}$. We conclude that
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\max\limits_{x\in\mathbb{R}}f^{\max}_{V_{1}V_{1}}(x)}$
$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\min\limits_{x\in\mathbb{R}}f^{\max}_{V_{1}V_{1}}(x)}$
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\max\limits_{x\in\mathbb{R}}f^{\min}_{V_{1}V_{1}}(x)}$
$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\min\limits_{x\in\mathbb{R}}f^{\min}_{V_{1}V_{1}}(x)}.$
Now, the theorem follows by Lemma 56. ∎
Remark 64.
  • 1.
    If $V_{1}$ is not close to $1$, then, for any $y_{0}$ and $\widehat{z}_{0}$, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ does not assume large or small values, as $t$ varies. In fact, we have
    $\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}}.$
  • 2.
    When $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$, we have
    $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1=\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$
    and then
    $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1,$
    as it can be directly verified by the definition of $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ and Proposition 52.
  • 3.
    When $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, we have
    $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1+V_{1}}{1-V_{1}}}$ (118)
    and
    $\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1-V_{1}}{1+V_{1}}}$ (119)
    with $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ at its maximum value and $\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ at its minimum value, as $\widehat{z}_{0}$ varies.
  • 4.
    If $V_{1}$ is close to $1$ and $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies. This follows by (118) and (119).
  • 5.
    The condition $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$ is equivalent to saying that the vectors $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$ are orthogonal in the complex plane, i.e.
    $(1+W_{1})c_{1}(\widehat{y}_{0})c_{1}(\widehat{z}_{0})+(1-W_{1})d_{1}(\widehat{y}_{0})d_{1}(\widehat{z}_{0})=0$
    by reminding (79).
    The condition $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$ is equivalent to saying that the vectors $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$ are parallel in the complex plane, i.e.
    $c_{1}(\widehat{y}_{0})d_{1}(\widehat{z}_{0})-d_{1}(\widehat{y}_{0})c_{1}(\widehat{z}_{0})=0.$
  • 6.
    Some specific situations are worth considering.
    • a)
      If $y_{0}$ is orthogonal to the first right singular vector of the matrix $R_{1}$ of Proposition 48 and $\widehat{z}_{0}$ is orthogonal to the second right singular vector (i.e. $c_{1}(\widehat{y}_{0})=0$ and $d_{1}(\widehat{z}_{0})=0$), then, by previous points 5 and 3, (118)-(119) hold. Moreover, by (115) we have
      $\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{1+W_{1}}{1-W_{1}}}\cdot\frac{|c_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$
      If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies. If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ large. Regarding the asymptotic condition number
      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0}),$
      it is expected that it assumes large values, as $t$ varies, if $V_{1}$ is close to $1$ or $W_{1}$ is close to $1$: we have
      $\max\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})=\sqrt{\frac{(1+W_{1})(1+V_{1})}{(1-W_{1})(1-V_{1})}}\cdot\frac{|c_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$ (120)
    • b)
      If $y_{0}$ is orthogonal to the second right singular vector of $R_{1}$ and $\widehat{z}_{0}$ is orthogonal to the first right singular vector (i.e. $d_{1}(\widehat{y}_{0})=0$ and $c_{1}(\widehat{z}_{0})=0$), then (118)-(119) hold again. Moreover, we have
      $\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{1-W_{1}}{1+W_{1}}}\cdot\frac{|d_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$
      If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies. If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ small. Regarding the asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$, it is expected that it assumes large values, as $t$ varies, if the ratio $\frac{1-W_{1}}{1-V_{1}}$ is large: we have
      $\max\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})=\sqrt{\frac{(1-W_{1})(1+V_{1})}{(1+W_{1})(1-V_{1})}}\cdot\frac{|d_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$ (121)
    • c)
      If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the first right singular vector of $R_{1}$ (i.e. $c_{1}(\widehat{y}_{0})=0$ and $c_{1}(\widehat{z}_{0})=0$), then, by previous point 2 and 5, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$. Moreover, we have
      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})=\frac{|d_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$
    • d)
      If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the second right singular vector of $R_{1}$ (i.e. $d_{1}(\widehat{y}_{0})=0$ and $d_{1}(\widehat{z}_{0})=0$), then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$. Moreover, we have
      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})=\frac{|c_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$
9.3. The oscillation scale factor $\mathrm{OSF}(y_{0})$
The oscillating scale factor $\mathrm{OSF}(y_{0})$ in (112) is the reciprocal of the modulus of the complex number $\widehat{w}^{(1)}\widehat{y}_{0}$. In case of the Euclidean norm as vector norm, such complex number can be an arbitrary non-zero point in the set $\mathcal{W}_{1}$ of Proposition 48.
Thus, by varying $y_{0}$, $\mathrm{OSF}(y_{0})$ can be any number in
$\left[\frac{1}{\sigma_{1}},+\infty\right)=\left[\sqrt{\frac{2}{1+W_{1}}},+\infty\right)$
when $n>2$ and any number in
$\left[\frac{1}{\sigma_{1}},\frac{1}{\mu_{1}}\right]=\left[\sqrt{\frac{2}{1+W_{1}}},\sqrt{\frac{2}{1-W_{1}}}\right].$
when $n=2$.
Remark 65.
By (80), we can express $\mathrm{OSF}(y_{0})$ in terms of the components $c_{1}(\widehat{y}_{0})$ and $d_{1}(\widehat{y}_{0})$ of $\widehat{y}_{0}$ along the first two right singular vectors of the matrix $R_{1}$. We have
$\mathrm{OSF}(y_{0})=\sqrt{\frac{2}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$ (122)
9.4. The oscillating term $\mathrm{OT}(t,y_{0})$
In case of the Euclidean norm as vector norm, by reminding Propositions 52 and 58, we can write the oscillating term (114) as
$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\frac{f_{V_{1}W_{1}}^{\max}(x_{1}(t))}{1+V_{1}\cos(x_{1}(t)+\Delta_{1}(\widehat{y_{0}}))}},$ (123)
where $x_{1}(t)$ and $\Delta_{1}(\widehat{y}_{0})$ are defined in (81) and (82).
In the case $V_{1}=0$, we have
$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1+W_{1}}{2}}$
by reminding (95). Therefore, we assume $V_{1}\neq 0$. We are interested to the maximum and minimum values
$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\max\limits_{x\in\mathbb{R}}\frac{f_{V_{1}W_{1}}^{\max}(x)}{1+V_{1}\cos(x+\Delta_{1}(\widehat{y_{0}}))}}$ (124)
and
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\min\limits_{x\in\mathbb{R}}\frac{f_{V_{1}W_{1}}^{\max}(x)}{1+V_{1}\cos(x+\Delta_{1}(\widehat{y_{0}}))}}.$ (125)
of the oscillating term. Observe that (124) and (125) hold since $x_{1}(t)$ varies in $\mathbb{R}$, as $t$ varies in $\mathbb{R}$. In the case $V_{1}\neq 0$ and $W_{1}=0$, we have
$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1+V_{1}}{2(1-V_{1})}}$
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1}{2}},$
by reminding (95). Observe that, for $V_{1}=0$ or $W_{1}=0$, such maximum and minimum are independent of $y_{0}$.
We proceed by assuming $V_{1}\neq 0$ and $W_{1}\neq 0$. We do not determine (124) and (125) for any $y_{0}$, but we look only at their maximum and minimum values as $y_{0}$ varies. The next theorem gives such maximum and minumum values.
Theorem 66.
Assume that $A$ is a real matrix and $\Lambda_{1}$ is complex. Assume $V_{1}\neq 0$ and $W_{1}\neq 0$.
We have
$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\max}(V_{1},W_{1}):=\sqrt{\frac{(1+W_{1})(1+V_{1})}{2(1-V_{1})}}$ (126)
and
$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\left\{\begin{array}[]{l}\sqrt{\frac{1+W_{1}}{2}\cdot\frac{1-V_{1}^{2}}{1-Q_{1}V_{1}}}\text{\ if\ }V_{1}\leq\frac{2W_{1}}{1+W_{1}}\\ \\ \sqrt{\frac{(1-W_{1})(1+V_{1})}{2(1-V_{1})}}\text{\ if\ }\frac{2W_{1}}{1+W_{1}}\leq V_{1},\end{array}\right.$ (130)
where
$Q_{1}:=\frac{1}{2}\cdot\frac{V_{1}}{W_{1}}(1+W_{1}).$ (131)
The maximum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$ and the minimum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is a multiple of $\pi$. The angle $\theta_{1}$ is defined in Subsection 7.4. In the case $V_{1}\leq\frac{2W_{1}}{1+W_{1}}$, we have
$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)\leq\sqrt{\frac{(1+W_{1})(1+V_{1})}{2}}.$ (132)
Moreover, we have
$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1+W_{1}}{2}}$ (133)
and
$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\min}(V_{1},W_{1}):=\left\{\begin{array}[]{l}\sqrt{\frac{(1+W_{1})(1-V_{1})}{2(1+V_{1})}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \sqrt{\frac{1-W_{1}}{2}}\text{\ if\ }V_{1}\geq W_{1}.\end{array}\right.$ (137)
The maximum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is a multiple of $\pi$ and the minimum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$.
Proof.
We introduce the function $H:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by
$H(x,\beta)=\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)},\ (x,\beta)\in\mathbb{R}^{2}.$ (138)
Thus, by (123), (124) and (125) we can write
$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H(x_{1}(t),\Delta_{1}(\widehat{y}_{0}))}$ (139)
and
$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H^{\max}(\Delta_{1}(\widehat{y}_{0}))}$
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H^{\min}(\Delta_{1}(\widehat{y}_{0}))},$
where the functions $H^{\max},H^{\min}:\mathbb{R}\rightarrow\mathbb{R}$ are given by
$H^{\max}(\beta):=\max\limits_{x\in\mathbb{R}}H(x,\beta)\text{\ \ and\ \ }H^{\min}(\beta):=\min\limits_{x\in\mathbb{R}}H(x,\beta),\ \beta\in\mathbb{R}.$
Observe that (140) holds since, as $t$ varies, $x_{1}(t)$ in (81) varies in $\mathbb{R}$.
Therefore, we have
$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta))}$
$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta))}$
$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta))}$
$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\min\limits_{\beta\in\mathbb{R}}H^{\min}(\beta))}.$
Observe that (141) hold since, as $y_{0}\in\mathbb{R}^{n}$ with $w^{(1)}y_{0}\neq 0$ varies, $\widehat{w}^{(1)}\widehat{y}_{0}$ can be an arbitrary non-zero point in $\mathcal{W}_{1}$ and then $\gamma_{1}(\widehat{y}_{0})$ varies in $[0,2\pi]$ and then
$\Delta_{1}(\widehat{y}_{0})=2(\gamma_{1}(\widehat{y}_{0})-\theta_{1})$
varies in $[-2\theta_{1},-2\theta_{1}+4\pi]$, over two periods of the functions $H^{\max}$ and $H^{\min}$.
The theorem will follow once we have determined
$\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta),\ \min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta),\ \max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)\text{\ and\ }\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)$
and for which values $\beta$ the maximum and minimum values are achieved.
By Lemma 56, we immediately see that
$\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)$ $=$ $\max\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}H(x,\beta)=\max\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)}$
$=$ $\frac{1+V_{1}}{(1-W_{1})(1-V_{1})}$
with the maximum value achieved if and only if $x$ is an even multiple of $\pi$ (thus $f_{V_{1}W_{1}}^{\max}(x)$ achieves its maximum value
$\frac{1+V_{1}}{1-W_{1}}$
as $x$ varies) and $x+\beta$ is an odd multiple of $\pi$ (thus $1+V_{1}\cos(x+\beta)$ achieves its minimum value $1-V_{1}$ as $x$ and $\beta$ vary). In other words, the maximum is achieved if and only if $\beta$ is an odd multiple of $\pi$. Thus, by (141) and (LABEL:MAXMAX) we have (126) with the maximum achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}=\frac{\Delta_{1}(\widehat{y}_{0})}{2}=\frac{\beta}{2}$ is an odd multiple of $\frac{\pi}{2}$. Remind that $\Delta_{1}(\widehat{y}_{0})=\beta$ in (140).
By Lemma 56, we also immediately see that
$\min\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)$ $=$ $\min\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}H(x,\beta)=\min\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)}$ (146)
$=$ $\left\{\begin{array}[]{l}\frac{1-V_{1}}{(1-W_{1})(1+V_{1})}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \frac{1}{1+W_{1}}\text{\ if\ }V_{1}\geq W_{1},\end{array}\right.$
with the minimum value achieved if and only $x$ is an odd multiple of $\pi$ (thus $f_{V_{1}W_{1}}^{\max}(x)$ achieves its minimum value
$\left\{\begin{array}[]{l}\frac{1-V_{1}}{1-W_{1}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \frac{1+V_{1}}{1+W_{1}}\text{\ if\ }V_{1}\geq W_{1}\end{array}\right.$
as $x$ varies) and $x+\beta$ is an even multiple of $\pi$ (thus $1+V_{1}\cos(x+\beta)$ achieves its maximum value $1+V_{1}$ as $x$ and $\beta$ vary). In other words, the minimum is achieved if and only if $\beta$ is an odd multiple of $\pi$. Thus, by (141) and (LABEL:MINMIN) we have (137) with the minimum achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$.
It remains to determine
$\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)=\min\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}H(x,\beta)\text{\ and\ }\max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)=\max\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}H(x,\beta).$
By using the property
$f_{V_{1}W_{1}}^{\max}(-x)=f_{V_{1}W_{1}}^{\max}(x),\ x\in\mathbb{R},$
(this follows since $\alpha^{\max}_{V_{1}W_{1}}(-x)=-\alpha^{\max}_{V_{1}W_{1}}(x)$, with $\alpha^{\max}_{V_{1}W_{1}}(x)$ defined in Lemma 54), we obtain
$H^{\max}(\pi-\beta)=H^{\max}(\pi+\beta)\text{\ and\ }H^{\min}(\pi-\beta)=H^{\min}(\pi+\beta),\ \beta\in\mathbb{R}.$
This fact shows that we can confine our analysis to the case $\beta\in[0,\pi]$ (remind that $H^{\max}$ and $H^{\min}$ have period $2\pi$).
For the remainder of the proof, we assume $V_{1}\neq W_{1}$. This ensures that the function $\alpha_{V_{1}W}^{\max}$ is defined and differentiable at every point $x\in\mathbb{R}$, by reminding Remark 55. The annoying case $V_{1}=W_{1}$ is treated as a limiting case of the situation $V_{1}\neq W_{1}$. This thanks to the following proposition.
Proposition 67.
Let $F:A\times B\rightarrow\mathbb{R}$ be a continuous function, with $A$ and $B$ metric spaces. If
$\lim\limits_{n\rightarrow\infty}a_{n}=a,\ \lim\limits_{n\rightarrow\infty}b_{n}=b\text{\ \ and\ \ }F(a_{n},b_{n})=\max\limits_{x\in A}F(x,b_{n}),$
where $a_{n},a\in A$ and $b_{n},b\in B$, then
$F(a,b)=\max\limits_{x\in A}F(x,b).$
Proof.
Let $x\in A$. We have
$F(a,b)$ $=$ $F(x,b)+F(x,b_{n})-F(x,b)+F(a_{n},b_{n})-F(x,b_{n})+F(a,b)-F(a_{n},b_{n})$
$\geq$ $F(x,b)+F(x,b_{n})-F(x,b)+F(a,b)-F(a_{n},b_{n})$
since
$F(a_{n},b_{n})-F(x,b_{n})\geq 0.$
By letting $n\rightarrow\infty$, we obtain
$F(a,b)\geq F(x,b)$
for the continuity of $F$. ∎
For a given $\beta\in[0,\pi]$, the values $H^{\max}(\beta)$ and $H^{\min}(\beta)$ are values $H(x,\beta)$, where $x$ is stationary point of the function $H(\cdot,\beta):\mathbb{R}\rightarrow\mathbb{R}$ and satisfies
$\frac{\partial H}{\partial x}(x,\beta)=0.$ (148)
Since
$\frac{\partial H}{\partial x}(x,\beta)=\frac{\partial}{\partial x}\left(\frac{f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)}{1+V_{1}\cos(x+\beta)}\right)$
$=\frac{1}{(1+V_{1}\cos(x+\beta))^{2}}$
$\ \ \ \ \cdot\left(\left(\underset{=0}{\underbrace{\frac{\partial f_{V_{1}W_{1}}}{\partial\alpha}(\alpha^{\max}_{V_{1}W_{1}}(x),x)}}(\alpha^{\max}_{V_{1}W_{1}})^{\prime}(x)+\frac{\partial f_{V_{1}W_{1}}}{\partial x}(\alpha^{\max}_{V_{1}W_{1}}(x),x)\right)\right.$
$\left.\ \ \ \ \ \ \ \ \ \ \ \cdot(1+V_{1}\cos(x+\beta))\right.$
$\left.\ \ \ \ \ \ \ \ \ \ \ -f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)\left(-V_{1}\sin(x+\beta)\right)\right)$
$=\frac{1}{(1+V_{1}\cos(x+\beta))^{2}}$
$\ \ \ \ \cdot\left(\frac{\partial f_{V_{1}W_{1}}}{\partial x}(\alpha^{\max}_{V_{1}W_{1}}(x),x)(1+V_{1}\cos(x+\beta))+f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)V_{1}\sin(x+\beta)\right),$
we see (remind (87) and (98)) that (148) is equivalent to
$-\sin(x+\alpha^{\max}_{V_{1}W_{1}}(x))+\sin(x+\beta)-V_{1}\sin(\alpha^{\max}_{V_{1}W_{1}}(x)-\beta)=0.$ (149)
By solving this equation for $\beta=0$, we obtain (see Appendix B)
$H^{\max}(0)=\left\{\begin{array}[]{l}\frac{1}{1-W_{1}}\cdot\frac{1-V_{1}^{2}}{1-Q_{1}V_{1}}\text{\ if\ }V_{1}\leq\frac{2W_{1}}{1+W_{1}}\\ \\ \frac{1}{1+W_{1}}\cdot\frac{1+V_{1}}{1-V_{1}}\text{\ if\ }\frac{2W_{1}}{1+W_{1}}\leq V_{1},\end{array}\right.$ (150)
where $Q_{1}$ is defined in (131), and
$H^{\min}(0)=\frac{1}{1-W_{1}}.$ (151)
Moreover, in the case $V_{1}\leq\frac{2W_{1}}{1+W_{1}}$, we have
$H^{\max}(0)\leq\frac{1}{1-W_{1}}(1+V_{1}).$ (152)
The stationary points of $H(\cdot,\beta)$ (i.e. the solutions of (149)) form, as function of $\beta\in[0,\pi]$, a family of continuous branches
$x_{i}:D_{i}\rightarrow\mathbb{R},\ \beta\mapsto x_{i}(\beta),\ i\in I,$
where $I$ is a finite set of indices and $D_{i}$ a closed interval included in $[0,\pi]$. These branches satisfy the differential equation
$\frac{d}{d\beta}\frac{\partial H}{\partial x}(x_{i}(\beta),\beta)=\frac{\partial^{2}H}{\partial x^{2}}(x_{i}(\beta),\beta)x_{i}^{\prime}(\beta)+\frac{\partial^{2}H}{\partial\beta\partial x}(x_{i}(\beta),\beta)=0.$ (153)
The ends of the interval $D_{i}$ are 00, $\pi$ or points $\beta$ where
$\frac{\partial^{2}H}{\partial x^{2}}(x_{i}(\beta),\beta)=0.$
Each branch $x_{i}$, $i\in I$, is continuously differentiable in the interior of $D_{i}$.
The values of $H(\cdot,\beta)$ at the stationary points also form a family of continuous branches
$h_{i}:D_{i}\rightarrow\mathbb{R},\ \beta\mapsto H(x_{i}(\beta),\beta),\ i\in I,$
continuously differentiable in the interior of $D_{i}$.
Since $H^{\max}(\beta)$ and $H^{\min}(\beta)$, $\beta\in[0,\pi]$, are values $H(x_{i}(\beta),\beta)$ for some $i\in I$, we have
$H^{\max}(\beta)=\max\limits_{\begin{subarray}{c}i\in I\\ \beta\in D_{i}\end{subarray}}h_{i}(\beta)\text{\ \ and\ \ }H^{\min}(\beta)=\min\limits_{\begin{subarray}{c}i\in I\\ \beta\in D_{i}\end{subarray}}h_{i}(\beta),\ \beta\in[0,\pi].$ (154)
In Figure 8, we see examples of such branches.
Remark 68.
Observe that the equation (149) can be factorized as
$-2\sin\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}\left(\cos\left(x+\frac{\alpha^{\max}_{V_{1}W_{1}}(x)+\beta}{2}\right)+V_{1}\cos\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}\right)=0$
and then its solutions are the union of the solutions of
$\alpha^{\max}_{V_{1}W_{1}}(x)-\beta\text{\ is a multiple of $2\pi$}$ (155)
and
$\cos\left(x+\frac{\alpha^{\max}_{V_{1}W_{1}}(x)+\beta}{2}\right)+V_{1}\cos\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}=0.$
The branches $h_{i}$ corresponding to solutions of (155) are given by
$h_{i}(\beta)=\frac{1}{1-W_{1}\cos\beta}$
as can be immediately verified by the definition (138) of $H$. In Figure 9, we see examples of such branches $x_{i}$ and $h_{i}$ corresponding to solutions of (155).
Figure 8. Left: branches $x_{i}$ for $(\beta,x)\in[0,\pi]\times[-\pi,\pi]$. Right: branches $h_{i}$ for $\beta\in[0,\pi]$.
Figure 9. Left: branches $x_{i}$ for $(\beta,x)\in[0,\pi]\times[-\pi,\pi]$ corresponding to solutions of (155). Right: branches $h_{i}$ for $\beta\in[0,\pi]$ corresponding to solutions of (155).
Proposition 69.
The branches $h_{i}$, $i\in I$, are either monotonically increasing or decreasing.
Proof.
Let $i\in I$. We show that there are no interiors points of $D_{i}$ where the derivative $h_{i}^{\prime}$ is zero.
For $\beta$ interior point of $D_{i}$, we have
$h_{i}^{\prime}(\beta)$ $=$ $\frac{d}{d\beta}H(x_{i}(\beta),\beta)$
$=$ $\frac{\partial H}{\partial x}(x_{i}(\beta),\beta)x_{i}^{\prime}(\beta)+\frac{\partial H}{\partial\beta}(x_{i}(\beta),\beta)$
$=$ $\frac{\partial H}{\partial\beta}(x_{i}(\beta),\beta)$
$=$ $\frac{V_{1}\sin(x_{i}(\beta)+\beta)}{(1+V_{1}\cos(x_{i}(\beta)+\beta))^{2}}f^{\max}_{V_{1},W_{1}}(x_{i}(\beta)),$
where $f^{\max}_{V_{1},W_{1}}(x_{i}(\beta))>0$ (see Lemma 56, or the definition (87), by reminding $V_{1},W_{1}\in[0,1)$). Then, for $h^{\prime}_{i}(\beta)=0$, we have
$-V_{1}\sin(x+\alpha)-W_{1}\sin\alpha+V_{1}W_{1}\sin x=0$
$-\sin(x+\alpha)+\sin(x+\beta)-V_{1}\sin(\alpha-\beta)=0$
$\sin(x+\beta)=0,$
where $x=x_{i}(\beta)$ and $\alpha=\alpha^{\max}_{V_{1}W_{1}}(x_{i}(\beta))$.
Regarding the equations in (157), observe that:
  • The first equation is the equation for the stationary points $\alpha$ of $f_{V_{1}W_{1}}(\cdot,x_{i}(\beta))$ (recall (88)): $\alpha=\alpha^{\max}_{V_{1}W_{1}}(x_{i}(\beta))$ is a stationary point of $f_{V_{1}W_{1}}(\cdot,x_{i}(\beta))$.

### v2

@@
-Asymptotic condition numbers for linear ODEs
+Asymptotic condition numbers for linear ordinary differential equations
 Abstract.
-We are interested in the (relative) conditioning of the linear problem $y_{0}\mapsto\mathrm{e}^{tA}y_{0}$, i.e.
-the conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector.
+We are interested in the relative conditioning of the problem $y_{0}\mapsto\mathrm{e}^{tA}y_{0}$, i.e., the relative conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector.
 The present paper is a qualitative study of the long-time behavior of this conditioning.
-In other words, we are interested to study the propagation to the solution $y(t)$ of perturbations of the initial value for a linear ODE $y^{\prime}(t)=Ay(t)$, by measuring these perturbations with relative errors instead of absolute errors.
+In other words, we are interested in studying the propagation to the solution $y(t)$ of perturbations of the initial value for a linear ordinary differential equation $y^{\prime}(t)=Ay(t)$, by measuring these perturbations with relative errors.
 We introduce three condition numbers: the first considers a specific initial value and a specific direction of perturbation; the second considers a specific initial value and the worst case by varying the direction of perturbation; and the third considers the worst case by varying both the initial value and the direction of perturbation.
@@
 1. Introduction
-We are interested to understand how a perturbation of the initial value $y_{0}$ of the linear $n-$dimensional Ordinary Differential Equation (ODE)
+We are interested in understanding how a perturbation of the initial value $y_{0}$ of the linear $n$-dimensional Ordinary Differential Equation (ODE)
 $\left\{\begin{array}[]{l}y^{\prime}\left(t\right)=Ay\left(t\right),\ t\in\mathbb{R},\\ y\left(0\right)=y_{0},\end{array}\right$. (1)
-where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$, is propagated to the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time $t$.
-This perturbation, propagating along the solution, is measured by a relative error rather than an absolute error.
-In other words, we study the (relative) conditioning of the linear problem
-$y_{0}\mapsto\mathrm{e}^{tA}y_{0}$ (2)
+where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$, is propagated to the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time interval.
+This perturbation, propagating along the solution, is measured by a relative error.
+In other words, we study the relative conditioning of the problem
+$y_{0}\mapsto y(t)=\mathrm{e}^{tA}y_{0}$ (2)
 for large time $t$.
-The conditioning of the problem (2) has received little attention in literature despite of its significance in the context of ODEs.
-In contrast, the study of the conditioning of the matrix exponential $\mathrm{e}^{tA}$ is quite extensive; see, for example, [10], [15], [9], [11], [14], [2], [16], [3], [5], and [1].
-However, none of these papers examined how the conditioning depends on $t$, a crucial aspect for ODEs.
-This type of analysis, for $A$ normal, was addressed in [12] for the problem (2) and in [7] for the problem $A\mapsto\mathrm{e}^{tA}y_{0}$.
-In the present paper, for the problem (2), we consider a general matrix $A$.
-The present paper complements the work presented in [13].
-The paper [13] offers a non-technical exposition without proofs of the main results found in the present paper.
-It includes motivations, practical examples, and numerical tests.
-Consequently, these aspects are not addressed in the present paper, which focuses on providing the proofs of all results and the necessary theoretical development for obtaining them.
+The relative conditioning of the matrix exponential function, i.e., the relative conditioning of the problem $A\mapsto\mathrm{e}^{A}$, or the problem
+$A\mapsto\mathrm{e}^{tA}$ (3)
+involving the time $t$, has been extensively studied: see [9], [15], [8], [10], [14], [2], [16], [3], [5], and [1].
+In the study of time evolutions, an important aspect is to understand how the relative conditioning depends on $t$.
+Many of the papers cited above have examined this issue for the problem (3).
+For a normal matrix $A$, it is known that the relative condition number of (3) grows linearly with $t$.
+On the other hand, for a general matrix $A$, the exact order of growth with $t$ and its dependence on the matrix $A$ are not known: we have only polynomial lower bounds and exponential upper bounds in $t$ for the relative condition number.
+It is also of interest to study the relative conditioning of the action of the matrix exponential on a vector.
+This is particularly important in the context of ODEs (1), where the solution is given by the action of the matrix exponentials $\mathrm{e}^{tA}$ on a known initial value $y_{0}$.
+In this context, we can consider the relative conditioning of the problems (2) and
+$A\mapsto\mathrm{e}^{tA}y_{0},$ (4)
+which are more important than the problem (3), since $y_{0}$ is involved. Despite the importance, there has been little attention in the literature to the relative conditioning of these problems.
+For the case of $A$ normal, an analysis of the relative conditioning, focused on time $t$, was carried out in [6] for the problem (4).
+The relative conditioning of problem (2) can be perceived, at first glance, as a trivial issue: (2) is a linear problem and its condition number can be immediately determined and computed.
+This perception is especially reinforced when the relative conditioning of (2) is compared with the relative conditioning of the non-linear problems (3) and (4).
+However, the relative conditioning of (2) ceases to be trivial once the time $t$ is taken into account, and we want to understand how it depends on $t$.
+Analyzing how the relative conditioning of (2) depends on $t$ can fill a gap in our understanding of linear dynamics.
+In fact, while it is well understood how the absolute conditioning of (2) depends on $t$, i.e., how absolute errors due to perturbations of $y_{0}$ propagate to $y(t)$ (they are governed by the real parts of the rightmost eigenvalues of $A$ for large $t$, and by the pseudospectra of $A$ for non-large $t$), how the relative conditioning of (2) depends on $t$, i.e., how relative errors propagate, is far less understood, even for large $t$.
+Such an analysis was carried out in [11] for the simple case of $A$ normal.
+In the present paper, we carry out this analysis for a general ODE (1), by considering the relative conditioning for large $t$.
+The analysis is far from trivial, as evidenced by the length of this paper and its sequels [12] and [13].
 1.1. The condition numbers
-[…]
-$\varepsilon=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$
+Suppose that the initial value $y_{0}\neq 0$ in (1) is perturbed to $\widetilde{y}_{0}$ and, as a consequence, the solution $y$ is perturbed to $\widetilde{y}$.
+Let $\left\|\ \cdot\ \right\|$ be an arbitrary vector norm on $\mathbb{C}^{n}$.
+We introduce the normwise relative error
+$\varepsilon:=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$
 of $\widetilde{y}_{0}$ and the normwise relative error
-$\delta\left(t\right)=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$
+$\delta\left(t\right):=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$
 of $\widetilde{y}(t)$. By writing
 $\widetilde{y}_{0}=y_{0}+\varepsilon\|y_{0}\|\widehat{z}_{0},$
-where $\widehat{z}_{0}\in\mathbb{C}^{n}$ with $\|\widehat{z}_{0}\|=1$ is the direction of perturbation, we obtain
+where $\widehat{z}_{0}\in\mathbb{C}^{n}$, with $\|\widehat{z}_{0}\|=1$, is the direction of perturbation, we obtain
 $\delta\left(t\right)=K\left(t,y_{0},\widehat{z}_{0}\right)\cdot\varepsilon,$
@@
 $K\left(t,y_{0},\widehat{z}_{0}\right):=\frac{\left\|\mathrm{e}^{tA}\widehat{z}_{0}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$ (5)
-with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$, is called the directional condition number of the problem (2).
-Along with the condition number (3), we also introduce other two condition numbers: the condition number of the problem (2) (see [4] for the definition of condition number of a general problem) given by
+with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$ the normalized initial value.
+The number $K\left(t,y_{0},\widehat{z}_{0}\right)$ is called the directional pointwise condition number of the problem (2): it is called “directional” because it depends on $\widehat{z}_{0}$, and “pointwise” because it depends on $y_{0}$.
+Along with the condition number (5), we also introduce two other condition numbers:
+  • the pointwise condition number of the problem (2) given by
     $K\left(t,y_{0}\right):=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}K\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|\mathrm{e}^{tA}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$ (6)
-[…]
+    where $\left\|\mathrm{e}^{tA}\right\|$ is the matrix norm of $\mathrm{e}^{tA}$ induced by the vector norm $\left\|\ \cdot\ \right\|$ (see [4] for the definition of condition number of a general problem, which corresponds to the pointwise condition number);
+  • the global condition number of the problem (2) given by
     $K\left(t\right):=\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}K\left(t,y_{0}\right)=\left\|\mathrm{e}^{tA}\right\|\cdot\left\|\mathrm{e}^{-tA}\right\|=\kappa\left(\mathrm{e}^{tA}\right),$ (7)
     which equals the standard condition number $\kappa\left(\mathrm{e}^{tA}\right)$ of the matrix $\mathrm{e}^{tA}$.
-Observe that $K\left(t,y_{0}\right)$ is the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$ and $K(t)$ is the worst condition number $K(t,y_{0})$ by varying $y_{0}$, i.e.
-the worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying both $y_{0}$ and $\widehat{z}_{0}$.
-The paper [12] studied the condition numbers (3), (4) and (5) in case of $A$ normal and the Euclidean norm as vector norm on $\mathbb{R}^{n}$.
-The present paper study the general case where $A$ (complex) and the vector norm are arbitrary.
+Observe that $K\left(t,y_{0}\right)$ is the worst $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$, and $K(t)$ is the worst $K(t,y_{0})$ by varying $y_{0}$, i.e., the worst $K(t,y_{0},\widehat{z}_{0})$ by varying both $y_{0}$ and $\widehat{z}_{0}$.
+The paper [11] studied the condition numbers (5), (6) and (7) in the particular case of $A$ normal. The present paper studies the general case.
+The aim of the present paper is to analyze the asymptotic (long-time) behavior of the three condition numbers $K(t,y_{0},\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$, i.e., their behavior as $t$ approaches infinity (becomes large).
+1.2. Plan of the paper
+Besides this introduction, the paper contains five sections and three appendices.
+Section 2 develops notions and notations for understanding the asymptotic forms of the three condition numbers.
+Section 3 analyzes the asymptotic behaviors of the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$, by introducing the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$.
+Section 4 analyzes the asymptotic behavior of the condition number $K(t)$, by introducing the asymptotic condition numbers $K_{\infty}^{+}(t)$ and $K_{\infty}(t)$.
+Section 5 introduces the Rightmost Last Generalized Eigenvector (RLGE) condition and summarizes the most important results.
+Conclusions are in Section 6.
+The three appendices contain the more technical material.
+They should be consulted as needed while reading the paper, and read in full only by readers interested in the mathematical details.
+Appendix A develops a suitable formula for the matrix exponential $\mathrm{e}^{tA}$ in the general non-diagonalizable case along with other fundamental material related to this formula.
+Appendix B investigates the properties of the key matrices $Q_{jl}(t)$ that determine the asymptotic behavior of the condition numbers.
+Appendix C analyzes the matrices $Q_{j}^{e}(t)$, which are important for defining the asymptotic condition number $K_{\infty}(t)$ related to $K(t)$.
+The present paper has two sequels.
+The first is [12], which develops the results of this paper in depth, for a real ODE (1) in a generic case.
+The second is [13], which presents extensive experimental tests, applications to real-world systems, and other issues, such as the non-asymptotic behavior of the condition numbers and how rapidly the asymptotic behavior sets in, also in relation to the non-normality of the matrix $A$.
+As a consequence, we will not discuss such practical questions in the present paper nor in [12].
+In any case, the present paper reaches a conclusion by fully characterizing the asymptotic behavior of the condition numbers (5), (6) and (7).
+Moreover, it contains in Sections 4 and 5 some numerical examples illustrating the results obtained in this paper.
+Papers [12] and [13] are further developments that cannot be included here for obvious space constraints.
+In what follows, we often refer to a generic case for an element $v$ of a finite-dimensional space $V$.
+By this, we mean that $v$ satisfies a property which is not satisfied only on a manifold $M$ of $V$ with $\mathrm{dim}M<\mathrm{dim}V$.
+Equivalently, if $v$ is drawn at random from $V$ (with respect to any distribution absolutely continuous with Lebesgue measure), the generic case holds with probability $1$.
+2. Asymptotic forms
+In the next Sections 3 and 4, we analyze the asymptotic behavior of the three condition numbers $K(t,y_{0},\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$.
+This analysis determines the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (5), of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (6), and of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ in (7).
+Then, the asymptotic behavior of the condition numbers is described by inserting in (5), (6) and (7) these asymptotic forms and by defining as asymptotic condition numbers the new expressions obtained by these substitutions.
+All of this might seem straightforward, but from a rigorous mathematical perspective it is not a simple task. Specifically, the following points should be remarked.
+  • It is necessary to precisely define what we mean by asymptotic form and how to determine it.
+The determination of the asymptotic forms is complicated in the non-diagonalizable case, where the Jordan Canonical Form of $A$ and generalized eigenvectors are involved.
+Moreover, proving that a given candidate for an asymptotic form is indeed an asymptotic form in our definition requires a certain mathematical effort: see Remark 6 at the end of this section.
+  • We also want to quantify how dominant the asymptotic forms are at finite times (these quantifications are used in the sequel papers [12] and [13]).
+  • The core of our analysis is the qualitative study of the asymptotic condition numbers.
+Understanding them requires significant mathematical effort, in particular in the case of rightmost complex eigenvalues (see the sequel paper [12]).
+  • If we are interested in defining asymptotic condition numbers for the problem (2), the asymptotic condition numbers derived from (6) and (7) as $t\rightarrow+\infty$ may not be appropriate, since they represent the asymptotic worst cases of (5), by varying $\widehat{z}_{0}$ only and both $\widehat{z}_{0}$ and $y_{0}$, respectively.
+Instead, we may be concerned with the worst asymptotic cases of (5).
+In other words: do “asymptotic” and “worst” commute?
+In this section, we determine the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, where $u\in\mathbb{C}^{n}$.
+In order to define them, some preliminary notions and results need to be introduced.
+2.1. Notations $\sim$ and $\approx$
+In this subsection, we make precise what we mean by asymptotic form.
+Let $f(t)$ and $g(t)$ be scalar, vector or matrix functions of $t\in\mathbb{R}$. For $g$ such that $g(t)\neq 0$ for $t$ in a neighborhood of $\infty$, we write
+$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$ (8)
+when
+$\lim\limits_{t\rightarrow+\infty}\frac{\|f(t)-g(t)\|}{\|g(t)\|}=0.$
+In case of scalars, vectors and matrices, $\|\;\cdot\;\|$ denotes, respectively, the modulus, a vector norm and a matrix norm.
+We interpret (8) as indicating that $g$ is an asymptotic form of $f$.
+Observe that (8) means that the relative error of $f$ with respect to its asymptotic form $g$ asymptotically vanishes.
+At a finite time, one may ask how dominant the asymptotic form is, i.e., how large the relative error of $f$ with respect to $g$ is.
+Therefore, we introduce the following notation.
+For $t\in\mathbb{R}$ such that $g(t)\neq 0$ and $\epsilon\geq 0$, we write
+$f\left(t\right)\approx g\left(t\right)\textit{\ with precision\ }\epsilon$ (9)
+when
+$\frac{\|f\left(t\right)-g\left(t\right)\|}{\|g\left(t\right)\|}\leq\epsilon.$
+Observe that (9) means that the relative error of $f$ with respect to $g$ at the time $t$ is not larger than $\epsilon$.
+The notation (9) serves to quantify how dominant the asymptotic form $g$ of $f$ is at the finite time $t$.
 Remark 1.
-We also consider the case of a real ODE (1).
-In this case, we replace $\mathbb{C}$ by $\mathbb{R}$ and we have $A\in\mathbb{R}^{n\times n}$ and $y_{0},y(t),\widetilde{y}_{0},\widetilde{y}(t)\in\mathbb{R}^{n}$.
-Moreover, the condition numbers (4) and (5) are defined by maxima over $\widehat{z}_{0}\in\mathbb{R}^{n}$ with $\|\widehat{z}_{0}\|=1$ and $y_{0}\in\mathbb{R}^{n}$ with $y_{0}\neq 0$, respectively.
-1.2. Plan of the paper
-The present paper is quite long and technical.
-For this reason, it should be read after or alongside the previously cited expository work [13], which presents the main results given here without proofs and extensive technical details.
-Here is the plan of the paper.
-Section 2 introduces some necessary technical linear algebra results concerning the Jordan Canonical Form (JCF) of $A$.
-Section 3 developes notions and notations for understanding the asymptotic (long-time) behavior considered in the next sections.
-Section 4 analyzes the asymptotic behaviors of the three conditions numbers $K(t,y_{0},$ $\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$, by introducing the three asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,y_{0})$ and $K_{\infty}(t)$.
-Section 5 summarizes the most important result of Section 4.
-From Section 6 to the last Section 10, the results of the asymptotic behavior analysis are applied in a generic situation for a real matrix $A$.
-2. Linear Algebra Results
+Note that, for vector or matrix functions $f(t)$ and $g(t)$,
+$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$
+implies
+$\|f\left(t\right)\|\sim\|g\left(t\right)\|,\ t\rightarrow+\infty,$
+and, for $\epsilon\geq 0$,
+$f\left(t\right)\approx g\left(t\right)\text{\ with precision\ }\epsilon$
+implies
+$\|f\left(t\right)\|\approx\|g\left(t\right)\|\text{\ with precision\ }\epsilon.$
+This follows by
+$\left|\left\|f(t)\right\|-\left\|g(t)\right\|\right|\leq\left\|f(t)-g(t)\right\|.$
+2.2. Partition of the spectrum and formula for the matrix exponential
+In this subsection, we introduce the tool for determining the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, $u\in\mathbb{C}^{n}$.
+The spectrum $\Lambda=\left\{\lambda_{1},\ldots,\lambda_{p}\right\}$ of $A$, where $\lambda_{1},\ldots,\lambda_{p}$ are the distinct eigenvalues of $A$, is partitioned by decreasing real parts (see Figure 1) in the subsets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, given by
+$\Lambda_{j}:=\{\lambda_{i_{j-1}+1},\lambda_{i_{j-1}+2},\ldots,\lambda_{i_{j}}\}$
+$\mathrm{Re}\left(\lambda_{i_{j-1}+1}\right)=\mathrm{Re}\left(\lambda_{i_{j-1}+2}\right)=\cdots=\mathrm{Re}\left(\lambda_{i_{j}}\right)=r_{j},$
+where the $q$ distinct real parts $r_{j}$, $j\in\{1,\ldots,q\}$, of the eigenvalues of $A$ satisfy $r_{1}>r_{2}>\cdots>r_{q}$.
+Observe that $\Lambda_{1}$ and $\Lambda_{q}$ are the sets of the rightmost and leftmost, respectively, eigenvalues of $A$.
+Figure 1. Partition of the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of AA by decreasing real part. The eigenvalues are marked by “x”.
+Now, we recall the formula (49) in Appendix A for the matrix exponential $\mathrm{e}^{tA}$:
+$\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$
+where $m_{i}$ is the ascent of the eigenvalue $\lambda_{i}$ (maximum dimension of the mini-blocks corresponding to $\lambda_{i}$ in the Jordan Canonical Form of $A$) and the matrices $P_{il}$ are defined and studied in Appendix A.
+By using the partition of the spectrum of $A$ given above, we have
+$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$ (10)
+where $\sqrt{-1}$ denotes the imaginary unit and $\omega_{i}$ denotes the imaginary part of the eigenvalue $\lambda_{i}$.
+By exchanging the two inner sums, we can rewrite (10) as
+$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t),$ (11)
+where
+$L_{j}:=\max_{\lambda_{i}\in\Lambda_{j}}m_{i}-1,\ j\in\{1,\ldots,q\},$
+and
+$Q_{jl}(t):=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il},\ j\in\{1,\ldots,q\}\text{\ and\ }l\in\{0,\ldots,L_{j}\}.$ (12)
+The formula (11) is used to determine the asymptotic forms in (5), (6) and (7).
+2.3. The asymptotic form of $\mathrm{e}^{tA}$
+By looking at the formula (11), we immediately identify
+$\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$ (13)
+as the asymptotic form of $\mathrm{e}^{tA}$, since it is the dominant term as $t\rightarrow+\infty$.
+Next proposition makes precise that (13) is indeed an asymptotic form of $\mathrm{e}^{tA}$ in our definition (8) of asymptotic form and quantifies its dominance at a finite time.
+Proposition 2.
+We have
+$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
+with precision
+$\epsilon(t):=\sum\limits_{l=0}^{L_{1}-1}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{1l}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}+\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\sum\limits_{l=0}^{L_{j}}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{jl}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}.$
+Moreover, we have
+$\epsilon(t)\rightarrow 0,\ t\rightarrow+\infty,$ (14)
+and then
+$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t),\ t\rightarrow+\infty.$
+Proof.
+By (11), we can write
+$\mathrm{e}^{tA}=\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)+\mathrm{e}^{r_{1}t}\sum\limits_{l=0}^{L_{1}-1}\frac{t^{l}}{l!}Q_{1l}(t)+\sum\limits_{j=2}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t).$
+The first part of the proposition regarding $\approx$ follows. By 1) and 3) in Proposition 41 of Appendix B, we obtain (14) and then the second part regarding $\sim$ follows. ∎
+Remark 3.
+If $A$ is diagonalizable, then we have
+$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}Q_{10}(t)$
+with precision
+$\epsilon(t)=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{\left\|Q_{j0}(t)\right\|}{\left\|Q_{10}(t)\right\|}$
+and
+$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}Q_{10}(t),\ t\rightarrow+\infty,$
+where
+$Q_{j0}(t)=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{i0},\ j\in\{1,\ldots,q\},$
+with $P_{i0}$ the projection onto the eigenspace corresponding to $\lambda_{i}$.
+2.4. Notations for the asymptotic form of $\mathrm{e}^{tA}u$
+The following notations are crucial for determining the asymptotic form of $\mathrm{e}^{tA}u$.
+For $j\in\{1,\ldots,q\}$ and $u\in\mathbb{C}^{n}$, we define $\Lambda_{j}(u)$ and $L_{j}(u)$.
+  • Let
+    $\Lambda_{j}(u):=\{\lambda_{i}\in\Lambda_{j}:\alpha^{(i)}(u)\neq 0\},$
+    where $\alpha^{(i)}(u)$, defined in Appendix A, is the vector of the Jordan basis components of $u$ along the generalized eigenvectors corresponding to $\lambda_{i}$.
+In other words, $\Lambda_{j}(u)$ is obtained from $\Lambda_{j}$ by including only the eigenvalues $\lambda_{i}$ for which $u$ has a non-zero projection onto the generalized eigenspace corresponding to $\lambda_{i}$.
+  • When $\Lambda_{j}(u)\neq\emptyset$, let
+    $L_{j}(u):=\max\{l_{i}(u):\lambda_{i}\in\Lambda_{j}(u)\},$
+    where
+    $l_{i}(u):=\max\{l\in\{0,\ldots,m_{i}-1\}:P_{il}u\neq 0\}$
+    (the matrices $P_{il}$ appear in the previous Subsection 2.2) is defined and studied in Appendix A: it is such that $l_{i}(u)+1$ is the maximum index $k$ such that $u$ has non-zero component along the $k$-th generalized eigenvector $v^{(i,j^{\prime},k)}$ of some Jordan chain
+    $(v^{(i,j^{\prime},k)})_{k=1,\ldots,m_{ij^{\prime}}},\ j^{\prime}\in\{1,\ldots,d_{i}\},$
+    corresponding to the eigenvalue $\lambda_{i}$.
+In other words, $L_{j}(u)+1$ is the maximum index $k$ such that $u$ has non-zero component along the $k$-th generalized eigenvector of some Jordan chain corresponding to an eigenvalue in $\Lambda_{j}(u)$.
+2.4.1. Indices of dominance
+For $u\in\mathbb{C}^{n}\setminus\{0\}$, we define the indices of dominance $j(u)$ and $L(u)$.
+  • Let
+    $j(u):=\min\{j\in\left\{1,\ldots,q\right\}:\Lambda_{j}(u)\neq\emptyset\}.$
+    In other words, $\Lambda_{j(u)}$ is the rightmost set $\Lambda_{j}$ (the sets $\Lambda_{j}$ are defined in the previous Subsection 2.2) such that $u$ has, for some $\lambda_{i}\in\Lambda_{j}$, non-zero projection onto the generalized eigenspace corresponding to $\lambda_{i}$.
+We call $j(u)$ the primary index of dominance of $u$.
+  • Let
+    $L(u):=L_{j(u)}(u).$
+    In other words, $L(u)+1$ is the maximum index $k$ such that $u$ has non-zero component along the $k$-th generalized eigenvector of some Jordan chain corresponding to an eigenvalue in $\Lambda_{j(u)}(u)$.
+We call $L(u)$ the secondary index of dominance of $u$.
+2.5. The asymptotic form of $\mathrm{e}^{tA}u$
+Let $u\in\mathbb{C}^{n}\setminus\{0\}$. When we use the formula (11) for $\mathrm{e}^{tA}$ in $\mathrm{e}^{tA}u$, we obtain
+$\mathrm{e}^{tA}u=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u.$ (15)
+By 1) and 2) in Proposition 40 of Appendix B, we identify
+$\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$ (16)
+as the asymptotic form of $\mathrm{e}^{tA}u$, since it is the dominant term as $t\rightarrow+\infty$.
+Next proposition states precisely that (16) is indeed an asymptotic form in our definition and quantifies its dominance.
+Proposition 4.
+For $u\in\mathbb{C}^{n}\setminus\{0\}$, we have
+$\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$
+with precision
+$\epsilon(t,u)$ $:=$ $\sum\limits_{l=0}^{L\left(u\right)-1}\frac{L\left(u\right)!}{l!}t^{l-L\left(u\right)}\frac{\left\|Q_{j(u)l}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}$
+$+\sum\limits_{j=j\left(u\right)+1}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\sum\limits_{l=0}^{L_{j}}\frac{L(u)!}{l!}t^{l-L(u)}\frac{\left\|Q_{jl}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}.$
+Moreover, we have
+$\epsilon(t,u)\rightarrow 0,\ t\rightarrow+\infty,$ (17)
+and then
+$\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u,\ t\rightarrow+\infty.$
+Proof.
+By (15) and Proposition 40 in Appendix B, we can write
+$\mathrm{e}^{tA}u$ $=$ $\sum\limits_{j=j(u)}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u$
+$=$ $\mathrm{e}^{r_{j(u)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u+\mathrm{e}^{r_{j(u)}t}\sum\limits_{l=0}^{L(u)-1}\frac{t^{l}}{l!}Q_{j(u)l}(t)u$
+$+\sum\limits_{j=j(u)+1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u.$
+The first part of the proposition regarding $\approx$ follows. By 2) and 4) in Proposition 41 of Appendix B, we obtain (17) and then the second part regarding $\sim$ follows. ∎
+Remark 5.
+If $A$ is diagonalizable, then we have
+$\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u$
+with precision
+$\epsilon(t,u)=\sum\limits_{j=j\left(u\right)+1}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\frac{\left\|Q_{j0}(t)u\right\|}{\left\|Q_{j(u)0}(t)u\right\|}$
+and
+$\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u,\ t\rightarrow+\infty.$
+The asymptotic form (16) of $\mathrm{e}^{tA}u$ is determined by the primary and secondary indices of dominance of $u$.
+In particular, the smaller the primary index of dominance $j(u)$, the higher the order of the dominant term (16), and for a fixed primary index, the larger the secondary index of dominance $L(u)$, the higher the order of (16).
+Remark 6.
+Recall the first point in Section 2.
+The contents of Propositions 2 and 4 are deeper than what might be apparent at first glance.
+We are not simply stating the obvious facts that (13) and (16) are the dominant terms in $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, respectively.
+Rather, we are asserting that they represent the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ according to our definition (8).
+In other words, the relative errors of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ with respect to asymptotic form (13) and (16), respectively, approach zero asymptotically.
+Proving these deeper conclusions is complicated: to obtain (14) and (17), one needs the key facts $\inf\limits_{t\in\mathbb{R}}\|Q_{1L_{1}}(t)\|>0$ and $\inf\limits_{t\in\mathbb{R}}\|Q_{j(u)L(u)}(t)u\|>0$, which follow from Propositions 37 and 38 and Lemma 39 in Appendix A, developed through the work in that appendix.
+3. The asymptotic behaviors of $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$
+This section and the next one form the core of our analysis. Building on the preparatory work of the previous section, we can now easily describe the asymptotic behavior of the condition numbers.
+In particular, in this section we study the asymptotic behavior of the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$.
+Their asymptotic forms are the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$.
+We also show that $K_{\infty}(t,y_{0})$ coincides with the worst $K_{\infty}(t,y_{0},\widehat{z}_{0})$, by varying $\widehat{z}_{0}$.
+3.1. The asymptotic condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$
+We set
+$j^{\ast}:=j\left(\widehat{y}_{0}\right)=j\left(y_{0}\right)\text{\ \ and\ \ }L^{\ast}:=L\left(\widehat{y}_{0}\right)=L\left(y_{0}\right)$
+as well as
+[…]
+i.e. $j^{\ast}$ and $j^{\ast\ast}$ are the primary indices of dominance, and $L^{\ast}$ and $L^{\ast\ast}$ are the secondary indices of dominance, of $y_{0}$ and $\widehat{z}_{0}$, respectively.
+The next theorem describes the asymptotic form of $K\left(t,y_{0},\widehat{z}_{0}\right)$.
+Theorem 7.
+We have
+$K\left(t,y_{0},\widehat{z}_{0}\right)\approx K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),$
+where
+$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right):=\frac{L^{\ast}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{j^{\ast}}\right)t}t^{L^{\ast\ast}-L^{\ast}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
+with precision
+$\frac{\epsilon(t,\widehat{z}_{0})+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})},$
+whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t,\widehat{z}_{0})$ and $\epsilon(t,\widehat{y}_{0})$ are defined in Proposition 4). Moreover, we have
+$K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\rightarrow+\infty.$
+Proof.
+Proposition 4 states the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (5), and quantifies how dominant they are at a finite time $t$: we have
+$\mathrm{e}^{tA}\widehat{z}_{0}\approx\mathrm{e}^{r_{j^{\ast\ast}}t}\frac{t^{L^{\ast\ast}}}{L^{\ast\ast}!}Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}$
+with precision $\epsilon(t,\widehat{z}_{0})$ and
+$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
+with precision $\epsilon(t,\widehat{y}_{0})$.
+The first part of the theorem regarding $\approx$ follows by using Remark 1 and by bounding the relative error of the ratio $\frac{\|\mathrm{e}^{tA}\widehat{z}_{0}\|}{\|\mathrm{e}^{tA}\widehat{y}_{0}\|}$in terms of the bounds $\epsilon(t,\widehat{z}_{0})$ and $\epsilon(t,\widehat{y}_{0})$ of the relative errors of $\|\mathrm{e}^{tA}\widehat{z}_{0}\|$ and $\|\mathrm{e}^{tA}\widehat{y}_{0}\|$ , respectively.
+These are relative errors with respect to the norms of the asymptotic forms.
+The second part regarding $\sim$ follows since $\epsilon(t,\widehat{y}_{0})\rightarrow 0$ and $\epsilon(t,\widehat{z}_{0})\rightarrow 0$, $t\rightarrow+\infty$.
+∎
+We define the function
+$t\rightarrow K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\in\mathbb{R},$
+as the asymptotic directional pointwise condition number of the problem (2).
+Remark 8.
+  • 1.
+    The asymptotic directional pointwise condition number of the problem (2):
+    • is bounded and away from zero, as $t$ varies, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}=L^{\ast\ast}$ (recall 2) and 4) in Proposition 41 of Appendix B) ;
+    • decays polynomially to zero, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}>L^{\ast\ast}$;
+    • diverges polynomially to infinity, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}<L^{\ast\ast}$;
+    • decays exponentially to zero, as $t\rightarrow+\infty$, if $j^{\ast}<j^{\ast\ast}$;
+    • diverges exponentially to infinity, as $t\rightarrow+\infty$, if $j^{\ast}>j^{\ast\ast}$.
+    Hence, whether the asymptotic condition number decreases to zero, diverges to infinity, or exhibits different behavior depends on which between $y_{0}$ and $\widehat{z}_{0}$ is more dominant.
+This is determined by which of $y_{0}$ and $\widehat{z}_{0}$ possesses the smaller primary index of dominance, or, in the case of equal primary indices, the larger secondary index of dominance.
+  • 2.
+    The case $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$ is generic for $y_{0}$ and $\widehat{z}_{0}$. In this generic case, we have
+    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
+3.2. The asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$
+The next theorem describes the asymptotic form of $K\left(t,y_{0}\right)$, worst $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$.
+Theorem 9.
+We have
+$K\left(t,y_{0}\right)\approx K_{\infty}\left(t,y_{0}\right),$
+where
+$K_{\infty}\left(t,y_{0}\right):=\frac{L^{\ast}!}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{j^{\ast}}\right)t}t^{L_{1}-L^{\ast}}\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
+with precision
+$\frac{\epsilon(t)+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
+whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t)$ and $\epsilon(t,\widehat{y}_{0})$ are defined in Propositions 2 and 4, respectively). Moreover, we have
+$K\left(t,y_{0}\right)\sim K_{\infty}\left(t,y_{0}\right),\ t\rightarrow+\infty.$
+Proof.
+Proposition 2 states the asymptotic form of $\mathrm{e}^{tA}$, and quantifies how dominant it is at a finite time $t$: we have
+$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
+with precision $\epsilon(t)$. Proposition 4 states the asymptotic form of $\mathrm{e}^{tA}\widehat{y}_{0}$, and quantifies how it is dominant at a finite time $t$: we have
+$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
+with precision $\epsilon(t,\widehat{y}_{0})$.
+The first part of the theorem regarding $\approx$ follows by using Remark 1 and by bounding the relative error of the ratio $\frac{\|\mathrm{e}^{tA}\|}{\|\mathrm{e}^{tA}\widehat{y}_{0}\|}$.
+The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,\widehat{y}_{0})\rightarrow 0$, $t\rightarrow+\infty$.
+∎
+We define the function
+$t\rightarrow K_{\infty}(t,y_{0}),\ t\in\mathbb{R},$
+as the asymptotic pointwise condition number of the problem (2).
+Remark 10.
+  • 1.
+    The asymptotic pointwise condition number of the problem (2):
+    • is bounded and away from zero as $t$ varies if $j^{\ast}=1$ and $L^{\ast}=L_{1}$;
+    • diverges polynomially to infinity, as $t\rightarrow+\infty$, if $j^{\ast}=1$ and $L^{\ast}<L_{1}$;
+    • diverges exponentially to infinity, as $t\rightarrow+\infty$, if $j^{\ast}>1$.
+    Therefore, the asymptotic condition number does not diverge to infinity if and only if $y_{0}$ is as dominant as possible, meaning that $y_{0}$ has the smallest possible primary index of dominance (i.e., $j^{\ast}=1$) and simultaneously the largest possible secondary index of dominance (i.e., $L^{\ast}=L_{1}$).
+  • 2.
+    The case $j^{\ast}=1$ and $L^{\ast}=L_{1}$ is generic for $y_{0}$. In this generic case, we have
+    $K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
+3.3. Is $K_{\infty}\left(t,y_{0}\right)$ the worst $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$?
+By definition, $K\left(t,y_{0}\right)$ is the worst $K\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$.
+Hence, an interesting question is the following.
+Does this fact hold asymptotically as $t\rightarrow+\infty$?
+Specifically, is $K_{\infty}\left(t,y_{0}\right)$ the worst $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$?
+In other words, do
+$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\ \text{(worst case) and }\ t\rightarrow+\infty\ \text{(asymptotic behavior)}$
+commute? The answer is YES and it is given by the next theorem, which considers the ratio
+$\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=\frac{L_{1}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{1}\right)t}t^{L^{\ast\ast}-L_{1}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\right\|},$ (18)
+which is independent of $y_{0}$.
+Theorem 11.
+We have
+$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=1.$ (19)
+In particular:
+  • a)
+    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}>1$ or $j^{\ast\ast}=1$ and $L^{\ast\ast}<L_{1}$, we have
+    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=0.$
+  • b)
+    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, we have
+    $\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}\leq 1.$
+  • c)
+    There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a sequence $\{t_{m}\}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
+    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m},y_{0}\right)}=1.$
+Proof.
+Points a), b) and c) imply (19).
+Points a) and b) immediately follow by (18).
+Now, we prove c).
+Consider a sequence $\{t_{k}\}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{z}_{0k}\}$ such that $\widehat{z}_{0k}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0k}\right\|=1$ and
+$\left\|Q_{1L_{1}}(t_{k})\right\|=\left\|Q_{1L_{1}}(t_{k})\widehat{z}_{0k}\right\|.$
+By the compactness of the unit sphere in $\mathbb{C}^{n}$, there exists a subsequence $\{\widehat{z}_{0k_{m}}\}$ of $\{\widehat{z}_{0k}\}$ converging to some $\widehat{z}_{0\infty}\in\mathbb{C}^{n}$ with $\left\|\widehat{z}_{0\infty}\right\|=1$.
+We have $j\left(\widehat{z}_{0\infty}\right)=1$ and $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$.
+In fact, for any index $m$, we have
+$\left|\ \left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|-\left\|Q_{1L_{1}}(t_{k_{m}})\right\|\ \right|\leq\sup\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|,$
+where the right-hand side goes to zero as $m\rightarrow\infty$ (remind point 1) in Proposition 41 of Appendix B). Therefore, there exists an index $m$ such that
+$\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|\geq\frac{1}{2}\inf\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|,$
+where the right-hand side is positive (remind point 3) in Proposition 41 of Appendix B).
+Since $Q_{1L_{1}}(t_{k_{m}})$ $\widehat{z}_{0\infty}\neq 0,$ we cannot have $j\left(\widehat{z}_{0\infty}\right)>1$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$ (remind point 1) in Proposition 40 of Appendix B).
+Hence, $j\left(\widehat{z}_{0\infty}\right)=1$.
+Moreover, since $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\neq 0,$ we cannot have $L_{1}\left(\widehat{z}_{0\infty}\right)<L_{1}$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$ (remind point 2) in Proposition 40 of Appendix B).
+Hence, $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$.
+By using as a direction of perturbation $\widehat{z}_{0\infty}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$. Thus, for any index $m$,
+$\left|\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}-1\right|$ $=$ $\left|\frac{\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|}{\left\|Q_{1L_{1}}(t_{k_{m}})\right\|}-1\right|$
+$\leq$ $\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|.$
+We conclude that
+$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}=1.$
+Remark 12.
+  • 1.
+    When $\Lambda_{1}$ consists of a real eigenvalue, the point c) is modified to:
+    • $c^{\prime}$)
+      There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, such that
+      $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right).$ (20)
+    In fact, in this case $Q_{1L_{1}}(t)=Q_{1L_{1}}$ is independent of $t$ (see (12)); hence, there exists $\widehat{z}_{0}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0}\right\|=1$, independent of $t$ such that
+    $\left\|Q_{1L_{1}}\right\|=\left\|Q_{1L_{1}}\widehat{z}_{0}\right\|.$
+    By using as a direction of perturbation $\widehat{z}_{0}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ and then we obtain (20) by (18).
+  • 2.
+    Observe that
+    $\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K\left(t,y_{0},\widehat{z}_{0}\right)}{K\left(t,y_{0}\right)}=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}$
+    and then
+    $\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K\left(t,y_{0},\widehat{z}_{0}\right)}{K\left(t,y_{0}\right)}=1,$
+    which, more clearly than (19), shows that
+    $\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\ \text{(worst case) and }\limsup\limits_{t\rightarrow+\infty}\text{(asymptotic behavior)}$
+    commute.
+4. The asymptotic behavior of $K(t)$
+In this section, we study the asymptotic behavior of the global condition number $K(t)$.
+Its asymptotic form is the asymptotic condition number $K^{+}_{\infty}(t)$.
+We also show that $K^{+}_{\infty}(t)$ does not coincide with the worst $K_{\infty}(t,y_{0})$, by varying $y_{0}$, i.e., in light of Theorem 11, it does not coincide with the worst $K_{\infty}(t,y_{0},\widehat{z}_{0})$, by varying $y_{0}$ and $\widehat{z}_{0}$.
+The worst $K_{\infty}(t,y_{0})$ is the asymptotic condition number $K_{\infty}(t)$.
+4.1. The asymptotic global condition number $K_{\infty}^{+}\left(t\right)$
+Next theorem describes the asymptotic form of $K\left(t\right)$, the worst $K\left(t,y_{0}\right)$ by varying $y_{0}$.
+Theorem 13.
+We have
+$K\left(t\right)\approx K_{\infty}^{+}(t),$
+where
+$K_{\infty}^{+}\left(t\right):=\frac{1}{L_{1}!L_{q}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}+L_{q}}\left\|Q_{1L_{1}}(t)\right\|\ \cdot\left\|Q_{qL_{q}}(-t)\right\|$ (21)
+with precision
+$\epsilon(t)+\epsilon(t,-A)+\epsilon(t)\epsilon(t,-A),$
+where $\epsilon(t,-A)$ is $\epsilon(t)$ for the matrix A-A. Moreover, we have
+$K\left(t\right)\sim K^{+}_{\infty}(t),\ t\rightarrow+\infty.$
+Proof.
+Proposition 2 states the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ and how they are dominant at a finite time $t$: we have
+$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
+with precision $\epsilon(t)$ and
+$\mathrm{e}^{-tA}=\mathrm{e}^{t(-A)}\approx\mathrm{e}^{r_{1}(-A)t}\frac{t^{L_{1}(-A)}}{L_{1}(-A)!}Q_{1L_{1}(-A)}(t,-A)=\mathrm{e}^{-r_{q}t}\frac{t^{L_{q}}}{L_{q}!}(-1)^{L_{q}}Q_{qL_{q}}(-t)$
+with precision $\epsilon(t,-A)$.
+For the latter, recall (62) and Proposition 42 in Appendix B.
+The first part of the theorem regarding $\approx$ follows by using Remark 1 and by bounding the relative error of the product $\|\mathrm{e}^{tA}\|\|\mathrm{e}^{t(-A)}\|$ in terms of the bounds $\epsilon(t)$ and $\epsilon(t,-A)$ of the relative errors of $\|\mathrm{e}^{tA}\|$ and $\|\mathrm{e}^{t(-A)}\|$, respectively.
+The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,-A)\rightarrow 0$, $t\rightarrow+\infty$.
+∎
+We define the function
+$t\rightarrow K^{+}_{\infty}(t),\ t\in\mathbb{R},$
+as the asymptotic global condition number of the problem (2).
+4.2. Is $K^{+}_{\infty}\left(t\right)$ the worst $K_{\infty}\left(t,y_{0}\right)$?
+By definition, $K\left(t\right)$ is the worst $K\left(t,y_{0}\right)$, by varying $y_{0}$.
+Hence, as in Subsection 3.3, an interesting question is the following.
+Does this fact hold asymptotically as $t\rightarrow+\infty$?
+Specifically, is $K^{+}_{\infty}\left(t\right)$ the worst $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$?
+In other words, do
+$\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}\ \text{(worst case) and }\ t\rightarrow+\infty\ \text{(asymptotic behavior)}$
+commute? Unlike the similar question in Subsection 3.3, here the answer is NO and it is given by the next theorem, which considers
+$K_{\infty}\left(t\right):=\frac{1}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}}\left\|Q_{1L_{1}}(t)\right\|\cdot\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|$ (22)
+and the ratio
+$\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=L^{\ast}!\mathrm{e}^{\left(r_{q}-r_{j^{\ast}}\right)t}t^{-L^{\ast}}\frac{1}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|},$ (23)
+where the subspace $U^{e}_{q}$ of $\mathbb{C}^{n}$ and the linear operator $Q^{e}_{q}(-t)|_{U^{e}_{q}}$ are defined in Appendix C.
+Theorem 14.
+We have
+$\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=1.$ (24)
+In particular:
+  • a)
+    For any initial value $y_{0}$ such that $j^{\ast}<q$ or $j^{\ast}=q$ and $L^{\ast}>0$, we have
+    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=0.$
+  • b)
+    For any initial value $y_{0}$ such that $j^{\ast}=q$ and $L^{\ast}=0$, we have
+    $\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}\leq 1.$
+  • c)
+    There exists an initial value $y_{0}\in U^{e}_{q}$, i.e., an initial value $y_{0}$ with $j^{\ast}=q$ and $L^{\ast}=0$, and a sequence $\{t_{m}\}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
+    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0}\right)}{K_{\infty}\left(t_{m}\right)}=1.$
+  • d)
+    For the initial value $y_{0}$ at point c), there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}$ of the sequence $\{t_{m}\}$ at point c), such that
+    $\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=1.$
+Proof.
+Points a), b) and c) imply (24).
+Point a) follows by (23): recall point 4) in Proposition 41 of Appendix B and Remark 50 of Appendix C.
+Point b) follows by Remark 48 of Appendix C and by observing that, for such $y_{0}$, we have $y_{0}\in U^{e}_{q}$ and then
+$Q_{q0}(t)\widehat{y}_{0}=Q^{e}_{q}(t)\widehat{y}_{0}$
+by Proposition 46 of Appendix C.
+For the point c), we proceed as in the proof of Theorem 11.
+Consider a sequence $\{t_{k}\}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{y}_{0k}\}$, where $\widehat{y}_{0k}\in U^{e}_{q}$ and $\left\|\widehat{y}_{0k}\right\|=1$, such that
+$\left\|Q_{q0}(t_{k})\widehat{y}_{0k}\right\|=\left\|Q^{e}_{q}(t_{k})\widehat{y}_{0k}\right\|=\frac{1}{\left\|Q^{e}_{q}(-t_{k})|_{U^{e}_{q}}\right\|}$ (25)
+(recall, in Appendix C, Proposition 46 and Remark 48).
+There exists a subsequence $\{\widehat{y}_{0k_{m}}\}$ of $\{\widehat{y}_{0k}\}$ converging to some $\widehat{y}_{0\infty}\in U^{e}_{q}$ with $\left\|\widehat{y}_{0\infty}\right\|=1$.
+By using $\widehat{y}_{0\infty}$ as initial value, we have $j^{\ast}=q$ and $L^{\ast}=0$: for the latter, observe that $l_{i}(\widehat{y}_{0\infty})=0$ for any $\lambda_{i}\in\Lambda_{q}(\widehat{y}_{0\infty})$ (see Proposition 36 in Appendix A).
+Therefore,
+$\left|\frac{K_{\infty}\left(t_{k_{m}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}}\right)}-1\right|$ $=$ $\left|\frac{1}{\left\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\right\|\left\|Q^{e}_{q}(-t_{k_{m}})|_{U^{e}_{q}}\right\|}-1\right|$
+[…]
+$\leq$ $\frac{\left\|Q_{q0}(t_{k_{m}})\right\|}{\left\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\right\|}\left\|\widehat{y}_{0k_{m}}-\widehat{y}_{0\infty}\right\|$
+$\rightarrow$ $0,\ m\rightarrow\infty$
+(use 2) and 4) in Proposition 41 of Appendix B).
+We conclude that
+$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{k_{m}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{t_{k_{m}}}\right)}=1.$
+For the point d), repeat the proof of point c) in Theorem 11 with the sequence $\{t_{k}\}$ replaced by the sequence $\{t_{m}\}$ at point c) of this theorem.
+In this way, we show that there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}$ of the sequence $\{t_{m}\}$, such that
+$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1$
+and then
+$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{m_{s}}\right)}\cdot\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1.$
+Remark 15.
+When $\Lambda_{q}$ consists of a real eigenvalue, the point c) is modified to:
+  • $c^{\prime}$)
+    For any initial value $y_{0}\in U^{e}_{q}$, we have
+    $K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
+See (23) and Remark 45 and Proposition 46 of Appendix C.
+In addition, if $\Lambda_{1}$ also consists of a real eigenvalue, the point d) is modified to:
+  • $d^{\prime}$)
+    There exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ such that, for any initial value $y_{0}\in U^{e}_{q}$, we have
+    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
+See point 1 in Remark 12.
+4.3. The global asymptotic condition number $K_{\infty}\left(t\right)$
+Theorem 14 says that $K_{\infty}\left(t\right)$ in (22), not $K^{+}_{\infty}\left(t\right)$ in (21), is the worst $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$, i.e., the worst $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying both $y_{0}$ and $\widehat{z}_{0}$.
+We define the function
+$t\rightarrow K_{\infty}(t),\ t\in\mathbb{R},$
+as the global asymptotic condition number of the problem (2).
+Observe that the global asymptotic condition number $K_{\infty}(t)$ is the worst asymptotic form of $K(t,y_{0},\widehat{z}_{0})$, as $y_{0}$ and $\widehat{z}_{0}$ vary.
+In contrast, the asymptotic global condition number $K_{\infty}^{+}(t)$ is the asymptotic form of the worst $K(t,y_{0},\widehat{z}_{0})$, as $y_{0}$ and $\widehat{z}_{0}$ vary.
+Moreover, observe that $K^{+}_{\infty}(t)$ can be significantly larger than $K_{\infty}(t)$: we have
+$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=L_{q}!t^{-L_{q}}\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{qL_{q}}(-t)\right\|}$
+and then
+$\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=0$
+for $L_{q}>0$ and
+$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}=\frac{\left\|Q_{q0}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}\leq 1$
+for $L_{q}=0$ (recall Proposition 46 in Appendix C).
+Remark 16.
+Both the asymptotic global condition number and the global asymptotic condition number of the problem (2):
+  • are bounded and away from zero as $t$ varies if $q=1$ and $L_{1}=0$;
+  • diverge polynomially to infinity, as $t\rightarrow+\infty$, if $q=1$ and $L_{1}>0$;
+  • diverge exponentially to infinity, as $t\rightarrow+\infty$, if $q>1$.
+Therefore, they do not diverge to infinity if and only if all the eigenvalues of $A$ have the same real part, i.e., they lie in a vertical line of the complex plane, and have ascent $1$, i.e., $A$ is diagonalizable.
+In the next example, we illustrate the difference between $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$.
+Example 17.
+Consider the non-diagonalizable matrix
+$A_{1}=V\left[\begin{array}[]{rr}1&1\\ 0&1\end{array}\right]W$
+and the diagonalizable matrix
+$A_{2}=V\left[\begin{array}[]{rr}1&0\\ 0&-1\end{array}\right]W,$
+where
+$V=\left[\begin{array}[]{rr}1&\frac{1}{2}\\ 1&1\end{array}\right]\text{\ \ and\ \ }W=V^{-1}=2\left[\begin{array}[]{rr}1&-\frac{1}{2}\\ -1&1\end{array}\right].$
+The matrix $A_{1}$
+For the matrix $A_{1}$, we have
+$\Lambda=\{1\},\ q=1{\ \ and\ \ }L_{1}=L_{q}=1.$
+By using the notation for columns of $V$ and rows of $W$ in Appendix A, we write $V$ and $W$ as
+$V=\left[v^{(1,1,1)}\ v^{(1,1,2)}\right]\text{\ \ and\ \ }W=\left[\begin{array}[]{c}w^{(1,1,1)}\\ w^{(1,1,2)}\end{array}\right].$
+The global asymptotic condition number and the asymptotic global condition number are
+$K_{\infty}(t)=t\left\|Q_{11}(t)\right\|\left\|Q^{e}_{1}(-t)|_{U^{e}_{1}}\right\|\text{\ and\ \ }K^{+}_{\infty}(t)=t^{2}\left\|Q_{11}(t)\right\|\left\|Q_{11}(-t)\right\|,$
+where
+$Q_{11}(t)=Q_{11}(-t)=P_{11}=v^{(1,1,1)}w^{(1,1,2)}=\left[\begin{array}[]{rr}-2&2\\ -2&2\end{array}\right]$
+(recall point 2 in Remark 28 of Appendix A) and
+$Q^{e}_{1}(-t)|_{U^{e}_{1}}=I|_{U^{e}_{1}}\\$
+(recall Remark 45 of Appendix C). Thus, for the Euclidean norm as vector norm, we have, since $\|P_{11}\|_{2}=4$,
+$K_{\infty}(t)=4t\text{\ and\ \ }K^{+}_{\infty}(t)=16t^{2}.$
+In Figure 2, we see $K(t)$ (blue dashed line), $K^{+}_{\infty}(t)$ (red dashed line) and $K_{\infty}(t)$ (red dashed line), $t\in[0,100]$, in logarithmic scale.
+From the beginning, $K(t)$ is not distinguishable from its asymptotic form $K^{+}_{\infty}(t)$.
+Figure 2. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{1}$. The vector norm is the Euclidean norm.
+Figure 2 confirms that the asymptotic global condition number $K^{+}_{\infty}(t)$ is not the worst asymptotic form of $K(t,y_{0},\widehat{z}_{0})$, by varying $y_{0}$ and $\widehat{z}_{0}$.
+In fact, in the figure we also see, for each $\tau\in\{10,20,30,40\}$, $K(t,y_{0\tau},\widehat{z}_{0\tau})$ (black solid lines) for an initial value $y_{0\tau}$ and a direction of perturbation $\widehat{z}_{0\tau}$ such that $K(\tau,y_{0\tau},\widehat{z}_{0\tau})=K(\tau)$.
+To obtain this, we take:
+  • the initial value $y_{0\tau}$ such that
+    $\|\mathrm{e}^{-\tau A}\|=\frac{1}{\|\mathrm{e}^{\tau A}\widehat{y}_{0\tau}\|};$
+    this is obtained with $y_{0\tau}=\mathrm{e}^{-\tau A}x_{0\tau}$, where $x_{0\tau}$ is such that
+    $\|\mathrm{e}^{-\tau A}\|=\frac{\|\mathrm{e}^{-\tau A}x_{0\tau}\|}{\|x_{0\tau}\|};$
+  • the direction of perturbation $\widehat{z}_{0\tau}$ such that
+    $\|\mathrm{e}^{\tau A}\|=\|\mathrm{e}^{\tau A}\widehat{z}_{0\tau}\|.$
+Observe that $K(t,y_{0\tau},\widehat{z}_{0\tau})$ is the worst $K(t,y_{0},\widehat{z}_{0})$, by varying $y_{0}$ and $\widehat{z}_{0}$, only at $t=\tau$, where we have $K(t,y_{0\tau},\widehat{z}_{0\tau})=K(t)$.
+But, $K(t,y_{0\tau},\widehat{z}_{0\tau})$ asymptotically falls below $K_{\infty}(t)$, which is much smaller than $K^{+}_{\infty}(t)$.
+This last fact confirms that $K_{\infty}(t)$ is the worst $K_{\infty}(t,y_{0},\widehat{z}_{0})$, by varying $y_{0}$ and $\widehat{z}_{0}$.
+Indeed, we have $K_{\infty}(t)=K_{\infty}(t,y_{0},\widehat{z}_{0})$ for $y_{0}\in U^{e}_{1}=\mathrm{span}(v^{(1,1,1)})$ and for $\widehat{z}_{0}$ such that
+$\left\|Q_{11}(t)\right\|=\left\|P_{11}\right\|=\left\|P_{11}\widehat{z}_{0}\right\|$
+(see d’) in Remark 15 and point 1 in Remark 12).
+The matrix $A_{2}$
+For the matrix $A_{2}$, we have
+$\Lambda=\{1,-1\},\ q=2\text{\ \ and\ \ }L_{1}=L_{q}=0.$
+By using the notation for columns and rows in Appendix A, we write
+$V=\left[v^{(1,1,1)}\ v^{(2,1,1)}\right]\text{\ \ and\ \ }W=\left[\begin{array}[]{c}w^{(1,1,1)}\\ w^{(2,1,1)}\end{array}\right].$
+The global asymptotic condition number and asymptotic global condition number are
+$K_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q^{e}_{2}(-t)|_{U^{e}_{2}}\right\|\text{\ and\ \ }K^{+}_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q_{20}(-t)\right\|,$
+where
+$Q_{10}(t)=P_{10}=v^{(1,1,1)}w^{(1,1,1)}=\left[\begin{array}[]{rr}2&-1\\ 2&-1\end{array}\right]$
+and
+$Q_{20}(-t)=P_{20}=v^{(2,1,1)}w^{(2,1,1)}=\left[\begin{array}[]{rr}-1&1\\ -2&2\end{array}\right]$
+(see Subsection A.3.2 of Appendix A) and
+$Q^{e}_{2}(-t)|_{U^{e}_{2}}=I|_{U^{e}_{2}}$
+(see Remark 45 of Appendix C). Thus, for the Euclidean norm as vector norm, we have, since $\|P_{10}\|_{2}=\|P_{20}\|_{2}=\sqrt{10}$,
+[…]
+In Figure 3, for the matrix $A_{2}$, we reproduce everything shown in Figure 2 for the matrix $A_{1}$.
+Now, $t\in[0,5]$ and $\tau\in\{1,2,3,4\}$.
+The same behaviour of $K(t,y_{0\tau},\widehat{z}_{0\tau})$ appears.
+Figure 3. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{2}$. The vector norm is the Euclidean norm.
+5. Final considerations about the asymptotic condition numbers
+Suppose that the matrix $A$ in the ODE (1) does not have all eigenvalues with same real part, meaning $q>1$ (recall that $q$ is the number of distinct real parts of the eigenvalues of $A$).
+In this case, recall Remark 16, the global asymptotic condition number $K_{\infty}\left(t\right)$ exponentially diverges.
+Hence, the relative error $\varepsilon$ of the initial value is exponentially magnified in the relative error $\delta(t)$ of the solution, in the worst case for the initial value and the perturbation of the initial value.
+However, this conclusion is too pessimistic, since this exponential magnification of $\varepsilon$ in $\delta(t)$ appears only in a non-generic case.
+In fact, as illustrated in point 2 of Remark 8, in the generic case $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$ for $y_{0}$ and $\widehat{z}_{0}$, we have
+$\frac{\delta(t)}{\varepsilon}=K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|},\ t\rightarrow+\infty.$ (26)
+In this generic case, $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$ remains bounded as well as away from zero by varying $t$ (see point 1 of Remark 8).
+We have introduced the three asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,$ $y_{0})$ and $K_{\infty}(t)$.
+The most important is $K_{\infty}(t,y_{0})$.
+In fact, $K_{\infty}(t,y_{0},\widehat{z}_{0})$ is given in terms of the (in general) unknown direction of perturbation $\widehat{z}_{0}$ and $K_{\infty}(t)$ is a worst asymptotic magnification factor $\frac{\delta(t)}{\varepsilon}$ which applies in a non-generic case.
+As illustrated in point 2 of Remark 10, in the generic case $j^{\ast}=1$ and $L^{\ast}=L_{1}$ for $y_{0}$, this most important asymptotic condition number $K_{\infty}(t,y_{0})$, representing the asymptotic magnification factor $\frac{\delta(t)}{\varepsilon}$ for the worst perturbation of $y_{0}$, is given by
+$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$ (27)
+In this generic case, $K_{\infty}\left(t,y_{0}\right)$ remains bounded as well as away from zero by varying $t$ (see point 1 of Remark 10).
+5.1. The RLGE condition
+In this final subsection, we consolidate the conclusions (26) and (27) in a theorem that provides further details.
+Let $u\in\mathbb{C}^{n}$.
+We say that $u$ satisfies the Rightmost Last Generalized Eigenvector (RLGE) condition if $j(u)=1$ and $L(u)=L_{1}$, i.e., there is a non-zero component of $u$ along the last generalized eigenvector in some of the longest Jordan chains of the rightmost eigenvalues of $A$.
+Observe that satisfying the RLGE condition is a generic case for $u$.
+Here is the theorem regarding the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$, where we set $Q_{1}(t):=Q_{1L_{1}}(t)$ to simplify the notation.
+Theorem 18.
+If $y_{0}$ and $\widehat{z}_{0}$ satisfy the RLGE condition, then
+$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|}$
+and
+$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1}(t)\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|}$
+with
+$Q_{1}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)}$
+and
+$Q_{1}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)},\ u\in\mathbb{C}^{n},$
+where:
+  • the sum
+    $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}$
+    is over the rightmost eigenvalues of $A$ with the longest Jordan chains of length
+    $M_{1}:=L_{1}+1=\max\limits_{\lambda_{i}\in\Lambda_{1}}m_{i}=\max\limits_{\lambda_{i}\in\Lambda_{1}}\max\limits_{j\in\{1,\ldots,d_{i}\}}m_{ij};$
+  • $v^{\left(i,j,1\right)}$ are the eigenvectors of these longest Jordan chains;
+  • $\alpha_{ijM_{1}}(u)$, $u\in\mathbb{C}^{n}$, are the components of $u$ along the last generalized eigenvectors of these longest Jordan chains;
+  • $w^{(i,j,M_{1})}$ are the last left generalized eigenvectors corresponding to these longest Jordan chains.
+Proof.
+By using (12) and Proposition 27 in Appendix A, we have
+$Q_{1}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)}=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)}$
+and, for $u\in\mathbb{C}^{n}$,
+$Q_{1}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}.$
+The previous theorem has the following corollary, which considers a particular situation.
+Corollary 19.
+Suppose $\Lambda_{1}$ consists of only a real eigenvalue $\lambda_{1}$ and there is a unique Jordan mini-block $J^{(1,j)}$, $j\in\{1,\ldots,d_{1}\}$, corresponding to $\lambda_{1}$ of maximum order $M_{1}$.
+If $w^{(1,j,M_{1})}y_{0}\neq 0$ and $w^{(1,j,M_{1})}\widehat{z}_{0}\neq 0$, then
+$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(y_{0},\widehat{z}_{0}\right)=\frac{|w^{(1,j,M_{1})}\widehat{z}_{0}|}{|w^{(1,j,M_{1})}\widehat{y}_{0}|}$
+and
+$K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(y_{0}\right)=\frac{\|w^{(1,j,M_{1})}\|}{|w^{(1,j,M_{1})}\widehat{y}_{0}|},$
+where $\|w^{(1,j,M_{1})}\|$ is the induced norm of the row $w^{(1,j,M_{1})}$.
+Proof.
+In this particular situation regarding $\Lambda_{1}$, the RLGE condition for $u$ is $\alpha_{1jM_{1}}(u)=w^{(1,j,M_{1})}u\neq 0$. Moreover, we have
+$Q_{1}(t)=v^{(1,j,1)}w^{(1,j,M_{1})}$
+and
+$Q_{1}(t)u=\alpha_{1jM_{1}}(u)v^{(1,j,1)}.$
+Therefore, if $w^{(1,j,M_{1})}y_{0}\neq 0$ and $w^{(1,j,M_{1})}\widehat{z}_{0}\neq 0$, then
+$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\|\alpha_{1jM_{1}}(\widehat{z}_{0})v^{(1,j,1)}\|}{\|\alpha_{1jM_{1}}(\widehat{y}_{0})v^{(1,j,1)}\|}=\frac{|\alpha_{1jM_{1}}(\widehat{z}_{0})|}{|\alpha_{1jM_{1}}(\widehat{y}_{0})|}=\frac{|w^{(1,j,M_{1})}\widehat{z}_{0}|}{|w^{(1,j,M_{1})}\widehat{y}_{0}|}.$
+and
+$K_{\infty}\left(t,y_{0}\right)$ $=$ $\frac{\|v^{(1,j,1)}w^{(1,j,M_{1})}\|}{\|\alpha_{1jM_{1}}(\widehat{y}_{0})v^{(1,j,1)}\|}=\frac{\|v^{(1,j,1)}\|\|w^{(1,j,M_{1})}\|}{\|\alpha_{1jM_{1}}(\widehat{y}_{0})v^{(1,j,1)}\|}$
+$=$ $\frac{\|w^{(1,j,M_{1})}\|}{|\alpha_{1jM_{1}}(\widehat{y}_{0})|}=\frac{\|w^{(1,j,M_{1})}\|}{|w^{(1,j,M_{1})}\widehat{y}_{0}|}$
+since the induced norm of the rank-one matrix $v^{(1,j,1)}w^{(1,j,M_{1})}$ is $\|v^{(1,j,1)}\|\|w^{(1,j,M_{1})}\|$. ∎
+The next example considers the asymptotic condition number $K_{\infty}(y_{0})$ in the particular situation considered in the corollary.
+Example 20.
+We consider a real non-diagonalizable matrix $A$ of order $3$ with two distinct real eigenvalues: the defective eigenvalue $r_{1}$ with one Jordan mini-block of dimension $2$ and the non-defective eigenvalue $r_{2}$, where $r_{2}<r_{1}$.
+The Jordan Canonical form of $A$ is
+$J=\left[\begin{array}[]{rrr}r_{1}&1&0\\ 0&r_{1}&0\\ 0&0&r_{2}\end{array}\right].$
+Under the RLGE condition $w^{\left(1,1,2\right)}y_{0}\neq 0$, we have
+$K_{\infty}(y_{0})=\frac{\|w^{(1,1,2)}\|}{|w^{(1,1,2)}\widehat{y}_{0}|}=\frac{\|w^{(1,1,2)}\|\|y_{0}\|}{|w^{(1,1,2)}y_{0}|}.$
+We consider the case where $r_{1}=0$, $r_{2}=-1$,
+$V=\left[\begin{array}[]{rrr}0&0&1\\ 1&0&-1\\ -1&1&0\end{array}\right]\text{\ \ and\ then\ \ }W=V^{-1}=\left[\begin{array}[]{ccc}1&1&0\\ 1&1&1\\ 1&0&0\end{array}\right]=\left[\begin{array}[]{c}w^{(1,1,1)}\\ w^{(1,1,2)}\\ w^{(2,1,1)}\end{array}\right].$
+For the rows of $W$, we are using the notation of Appendix A. The matrix $A$ is
+$A=VJW=\left[\begin{array}[]{rrr}-1&0&0\\ 2&1&1\\ -1&-1&-1\end{array}\right].$ (28)
+The $1$-norm is used as vector norm. We have $w^{(1,1,2)}=[\ 1\ \ 1\ \ 1\ ]$ and then
+$\|w^{(1,1,2)}\|=\left\|\left(w^{(1,1,2)}\right)^{T}\right\|_{\infty}=1.$
+In Figure 4, we see $K(t,y_{0})$ and
+$K_{\infty}(y_{0})=\frac{\|y_{0}\|_{1}}{|y_{01}+y_{02}+y_{03}|}=\frac{|y_{01}|+|y_{02}|+|y_{03}|}{|y_{01}+y_{02}+y_{03}|}$
+for $t\in\left[0,50\right]$, for three different initial values $y_{0}$.
+We can observe a slow approach of $K(t,y_{0})$ to $K_{\infty}(y_{0})$, as $t\rightarrow+\infty$.
+The curves $K(t,y_{0})$ appear, for $t$ close to $50$, nearly parallel to the horizontal lines representing their asymptotic forms, with a small gap remaining between them.
+Indeed, we have a $\frac{1}{t}$-convergence of $K(t,y_{0})$ to $K_{\infty}(y_{0})$ as $t\rightarrow+\infty$, instead of the exponential convergence valid for $L_{1}=0$ (see Theorem 9 and Propositions 2 and 4 for $\epsilon(t)$ and $\epsilon(t,\widehat{y}_{0})$).
+(a) $y_{0}=(-0.8,-2.9,1.4)$
+(b) $y_{0}=(-1,1,0.05)$
+(c) $y_{0}=(1,2,3)$
+Figure 4. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$, $t\in[0,50]$ for the non-diagonalizable matrix (28). The vector norm is the 11-norm.
+The previous Corollary 19 addresses the case in which $\Lambda_{1}$ consists of a real eigenvalue.
+The more intricate case, where $\Lambda_{1}$ consists of a complex conjugate pair of eigenvalues, is treated in [12].
+6. Conclusion
+In the present paper, we have considered how a perturbation in the initial value of the ODE (1) is propagated to the solution over a long time, by measuring the perturbation with a normwise relative error.
+In other words, we have studied the long time relative conditioning of the problem (2).
+We have defined:
+  • a directional pointwise condition number $K(t,y_{0},\widehat{z}_{0})$ such that
+    $K(t,y_{0},\widehat{z}_{0})=\frac{\delta(t)}{\varepsilon},$
+    where $\varepsilon$ is the normwise relative error of the perturbed initial value and $\delta(t)$ is the normwise relative error of the perturbed solution;
+  • a pointwise condition number $K(t,y_{0})$, worst $K(t,y_{0},\widehat{z}_{0})$ by varying the direction of perturbation $\widehat{z}_{0}$;
+  • a global condition number $K(t)$, worst $K(t,y_{0},\widehat{z}_{0})$ by varying both $\widehat{z}_{0}$ and the initial value $y_{0}$.
+The asymptotic forms $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,y_{0})$ and $K_{\infty}^{+}(t)$ of $K(t,y_{0},\widehat{z}_{0})$, $K(t,y_{0})$ and $K(t)$ have been determined in Theorems 7, 9 and 13, respectively.
+For $y_{0}$ and $\widehat{z}_{0}$ satisfying the Rightmost Last Generalized Eigenvector (RLGE) condition, the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$ remain bounded as well as away from zero by varying $t$.
+Expressions for $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$ valid in the RLGE condition are presented in Theorem 18 and Corollary 19.
+Satisfying the RLGE condition is a generic case for $y_{0}$ and $\widehat{z}_{0}$.
+On the contrary, the asymptotic condition number $K_{\infty}^{+}(t)$ is an exponentially diverging function of $t$.
+Moreover, it has been proved in Theorem 11 that $K_{\infty}(t,y_{0})$, the asymptotic form of the worst $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$, coincides with the worst asymptotic form $K_{\infty}(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$.
+On the contrary, it has also been proved that $K_{\infty}^{+}(t)$, the asymptotic form of the worst $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$ and $y_{0}$, does not coincide with the worst asymptotic form $K_{\infty}(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$ and $y_{0}$.
+This worst asymptotic $K_{\infty}(t,y_{0},\widehat{z}_{0})$, denoted by $K_{\infty}(t)$, has been determined in Theorem 14.
+The topic of the present paper is further investigated in the subsequent papers [12] and [13].
+In [12], the case of a generic real ODE (1) is explored in depth, while [13] is less theoretical and deals with more practical issues.
+Acknowledgements: the research was supported by the INdAM Research group GNCS (Gruppo Nazionale di Calcolo Scientifico).
+Appendix A Linear Algebra Results
 We introduce linear algebra notations and results necessary for the analysis developed in this paper.
-Although the content pertains to the well-known topic of the JCF of a matrix, it addresses very specific aspects that are either not known or not sufficiently detailed for the purposes of this paper.
-This section can be skipped and referred back to as needed while reading the subsequent sections.
+Although the content pertains to the well-known topic of the Jordan Canonical Form (JCF) of a matrix, it addresses very specific aspects that are either not known or not sufficiently detailed for the purposes of this paper.
 The main goal of the section is to derive a formula for the matrix exponential $\mathrm{e}^{tA}$, where it is explicitly identified how $\mathrm{e}^{tA}$ depends on $t$.
 Since it is based on the JCF of $A$, first we revise such a form.
-2.1. The JCF of the matrix $A$
+A.1. The JCF of the matrix AA
 Let $A\in\mathbb{C}^{n\times n}$ and let $\lambda_{1},\ldots,\lambda_{p}$ be the distinct eigenvalues of $A$.
@@
     $J^{(i)}=\mathrm{diag}\left(J^{\left(i,1\right)},\ldots,J^{\left(i,d_{i}\right)}\right)\in\mathbb{C}^{\nu_{i}\times\nu_{i}}.$
-  • For any $i\in\{1,\ldots,p\}$ and for any $j\in\{1,\ldots,d_{i}\}$, the Jordan mini-block $J^{\left(i,j\right)}$ is upper bi-diagonal:
+  • For any $i\in\{1,\ldots,p\}$ and for any $j\in\{1,\ldots,d_{i}\}$, the Jordan mini-block $J^{\left(i,j\right)}$ is upper bidiagonal:
     $J^{\left(i,j\right)}=\left[\begin{array}[]{ccccc}\lambda_{i}&1&&&\\ &.&1&&\\ &&.&1&\\ &&&.&1\\ &&&&\lambda_{i}\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}.$
@@
 $J^{(i)}$ is diagonal.
-The matrix $A$ is diagonalizable if and only if all the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ are non-defective, i.e.
-$J$ is diagonal.
+The matrix $A$ is diagonalizable if and only if all the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ are non-defective, i.e., $J$ is diagonal.
 An eigenvalue $\lambda_{i}$, $i\in\{1,\ldots,p\}$, is called simple if $d_{i}=\nu_{i}=1$.
-2.1.1. Three-index notation
+A.1.1. Three-index notation
 In the following, we denote an index $l\in\{1,\ldots,n\}$ by three indices $\left(i,j,k\right)$, where (see Figure 5):
-  • $i\in\{1,\ldots,p\}$ is the index of the block $J^{(i)}$ traversed by the $l$-th column (or row) of $J$.;
+  • $i\in\{1,\ldots,p\}$ is the index of the block $J^{(i)}$ traversed by the $l$-th column (or row) of $J$;
   • $j\in\left\{1,\ldots,d_{i}\right\}$ is the index of the mini-block $J^{(i,j)}$ of $J^{(i)}$ traversed by the $l$-th column (or row) of $J$;
-  • $k\in\left\{1,\ldots,m_{ij}\right\}$ is the index of the column (or row) of the mini-block $J^{(i,j)}$ included in the $l$-th column (row) of $J$.
+  • $k\in\left\{1,\ldots,m_{ij}\right\}$ is the index of the column (or row) of the mini-block $J^{(i,j)}$ included in the $l$-th column (or row) of $J$.
 Observe that the triples $(i,j,k)$ appear lexicographically ordered when the index $l$ moves from $1$ to $n$.
 Figure 5. The index $l$ and the three indices $(i,j,k)$.
-2.1.2. The Jordan basis
+A.1.2. The Jordan basis
 Since $A$ is similar to $J$, there exists $V\in\mathbb{C}^{n\times n}$ non-singular such that
@@
 $v^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\}.$ (31)
+The vectors (31) appear as columns of $V$ in lexicographic order.
 For $i\in\{1,\ldots,p\}$, the vectors
@@
 are generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$: they constitute a basis for the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$.
-2.1.3. Jordan chains
-The Jordan basis (8) is partitioned in the Jordan chains
+A.1.3. Jordan chains
+The Jordan basis (31) is partitioned into the Jordan chains
 $\left(v^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}}=\left(v^{\left(i,j,1\right)},v^{\left(i,j,2\right)},\ldots,v^{\left(i,j,m_{ij}\right)}\right)$
@@
 $\left(A-\lambda_{i}I\right)x=v^{\left(i,j,m_{ij}\right)}$
-has not a solution.
-Observe that (7) is equivalent to have (2.1.3), for $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, and that there is a correspondence one-to-one between mini-blocks and Jordan chains.
-2.1.4. The matrix $zA$
+has no solution.
+Observe that (30) is equivalent to have (A.1.3), for all $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, and that there is a correspondence one-to-one between mini-blocks and Jordan chains.
+A.1.4. The matrix zAzA
 The next proposition describes the JCF and a Jordan basis of the matrix $zA$, where $z\in\mathbb{C}\setminus\{0\}$, in terms of the JCF and a Jordan basis of $A$.
@@
 Let $z\in\mathbb{C}\setminus\{0\}$.
-The distinct eigenvalues of $zA$ are $z\lambda_{i}$$i\in\{1,\ldots,p\}$.
+The distinct eigenvalues of $zA$ are $z\lambda_{i}$, $i\in\{1,\ldots,p\}$.
 For any $i\in\{1,\ldots,p\}$, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ in the JCF of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ in the JCF of $A$.
@@
 is a Jordan chain of $zA$ corresponding to the eigenvalue $z\lambda_{i}$.
-In fact, we have (see (2.1.3))
+In fact, by (A.1.3) we have
 $\left(zA-z\lambda_{i}I\right)v^{\left(i,j,1\right)}=0$
@@
 would have the solution $y=z^{m_{ij}}x$.
-By exchanging the role of $A$ and $zA$, i.e. we consider $zA$ and $z^{-1}(zA)$, we also see that given a Jordan chain
+By exchanging the role of $A$ and $zA$, i.e., we consider $zA$ and $z^{-1}(zA)$, we also see that, given a Jordan chain
 $\left(u^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}(zA)}$
@@
 $\left(z^{k-1}u^{\left(i,j,k\right)}\right)_{k=1,\ldots,m_{ij}(zA)}$
-is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i}$ (here, $m_{ij}(zA)$ denotes the length of a Jordan chain of $zA$).
+is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i}$.
+Here, $m_{ij}(zA)$ denotes the length of a Jordan chain of $zA$.
 This shows that the Jordan chains of $zA$ are all of type (34).
-Consequently, the number and the lenghts of the Jordan chains corresponding to the eigenvalue $z\lambda_{i}$ of $\alpha A$ are equal to the number and the lengths of the Jordan chains corresponding to the eigenvalue $\lambda_{i}$ of $A$.
+Consequently, the number and the lenghts of the Jordan chains corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the lengths of the Jordan chains corresponding to the eigenvalue $\lambda_{i}$ of $A$.
 Therefore, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ of $A$.
 A Jordan basis for $zA$ is given by collecting all the Jordan chains (34). Thus, we obtain the Jordan basis (33) for $zA$. ∎
-2.1.5. The real case
+A.1.5. The real case
 When $A$ is a real matrix, the distinct complex eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are divided in real eigenvalues and complex conjugate pairs of eigenvalues.
@@
 is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
-In fact, by conjugating both sides in (2.1.3), we have
+In fact, by conjugating both sides in (A.1.3), we have
 $\left(A-\overline{\lambda_{i_{1}}}I\right)\overline{v^{\left(i_{1},j,1\right)}}=0$
@@
 would have the solution $y=\overline{x}$.
-By exchanging the role of $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e. we consider $\lambda_{i_{2}}$ and $\overline{\lambda_{i_{2}}}$, we also see that given a Jordan chain
+By exchanging the role of $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e., we consider $\lambda_{i_{2}}$ and $\overline{\lambda_{i_{2}}}$, we also see that, given a Jordan chain
 $\left(u^{\left(i_{2},j,k\right)}\right)_{k=1,\ldots,m_{i_{2}j}}$
@@
 This implies (35).
-By collecting all the Jordan chains (14) corresponding to $\lambda_{i_{1}}$, we obtain a Jordan basis satisfying (13).
+By collecting all the Jordan chains (37) corresponding to $\lambda_{i_{2}}$, we obtain a Jordan basis satisfying (36).
 The result about a real eigenvalue $\lambda_{i}$ follows from the next two facts:
-  • the eigenspace of $\lambda_{i}$ in $\mathbb{C}^{n}$ has a basis of real eigenvalues;
+  • the eigenspace of $\lambda_{i}$ in $\mathbb{C}^{n}$ has a basis of real eigenvectors;
   • if $v^{(i,j,k)}$ is real and the linear system
@@
 In the following, in case of a real matrix $A$, we assume to have a Jordan basis as that described in the previous proposition.
-2.2. The matrices $V^{(i)}$, $W^{(i)}$, $N^{(i,l)}$ and the vector $\alpha^{(i)}(u)$
+A.2. The matrices $V^{(i)}$, $W^{(i)}$, $N^{(i,l)}$ and the vector $\alpha^{(i)}(u)$
 The formula of our interest for the matrix exponential $\mathrm{e}^{tA}$ is constructed by using the matrices $V^{(i)}$, $W^{(i)}$ and $N^{(i,l)}$ now introduced.
@@
 We also introduce the vector $\alpha^{(i)}(u)$ of components of $u$ in the Jordan basis.
-Remind that $V$ is the matrix whose columns (8) constitute a Jordan basis.
+Recall that $V$ is the matrix whose columns (31) constitute a Jordan basis.
   • For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
@@
     $V=\left[V^{(1)}\ \cdots\ V^{(p)}\right].$
-  • Let $W=V^{-1}$ and let
+  • Let $W:=V^{-1}$ and let
     $w^{\left(i,j,k\right)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$ (39)
     be the rows of $W$ in the three-index notation.
-The $n$ row vectors in (16) are called left generalized eigenvectors of $A$, whereas, as we have already seen, the $n$ columns vectors in (8) are called (right) generalized eigenvectors of $A$.
+They appear in $W$ in lexicographic order.
+The $n$ row vectors in (39) are called left generalized eigenvectors of $A$, whereas, as we have already seen, the $n$ column vectors in (31) are called (right) generalized eigenvectors of $A$.
 For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
@@
     $\alpha_{ijk}(u)$ is the component of $u$ along $v^{(i,j,k)}$. For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
-    $\alpha^{(i,j)}(u):=\left[\begin{array}[]{c}\alpha_{ij1}(u)\\ \vdots\\ \alpha_{ijm_{ij}}(u)\end{array}\right]\in\mathbb{C}^{m_{ij}}.$
+    $\alpha^{(i,j)}(u):=\left[\begin{array}[]{c}\alpha_{ij1}(u)\\ \vdots\\ \alpha_{ijm_{ij}}(u)\end{array}\right]\in\mathbb{C}^{m_{ij}}$
     and let
@@
   • For $i\in\{1,\ldots,p\}$, $l\in\{0,\ldots,m_{i}-1\}$ and $j\in\{1,\ldots,d_{i}\}$, let
-    $N^{(i,l,j)}:=\left\{\begin{array}[]{l}\left[\begin{array}[]{ccccccc}0&.&0&1&0&.&0\\ &.&.&.&.&.&.\\ &&.&.&.&.&0\\ &&&.&.&.&1\\ &&&&.&.&0\\ &&&&&.&.\\ &&&&&&0\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }l\leq m_{ij}-1\\ \\ 0\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }m_{ij}\leq l\leq m_{i}-1,\end{array}\right.$ (18)
+    $N^{(i,l,j)}:=\left\{\begin{array}[]{l}\left[\begin{array}[]{ccccccc}0&.&0&1&0&.&0\\ &.&.&.&.&.&.\\ &&.&.&.&.&0\\ &&&.&.&.&1\\ &&&&.&.&0\\ &&&&&.&.\\ &&&&&&0\end{array}\right]\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }l\leq m_{ij}-1\\ \\ 0\in\mathbb{C}^{m_{ij}\times m_{ij}}\text{ if }m_{ij}\leq l\leq m_{i}-1,\end{array}\right$.
+(41)
     where the upper diagonal of elements equal to $1$ is the $l$-th upper diagonal. For $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$, let
     $N^{(i,l)}:=\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\in\mathbb{C}^{\nu_{i}\times\nu_{i}}.$ (42)
-    Observe that the matrix $N^{(i,l)}$ has the same dimensions $\nu_{i}\times\nu_{i}$ of the Jordan block $J^{(i)}=\mathrm{diag}\left(J^{(i,1)},\ldots,J^{(i,d_{i})}\right)$ and the diagonal blocks $N^{(i,l,1)},$ $\ldots$,… , $N^{(i,l,d_{i})}$ have the same dimensions of the Jordan mini-blocks $J^{(i,1)},\ldots,$ $J^{(i,d_{i})}$, respectively.
-Thus, as it is illustrated in Figure 2, the matrix $N^{(i,l)}$ has $1$ in the intersection of the $l$-th upper diagonal with the (frame of the) Jordan mini-blocks and 00 in all the other places.
+    Observe that the matrix $N^{(i,l)}$ has the same dimensions $\nu_{i}\times\nu_{i}$ of the Jordan block $J^{(i)}=\mathrm{diag}\left(J^{(i,1)},\ldots,J^{(i,d_{i})}\right)$ and the diagonal blocks $N^{(i,l,1)},$ $\ldots,$ $N^{(i,l,d_{i})}$ have the same dimensions of the Jordan mini-blocks $J^{(i,1)},\ldots,$ $J^{(i,d_{i})}$, respectively.
+Thus, as it is illustrated in Figure 6, the matrix $N^{(i,l)}$ has $1$ in the intersection of the $l$-th upper diagonal with the (frame of the) Jordan mini-blocks and 0 in all the other places.
+Remark 23.
+Regarding the matrices $N^{(i,l,j)}$ and $N^{(i,l)}$, where $i\in\{1,\dots,p\}$, $l\in\{1,\ldots,m_{i}-1\}$ and $j\in\{1,\ldots,d_{i}\}$, note that
+$N^{(i,l,j)}=\left(N^{(i,1,j)}\right)^{l}\text{\ \ and\ \ }N^{(i,l)}=\left(N^{(i,1)}\right)^{l}.$
+Viewing $N^{(i,l,j)}$ and $N^{(i,l)}$ as powers of the nilpotent matrices $N^{(i,1,j)}$ and $N^{(i,1)}$ is how the matrices $N^{(i,l,j)}$ and $N^{(i,l)}$ are presented in usual expositions of the JCF.
 Figure 6. The matrix $N^{(i,l)}$.
-2.2.1. The matrix $zA$
-When we replace $A$ by $zA$, $z\in\mathbb{C}\setminus\{0\}$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2).
+A.2.1. The matrix zAzA
+When we replace $A$ by $zA$, $z\in\mathbb{C}\setminus\{0\}$, the number and the dimensions of blocks and mini-blocks remain the same (recall Proposition 21).
 Moreover, a Jordan basis for $zA$ is given in (33).
-Therefore, we know how it is transformed the matrix $V$ by replacing $A$ by $zA$.
-The next proposition says this and, in addition, how it is transformed the matrix $W=V^{-1}$.
+Therefore, we know how the matrix $V$ is transformed by replacing $A$ by $zA$.
+The next proposition says this and, in addition, how the matrix $W=V^{-1}$ is transformed.
 Proposition 24.
@@
 since $W$, whose rows are (39), is the inverse of $V$, whose columns are (31). This shows that the matrix of rows (44) is the inverse of the matrix of columns (43). ∎
-2.2.2. The real case
+A.2.2. The real case
 The next proposition describes, when the matrix $A$ and the vector $u$ are real, the matrices $V^{(i)}$ and $W^{(i)}$ and the vector $\alpha^{(i)}(u)$ in terms of the real eigenvalues and the complex conjugate pairs of eigenvalues of $A$.
 Proposition 25.
-Assume $A\in\mathbb{R}^{n\times n}$ and $u\in\mathbb{R}^{n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
+Assume $A\in\mathbb{R}^{n\times n}$ and $u\in\mathbb{R}^{n}$. Moreover, assume we have a Jordan basis as that described in Proposition 22.
 For $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$, we have
@@
 Proof.
-Since we are assuming to have a Jordan basis as that described in Proposition 3, we have $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
-Now, we prove that for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ we have $\alpha^{(i)}(u)\in\mathbb{R}$, and for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$ we have $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$.
+Since we have a Jordan basis as that described in Proposition 22, we have $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
+Now, we prove that for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ we have $\alpha^{(i)}(u)\in\mathbb{R}$, and for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ with $\mathrm{Im}\left(\lambda_{i_{1}}\right)>0$ we have $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$.
 By conjugating both sides of
@@
 $u=\sum\limits_{i=1}^{p}\overline{V^{(i)}}\ \overline{\alpha^{(i)}(u)}.$
-So, by separating real eigenvalues and complex conjugate pairs of eigenvalues, we have
+Therefore, by separating real eigenvalues and complex conjugate pairs of eigenvalues, we have
 $u=\sum\limits_{i=1}^{p}V^{(i)}\alpha^{(i)}(u)=\sum\limits_{\lambda_{i}\in\mathbb{R}}V^{(i)}\alpha^{(i)}(u)+\sum\limits_{\begin{subarray}{c}\lambda_{i_{1}}\in\mathbb{C}\\ \mathrm{Im}\left(\lambda_{i_{1}}\right)>0\end{subarray}}\left(V^{(i_{1})}\alpha^{(i_{1})}(u)+V^{(i_{2})}\alpha^{(i_{2})}(u)\right),$
@@
 $u=\sum\limits_{i=1}^{p}\overline{V^{(i)}}\ \ \overline{\alpha^{(i)}(u)}=\sum\limits_{\lambda_{i}\in\mathbb{R}}V^{(i)}\ \overline{\alpha^{(i)}(u)}+\sum\limits_{\begin{subarray}{c}\lambda_{i_{1}}\in\mathbb{C}\\ \mathrm{Im}\left(\lambda_{i_{1}}\right)>0\end{subarray}}\left(V^{(i_{2})}\overline{\alpha^{(i_{1})}(u)}+V^{(i_{1})}\overline{\alpha^{(i_{2})}(u)}\right)$
-by reminding that $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$.
+by recalling that $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$.
 Since $u$ can be expressed as a linear combination of the Jordan basis in a unique manner, we obtain $\alpha^{(i)}(u)=\overline{\alpha^{(i)}(u)}$, i.e.
@@
 $=$ $\overline{W^{(i_{1})}\left[e^{\left(1\right)}\ \cdots\ e^{\left(n\right)}\right]}=\overline{W^{(i_{1})}}.$
-2.3. The formula for $\mathrm{e}^{tA}$
+A.3. The formula for $\mathrm{e}^{tA}$
 Next proposition provides the announced formula for the matrix exponential $\mathrm{e}^{tA}$.
-For taste of generality, we consider the general situation of a function of matrix $f(A)$, where $f(z)$, $z\in\mathcal{D}\subseteq\mathbb{C}$, is an analytic complex function of $z$.
+For sake of generality, we consider a matrix function $f(A)$, where $f(z)$, $z\in\mathcal{D}\subseteq\mathbb{C}$, is an analytic complex function of $z$.
 The domain $\mathcal{D}$ of $f$ is an open subset of $\mathbb{C}$ and we assume that the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are contained in $\mathcal{D}$.
@@
 Proof.
-By recalling the definition of function of matrix by the JCF (see, e.g., [8]), we have
+By recalling the definition of matrix function by the JCF (see, e.g., [7]), we have
 $f(A)=Vf(J)W=\sum\limits_{i=1}^{p}V^{\left(i\right)}f(J^{(i)})W^{\left(i\right)},$ (47)
@@
 $\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il}.$ (49)
-2.3.1. The matrices $P_{i0}$
-For $i\in\{1,\ldots,p\}$, since $N^{(i,0)}=I_{n}$, we have
+A.3.1. The matrices $P_{i0}$
+For $i\in\{1,\ldots,p\}$, since $N^{(i,0)}=I_{\nu_{i}}$, we have
 $P_{i0}=V^{\left(i\right)}W^{\left(i\right)}=\sum\limits_{j=1}^{d_{i}}\sum\limits_{k=1}^{m_{ij}}v^{(i,j,k)}w^{(i,j,k)}.$ (50)
-The matrix $P_{i0}$ is the projection onto the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e.
-the subspace spanned by the generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
+The matrix $P_{i0}$ is the projection onto the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e., the subspace spanned by the generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
 In fact, we have, for $u\in\mathbb{C}^{n}$,
 $P_{i0}u=V^{\left(i\right)}W^{\left(i\right)}u=V^{\left(i\right)}\alpha^{\left(i\right)}(u)=\sum\limits_{j=1}^{d_{i}}\sum\limits_{k=1}^{m_{ij}}\alpha_{ijk}(u)v^{(i,j,k)}.$
-2.3.2. The case $A$ diagonalizable
+A.3.2. The case $A$ diagonalizable
 When $A$ is diagonalizable, the formula (45) simplifies to the well-known formula
@@
 $P_{i0}=\sum\limits_{j=1}^{d_{i}}v^{(i,j,1)}w^{(i,j,1)}.$
-In this case of $A$ diagonalizable, the matrix $P_{i0}$ is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e.
-the subspace spanned by the eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
+In this case of $A$ diagonalizable, the matrix $P_{i0}$ is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e., the subspace spanned by the eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
 For the case $f(A)=\mathrm{e}^{tA}$ of our interest, formula (51) becomes
 $\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}P_{i0}.$
-2.4. The matrices $P_{il}$
+A.4. The matrices $P_{il}$
 In this subsection, we see some properties of the matrices $P_{il}$ defined in (46).
-2.4.1. An alternative expression for $P_{il}$
+A.4.1. An alternative expression for $P_{il}$
 Next proposition gives an alternative expression for $P_{il}$, more informative than the definition (46).
@@
 $P_{il}u=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}\alpha_{ij\,l+k}(u)v^{\left(i,j,k\right)}.$ (53)
-The indeces $j$ in the outer sum
+The indices $j$ in the outer sum
 $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}$
-in (29) and (30) are the indeces $j$ of the mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal. The indeces $k$ in the inner sum
+in (52) and (53) are the indices $j$ of the mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal. The indices $k$ in the inner sum
 $\sum\limits_{k=1}^{m_{ij}-l}$
-are the row indeces of the elements of the mini-block $J^{(i,j)}$ on this $l$-th upper diagonal (see Figure 3).
-The column indeces $l+k$ of these elements appear as third indeces in $w^{(i,j,l+k)}$ in (29) and $\alpha_{ij\,l+k}(u)$ in (30).
-In other words, with reference to the previous Figure 2, in the double sum in (29) and (30) we are summing over all elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with the mini-blocks: the three-index notation $(i,j,k)$ of the row of these elements appears in $v^{(i,j,k)}$ and the three-index notation $(i,j,l+k)$ of the columns appears in $w^{(i,j,l+k)}$ for (29) and in $\alpha_{ijl+k}(u)$ for (30).
+are the row indices of the elements of the mini-block $J^{(i,j)}$ on this $l$-th upper diagonal (see Figure 7).
+The column indices $l+k$ of these elements appear as third indices in $w^{(i,j,l+k)}$ in (52) and $\alpha_{ij\,l+k}(u)$ in (53).
+In other words, with reference to the previous Figure 6, in the double sum in (52) and (53) we are summing over all elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with the mini-blocks: the three-index notation $(i,j,k)$ of the row index of these elements appears in $v^{(i,j,k)}$ and the three-index notation $(i,j,l+k)$ of the column index appears in $w^{(i,j,l+k)}$ for (52) and in $\alpha_{ijl+k}(u)$ for (53).
 Proof.
-As for (29), by the definition (23) we have
-$P_{il}$ $=$ $\sum\limits_{j=1}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
-$=$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
+Regarding (52), by the definition (46) we have
+$P_{il}$ $=$ $\sum\limits_{j=1}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
 $=$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}v^{\left(i,j,k\right)}w^{\left(i,j,l+k\right)}$
-by reminding the form (18) of $N^{(i,l,j)}$.
-As for (30), we have
+by recalling the form (41) of $N^{(i,l,j)}$.
+Regarding (53), we have
 $P_{il}u=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}v^{\left(i,j,k\right)}\underset{=\alpha_{ij\,l+k}(u)}{\underbrace{w^{\left(i,j,l+k\right)}u}}$.
 Figure 7.
-Indeces $j$ and $k$ in (29) and (30): $j$ is an index of mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal and $k$ is an index of row of elements in the $l$-th upper diagonal.
+Indices $j$ and $k$ in (52) and (53): $j$ is an index of mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal and $k$ is an index of row of elements in the $l$-th upper diagonal.
 Remark 28.
@@
   • 2.
-    If there exists an unique $j\in\{1,\ldots,d_{i}\}$ such that $m_{i}=m_{ij}$, then
+    If there exists a unique $j\in\{1,\ldots,d_{i}\}$ such that $m_{i}=m_{ij}$, then
     $P_{i\,m_{i}-1}=v^{\left(i,j,1\right)}w^{\left(i,j,m_{i}\right)}$
@@
     since these vectors are multiplied by coefficients different from the corresponding components $\alpha_{ijk}(u)$.
-The coefficients are the components $\alpha_{ijl+k}(u)$ shifted by $l$ in the third index.
-2.4.2. The matrix $zA$
+These coefficients are the components $\alpha_{ijl+k}(u)$ shifted by $l$ in the third index.
+A.4.2. The matrix zAzA
 The next proposition shows how the matrices $P_{il}$ are transformed when the matrix $A$ is replaced by $zA$, $z\in\mathbb{C}\setminus\{0\}$.
-Observe that, when $A$ is replaced by $zA$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2).
-Thus, the indeces $i$ and $l$ for the matrices $P_{il}(zA)$ corresponding to $zA$ are the same indeces $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ for the matrices $P_{il}$ corresponding to $A$.
+Observe that, when $A$ is replaced by $zA$, the number and the dimensions of blocks and mini-blocks remain the same (see Proposition 21).
+Thus, the indices $i$ and $l$ for the matrices $P_{il}(zA)$ corresponding to $zA$ are the same indices $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ for the matrices $P_{il}$ corresponding to $A$.
 Proposition 29.
@@
 Proof.
-The columns of the matrix $V(zA)$ corresponding to $zA$ are given in (20) and the rows of the matrix $W(zA)$ corresponding to $zA$ are given in (21). Now, use the formula (29). ∎
-2.4.3. The real case
+Recall Proposition 24.
+The columns of the matrix $V(zA)$ corresponding to $zA$ are given in (43) and the rows of the matrix $W(zA)$ corresponding to $zA$ are given in (44).
+Now, use the formula (52).
+∎
+A.4.3. The real case
 The next proposition describes the matrices $P_{il}$, when $A$ is real, in terms of real eigenvalues and complex conjugate pairs of eigenvalues of $A$.
 Proposition 30.
-Assume $A\in\mathbb{R}^{n\times n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
+Assume $A\in\mathbb{R}^{n\times n}$. Moreover, assume we have a Jordan basis as that described in Proposition 22.
 For $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$, we have
@@
 $P_{i_{2}l}=V^{\left(i_{2}\right)}N^{(i_{2},l)}W^{\left(i_{2}\right)}=\overline{V^{\left(i_{1}\right)}}N^{(i_{1},l)}\overline{W^{\left(i_{1}\right)}}=\overline{V^{\left(i_{1}\right)}N^{(i_{1},l)}W^{\left(i_{1}\right)}}=\overline{P_{i_{1}l}}.$
-2.4.4. The kernel of the matrices $P_{il}$
+A.4.4. The kernel of the matrices $P_{il}$
 Next three propositions concern the kernel of the matrices $P_{il}$. The first proposition describes this kernel.
@@
 $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.\ \text{and for all }k\in\{l+1,\ldots,m_{ij}\}\right\}.$
-In other words, with reference to the previous Figure 2, the proposition says that the kernel of $P_{il}$ is constituted by the vectors $u\in\mathbb{C}^{n}$ with zero component along all the generalized eigenvectors $v^{(i,j,k)}$ whose third index $k$ (in the three-index notation) is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
+Remark 32.
+With reference to the previous Figure 6 and 7, Proposition 31 states that the kernel of $P_{il}$ is constituted by the vectors $u\in\mathbb{C}^{n}$ with zero component $\alpha_{ijk}(u)$ along all the generalized eigenvectors $v^{(i,j,k)}$ such that, in the three-index notation, $(i,j,k)$ is an index column of elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with mini-blocks.
 Proof.
-For $u\in\mathbb{C}^{n}$, we have
-$P_{il}u=0$$\Leftrightarrow$ $V^{\left(i\right)}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
-$\Leftrightarrow$ $N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=0$
-by the linear independence of the vectors $v^{(i,j,k)}$ in (8)
-$\Leftrightarrow$ $N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0\ \text{for\ all\ }j=1,\ldots,d_{i},$
+For $u\in\mathbb{C}^{n}$, by Proposition 27 we have
+$P_{il}u=0$ $\Leftrightarrow$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}\sum\limits_{k=1}^{m_{ij}-l}\alpha_{ij\,l+k}(u)v^{\left(i,j,k\right)}=0$
 $\Leftrightarrow$ $\alpha_{ijk}\left(u\right)=0\ \text{for\ all }j\in\{1,\ldots,d_{i}\}\ \text{with $m_{ij}\geq l+1$}$
- and for all $\text{\ and\ for all\ }k\in\{l+1,\ldots,m_{ij}\}.$
+and for all $k\in\{l+1,\ldots,m_{ij}\}$,
+where the second $\Leftrightarrow$ follows by the linear independence of the vectors $v^{(i,j,k)}$ in (31). ∎
 Next two propositions are immediate consequences of the previous proposition.
@@
 $u\in\mathrm{ker}\left(P_{i0}\right)\ \ \Leftrightarrow\ \ \alpha^{\left(i\right)}\left(u\right)=0.$
-The Proposition 13 confirms our previous observation in Subsection 2.4 that $P_{i0}$ is the projection on the generalized eigenspace corresponding to $\lambda_{i}$.
-2.5. The index $l_{i}(u)$
-In view of the Propositions 12 and 13, for $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{(i)}(u)\neq 0$, we define the index
-$l_{i}\left(u\right):=\max\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}$
+Proposition 34 confirms our previous observation in Subsection A.3.1 that $P_{i0}$ is the projection on the generalized eigenspace corresponding to $\lambda_{i}$.
+A.5. The index $l_{i}(u)$
+In view of the Propositions 33 and 34, for $i\in\{1,\ldots,p\}$ and $u\in\mathbb{C}^{n}$ such that $\alpha^{(i)}(u)\neq 0$, we define the index
+$l_{i}\left(u\right):=\max\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}$ (54)
 (see Figure 8).
@@
 $\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}=\emptyset$
-by Propositions 12 and 13. In this case, the index $l_{i}(u)$ cannot be defined. On the other hand, if $\alpha^{(i)}(u)\neq 0$, then
+by Propositions 33 and 34. In this case, the index $l_{i}(u)$ cannot be defined. On the other hand, if $\alpha^{(i)}(u)\neq 0$, then
 $\left\{l\in\left\{0,\ldots,m_{i}-1\right\}:u\notin\mathrm{ker}\left(P_{il}\right)\right\}\neq\emptyset,$
 since $u\notin\mathrm{ker}(P_{i0})$ by Proposition 34. In this case, the index $l_{i}(u)$ can be defined. By Proposition 33, we have
-$u\notin\mathrm{ker}(P_{il}),\ l\in\{0,\ldots,l_{i}(u)\},$
+$u\notin\mathrm{ker}(P_{il}),\ l\in\{0,\ldots,l_{i}(u)\}$
 and
-$u\in\mathrm{ker}(P_{il}),\ l\in\{l_{i}(u)+1,\ldots,m_{i}\}.$
-Observe that if $\lambda_{i}$ is a non-defective eigenvalue, i.e. $m_{i}=0$, then $l_{i}\left(u\right)=0$.
+$u\in\mathrm{ker}(P_{il}),\ l\in\{l_{i}(u)+1,\ldots,m_{i}-1\},$
+as illustrated in Figure 8.
+Observe that if $\lambda_{i}$ is a non-defective eigenvalue, i.e., $m_{i}=1$, and $\alpha^{(i)}(u)\neq 0$, then $l_{i}\left(u\right)=0$.
 Figure 8.
@@
 In this case $l_{i}(u)=2$, since $u\in\mathrm{ker}\left(P_{i2}\right)^{c}$ but $u\notin\mathrm{ker}\left(P_{i3}\right)^{c}$.
-The following proposition gives $l_{i}\left(u\right)$ in terms of the zero components of $\alpha^{\left(i\right)}\left(u\right)$.
+Remark 35.
+By recalling the definition (54) of $l_{i}(u)$ and Remark 32 regarding $\mathrm{ker}(P_{il})$ and referencing Figures 6 and 7, we can say that $l_{i}(u)$ is the maximum index $l$ for which $u$ has a non-zero component $\alpha_{ijk}(u)$ along some generalized eigenvector $v^{(i,j,k)}$ such that, in the three-index notation, $(i,j,k)$ is an index column of elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with mini-blocks.
+The following proposition relates $l_{i}(u)$ to Jordan chains.
 Proposition 36.
@@
 $=\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\}.$
-In other words, the proposition says that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
+In other words, the proposition states that $l_{i}(u)+1$ is the maximum index $k$ such that $u$ has non-zero component $\alpha_{ijk}(u)$ along the $k$-th generalized eigenvector $v^{(i,j,k)}$ of some Jordan chain
 $(v^{(i,j,k)})_{k=1,\ldots,m_{ij}},\ j\in\{1,\ldots,d_{i}\},$
-corresponding to the eigenvalue $\lambda_{i}$, such that $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$.
-Alternatively, with reference to Figure 2, the proposition says that $l_{i}(u)$ is the maximum index $l$ such that $u$ has a non-zero component along some generalized eigenvector $v^{(i,j,k)}$ whose third index $k$ is an index column of elements in the intersection of $l$-th diagonal with mini-blocks.
+corresponding to the eigenvalue $\lambda_{i}$.
 Proof.
-The proof is straightforward by recalling the interpretation of the statement of Proposition 14 from a few lines above, referencing Figure 2, and the observation about $\mathrm{ker}(P_{il})$, also referencing Figure 2, immediately following Proposition 11.
-However, a more formal proof is now presented below.
+The proof should be straightforward by recalling Remark 35. Nevertheless, a more formal proof is presented below.
 Suppose $l_{i}\left(u\right)=m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{im_{i}-1}\right)$.
@@
 Suppose $l_{i}\left(u\right)<m_{i}-1$ and then $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$ and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$.
-By Proposition 11 and $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$, there exist $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}(u)+1$ and $\overline{k}\in\{l_{i}\left(u\right)+1,\ldots,m_{i\overline{j}}\}$ such that $\alpha_{i\overline{j}\ \overline{k}}\neq 0$.
+By Proposition 31 and $u\notin\mathrm{ker}\left(P_{il_{i}(u)}\right)$, there exist $\overline{j}\in\{1,\ldots,d_{i}\}$ with $m_{i\overline{j}}\geq l_{i}(u)+1$ and $\overline{k}\in\{l_{i}\left(u\right)+1,\ldots,m_{i\overline{j}}\}$ such that $\alpha_{i\overline{j}\ \overline{k}}(u)\neq 0$.
 Then
@@
 $\leq\max\left\{k\in\left\{1,\ldots,m_{i}\right\}:\alpha_{ijk}\left(u\right)\neq 0\text{ for some }j\in\{1,\ldots,d_{i}\}\text{\ with $m_{ij}\geq k$}\right\}.$
-On the other hand, by Proposition 11 and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$, we have $\alpha_{ijk}\left(u\right)=0$for all $j\in\{1,\ldots,d_{i}\}$ with $m_{ij}\geq l_{i}(u)+2$ and for all $k\in\{l_{i}\left(u\right)+2,\ldots,m_{ij}\}$.
+On the other hand, by Proposition 31 and $u\in\mathrm{ker}\left(P_{i\,l_{i}(u)+1}\right)$, we have $\alpha_{ijk}\left(u\right)=0\$for all $j\in\{1,\ldots,d_{i}\}$ with $m_{ij}\geq l_{i}(u)+2$ and for all $k\in\{l_{i}\left(u\right)+2,\ldots,m_{ij}\}$.
 Then
@@
 Thus, (55) holds. ∎
-2.6. Linear independence
+A.6. Linear independence
 This subsection deals with the linear independence of the vectors $P_{il}u$ and the matrices $P_{il}$.
@@
 $0=\sum\limits_{\begin{subarray}{c}i=1\\ \alpha^{(i)}(u)\neq 0\end{subarray}}^{p}\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}P_{il}u=\sum\limits_{\begin{subarray}{c}i=1\\ \alpha^{(i)}(u)\neq 0\end{subarray}}^{p}V^{\left(i\right)}\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right),$ (57)
-where the second inequality follows by (23). Since $V$ has linearly independent columns, we obtain, for $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$,
-00 0 $=$ $\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\alpha^{\left(i\right)}\left(u\right)$
+where the second equality follows by (46). Since $V$ has linearly independent columns, we obtain, for $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$,
+$0$ $=$ $\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l)}\alpha^{\left(i\right)}\left(u\right)=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}\mathrm{diag}\left(N^{(i,l,1)},\ldots,N^{(i,l,d_{i})}\right)\alpha^{\left(i\right)}\left(u\right)$
 $=$ $\left(\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,1)}\alpha^{\left(i,1\right)}\left(u\right),\ldots,\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,d_{i})}\alpha^{\left(i,d_{i}\right)}\left(u\right)\right)$
-and then, for $j\in\{1,\ldots,d_{i}\}$,
-$\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0.$
+and then
+$\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,j)}\alpha^{\left(i,j\right)}\left(u\right)=0,\ j\in\{1,\ldots,d_{i}\}.$
 Now, fix $i\in\{1,\ldots,p\}$ such that $\alpha^{(i)}(u)\neq 0$.
@@
 $0=\sum\limits_{l=0}^{l_{i}\left(u\right)}c_{il}N^{(i,l,\overline{j})}\alpha^{\left(i,\overline{j}\right)}\left(u\right)$
-$=\left[\begin{array}[]{ccccccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}&|&0&\cdot&0\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&|&c_{il_{i}\left(u\right)}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{il_{i}\left(u\right)-1}&\cdot&0\\ \cdot&\cdot&\cdot&c_{i0}&c_{i1}&|&\cdot&\cdot&c_{il_{i}\left(u\right)}\\ \cdot&\cdot&\cdot&\cdot&c_{i0}&|&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ -&-&-&-&-&&-&-&-\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{i0}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&\cdot&\cdot&c_{i1}\\ 0&\cdot&\cdot&\cdot&\cdot&|&\cdot&0&c_{i0}\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}l_{i}\left(u\right)+1}\left(u\right)\\ -\\ 0\\ \cdot\\ 0\end{array}\right]$
+${\footnotesize=\left[\begin{array}[]{ccccccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}&|&0&\cdot&0\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&|&c_{il_{i}\left(u\right)}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{il_{i}\left(u\right)-1}&\cdot&0\\ \cdot&\cdot&\cdot&c_{i0}&c_{i1}&|&\cdot&\cdot&c_{il_{i}\left(u\right)}\\ \cdot&\cdot&\cdot&\cdot&c_{i0}&|&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ -&-&-&-&-&&-&-&-\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&c_{i0}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&|&\cdot&\cdot&c_{i1}\\ 0&\cdot&\cdot&\cdot&\cdot&|&\cdot&0&c_{i0}\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}l_{i}\left(u\right)+1}\left(u\right)\\ -\\ 0\\ \cdot\\ 0\end{array}\right]}$
 and then
 $0=\left[\begin{array}[]{ccccc}c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}&c_{il_{i}\left(u\right)}\\ 0&c_{i0}&c_{i1}&\cdot&c_{il_{i}\left(u\right)-1}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 0&0&\cdot&c_{i0}&c_{i1}\\ 0&0&\cdot&0&c_{i0}\\ &&&&\end{array}\right]\left[\begin{array}[]{c}\alpha_{i\overline{j}1}\left(u\right)\\ \alpha_{i\overline{j}2}\left(u\right)\\ \cdot\\ \alpha_{i\overline{j}l_{i}\left(u\right)}\left(u\right)\\ \alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\\ \end{array}\right].$
-Now, since $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$, we obtain $c_{il}=0$ successively for $l=0,1,\ldots,l_{i}\left(u\right)$.
+Since $\alpha_{i\overline{j}\,l_{i}\left(u\right)+1}\left(u\right)\neq 0$, we obtain $c_{il}=0$ successively for $l=0,1,\ldots,l_{i}\left(u\right)$.
 We have proved that the coefficients of the zero linear combination (57) are all zero. ∎
@@
 We have $l_{i}(u)=m_{i}-1$ for $i\in\{1,\ldots,p\}$.
-Since
+Since we have
 $0=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{m_{i}-1}c_{il}P_{il}u=\sum\limits_{i=1}^{p}\sum\limits_{l=0}^{l_{i}(u)}c_{il}P_{il}u$
-and $P_{il}u$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,l_{i}(u)\}$, are linearly independent (see previous Proposition 15), we have $c_{il}=0$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$. ∎
-2.6.1. Linear combinations depending on $t$
-In the present paper, we deal with linear combinations of some of the vectors $P_{il}u$ in (32) or some of the matrices $P_{il}$ in (36), whose coefficients depend on time $t$ and they are not all zero.
-For a fixed $t$, such a linear combination is not zero since the vectors and the matrices are linearly independent (recall the two previous Propositions 15 and 16).
+and $P_{il}u$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,l_{i}(u)\}$, are linearly independent by the previous Proposition 37, we obtain $c_{il}=0$, $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$.
+∎
+A.6.1. Linear combinations depending on $t$
+In the present paper, we deal with linear combinations of some of the vectors $P_{il}u$ in (56), or some of the matrices $P_{il}$ in (60), whose coefficients depend on time $t$ and they are not all zero.
+For a fixed $t$, such a linear combination is not zero since the vectors and the matrices are linearly independent: recall the two previous Propositions 37 and 38.
 However, in the analysis of the asymptotic behavior of the condition numbers, we require that the norm of the linear combination is away from zero, uniformly with respect to $t$.
@@
 Lemma 39.
-[…]
+Let $\mathcal{V}$ be a vector space over $\mathbb{C}$ equipped with the norm $\left\|\ \cdot\ \right\|$, let $a_{1},\ldots,a_{K}\in\mathcal{V}$ and let $f:I\rightarrow\mathbb{C}^{K}$, where $I$ is an arbitrary set.
+For any $t\in I$, the components $f_{1}(t),\ldots,f_{K}(t)$ of $f(t)$ serve as coefficients in the linear combination
 $\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}$
@@
 $\inf_{t\in I}\left\|\sum\limits_{k=1}^{K}f_{k}\left(t\right)a_{k}\right\|\geq m\inf_{t\in I}\|f(t)\|_{\infty}>0.$
-In the paper, we apply this lemma to the case where $a_{1},\ldots,a_{K}$ are some of the linearly independent vectors $P_{il}u$ in (32), or some of the linearly independent matrices $P_{il}$ in (36), and $f_{1}(t),\ldots,f_{K}(t)$ are of the form $\mathrm{e}^{\sqrt{-1}\omega t}$, with $\sqrt{-1}$ the imaginary unit and $\omega\in\mathbb{R}$, and then $\|f(t)\|_{\infty}=1$.
-3. Asymptotic forms
-In the next Section 4, we analyze the asymptotic (long-time) behavior of the three condition numbers (3), (4) and (5), i.e. what happens to (3), (4) and (5) as $t$ approaches infinity (becomes large).
-This analysis determines the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (3), of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (4) and of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ in (5).
-Then, the asymptotic behavior of the condition numbers is described by substituting in (3), (4) and (5) these asymptotic forms and by defining asymptotic condition numbers the new expressions obtained by these substitutions.
-All of this might seem straightforward, but from a mathematical perspective it is not a simple task. Specifically:
-  • It is necessary to precisely define what we mean by asymptotic form and how to determine it.
-This is particularly complicated in the non-diagonalizable case, where generalized eigenvectors are involved.
-  • We also want to quantify how dominant the asymptotic form is at finite times.
-  • The core of our analysis is the qualitative study of the asymptotic condition numbers. Understanding them requires significant mathematical effort in case of complex eigenvalues.
-  • If we are interested in defining asymptotic condition numbers for the problem (2), the asymptotic condition numbers derived from (4) and (5) as $t\rightarrow+\infty$ may not be appropriate, since they represent the asymptotic worst cases of (3) and (4), respectively.
-Instead, we are concerned with the worst asymptotic cases of (3) and (4) by varying $\widehat{z}_{0}$ and $y_{0}$.
-In other words: do “asymptotic” and “worst” commute?
-In this section, we determine the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, where $u$ is a vector.
-In order to define them some preliminary notions and results need to be introduced.
-3.1. Notations $\sim$ and $\approx$
-In this subsection, we precise what we mean for asymptotic form.
-Let $f(t)$ and $g(t)$ be scalar, vector or matrix functions of $t\in\mathbb{R}$. For $g$ such that $g(t)\neq 0$ for $t$ in a neighborhood of +$\infty$+ ∞, we write
-$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$ (38)
-when
-$\lim\limits_{t\rightarrow+\infty}\frac{\|f(t)-g(t)\|}{\|g(t)\|}=0.$
-[…]
-We interpret (38) as indicating that $g$ is an asymptotic form of $f$.
-Observe that (38) means that the relative error of $f$ with respect to its asymptotic form $g$ asymptotically vanishes.
-At a finite time, one may ask how dominant the asymptotic form is, i.e., how large the relative error of $f$ with respect to $g$ is.
-Therefore, we introduce the following notation.
-For $t\in\mathbb{R}$ such that $g(t)\neq 0$ and $\epsilon\geq 0$, we write
-$f\left(t\right)\approx g\left(t\right)\textit{\ with precision\ }\epsilon$ (39)
-when
-$\frac{\|f\left(t\right)-g\left(t\right)\|}{\|g\left(t\right)\|}\leq\epsilon.$
-Observe that (39) means that the relative error of $f$ with respect to $g$ at the time $t$ is not larger than $\epsilon$.
-The notation (39) serves to quantify how much dominant the asymptotic form $g$ of $f$ is at the finite time $t$.
-Remark 18.
-Note that, for vector or matrix functions $f(t)$ and $g(t)$,
-$f\left(t\right)\sim g\left(t\right),\ t\rightarrow+\infty,$
-implies
-$\|f\left(t\right)\|\sim\|g\left(t\right)\|,\ t\rightarrow+\infty,$
-and, for $\epsilon\geq 0$,
-$f\left(t\right)\approx g\left(t\right)\text{\ with precision\ }\epsilon$
-implies
-$\|f\left(t\right)\|\approx\|g\left(t\right)\|\text{\ with precision\ }\epsilon.$
-This follows by
-$\left|\left\|f(t)\right\|-\left\|g(t)\right\|\right|\leq\left\|f(t)-g(t)\right\|.$
-3.2. Partition of the spectrum and formula for the matrix exponential
-In this subsection, we introduce the tool for determining the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, $u\in\mathbb{C}^{n}$.
-The spectrum $\Lambda=\left\{\lambda_{1},\ldots,\lambda_{p}\right\}$ of $A$, where $\lambda_{1},\ldots,\lambda_{p}$ are the distinct eigenvalues of $A$, is partitioned by decreasing real parts (see Figure 5) in the subsets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, given by
-$\Lambda_{j}:=\{\lambda_{i_{j-1}+1},\lambda_{i_{j-1}+2},\ldots,\lambda_{i_{j}}\}$
-$\mathrm{Re}\left(\lambda_{i_{j-1}+1}\right)=\mathrm{Re}\left(\lambda_{i_{j-1}+2}\right)=\cdots=\mathrm{Re}\left(\lambda_{i_{j}}\right)=r_{j},$
-where the $q$ distinct real parts $r_{j}$, $j\in\{1,\ldots,q\}$, of the eigenvalues of $A$ satisfy $r_{1}>r_{2}>\cdots>r_{q}$.
-Figure 5. Partition of the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ by decreasing real part. The eigenvalues are marked by “x”.
-Now, we recall the formula (26) for the matrix exponential $\mathrm{e}^{tA}$:
-$\mathrm{e}^{tA}=\sum\limits_{i=1}^{p}\mathrm{e}^{\lambda_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$
-where $m_{i}$ is the ascent of the eigenvalue $\lambda_{i}$ (maximum dimension of the mini-blocks corresponding to $\lambda_{i}$ in the JCF of $A$) and the matrices $P_{il}$ are defined and studied in Section 2.
-By using the partition of the spectrum given above, we have
-$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{l=0}^{m_{i}-1}\frac{t^{l}}{l!}P_{il},$ (40)
-where $\sqrt{-1}$ denotes the imaginary unit and $\omega_{i}$ denotes the imaginary part of the eigenvalue $\lambda_{i}$.
-Now, by exchanging the two inner sums we can rewrite (40) as
-$\mathrm{e}^{tA}=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t),$ (41)
-where
-$L_{j}:=\max_{\lambda_{i}\in\Lambda_{j}}m_{i}-1,\ j\in\{1,\ldots,q\},$
-and
-$Q_{jl}(t):=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il},\ j\in\{1,\ldots,q\}\text{\ and\ }l\in\{0,\ldots,L_{j}\}.$ (42)
-The formula (41) is used to determine the asymptotic forms in (3), (4) and (5).
-3.3. Notations for the asymptotic form of $\mathrm{e}^{tA}u$
-The following notations are crucial for determining the asymptotic form of $\mathrm{e}^{tA}u$.
-For $j\in\{1,\ldots,q\}$ and $u\in\mathbb{C}^{n}$:
-  • Let
-    $\Lambda_{j}(u):=\{\lambda_{i}\in\Lambda_{j}:\alpha^{(i)}(u)\neq 0\},$
-    where $\alpha^{(i)}(u)$, defined in Section 2, is the vector of the Jordan basis components of $u$ along the generalized eigenvectors corresponding to $\lambda_{i}$.
-In other words, $\Lambda_{j}(u)$ is obtained from $\Lambda_{j}$ by including only the eigenvalues for which $u$ has a non-zero projection onto the corresponding generalized eigenspace.
-  • In the case where $\Lambda_{j}(u)\neq\emptyset$, let
-    $L_{j}(u):=\max\{l_{i}(u):\lambda_{i}\in\Lambda_{j}(u)\},$
-    where
-    $l_{i}(u)=\max\{l\in\{0,\ldots,m_{i}-1\}:P_{il}u\neq 0\}$
-    is defined and studied in Section 2, and it is such that $l_{i}(u)+1$ is the maximum index $k$, in a Jordan chain
-    $(v^{(i,j,k)})_{k=1,\ldots,m{ij}},\ j\in\{1,\ldots,d_{j}\},$
-    corresponding to the eigenvalue $\lambda_{i}$, for which $u$ has a non-zero component along the generalized eigenvector $v^{(i,j,k)}$.
-In other words, $L_{j}(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j}$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain.
-3.3.1. Indeces of dominance
-For $u\in\mathbb{C}^{n}\setminus\{0\}$:
-  • Let
-    $j(u):=\min\{j\in\left\{1,\ldots,q\right\}:\Lambda_{j}(u)\neq\emptyset\}.$
-    In other words, $r_{j(u)}$ is the rightmost real part of an eigenvalue of $A$ such that $u$ has a non-zero projection on the corresponding generalized eigenspace.
-We call $j(u)$ the primary index of dominance of $u$.
-  • Let
-    $L(u):=L_{j(u)}(u).$
-    In other words, $L(u)+1$ is the maximum index $k$, in a Jordan chain corresponding to an eigenvalue in $\Lambda_{j(u)}(u)$, for which $u$ has a non-zero component along the $k$-th generalized eigenvector of the chain.
-We call $L(u)$ the secondary index of dominance of $u$.
-Similarly to Section 2, the next two subsections, 3.4 and 3.5, can be skipped and referred back to when necessary.
-3.4. Properties of the matrices $Q_{jl}(t)$
-In our study of asymptotic forms and asymptotic condition numbers, the following properties of the matrices $Q_{jl}(t)$ in (42) are fundamental. We collect these properties in four propositions.
-The first proposition relates the indices $j(u)$ and $L_{j}(u)$ to the condition $Q_{jl}(t)u=0$.
+In the present paper, we apply this lemma to the case where $a_{1},\ldots,a_{K}$ are some of the linearly independent vectors $P_{il}u$ in (56), or some of the linearly independent matrices $P_{il}$ in (60), and $f_{1}(t),\ldots,f_{K}(t)$ are of the form $\mathrm{e}^{\sqrt{-1}\omega t}$, with $\sqrt{-1}$ the imaginary unit and $\omega\in\mathbb{R}$, and then $\|f(t)\|_{\infty}=1$.
+Appendix B Properties of the matrices $Q_{jl}(t)$
+In our study of asymptotic forms and asymptotic condition numbers, the following properties of the matrices $Q_{jl}(t)$ defined in (12) are fundamental.
+We collect these properties in four propositions.
+The first proposition relates $\Lambda_{j}(u)$ and $L_{j}(u)$, defined in Subsection 2.4, to the condition $Q_{jl}(t)u=0$.
 Proposition 40.
@@
   • 2)
-    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,j(u)-1\}$ and $l\in\{0,\ldots,L_{j}\}$;
-  • 3)
-    $Q_{jl}(t)u=0$ for $j\in\{j(u),\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
-In 2) and 3), we assume $u\neq 0$ in order to have $j(u)$ defined.
+    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
 Proof.
-For $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{j}\}$, we have
-$Q_{jl}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u$
-by recalling (42).
 Proof of point 1).
 Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)=\emptyset$ and $l\in\{0,\ldots,L_{j}\}$.
-$\Lambda_{j}(u)=\emptyset$ implies, for any $\lambda_{i}\in\Lambda_{j}$, $\alpha^{(i)}(u)=0$ and then, by Propositions 12 and 13, for any $\lambda_{i}\in\Lambda_{j}$ we have $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$.
-In particular, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, we have $P_{il}u=0$.
-Point 2) follows by point 1).
-Proof of point 3).
+$\Lambda_{j}(u)=\emptyset$ implies, for any $\lambda_{i}\in\Lambda_{j}$, $\alpha^{(i)}(u)=0$ and then, for any $\lambda_{i}\in\Lambda_{j}$ we have $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$, by Propositions 33 and 34 in Appendix A.
+In particular, we have $P_{il}u=0$ in (12), for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$.
+Proof of point 2).
 Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
-For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$.
-For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)\neq 0$, we also have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{l_{i}(u)+1,\ldots,m_{i}-1\}$ by the definition of $l_{i}(u)$, and $l\geq L_{j}(u)+1\geq l_{i}(u)+1$.
+In (12), for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$.
+Moreover, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)\neq 0$, we also have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{l_{i}(u)+1,\ldots,m_{i}-1\}$ by the definition of $l_{i}(u)$ given in Appendix A, and $l\geq L_{j}(u)+1\geq l_{i}(u)+1$.
 ∎
@@
 Proof.
-Since (see (42))
-$\left\|Q_{jl}(t)\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}\right\|\text{\ \ and\ \ }\left\|Q_{jl}(t)u\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}u\right\|$
-hold, the points 1) and 2) follow.
-Proposition 16 and Lemma 17, as applied to the linear combination
+The points 1) and 2) are trivial: by (12), we have
+[…]
+Proposition 38 and Lemma 39 in Appendix A, as applied to the linear combination
 $Q_{jl}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il},$
@@
 $Q_{jl}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}(u)\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}(u)\\ m_{i}\geq l+1\\ l_{i}(u)\geq l\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u,$
-where the second equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$ we have $P_{il}u=0$; and the third equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, $\alpha^{(i)}(u)\neq 0$ and $l_{i}(u)<l$ we have $P_{il}u=0$ by the definition of $l_{i}(u)$.
-Now, Proposition 15 and Lemma 17 imply 4).
+where the second equality holds since, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$ by Propositions 33 and 34 in Appendix A; and the third equality holds since, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, $\alpha^{(i)}(u)\neq 0$ and $l_{i}(u)<l$, we have $P_{il}u=0$ by the definition of $l_{i}(u)$ in Appendix A.
+Now, Proposition 37 and Lemma 39 in Appendix A imply 4).
 ∎
-The third proposition says how the matrices $Q_{jl}(t)$ are transformed when we replace the matrix $A$ by $-A$.
-Observe that the matrix $-A$ has opposite eigenvalues, i.e.
-eigenvalues with opposite imaginary and real parts, with respect to the matrix $A$ and the dimensions of blocks and mini-blocks in the JCF of $-A$ are the same as in the JCF of $A$ (see Proposition 2 with $z=-1$).
-Therefore, we have the same number $q$ of different real parts for the eigenvalues of $-A$ and $A$.
-Moreover, the set $\Lambda_{j}(-A)$ and the numbers $r_{j}(-A)$ and $L_{j}(-A)$, $j\in\{1,\ldots,q\}$, corresponding to $-A$ are
+The third proposition says how the matrices $Q_{jl}(t)$ are transformed when we replace the matrix $A$ by A-A.
+Observe that the matrix A-A has opposite eigenvalues, i.e., eigenvalues with opposite imaginary and real parts, with respect to the matrix $A$ and the dimensions of blocks and mini-blocks in the JCF of A-A are the same as in the JCF of $A$: see Proposition 21 in Appendix A with z=1z=-1.
+Therefore, we have the same number $q$ of different real parts for the eigenvalues of A-A and $A$.
+Moreover, the set $\Lambda_{j}(-A)$ and the numbers $r_{j}(-A)$ and $L_{j}(-A)$, $j\in\{1,\ldots,q\}$, corresponding to A-A are
 $\Lambda_{j}(-A)=-\Lambda_{q+1-j},\ r_{j}(-A)=-r_{q+1-j}\text{\ \ and\ \ }L_{j}(-A)=L_{q+1-j},$ (62)
 where $\Lambda_{q+1-j}$, $r_{q+1-j}$ and $L_{q+1-j}$ correspond to $A$.
-The indeces $j$ and $l$ for the matrices $Q_{jl}(t,-A)$ corresponding to $-A$ range over $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, respectively.
+The indices $j$ and $l$ for the matrices $Q_{jl}(t,-A)$ corresponding to A-A range over $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, respectively.
 Proposition 42.
@@
 $Q_{jl}(t,-A)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{q+1-j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ (-\omega_{i})t}P_{il}(-A).$
-Now, use Proposition 9 with $z=-1$. ∎
+Now, use Proposition 29 in Appendix A with z=1z=-1. ∎
 In the case of a real matrix $A$, the fourth proposition explains how to rewrite the expression (12) that defines $Q_{jl}(t)$ in terms of the real eigenvalues and complex conjugate pairs of eigenvalues of $A$.
-For a matrix $Z$, we denote by $\mathrm{Re}(Z)$ the matrix given by the real parts of the elements of $Z$.
+For a matrix $Z$, we denote by $\mathrm{Re}(Z)$ and $\mathrm{Im}(Z)$ the matrices given by the real parts and imaginary parts, respectively, of the elements of $Z$.
 Proposition 43.
@@
 For $\lambda_{i}\in\Lambda_{j}$ such that $\lambda_{i}$ is real, we have $\omega_{i}=0$.
-For a complex conjugate pairs $\lambda_{i_{1}},\lambda_{i_{2}}\in\Lambda_{j}$ with $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ and $\omega_{i_{1}}>0$, by Proposition 10 we have
+For a complex conjugate pair $\lambda_{i_{1}},\lambda_{i_{2}}\in\Lambda_{j}$ with $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ and $\omega_{i_{1}}>0$, by Proposition 30 in Appendix A we have
 $\mathrm{e}^{\sqrt{-1}\omega_{i_{2}}t}P_{i_{2}l}$ $=$ $\mathrm{e}^{-\sqrt{-1}\ \omega_{i_{1}}t}\overline{P_{i_{1}l}}=\overline{\mathrm{e}^{\sqrt{-1}\ \omega_{i_{1}}t}P_{i_{1}l}}$
@@
 $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ \lambda_{i}\text{ is real}\\ m_{i}\geq l+1\end{subarray}}P_{il}$
-has zero or one term $P_{il}$, which is real (see Proposition 10). Moreover, in the other sum, each term
+has zero or one term $P_{il}$, which is real: see Proposition 30 in Appendix A. Moreover, in the other sum, each term
 $\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{i}t}P_{il}\right)=\cos\omega_{i}t\cdot\mathrm{Re}(P_{il})-\sin\omega_{i}t\cdot\mathrm{Im}(P_{il})$
 is a periodic function of $t$ of period $\frac{2\pi}{\omega_{i}}$.
-3.5. The matrices $Q^{e}_{j}(t)$
-The contents of this section are used for defining in Subsection 4.6 the global asymptotic condition number of the problem (2).
+Appendix C The matrices $Q^{e}_{j}(t)$
+The contents of this section are used for defining in Subsections 4.2 and 4.3 the global asymptotic condition number $K_{\infty}(t)$ of the problem (2).
 For $j\in\{1,\ldots,q\}$, we introduce
@@
 where
-$P_{i}^{e}:=V^{(i)}_{e}W^{(i)}_{e},$
+$P_{i}^{e}:=V^{(i)}_{e}W^{(i)}_{e}$
 is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, with
 $V^{(i)}_{e}:=\left[v^{(i,1,1)}\ \cdots\ v^{(i,d_{i},1)}\right]\in\mathbb{C}^{n\times d_{i}}\text{\ \ and\ \ }W^{(i)}_{e}:=\left[\begin{array}[]{c}w^{(i,1,1)}\\ \vdots\\ w^{(i,d_{i},1)}\end{array}\right]\in\mathbb{C}^{d_{i}\times n}$
-the matrices of the (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$ (see Section 2). The superscript or subscript $e$ in the notation denotes “eigen”.
+the matrices of the (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$ (see Appendix A). The superscript, or subscript, $e$ in the notation stands for “eigen”.
 Let $U^{e}_{j}$ be the sum of the eigenspaces corresponding to eigenvalues in $\Lambda_{j}$.
-Observe that the linear operator $Q^{e}_{j}(t)$ maps $\mathbb{C}^{n}$ in the subspace $U^{e}_{j}$.
-Therefore, $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$.
+Observe that the linear operators $P_{i}^{e}$, $\lambda_{i}\in\Lambda_{j}$, and $Q^{e}_{j}(t)$ map $\mathbb{C}^{n}$ into the subspace $U^{e}_{j}$.
+Therefore, $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$ and $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$.
+Remark 45.
+When $\Lambda_{j}$ consists of a real eigenvalue $\lambda_{i}$, we have
+$Q^{e}_{j}(t)|_{U^{e}_{j}}=P_{i}^{e}|_{U^{e}_{j}}=I_{U^{e}_{j}}.$
+The next proposition establishes that $P_{i}^{e}$ and $Q_{j}^{e}(t)$ coincide with $P_{i0}$ and $Q_{j0}(t)$ when restricted to $U_{j}^{e}$.
+$P_{i0}$ is defined in (46) of Appendix A (see also Subsection A.3.1 in Appendix A) and $Q_{j0}$ is defined in (12).
+Proposition 46.
+Let $j\in\{1,\ldots,q\}$. We have $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$, $\lambda_{i}\in\Lambda_{j}$, and $Q^{e}_{j}(t)|_{U^{e}_{j}}=Q_{j0}(t)|_{U^{e}_{j}}$.
+Proof.
+For $\lambda_{i}\in\Lambda_{j}$, $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$ follows by Subsection A.3.1 in Appendix A and the fact that, for $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$, for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$.
+$Q^{e}_{j}(t)|_{U^{e}_{j}}=Q_{j0}(t)|_{U^{e}_{j}}$ follows by $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$, $\lambda_{i}\in\Lambda_{j}$.
+∎
 The next proposition states that the restriction $Q^{e}_{j}(t)|_{U^{e}_{j}}$ is invertible.
@@
 We have
-$Q^{e}_{j}(-t)Q^{e}_{j}(t)$ $=$ $\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{-\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)$
-$=$ $\sum\limits_{\lambda_{i}\in\Lambda_{j}}P^{e}_{i}$
-and then
+$Q^{e}_{j}(-t)Q^{e}_{j}(t)$ $=$ $\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{-\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)\left(\sum\limits_{\lambda_{k}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{k}t}P^{e}_{k}\right)$
+$=$ $\sum\limits_{\lambda_{i},\lambda_{k}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ (-\omega_{i}+\omega_{k})t}P^{e}_{i}P^{e}_{k}=\sum\limits_{\lambda_{i}\in\Lambda_{j}}P^{e}_{i},$
+since $P^{e}_{i}P^{e}_{k}=0$ for $\lambda_{i}\neq\lambda_{k}$ and $P^{e}_{i}P^{e}_{k}=P^{e}_{i}$ for $\lambda_{i}=\lambda_{k}$. Thus,
 $Q_{j}^{e}(-t)|_{U_{j}^{e}}Q_{j}^{e}(t)|_{U_{j}^{e}}=I_{U_{j}^{e}}.$
 Remark 48.
-  • 1.
 As a consequence of the previous proposition, we have
 $\min\limits_{\begin{subarray}{c}\widehat{u}\in U^{e}_{j}\\ \left\|\widehat{u}\right\|=1\end{subarray}}\left\|Q^{e}_{j}(t)\widehat{u}\right\|=\frac{1}{\left\|Q^{e}_{j}(-t)|_{U^{e}_{j}}\right\|}.$
-  • 2.
-    In the case where $\Lambda_{j}$ consists of a real eigenvalue $\lambda_{i}$, we have
-    $Q^{e}_{j}(t)|_{U^{e}_{j}}=P_{i}^{e}|_{U^{e}_{j}}=I_{U^{e}_{j}}.$
-Next proposition considers the linear operators $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$, $\lambda_{i}\in\Lambda_{j}$.
+Next proposition considers the linear independence of the linear operators $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$, $\lambda_{i}\in\Lambda_{j}$.
 Proposition 49.
@@
 Consider a zero linear combination of the linear operators (64):
-$0=\sum\limits_{i=1}^{p}c_{i}P_{i}^{e}|_{U^{e}_{j}}.$
+$0=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i}^{e}|_{U^{e}_{j}}.$
 Consider $u\in U^{e}_{j}$ such that $\alpha_{ij^{\prime}1}(u)\neq 0$ for $\lambda_{i}\in\Lambda_{j}$ and $j^{\prime}\in\{1,\ldots,d_{i}\}$.
 Since $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$.
-Thus $l_{i}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, by Proposition 14.
-Since (recall Subsection 2.3.1)
-$P_{i}^{e}u=P_{i0}u,\ \lambda_{i}\in\Lambda_{j},$
+Thus $l_{i}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, by Proposition 36 in Appendix A.
 We have the zero linear combination
 $0=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i}^{e}u=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i0}u=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\sum\limits_{l=0}^{l_{i}(u)=0}c_{il}P_{il}u,$
-where we set $c_{i0}=c_{i}$, $i\in\{1,\ldots,p\}$. By using Proposition 15, we obtain $c_{i}=c_{i0}=0$, $\lambda_{i}\in\Lambda_{j}$. ∎
+where the second equality follows by Proposition 46 in this appendix, and we set $c_{i0}=c_{i}$, $\lambda_{i}\in\Lambda_{j}$.
+By using Proposition 37 in Appendix A, we obtain $c_{i}=c_{i0}=0$, $\lambda_{i}\in\Lambda_{j}$.
+∎
 Remark 50.
-As a consequence of the previous proposition and Lemma 17, we obtain
+As a consequence of the previous Proposition 49 and Lemma 39 in Appendix A, we obtain
 $\inf\limits_{t\in\mathbb{R}}\left\|Q^{e}_{j}(t)|_{U^{e}_{j}}\right\|>0.$
-3.6. The asymptotic form of $\mathrm{e}^{tA}$
-By looking at the formula (41), we immediately identify
-$\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$ (46)
-as the asymptotic form of $\mathrm{e}^{tA}$, since it is the dominant term as $t\rightarrow+\infty$: it is the term with the highest order of infinite.
-Next proposition precises that (46) is indeed an asymptotic form of $\mathrm{e}^{tA}$ in our definition (38) of asymptotic form and quantifies its dominance at a finite time.
-Proposition 28.
-We have
-$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
-with precision
-$\epsilon(t):=\sum\limits_{l=0}^{L_{1}-1}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{1l}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}+\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\sum\limits_{l=0}^{L_{j}}\frac{L_{1}!}{l!}t^{l-L_{1}}\frac{\left\|Q_{jl}(t)\right\|}{\left\|Q_{1L_{1}}(t)\right\|}.$
-Moreover, we have
-$\epsilon(t)\rightarrow 0,\ t\rightarrow+\infty,$ (47)
-and then
-$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t),\ t\rightarrow+\infty.$
-Proof.
-By (41), we can write
-$\mathrm{e}^{tA}$ $=$ $\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)$
-$=$ $\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)+\mathrm{e}^{r_{1}t}\sum\limits_{l=0}^{L_{1}-1}\frac{t^{l}}{l!}Q_{1l}(t)+\sum\limits_{j=2}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t).$
-The first part of the proposition regarding $\approx$ follows. By 1) and 3) in Proposition 20, we obtain (47) and then the second part regarding $\sim$ follows. ∎
-Remark 29.
-If $A$ is diagonalizable, then we have
-$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}Q_{10}(t)$
-with precision
-$\epsilon(t)=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{\left\|Q_{j0}(t)\right\|}{\left\|Q_{10}(t)\right\|}$
-and
-$\mathrm{e}^{tA}\sim\mathrm{e}^{r_{1}t}Q_{10}(t),\ t\rightarrow+\infty,$
-where
-$Q_{j0}(t)=\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{i0},\ j\in\{1,\ldots,q\},$
-with $P_{i0}$ the projection onto the eigenspace corresponding to $\lambda_{i}$.
-3.7. The asymptotic form of $\mathrm{e}^{tA}u$
-Let $u\in\mathbb{C}^{n}\setminus\{0\}$. When we use the formula (41) for $\mathrm{e}^{tA}$ in $\mathrm{e}^{tA}u$, we obtain
-$\mathrm{e}^{tA}u=\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u.$
-By 2) and 3) in Proposition 19 we identify
-$\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$ (48)
-as the asymptotic form of $\mathrm{e}^{tA}u$, since it is the dominant term as $t\rightarrow+\infty$.
-Next proposition precises that (48) is indeed an asymptotic form in our definition and quantifies its dominance.
-Proposition 30.
-For $u\in\mathbb{C}^{n}\setminus\{0\}$, we have
-$\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u$
-with precision
-$\epsilon(t,u)$ ::=:= $\sum\limits_{l=0}^{L\left(u\right)-1}\frac{L\left(u\right)!}{l!}t^{l-L\left(u\right)}\frac{\left\|Q_{j(u)l}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}$
-$+\sum\limits_{\begin{subarray}{c}j=j\left(u\right)+1\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\sum\limits_{l=0}^{L_{j}(u)}\frac{L(u)!}{l!}t^{l-L(u)}\frac{\left\|Q_{jl}(t)u\right\|}{\left\|Q_{j(u)L(u)}(t)u\right\|}.$
-Moreover, we have
-$\epsilon(t,u)\rightarrow 0,\ t\rightarrow+\infty,$ (49)
-and then
-$\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u,\ t\rightarrow+\infty.$
-Proof.
-By (41) and Proposition 19, we can write
-$\mathrm{e}^{tA}u$ $=$ $\sum\limits_{j=1}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u=\sum\limits_{j=j(u)}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}}\frac{t^{l}}{l!}Q_{jl}(t)u$
-$=$ $\sum\limits_{\begin{subarray}{c}j=j(u)\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}(u)}\frac{t^{l}}{l!}Q_{jl}(t)u$
-$=$ $\mathrm{e}^{r_{j(u)}t}\frac{t^{L(u)}}{L(u)!}Q_{j(u)L(u)}(t)u+\mathrm{e}^{r_{j(u)}t}\sum\limits_{l=0}^{L(u)-1}\frac{t^{l}}{l!}Q_{j(u)l}(t)u$
-$+\sum\limits_{\begin{subarray}{c}j=j(u)+1\\ \Lambda_{j}(u)\neq\emptyset\end{subarray}}^{q}\mathrm{e}^{r_{j}t}\sum\limits_{l=0}^{L_{j}(u)}\frac{t^{l}}{l!}Q_{jl}(t)u.$
-The first part of the proposition regarding $\approx$ follows. By 2) and 4) in Proposition 20, we obtain (49) and then the second part regarding $\sim$ follows. ∎
-Remark 31.
-  • 1.
-    The asymptotic form (48) of $\mathrm{e}^{tA}u$ is determined by the primary and secondary indices of dominance of $u$.
-In particular, the smaller the primary index of dominance $j(u)$, the higher the order of infinity of (48), and for a fixed primary index, the larger the secondary index of dominance $L(u)$, the higher the order of infinity of (48).
-  • 2.
-    The contents of Propositions 28 and 30 are deeper than what might be apparent at first glance.
-We are not simply stating the obvious facts that (46) and (48) are the dominant terms in $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$, respectively.
-Rather, we are asserting that they represent the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ according to our definition (38).
-In other words, the relative errors of $\mathrm{e}^{tA}$ and $\mathrm{e}^{tA}u$ with respect to (46) and (48), respectively, approach zero asymptotically.
-Proving these deeper conclusions requires Propositions 15 and 16 and Lemma 17, which were developed through the work in Section 2.
-  • 3.
-    If $A$ is diagonalizable, then we have
-    $\mathrm{e}^{tA}u\approx\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u$
-    with precision
-    $\epsilon(t,u)=\sum\limits_{j=j\left(u\right)+1}^{q}\mathrm{e}^{\left(r_{j}-r_{j\left(u\right)}\right)t}\frac{\left\|Q_{j0}(t)u\right\|}{\left\|Q_{j(u)0}(t)u\right\|}$
-    and
-    $\mathrm{e}^{tA}u\sim\mathrm{e}^{r_{j\left(u\right)}t}Q_{j(u)0}(t)u,\ t\rightarrow+\infty.$
-4. The asymptotic behavior of the condition numbers
-This section forms the core of our analysis. Building on the preparatory work of the previous two sections, we can now easily describe the asymptotic behavior of the condition numbers.
-4.1. The asymptotic directional condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$
-We set
-$j^{\ast}:=j\left(\widehat{y}_{0}\right)=j\left(y_{0}\right)\text{\ \ and\ \ }L^{\ast}:=L\left(\widehat{y}_{0}\right)=L\left(y_{0}\right)$
-as well as
-$j^{\ast\ast}:=j\left(\widehat{z}_{0}\right)\text{\ \ and\ \ }L^{\ast\ast}:=L\left(\widehat{z}_{0}\right),$
-i.e. $j^{\ast}$ and $j^{\ast\ast}$ are the primary indeces of dominance, and $L^{\ast}$ and $L^{\ast\ast}$ are the secondary indeces of dominance, of $y_{0}$ and $\widehat{z}_{0}$, respectively.
-The next theorem describes the long-time behavior of the directional condition number $K\left(t,y_{0},\widehat{z}_{0}\right)$.
-Theorem 32.
-We have
-$K\left(t,y_{0},\widehat{z}_{0}\right)\approx K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),$
-where
-$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right):=\frac{L^{\ast}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{j^{\ast}}\right)t}t^{L^{\ast\ast}-L^{\ast}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
-with precision
-$\frac{\epsilon(t,\widehat{z}_{0})+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})},$
-whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t,u)$, $u=\widehat{z}_{0},\widehat{y}_{0}$, defined in Proposition 30). Moreover, we have
-$K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\rightarrow+\infty.$
-Proof.
-Proposition 30 says what are the asymptotic forms of $\mathrm{e}^{tA}\widehat{z}_{0}$ and $\mathrm{e}^{tA}\widehat{y}_{0}$ in (3) and how much they are dominant at a finite time $t$: we have
-$\mathrm{e}^{tA}\widehat{z}_{0}\approx\mathrm{e}^{r_{j^{\ast\ast}}t}\frac{t^{L^{\ast\ast}}}{L^{\ast\ast}!}Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}$
-with precision $\epsilon(t,\widehat{z}_{0})$ and
-$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
-with precision $\epsilon(t,\widehat{y}_{0})$.
-The first part of the theorem regarding $\approx$ follows by using Remark 18.
-The second part regarding $\sim$ follows since $\epsilon(t,\widehat{y}_{0})\rightarrow 0$ and $\epsilon(t,\widehat{z}_{0})\rightarrow 0$, $t\rightarrow+\infty$.
-∎
-We define the function
-$t\rightarrow K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right),\ t\in\mathbb{R},$
-as the asymptotic directional condition number of the problem (2).
-Remark 33.
-  • 1.
-    The asymptotic directional condition number of the problem (2):
-    • is bounded and away from zero as $t$ varies (remind 2) and 4) in Proposition 20) if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}=L^{\ast\ast}$;
-    • decays polynomially to zero, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}>L^{\ast\ast}$;
-    • diverges polynomially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}=j^{\ast\ast}$ and $L^{\ast}<L^{\ast\ast}$;
-    • decays exponentially to zero, as $t\rightarrow+\infty$, if $j^{\ast}<j^{\ast\ast}$;
-    • diverges exponentially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}>j^{\ast\ast}$.
-    Hence, whether the asymptotic condition number decreases to zero, diverges to infinity, or exhibits different behavior depends on which between $y_{0}$ and $\widehat{z}_{0}$ is more dominant.
-This is determined by which of $y_{0}$ and $\widehat{z}_{0}$ possesses the smaller primary index of dominance, or, in the case of equal primary indices, the larger secondary index of dominance.
-  • 2.
-    The generic case for $y_{0}$ and $\widehat{z}_{0}$ is $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$. In this generic case, we have
-    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
-4.2. The asymptotic condition number $K\left(t,y_{0}\right)$
-The next theorem describes the long-time behavior of the condition number $K\left(t,y_{0}\right)$, worst condition number $K(t,y_{0},\widehat{z}_{0})$ by varying $\widehat{z}_{0}$.
-Theorem 34.
-We have
-$K\left(t,y_{0}\right)\approx K_{\infty}\left(t,y_{0}\right),$
-where
-$K_{\infty}\left(t,y_{0}\right):=\frac{L^{\ast}!}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{j^{\ast}}\right)t}t^{L_{1}-L^{\ast}}\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|},$
-with precision
-$\frac{\epsilon(t)+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
-whenever $\epsilon(t,\widehat{y}_{0})<1$ ($\epsilon(t)$ defined in Proposition 28 and $\epsilon(t,\widehat{y}_{0})$ defined in Proposition 30). Moreover, we have
-$K\left(t,y_{0}\right)\sim K_{\infty}\left(t,y_{0}\right),\ t\rightarrow+\infty.$
-Proof.
-Proposition 28 says what is the asymptotic form of $\mathrm{e}^{tA}$ and how much it is dominant: we have
-$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
-with precision $\epsilon(t)$. Proposition 30 says what is the asymptotic form of $\mathrm{e}^{tA}\widehat{y}_{0}$ and how much it is dominant: we have
-$\mathrm{e}^{tA}\widehat{y}_{0}\approx\mathrm{e}^{r_{j^{\ast}}t}\frac{t^{L^{\ast}}}{L^{\ast}!}Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}$
-with precision $\epsilon(t,\widehat{y}_{0})$.
-The first part of the theorem regarding $\approx$ follows by using Remark 18.
-The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,\widehat{y}_{0})\rightarrow 0$, $t\rightarrow+\infty$.
-∎
-We define the function
-$t\rightarrow K_{\infty}(t,y_{0}),\ t\in\mathbb{R},$
-as the asymptotic condition number of the problem (2).
-Remark 35.
-  • 1.
-    The asymptotic condition number of the problem (2):
-    • is bounded and away from zero as $t$ varies if $j^{\ast}=1$ and $L^{\ast}=L_{1}$;
-    • diverges polynomially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}=1$ and $L^{\ast}<L_{1}$;
-    • diverges exponentially to infinite, as $t\rightarrow+\infty$, if $j^{\ast}>1$.
-    Therefore, the asymptotic condition number does not diverge to infinity if and only if $y_{0}$ is as dominant as possible, meaning that $y_{0}$ has the smallest possible primary index of dominance (i.e.
-$j^{\ast}=1$) and simultaneously the largest possible secondary index of dominance (i.e.
-$L^{\ast}=1$).
-  • 2.
-    The generic case for $y_{0}$ is $j^{\ast}=1$ and $L^{\ast}=L_{1}$. In this generic case, we have
-    $K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1L_{1}}(t)\right\|}{\left\|Q_{1L_{1}}(t)\widehat{y}_{0}\right\|}.$
-4.3. Is $K_{\infty}\left(t,y_{0}\right)$ the worst $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$?
-By definition, the condition number $K\left(t,y_{0}\right)$ is the worst condition number $K\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$.
-Hence, an interesting question is the following.
-Does this fact hold asymptotically as $t\rightarrow+\infty$?
-Specifically, is the asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$ the worst asymptotic condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying $\widehat{z}_{0}$?
-In other words, do
-[…]
-commute? The answer is YES and it is given by the next theorem, which considers the ratio
-$\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=\frac{L_{1}!}{L^{\ast\ast}!}\mathrm{e}^{\left(r_{j^{\ast\ast}}-r_{1}\right)t}t^{L^{\ast\ast}-L_{1}}\frac{\left\|Q_{j^{\ast\ast}L^{\ast\ast}}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1L_{1}}(t)\right\|}$ (50)
-independent of $y_{0}$.
-Theorem 36.
-We have
-$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=1.$ (51)
-In particular:
-  • a)
-    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}<1$ or $j^{\ast\ast}=1$ and $L^{\ast\ast}<L_{1}$, we have
-    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=0.$
-  • b)
-    For any direction of perturbation $\widehat{z}_{0}$ such that $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, we have
-    $\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}\leq 1.$
-  • c)
-    There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a sequence $\{t_{m}\}_{m}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
-    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m},y_{0}\right)}=1.$
-Proof.
-Points a), b) and c) imply (51). Points a) and b) immediately follows by (50). Now, we prove c).
-Consider a sequence $\{t_{k}\}_{k}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{z}_{0k}\}_{k}$ such that $\widehat{z}_{0k}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0k}\right\|=1$ and
-$\left\|Q_{1L_{1}}(t_{k})\right\|=\left\|Q_{1L_{1}}(t_{k})\widehat{z}_{0k}\right\|.$
-By the compactness of the unit sphere in $\mathbb{C}^{n}$, there exists a subsequence $\{\widehat{z}_{0k_{m}}\}_{m}$ of $\{\widehat{z}_{0k}\}_{k}$ converging to some $\widehat{z}_{0\infty}\in\mathbb{C}^{n}$ with $\left\|\widehat{z}_{0\infty}\right\|=1$.
-By Proposition 19 (points 2) and 3)) and Proposition 20 (points 1) and 3)), we obtain $j\left(\widehat{z}_{0\infty}\right)=1$ and $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$.
-In fact, for any index $m$, we have
-$\left|\ \left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|-\left\|Q_{1L_{1}}(t_{k_{m}})\right\|\ \right|\leq\sup\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|,$
-where the right-hand side goes to zero as $m\rightarrow\infty$. Therefore, there exists an index $m$ such that
-$\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|\geq\frac{1}{2}\inf\limits_{t\in\mathbb{R}}\left\|Q_{1L_{1}}(t)\right\|,$
-where the right-hand side is positive.
-Since $Q_{1L_{1}}(t_{k_{m}})$ $\widehat{z}_{0\infty}\neq 0,$ we cannot have $j\left(\widehat{z}_{0\infty}\right)>1$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$.
-Hence, $j\left(\widehat{z}_{0\infty}\right)=1$.
-Once again since $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\neq 0,$ we cannot have $L_{1}\left(\widehat{z}_{0\infty}\right)<L_{1}$, otherwise $Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}=0$.
-Hence, $L_{1}\left(\widehat{z}_{0\infty}\right)=L_{1}$.
-By using as direction of perturbation $\widehat{z}_{0\infty}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$. So
-$\left|\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}-1\right|$ $=$ $\left|\frac{\left\|Q_{1L_{1}}(t_{k_{m}})\widehat{z}_{0\infty}\right\|}{\left\|Q_{1L_{1}}(t_{k_{m}})\right\|}-1\right|$
-$\leq$ $\left\|\widehat{z}_{0\infty}-\widehat{z}_{0k_{m}}\right\|.$
-We conclude that
-$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{k_{m}},y_{0},\widehat{z}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}},y_{0}\right)}=1.$
-Remark 37.
-In the case where $\Lambda_{1}$ consists of a real eigenvalue, the point c) is modified to:
-  • $c^{\prime}$)
-    There exists a direction of perturbation $\widehat{z}_{0}$, independent of $y_{0}$ and with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, such that
-    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right).$ (52)
-In fact, in this case $Q_{1L_{1}}(t)=Q_{1L_{1}}$ is independent of $t$ and so there exists $\widehat{z}_{0}\in\mathbb{C}^{n}$, $\left\|\widehat{z}_{0}\right\|=1$, independent of $t$ such that
-$\left\|Q_{1L_{1}}\right\|=\left\|Q_{1L_{1}}\widehat{z}_{0}\right\|.$
-By using as direction of perturbation $\widehat{z}_{0}$, we have $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ and then we obtain (52) by (50).
-4.4. The asymptotic global condition number $K_{\infty}^{+}\left(t,\right)$
-Next theorem describes the long-time behavior of the global condition number $K\left(t\right)$, worst condition number $K\left(t,y_{0}\right)$ by varying $y_{0}$.
-Theorem 38.
-We have
-$K\left(t\right)\approx K_{\infty}^{+}(t),$
-where
-$K_{\infty}^{+}\left(t\right):=\frac{1}{L_{1}!L_{q}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}+L_{q}}\left\|Q_{1L_{1}}(t)\right\|\ \cdot\left\|Q_{qL_{q}}(-t)\right\|$ (53)
-with precision
-$\epsilon(t)+\epsilon(t,-A)+\epsilon(t)\epsilon(t,-A),$
-where $\epsilon(t,-A)$ is $\epsilon(t)$ for the matrix $-A$. Moreover, we have
-$K\left(t\right)\sim K^{+}_{\infty}(t),\ t\rightarrow+\infty.$
-Proof.
-Propositions 28 says what are the asymptotic forms of $\mathrm{e}^{tA}$ and $\mathrm{e}^{-tA}$ and how much they are dominant: we have
-$\mathrm{e}^{tA}\approx\mathrm{e}^{r_{1}t}\frac{t^{L_{1}}}{L_{1}!}Q_{1L_{1}}(t)$
-with precision $\epsilon(t)$ and
-$\mathrm{e}^{-tA}=\mathrm{e}^{t(-A)}\approx\mathrm{e}^{-r_{q}t}\frac{t^{L_{q}}}{L_{q}!}(-1)^{L_{q}}Q_{qL_{q}}(-t)$
-with precision $\epsilon(t,-A)$.
-For the latter, remind (43) and Proposition 21.
-The first part of the theorem regarding $\approx$ follows by using Remark 18.
-The second part regarding $\sim$ follows since $\epsilon(t)\rightarrow 0$ and $\epsilon(t,-A)\rightarrow 0$, $t\rightarrow+\infty$.
-∎
-We define the function
-$t\rightarrow K^{+}_{\infty}(t),\ t\in\mathbb{R},$
-as the asymptotic global condition number of the problem (2).
-4.5. Is $K^{+}_{\infty}\left(t\right)$ the worst $K_{\infty}\left(t,y_{0}\right)$?
-By definition, the condition number $K\left(t\right)$ is the worst condition number $K\left(t,y_{0}\right)$, by varying $y_{0}$.
-Hence, as in Subsection 4.3, an interesting question is the following.
-Does this fact hold asymptotically as $t\rightarrow+\infty$?
-Specifically, is the asymptotic condition number $K^{+}_{\infty}\left(t\right)$ the worst asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$?
-In other words, do
-[…]
-commute? Unlike Subsection 4.3, here the answer is NO and it is given by the next theorem, which considers
-$K_{\infty}\left(t\right):=\frac{1}{L_{1}!}\mathrm{e}^{\left(r_{1}-r_{q}\right)t}t^{L_{1}}\left\|Q_{1L_{1}}(t)\right\|\cdot\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|,$ (54)
-where $Q^{e}_{q}(-t)|_{U^{e}_{q}}$ is defined in Subsection 3.5, and the ratio
-$\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=L^{\ast}!\mathrm{e}^{\left(r_{q}-r_{j^{\ast}}\right)t}t^{-L^{\ast}}\frac{1}{\left\|Q_{j^{\ast}L^{\ast}}(t)\widehat{y}_{0}\right\|\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}.$ (55)
-Theorem 39.
-We have
-$\max\limits_{\begin{subarray}{c}y_{0}\in\mathbb{C}^{n}\\ y_{0}\neq 0\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=1.$ (56)
-In particular:
-  • a)
-    For any initial value $y_{0}$ such that $j^{\ast}<q$ or $j^{\ast}=q$ and $L^{\ast}>0$, we have
-    $\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}=0.$
-  • b)
-    For any initial value $y_{0}$ such that $j^{\ast}=q$ and $L^{\ast}=0$, we have
-    $\frac{K_{\infty}\left(t,y_{0}\right)}{K_{\infty}\left(t\right)}\leq 1.$
-  • c)
-    There exist an initial value $y_{0}$, with $j^{\ast}=q$ and $L^{\ast}=0$, and a sequence $\{t_{m}\}_{m}$ with $t_{m}\rightarrow+\infty$, as $m\rightarrow\infty$, such that
-    $\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},y_{0}\right)}{K_{\infty}\left(t_{m}\right)}=1.$
-  • d)
-    For the initial value $y_{0}$ at point c), there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}_{s}$ of the sequence $\{t_{m}\}_{m}$ at point c), such that
-    $\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=1.$
-Proof.
-Points a), b) and c) imply (56). Point a) follows by (55) (remind point 4 in Proposition 20 and Remark 27).
-Point b) follows by point 1 in Remark 25 and by observing that, for such $y_{0}$, we have $y_{0}\in U^{e}_{q}$ and then
-$Q_{q0}(t)\widehat{y}_{0}=Q^{e}_{q}(t)\widehat{y}_{0}.$
-For the point c), we proceed as in the proof of Theorem 36.
-Consider a sequence $\{t_{k}\}_{k}$ such that $t_{k}\rightarrow+\infty$, $k\rightarrow\infty$, and a sequence $\{\widehat{y}_{0k}\}_{k}$, where $\widehat{y}_{0k}\in U^{e}_{q}$ and $\left\|\widehat{y}_{0k}\right\|=1$, such that
-$\left\|Q_{q0}(t_{k})\widehat{y}_{0k}\right\|=\left\|Q^{e}_{q}(t_{k})\widehat{y}_{0k}\right\|=\frac{1}{\left\|Q^{e}_{q}(-t_{k})|_{U^{e}_{q}}\right\|}$ (57)
-(recall point 1 in Remark 25).
-There exists a subsequence $\{\widehat{y}_{0k_{m}}\}_{m}$ of $\{\widehat{y}_{0k}\}_{k}$ converging to some $\widehat{y}_{0\infty}\in U^{e}_{q}$ with $\left\|\widehat{y}_{0\infty}\right\|=1$.
-By using $\widehat{y}_{0\infty}$ as initial value, we have $j^{\ast}=q$ and $L^{\ast}=0$ (we have $l_{i}(\widehat{y}_{0\infty})=0$ for any $\lambda_{i}\in\Lambda_{q}$, see Proposition 14).
-Therefore,
-$\left|\frac{K_{\infty}\left(t_{k_{m}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{k_{m}}\right)}-1\right|$ $=$ $\left|\frac{1}{\left\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\right\|\left\|Q^{e}_{q}(-t_{k_{m}})|_{U^{e}_{q}}\right\|}-1\right|$
-$=$ $\left|\frac{\|Q_{q0}(t_{k_{m}})\widehat{y}_{0k_{m}}\|}{\|Q_{q0}(t_{k_{m}})\widehat{y}_{0\infty}\|}-1\right|$
-[…]
-[…]
-We conclude that
-$\lim\limits_{m\rightarrow\infty}\frac{K_{\infty}\left(t_{m},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{m}\right)}=1.$
-For the point d), repeat the proof of point c) in Theorem 36 with the sequence $\{t_{k}\}_{k}$ replaced by the sequence $\{t_{m}\}_{m}$ above in point c) of this proof.
-In this way, we show that there exists a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$, and a subsequence $\{t_{m_{s}}\}_{s}$ of the sequence $\{t_{m}\}_{m}$, such that
-$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1$
-and then
-$\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}}\right)}=\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}{K_{\infty}\left(t_{m_{s}}\right)}\cdot\lim\limits_{s\rightarrow\infty}\frac{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty},\widehat{z}_{0}\right)}{K_{\infty}\left(t_{m_{s}},\widehat{y}_{0\infty}\right)}=1.$
-Remark 40.
-In the case where $\Lambda_{q}$ consists of real eigenvalue, by recalling (55) and point 2 in Remark 25, the point c) is modified to:
-  • $c^{\prime}$)
-    For any initial value $y_{0}\in U^{e}_{q}$, we have
-    $K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
-In addition, in the case where $\Lambda_{1}$ consists of a real eigenvalue, the point d) is modified to:
-  • $d^{\prime}$)
-    There exist a direction of perturbation $\widehat{z}_{0}$ with $j^{\ast\ast}=1$ and $L^{\ast\ast}=L_{1}$ such that, for any initial value $y_{0}\in U^{e}_{q}$, we have
-    $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(t\right).$
-See Remark 37.
-4.6. The global asymptotic condition number $K_{\infty}\left(t\right)$
-Theorem 39 says that $K_{\infty}\left(t\right)$ in (54), not $K^{+}_{\infty}\left(t\right)$ in (53), is the worst asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$, by varying $y_{0}$, namely the worst asymptotic condition $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$, by varying both $y_{0}$ and $\widehat{z}_{0}$.
-We define the function
-$t\rightarrow K_{\infty}(t),\ t\in\mathbb{R},$
-as the global asymptotic condition number of the problem (2).
-Remark 41.
-  • 1.
-    Observe that $K_{\infty}(t)$ is the global asymptotic condition number, i.e.
-the worst-case asymptotic condition number $K_{\infty}(t,y_{0})$ as $y_{0}$ varies.
-In contrast, $K_{\infty}^{+}(t)$ is the asymptotic global condition number, i.e.
-the asymptotic form of the worst-case condition number $K(t,y_{0})$ as $y_{0}$ varies.
-  • 2.
-    Both the global asymptotic condition number and the asymptotic global condition number of the problem (2):
-    • are bounded and away from zero as $t$ varies if $q=1$ and $L_{1}=0$;
-    • diverge polynomially to infinite, as $t\rightarrow+\infty$, if $q=1$ and $L_{1}>0$;
-    • diverge exponentially to infinite, as $t\rightarrow+\infty$, if $q>1$.
-    Therefore, the they do not diverges to infinite if and only if all the eigenvalues of $A$ have the same real part, i.e. they lie in a vertical line of the complex plane, and are simple.
-Observe that $K^{+}_{\infty}(t)$ can be significantly larger than $K_{\infty}(t)$: we have
-$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=L_{q}!t^{-L_{q}}\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{qL_{q}}(-t)\right\|}$
-and then
-$\lim\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=0$
-for $L_{q}>0$ and
-$\frac{K_{\infty}\left(t\right)}{K^{+}_{\infty}\left(t\right)}=\frac{\left\|Q^{e}_{q}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}=\frac{\left\|Q_{q0}(-t)|_{U^{e}_{q}}\right\|}{\left\|Q_{q0}(-t)\right\|}\leq 1$
-for $L_{q}=0$.
-In the next example, we illustrate the difference between $K^{+}_{\infty}(A)$ and $K_{\infty}(A)$.
-Example 42.
-Consider the one Jordan block matrix
-$A_{1}=V\left[\begin{array}[]{rr}1&1\\ 0&1\end{array}\right]W$
-and the diagonalizable matrix
-$A_{2}=V\left[\begin{array}[]{rr}1&0\\ 0&-1\end{array}\right]W,$
-where
-$V=\left[\begin{array}[]{rr}1&\frac{1}{2}\\ 1&1\end{array}\right]\text{\ \ and\ \ }W=V^{-1}=2\left[\begin{array}[]{rr}1&-\frac{1}{2}\\ -1&1\end{array}\right].$
-For the matrix $A_{1}$, we have
-$\Lambda=\{1\},\ q=1\text{\ \ and\ \ }L_{1}=L_{q}=1$
-and then
-$K_{\infty}(t)=t\left\|Q_{11}(t)\right\|\left\|Q^{e}_{1}(-t)|_{U^{e}_{1}}\right\|$
-$K^{+}_{\infty}(t)=t^{2}\left\|Q_{11}(t)\right\|\left\|Q_{11}(-t)\right\|$
-where
-$Q_{11}(t)=Q_{11}(-t)=P_{11}=v^{(1,1,1)}w^{(1,1,2)}=\left[\begin{array}[]{rr}-2&2\\ -2&2\end{array}\right]$
-(remind point 2 in Remark 8)
-$Q^{e}_{1}(-t)|_{U^{e}_{1}}=I|_{U^{e}_{1}}$
-(remind point 2 in Remark 25).
-Thus, for the Euclidean norm as vector norm, we have, since $\|P_{11}\|_{2}=4$,
-$K_{\infty}(t)=4t$
-$K^{+}_{\infty}(t)=16t^{2}.$
-For the matrix $A_{2}$, we have
-$\Lambda=\{1,-1\},\ q=2\text{\ \ and\ \ }L_{1}=L_{q}=0$
-and then
-$K_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q^{e}_{2}(-t)|_{U^{e}_{2}}\right\|$
-$K^{+}_{\infty}(t)=\mathrm{e}^{2t}\left\|Q_{10}(t)\right\|\left\|Q_{20}(-t)\right\|,$
-where
-$Q_{10}(t)=P_{10}=v^{(1,1,1)}w^{(1,1,1)}=\left[\begin{array}[]{rr}2&-1\\ 2&-1\end{array}\right]$
-$Q_{20}(-t)=P_{20}=v^{(2,1,1)}w^{(2,1,1)}=\left[\begin{array}[]{rr}-1&1\\ -2&2\end{array}\right]$
-$Q^{e}_{2}(-t)|_{U^{e}_{2}}=I|_{U^{e}_{2}}.$
-Thus, for the Euclidean norm as vector norm, we have, since $\|P_{10}\|_{2}=\|P_{20}\|_{2}=\sqrt{10}$,
-$K_{\infty}(t)=\sqrt{10}\mathrm{e}^{2t}$
-$K^{+}_{\infty}(t)=10\mathrm{e}^{2t}.$
-In Figure 6 for the matrix $A_{1}$, we see in logarithmic scale the condition numbers $K(t)$ (blu dashed line) , $K^{+}_{\infty}(t)$ (red dashed line) and $K_{\infty}(t)$ (red dashed line), $t\in[0,100]$.
-From the beginning, $K(t)$ is not distinguishable from its asymptotic version $K^{+}_{\infty}(t)$.
-Figure 6. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{1}$.
-The asymptotic global condition number $K^{+}_{\infty}(t)$ is unrelated to the worst asymptotic behavior of the condition numbers $K(t,y_{0},\widehat{z}_{0})$.
-In fact, in Figure 6 we also see, for each $\tau\in\{10,20,30,40\}$, the condition number $K(t,y_{0\tau},\widehat{z}_{0\tau})$ (black solid lines) for an initial value $y_{0\tau}$ and a direction of perturbation $\widehat{z}_{0\tau}$ such that $K(\tau,y_{0\tau},\widehat{z}_{0\tau})=K(\tau)$.
-To obtain this, we take the initial value $y_{0\tau}$ such that
-$\|\mathrm{e}^{-\tau A}\|=\frac{1}{\|\mathrm{e}^{\tau A}\widehat{y}_{0\tau}\|}$
-(we take $y_{0\tau}=\mathrm{e}^{-\tau A}x_{0\tau}$, where $x_{0\tau}$ is such that
-$\|\mathrm{e}^{-\tau A}\|=\frac{\|\mathrm{e}^{-\tau A}x_{0\tau}\|}{\|x_{0\tau}\|})$
-and the direction of perturbation $\widehat{z}_{0\tau}$ such that
-$\|\mathrm{e}^{\tau A}\|=\|\mathrm{e}^{\tau A}\widehat{z}_{0\tau}\|.$
-The condition number $K(t,y_{0\tau},\widehat{z}_{0\tau})$ is the worst condition number $K(t,y_{0},\widehat{z}_{0})$, by varying the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$, only at $t=\tau$.
-But, asymptotically it goes below $K_{\infty}(t)$.
-On the other hand, the global asymptotic condition number $K_{\infty}(t)$ is the worst asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$, by varying the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$.
-Indeed, we have $K_{\infty}(t)=K_{\infty}(t,y_{0},\widehat{z}_{0})$ for $y_{0}\in U^{e}_{2}$ and for $\widehat{z}_{0}$ such that
-$\left\|Q_{11}(t)\right\|=\left\|P_{11}\right\|=\left\|P_{11}\widehat{z}_{0}\right\|$
-(see point d’) in Remark 40 and Remark 37).
-In Figure 7, we see the same for the matrix $A_{2}$. Here $\tau\in\{1,2,3,4\}$.
-Figure 7. The condition numbers $K(t)$, $K^{+}_{\infty}(t)$ and $K_{\infty}(t)$ for the matrix $A_{2}$.
-5. Some considerations about the asymptotic condition numbers
-Suppose that the matrix $A$ in the ODE (1) does not have all eigenvalues with the same real part, meaning $q>1$ (recall that $q$ is the number of distinct real parts).
-The global asymptotic condition number $K_{\infty}\left(t\right)$ grows exponentially with $t$.
-Hence, the relative error $\varepsilon$ in the initial value is exponentially magnified in the relative error $\delta(t)$ in the solution, in the worst case for the initial value and the perturbation of the initial value.
-However, this conclusion is too pessimistic, since this exponential magnification of $\varepsilon$ in $\delta(t)$ appears only in a non-generic situation for the initial value $y_{0}$.
-In fact, as illustrated in point 2 of Remark 33, in the generic situation $j^{\ast}=j^{\ast\ast}=1$ and $L^{\ast}=L^{\ast\ast}=L_{1}$ for $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$, we have
-$\frac{\delta(t)}{\varepsilon}=K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{1}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|},\ t\rightarrow+\infty,$ (61)
-where $Q_{1}(t):=Q_{1L_{1}}(t)$ to semplify the notation.
-In this generic situation, the magnification of the initial error remains bounded as well as away from zero by varying $t$ (see point 1 of Remark 33).
-We have introduced the three asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$, $K_{\infty}(t,$ $y_{0})$ and $K_{\infty}(t)$.
-The most important asymptotic condition number is $K_{\infty}(t,y_{0})$.
-In fact, $K_{\infty}(t,y_{0},\widehat{z}_{0})$ is given in terms of the (in general) unknown direction of perturbation $\widehat{z}_{0}$ and $K_{\infty}(t)$ is a worst asymptotic magnification factor which applies in a non-generic situation for $y_{0}$.
-As illustrated in point 2 of Remark 35, in the generic situation $j^{\ast}=1$ and $L^{\ast}=L_{1}$ for $y_{0}$, this most important asymptotic condition number $K_{\infty}(t,y_{0})$, representing the asympotic magnification factor $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$ in (61) for the worst perturbation of $y_{0}$, is given by
-$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{1}(t)\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|}.$ (62)
-Going into more detail, in (61) and (62), the vectors $Q_{1}(t)u$, with $u=\widehat{y}_{0}$ or $u=\widehat{z}_{0}$, and the matrix $Q_{1}(t)$ are given by (see (42), (29) and (30))
-$Q_{1}(t)u$ $=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}$
-$=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}$
-and
-$Q_{1}(t)$ $=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ m_{i}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}=M_{1}\end{subarray}}^{d_{i}}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)}$
-$=$ $\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)},$
-where
-$M_{1}:=L_{1}+1=\max\limits_{\lambda_{i}\in\Lambda_{1}}m_{i}=\max\limits_{\lambda_{i}\in\Lambda_{1}}\max\limits_{j\in\{1,\ldots,d_{i}\}}m_{ij}$
-to semplify the notation. This shows that the asymptotic behavior of the magnification in the solution of the initial relative error is determined by
-  • the imaginary parts $\omega_{i}$ of the rightmost eigenvalues of $A$ with the longest Jordan chains (chains of length $M_{1}$); in other words, with the largest mini-blocks in the JCF (mini-blocks of order $M_{1}$);
-  • the eigenvectors $v^{\left(i,j,1\right)}$ of these longest Jordan chains;
-  • the components $\alpha_{ijM_{1}}(u)$ of $u=\widehat{y}_{0},\widehat{z}_{0}$ along the last generalized eigenvectors of these longest Jordan chains;
-  • for the worst case of $\widehat{z}_{0}$, the last left generalized eigenvectors $w^{(i,j,M_{1})}$ corresponnding to these longest Jordan chains.
-For $u\in\mathbb{C}^{n}$, we say that $u$ satisfies the Rightmost Last Generalized Eigenvector (RLGE) condition if $j(u)=1$ and $L(u)=L_{1}$, i.e.
-there is a non-zero component $\alpha_{ijM_{1}}(u)$ along the last generalized eigenvctors of the longest Jordan chain of the rightmost eigenvalues of $A$.
-Therefore, we can say that (61) holds if $y_{0}$ (or $\widehat{y}_{0}$) and $\widehat{z}_{0}$ satisfy the RLGE condition and (62) holds if $y_{0}$ (or $\widehat{y}_{0}$) satisfies the RLGE condition.
-This terminology is utilized in [13].
-6. The generic real case. Part I: introduction
-In this section, Part I, and the following four sections, Parts II, III, IV, and V, we tailor the results from the previous Section 4 to a real ODE (1) in a generic situation for the matrix $A$, the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$.
-Since, as noted in the previous Section 5, the worst-case scenario of relative error magnification described by the asymptotic condition number $K_{\infty}(t)$ occurs in a non-generic situation for the initial value $y_{0}$, we focus our analysis on the other two more significant asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$.
-Parts I and II introduce concepts and present results essential for the developments in Parts III and IV, where asymptotic condition numbers are addressed.
-Parts I and III consider an arbitrary vector norm, while Parts II and IV explore the concepts introduced in Parts I and III, respectively, in the case of the Euclidean vector norm and a complex conjugate pairs of eigenvalues.
-The concluding section, Part V, investigates the onset of asymptotic behavior, specifically studying when a condition number begins to be close to its asymptotic form and identifying the factors that influence its rapid convergence towards this form.
-6.1. Real and complex induced norms
-[…]
-$\|B\|=\max_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \left\|u\right\|=1\end{subarray}}\left\|Bu\right\|,$ (63)
-instead of the complex induced norm
-$\max_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \left\|u\right\|=1\end{subarray}}\left\|Bu\right\|.$
-For a complex row vectors $b\in\mathbb{C}^{1\times n}$, we utilize both the real induced norm
-$\left\|b\right\|_{\mathbb{R}}:=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left|bu\right|$
-and the complex induced norm
-$\left\|b\right\|=\max_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \left\|u\right\|=1\end{subarray}}\left|bu\right|.$
-Observe that we have
-$\left\|b\right\|_{\mathbb{R}}\leq\left\|b\right\|.$ (64)
-Remark 43.
-[…]
-$\left\|b\right\|=\left\|b^{T}\right\|_{q},$
-where $\|\ \cdot\ \|_{q}$ is the vector $q$-norm with $\frac{1}{p}+\frac{1}{q}=1$.
-6.2. The generic $\Lambda_{j}$
-Assume $A\in\mathbb{R}^{n\times n}$ in (1).
-Let $j\in\{1,\ldots,q\}$ be such that the set $\Lambda_{j}$ consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues.
-We say that $\Lambda_{j}$ is real in the first case and complex in the second case.
-We consider a set $\Lambda_{j}$ of this type, since it is a generic situation for the real matrix $A$ to have all sets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, of such a form.
-When $\Lambda_{j}$ is real, we denote by $\lambda_{j}$ the real eigenvalue in $\Lambda_{j}$.
-When $\Lambda_{j}$ is complex, we denote by $\lambda_{j}$ and $\overline{\lambda_{j}}$ the complex conjugate pair in $\Lambda_{j}$, where $\lambda_{j}$ is the eigenvalue with positive imaginary part.
-This positive imaginary part is denoted by $\omega_{j}$.
-Let $V$ be the matrix whose columns form a Jordan basis for the matrix $A$ and let $W=V^{-1}$ (see Section 2).
-Let $v^{(j)}\in\mathbb{C}^{n}$ be the eigenvector of $A$ corresponding to the simple eigenvalue $\lambda_{j}$ (real eigenvalue or complex eigenvalue with positive imaginary part in the complex conjugate pair) appearing as a column in the matrix $V$.
-Let $w^{(j)}\in\mathbb{C}^{1\times n}$ be the corresponding row in the matrix $W$, i.e.
-the index of $v^{(j)}$ in $V$ matches the index of $w^{(j)}$ in $W$.
-Observe that $w^{(j)}$ is a left eigenvector of the matrix $A$ corresponding to the eigenvalue $\lambda_{j}$.
-If $\Lambda_{j}$ is real, then $v^{(j)}$ and $w^{(j)}$ are real vectors (see Proposition 5).
-The projection onto the eigenspace of the simple (real or complex) eigenvalue $\lambda_{j}$ is $v^{(j)}w^{(j)}$ (see (27)). Thus, by recalling Proposition 22, we have
-$Q_{j0}(t)=v^{(j)}w^{(j)}\in\mathbb{R}^{n\times n},$ (65)
-when $\Lambda_{j}$ is real and
-$Q_{j0}(t)=2\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}\right)\in\mathbb{R}^{n\times n},$ (66)
-when $\Lambda_{j}$ is complex.
-Remark 44.
-Recall Propostion 22.
-When $\Lambda_{j}$ is real, the projection on the eigenspace corresponding to $\lambda_{j}$ is $Q_{j0}(t)$, independent of $t$.
-When $\Lambda_{j}$ is complex, the projection onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$ is
-$Q_{j0}(0)=v^{(j)}w^{(j)}+\overline{v^{(j)}}\ \overline{w^{(j)}}=2\mathrm{Re}\left(v^{(j)}w^{(j)}\right).$
-6.3. The unit vectors $\widehat{v}^{\left(j\right)}$ and $\widehat{w}^{\left(j\right)}$ and the number $f_{j}$
-Let
-$\widehat{v}^{\left(j\right)}:=\frac{v^{\left(j\right)}}{\left\|v^{\left(j\right)}\right\|},\ \widehat{w}^{\left(j\right)}:=\frac{w^{\left(j\right)}}{\left\|w^{\left(j\right)}\right\|}\ \text{and}\ f_{j}:=\left\|w^{\left(j\right)}\right\|\cdot\left\|v^{\left(j\right)}\right\|.$ (67)
-In (67), when $\Lambda_{j}$ is real, $\left\|w^{\left(j\right)}\right\|$ denotes the real induced norm
-$\left\|w^{\left(j\right)}\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
-of the real row vector $w^{(j)}$. On the other hand, when $\Lambda_{j}$ is complex, $\left\|w^{\left(j\right)}\right\|$ denotes the complex induced norm
-$\left\|w^{\left(j\right)}\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \|u\|=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
-of the complex row vector $\widehat{w}^{(j)}$.
-6.4. The numbers $g_{j}(t,u)$ and $g_{j}(t)$, the vector $\widehat{\Theta}_{j}(t,u)$ and the matrix $\widehat{\Theta}_{j}(t)$.
-In the following, in both cases for $\Lambda_{j}$, real or complex, we write the norm of the real vector $Q_{j0}(t)u$, where $u\in\mathbb{R}^{n}$, as
-$\|Q_{j0}(t)u\|=f_{j}\left|\widehat{w}^{(j)}u\right|g_{j}(t,u)$ (68)
-for a suitable number $g_{j}(t,u)$. Analogously, we write the real induced norm (63) of the real matrix $Q_{j0}(t)$ as
-$\|Q_{j0}(t)\|=f_{j}g_{j}(t)$ (69)
-for a suitable number $g_{j}(t)$.
-Next two lemmas say what $g_{j}(t,u)$ and $g_{j}(t)$ are.
-Lemma 45.
-Let $j\in\{1,\ldots,q\}$. If $\Lambda_{j}$ is real, then
-$g_{j}(t,u)=1,\ u\in\mathbb{R}^{n},$
-$g_{j}(t)=1.$
-Proof.
-For $u\in\mathbb{R}^{n}$, by (65) we have
-$Q_{j0}(t)u=v^{(j)}w^{(j)}u=\left(w^{(j)}u\right)v^{(j)}=f_{j}\left(\widehat{w}^{(j)}u\right)\widehat{v}^{(j)}$
-and then
-$\left\|Q_{j0}(t)u\right\|=\left|w^{(j)}u\right|\left\|v^{(j)}\right\|=f_{j}\left|\widehat{w}^{(j)}u\right|.$ (70)
-By the second equality in (70), we obtain $g_{j}(t,u)=1$. By the first equality, we obtain
-$\left\|Q_{j0}(t)\right\|=\left\|w^{(j)}\right\|\left\|v^{(j)}\right\|=f_{j}$
-and then $g_{j}(t)=1$. ∎
-Lemma 46.
-Let $j\in\{1,\ldots,q\}$. If $\Lambda_{j}$ is complex, then
-$g_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|,\ u\in\mathbb{R}^{n},$
-$g_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|,$
-where
-$\widehat{\Theta}_{j}\left(t,u\right)=\left(\left|\widehat{v}_{k}^{\left(j\right)}\right|\cos\left(\omega_{j}t+\alpha_{jk}+\gamma_{j}\left(u\right)\right)\right)_{k=1,\ldots,n}\in\mathbb{R}^{n}$
-$\widehat{\Theta}_{j}\left(t\right)=\left(\left|\widehat{v}_{k}^{\left(j\right)}\right|\left|\widehat{w}_{l}^{\left(j\right)}\right|\cos\left(\omega_{j}t+\alpha_{jk}+\beta_{jl}\right)\right)_{k,l=1,\ldots,n}\in\mathbb{R}^{n\times n}$
-with $\alpha_{jk}$, $\beta_{jl}$ and $\gamma_{j}(u)$ defined by
-$\widehat{v}_{k}^{\left(j\right)}=\left|\widehat{v}_{k}^{\left(j\right)}\right|\mathrm{e}^{\sqrt{-1}\alpha_{jk}},\ k\in\{1,\ldots,n\},$
-$\widehat{w}_{l}^{\left(j\right)}=\left|\widehat{w}_{l}^{\left(j\right)}\right|\mathrm{e}^{\sqrt{-1}\beta_{jl}},\ l\in\{1,\ldots,n\},$
-$\widehat{w}^{\left(j\right)}u=\left|\widehat{w}^{\left(j\right)}u\right|\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)},$
-i.e.
-they are, respectively, the angles in the polar forms of the components of the complex vector $\widehat{v}^{(j)}$, the angles in the polar forms of the components of the complex vector $\widehat{w}^{(j)}$ and the angle in the polar form of the complex number $\widehat{w}^{\left(j\right)}u$.
-Moreover, we have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left(\left|\widehat{w}^{\left(j\right)}u\right|\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|\right).$ (71)
-In this lemma, $\left\|\widehat{\Theta}_{j}(t)\right\|$ is the real induced norm (63) of the real matrix $\widehat{\Theta}_{j}(t)$.
-Proof.
-By (66), we have
-$Q_{j0}(t)=2\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}\right)=2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right).$ (72)
-Thus, for $u\in\mathbb{R}^{n}$, we have
-$Q_{j0}(t)u$ $=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right)u$
-$=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}u\right)$
-$=$ $2f_{j}\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\left(\widehat{w}^{\left(j\right)}u\right)\widehat{v}^{(j)}\right)$
-$=$ $2f_{j}\left|\widehat{w}^{\left(j\right)}u\right|\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)}\widehat{v}^{(j)}\right)$
-and
-$\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\mathrm{e}^{\sqrt{-1}\gamma_{j}\left(u\right)}\widehat{v}^{(j)}\right)=\widehat{\Theta}_{j}\left(t,u\right).$
-We have obtained
-$Q_{j0}(t)u=2f_{j}\left|\widehat{w}^{\left(j\right)}u\right|\widehat{\Theta}_{j}\left(t,u\right).$ (73)
-and then $g_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|$.
-Moreover, in (72) we have
-$\mathrm{Re}\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}\widehat{v}^{(j)}\widehat{w}^{(j)}\right)=\widehat{\Theta}_{j}(t).$
-Therefore, we obtain
-$Q_{j0}(t)=2f_{j}\widehat{\Theta}_{j}\left(t\right)$ (74)
-and then $g_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|$.
-Finally, by (74) and (73), we have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|=\frac{\left\|Q_{j0}(t)\right\|}{2f_{j}}=\frac{1}{2f_{j}}\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left\|Q_{j0}(t)u\right\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}\left(\left|\widehat{w}^{\left(j\right)}u\right|\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|\right).$
-Remark 47.
-  • 1.
-    The functions $t\mapsto\widehat{\Theta}_{j}(t,u)$ and $t\mapsto\widehat{\Theta}_{j}(t)$ are periodic functions of period $\frac{2\pi}{\omega_{j}}$ and the functions $t\mapsto\|\widehat{\Theta}_{j}(t,u)\|$ and $t\mapsto\|\widehat{\Theta}_{j}(t)\|$ are periodic functions of period $\frac{\pi}{\omega_{j}}$.
-  • 2.
-    When the vector norm is a $p$-norm, we have
-    $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{p}=\left(\sum\limits_{k=1}^{n}\left|\widehat{v}_{k}^{\left(j\right)}\right|^{p}\left|\cos\left(\omega_{j}t+\alpha_{jk}+\gamma_{j}\left(u\right)\right)\right|^{p}\right)^{\frac{1}{p}}\leq\left\|\widehat{v}^{\left(j\right)}\right\|_{p}=1.$
-  • 3.
-    When the vector norm is a $p$-norm, we have $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{p}\leq 1$. This follows by (71),
-[…]
-    and the previous point 2.
-7. The generic real case. Part II: introduction (euclidean norm and complex conjugate eigenvalues)
-[…]
-In the following, for a complex row or column vector $z$, $z^{T}$ denotes the transpose of $z$, representing only transposition without conjugation.
-The conjugate transpose of $z$ is denoted by $z^{H}$.
-7.1. The set $\mathcal{W}_{j}$ and the ellipse $\mathcal{E}_{j}$.
-For $u\in\mathbb{R}^{n}$, we have (recall (68) and Lemma 46)
-$\left\|Q_{j0}(t)u\right\|_{2}=2f_{j}\left|\widehat{w}^{(j)}u\right|\left\|\widehat{\Theta}_{j}(t,u)\right\|_{2}.$ (75)
-The complex number $\widehat{w}^{(j)}u$ plays a crucial role in this formula: its modulus is part of it, and the vector $\widehat{\Theta}_{j}(t,u)$ is defined in Lemma 46 by means of the angle $\gamma_{j}(u)$ in the polar form of $\widehat{w}^{(j)}u$.
-For this reason, now we study the set
-$\mathcal{W}_{j}:=\{\widehat{w}^{(j)}u:u\in\mathbb{R}^{n}\text{\ and\ }\left\|u\right\|_{2}=1\}$
-of the complex numbers $\widehat{w}^{(j)}u$, by varying the unit real vector $u$.
-Next proposition describes the set $\mathcal{W}_{j}$ and states the real induced norm
-$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left|w^{\left(j\right)}u\right|$
-of the complex row vector $\widehat{w}^{(j)}$, whose complex induced norm is equal to $1$.
-Proposition 48.
-Assume that $A$ is a real matrix.
-Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex.
-Consider, in the complex plane, the ellipse $\mathcal{E}_{j}$ centered at the origin with semi-axes the vectors $\sigma_{j}\alpha^{(j)}$ and $\mu_{j}\beta^{(j)}$, where $\sigma_{j}\geq\mu_{j}\geq 0$ are the singular values of the $2\times n$ real matrix
-$R_{j}:=\left[\begin{array}[]{c}\mathrm{Re}(\widehat{w}^{(j)})\\ \mathrm{Im}(\widehat{w}^{(j)})\end{array}\right]$
-and $\alpha^{(j)},\beta^{(j)}\in\mathbb{R}^{2}$ are the corresponding left singular vectors. We have
-$\sigma_{j}=\sqrt{\frac{1+W_{j}}{2}}\text{\ \ and\ \ }\mu_{j}=\sqrt{\frac{1-W_{j}}{2}},$ (76)
-where $W_{j}$ is the modulus of the complex number $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$.
-If $n=2$, the set $\mathcal{W}_{j}$ is the ellipse $\mathcal{E}_{j}$. If $n>2$, the set $\mathcal{W}_{j}$ is the union of the ellipse $\mathcal{E}_{j}$ and its interior. Moreover, we have
-$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\sigma_{j}=\sqrt{\frac{1+W_{j}}{2}}.$
-Proof.
-We begin by proving (76). The singular values $\sigma_{j}$ and $\mu_{j}$ of the matrix $R_{j}$ are the square roots of the largest eigenvalue and the smallest eigenvalue, respectively, of the matrix
-$R_{j}R_{j}^{T}=\left[\begin{array}[]{c}\mathrm{Re}(w^{(j)})\\ \mathrm{Im}(w^{(j)})\end{array}\right]\left[\mathrm{Re}(w^{(j)})^{T}\ \mathrm{Im}(w^{(j)})^{T}\right]=\left[\begin{array}[]{cc}a_{j}^{2}&p_{j}\\ p_{j}&b_{j}^{2}\end{array}\right],$
-where
-$a_{j}=\left\|\mathrm{Re}\left(\widehat{w}^{(j)}\right)^{T}\right\|_{2},\ b_{j}=\left\|\mathrm{Im}\left(\widehat{w}^{(j)}\right)^{T}\right\|_{2}\text{\ \ and\ \ }p_{j}=\mathrm{Re}\left(\widehat{w}^{(j)}\right)\mathrm{Im}\left(\widehat{w}^{(j)}\right)^{T}.$
-Therefore, we obtain
-$\sigma_{j}^{2}=\frac{1+\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}}{2}\text{\ and\ }\mu_{j}^{2}=\frac{1-\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}}{2}.$
-by observing that $a_{j}^{2}+b_{j}^{2}=\left\|\widehat{w}^{(j)}\right\|_{2}^{2}=1$. Since
-$\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}=a_{j}^{2}-b_{j}^{2}+\sqrt{-1}\cdot 2p_{j},$
-we get
-$W_{j}=\sqrt{\left(a_{j}^{2}-b_{j}^{2}\right)^{2}+4p_{j}^{2}}$
-and (76) follows.
-Now, we pass to describe the set $\mathcal{W}_{j}$ in terms of the ellipse $\mathcal{E}_{j}$.
-By looking to a complex number as the pair constituted by the real and imaginary parts, the complex number $\widehat{w}^{(j)}u$, $u\in\mathbb{R}^{n}$, is $R_{j}u$.
-By introducing the SVD of $R_{j}$, we see that
-$R_{j}u=\sigma_{j}c_{j}(u)\alpha^{(j)}+\mu_{j}d_{j}(u)\beta^{(j)}$
-where $c_{j}(u)$ and $d_{j}(u)$ are the components of $u$ along the first and second right singular vectors of $R_{j}$, i.e.
-the right singular vectors corresponding to $\sigma_{j}$ and $\mu_{j}$, respectively.
-Thus,
-$\mathcal{W}_{j}=\{\sigma_{j}c\alpha^{(j)}+\mu_{j}d\beta^{(j)}:(c,d)\in\mathbb{R}^{2}\text{\ and\ }\left\|(c,d)\right\|_{2}=1\}$ (77)
-when $n=2$ and
-$\mathcal{W}_{j}=\{\sigma_{j}c\alpha^{(j)}+\mu_{j}d\beta^{(j)}:(c,d)\in\mathbb{R}^{2}\text{\ and\ }\left\|(c,d)\right\|_{2}\leq 1\}$ (78)
-when $n>2$.
-In the orthonormal basis of $\mathbb{R}^{2}$ given by the vector $\alpha^{(j)}$ and $\beta^{(j)}$, the coordinates $x=\sigma_{j}c$ and $y=\mu_{j}d$ of the elements of the set (77) satisfies the equation
-$\frac{x^{2}}{\sigma_{j}^{2}}+\frac{y^{2}}{\mu_{j}^{2}}=1,$
-which is the equation of an ellipse centered at the origin with lengths of the semi-axes $\sigma_{j}$ and $\mu_{j}$.
-Thus, we have $\mathcal{W}_{j}=\mathcal{E}_{j}$.
-On the other hand, the coordinates $x=\sigma_{j}c$ and $y=\mu_{j}d$ of the elements of the set (78) satisfies the inequality
-$\frac{x^{2}}{\sigma_{j}^{2}}+\frac{y^{2}}{\mu_{j}^{2}}\leq 1.$
-Thus, $\mathcal{W}_{j}$ is the union of the ellipse $\mathcal{E}_{j}$ and its interior.
-Finally, we have
-$\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left|\widehat{w}^{\left(j\right)}u\right|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|_{2}=1\end{subarray}}\left\|R_{j}u\right\|_{2}=\left\|R_{j}\right\|_{2}=\sigma_{j}.$
-Remark 49.
-  • 1.
-    By the Cauchy-Schwarz inequality and Remark 43, or only by Remark 43, we have
-    $W_{j}=\left|\widehat{w}^{\left(j\right)}\left(\widehat{w}^{\left(j\right)}\right)^{T}\right|=\left|\left(\left(\overline{\widehat{w}^{\left(j\right)}}\right)^{T}\right)^{H}\left(\widehat{w}^{\left(j\right)}\right)^{T}\right|\leq 1.$
-    Moreover, $W_{j}<1$ holds, otherwise $\widehat{w}^{\left(j\right)}$ and $\overline{\widehat{w}^{\left(j\right)}}$ were linearly dependent, which is impossible since $w^{\left(j\right)}$ and $\overline{w^{\left(j\right)}}$ are the rows of the non-singular matrix $W$ corresponding to the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively.
-  • 2.
-    Remind the inequality (64). By the previous point 1, we have
-    $\left\|\widehat{w}^{\left(j\right)}\right\|_{\mathbb{R}}=\sqrt{\frac{1+W_{j}}{2}}<1$
-    for the real induced norm of the row vector $\widehat{w}^{(j)}$. On the other hand,
-    $\left\|\widehat{w}^{(j)}\right\|_{2}=1.$
-    for the complex induced norm of $\widehat{w}^{(j)}$.
-7.2. Projection onto the eigenspaces of $\lambda_{j}$ and $\overline{\lambda}_{j}$
-The projection of $u\in\mathbb{R}^{n}$ onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda}_{j}$ is (see Remark 44)
-$f_{j}\left(\widehat{w}^{(j)}u\right)\widehat{v}^{(j)}+f_{j}\overline{\left(\widehat{w}^{(j)}u\right)}\ \overline{\widehat{v}^{(j)}},$
-where the first and the second terms are the projections onto the eigenspace corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively.
-Since $\widehat{v}^{(j)}$ and $\overline{\widehat{v}^{(j)}}$ are linearly independent, the projection onto the sum of the eigenspaces is zero if and only if the projection onto the eigenspace corresponding to $\lambda_{j}$ is zero, i.e.
-$\widehat{w}^{(j)}u=0$.
-The Euclidean norm of the projection onto the eigenspace corresponding to $\lambda_{j}$ is $f_{j}\left|\widehat{w}^{(j)}u\right|$, equal to the norm of the projection onto the eigenspace corresponding to $\overline{\lambda_{j}}$.
-The Euclidean norm of the projection on the sum of the eigenspaces is given in Remark 53 below.
-7.3. The right singular vectors of the matrix $R_{j}$
-Vectors in $\mathbb{R}^{n}$ can be expressed by real components in the orthonormal basis of $\mathbb{R}^{n}$ of the right singular vectors of the matrix $R_{j}$ of Proposition 48.
-As we saw in the proof of Proposition 48, for a vector $u\in\mathbb{R}^{n}$ we have
-$\widehat{w}^{(j)}u=a+\sqrt{-1}b$ (79)
-with
-$(a,b)=R_{j}u=\sigma_{j}c_{j}(u)\alpha^{(j)}+\mu_{j}d_{j}(u)\beta^{(j)},$
-where: $\sigma_{j}$ and $\mu_{j}$ are the two singular values of $R_{j}$; $c_{j}(u)$ and $d_{j}(u)$ are the components of $u$ along the first and second, respectively, right singular vectors of $R_{j}$; and $\alpha^{(j)}$ and $\beta^{(j)}$ are the left singular vectors of $R_{j}$.
-Therefore, $u$ has projection zero onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda}_{j}$, i.e.
-$\widehat{w}^{(j)}u=0$ by reminding the previous Subsection 7.2, if and only if $u$ is spanned by the last $n-2$ right singular vectors of $R_{j}$.
-By (79), we have
-$\left|\widehat{w}^{(j)}u\right|=\left\|R_{j}u\right\|_{2}=\sqrt{\sigma_{j}^{2}c_{j}(u)^{2}+\mu_{j}^{2}d_{j}(u)^{2}}.$ (80)
-Remark 50.
-Suppose $n>2$.
-By considering the components $c_{j}(u)$ and $d_{j}(u)$ of $u\in\mathbb{R}^{n}$ with $\left\|u\right\|_{2}=1$ along the first and the second right singular vectors of $R_{j}$, the set $\mathcal{W}_{j}$ can be partitioned by level curves of
-$\sigma_{j}^{2}c_{j}(u)^{2}+\mu_{j}^{2}d_{j}(u)^{2}=\left|\widehat{w}^{(j)}u\right|^{2},$
-whose set of possible values by varying $u$ is $[0,\sigma_{j}^{2}]$, or by level curves of
-$c_{j}(u)^{2}+d_{j}(u)^{2},$
-whose set of possible values is $[0,1]$.
-In the first case, the level curves are concentric circles with center the origin.
-In the second case, the level curves are ellipses concentric to the ellipse $\mathcal{E}_{j}$.
-7.4. Notations
-Before continuing with our analysis, we recall two important notations and introduce a few others.
-Here are the two previously introduced notations.
-  • The modulus $W_{j}$ of the complex number $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$, introduced in Proposition 48;
-  • The angle $\gamma_{j}(u)$ in the polar form of the complex number $\widehat{w}^{(j)}u$, where $u\in\mathbb{R}^{n}$, introduced In Lemma 46.
-Here are the new notations.
-  • The modulus $V_{j}$ of the complex number $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$.
-  • The angle $\delta_{j}$ in the polar form of the complex number $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$.
-  • The angle $\theta_{j}$ in the complex plane between the real axis and the major axis $\sigma_{j}\alpha^{(j)}$ of the ellipse $\mathcal{E}_{j}$ of Proposition 48.
-  • $x_{j}(t):=2(\omega_{j}t+\theta_{j})+\delta_{j}.$ (81)
-  • For $u\in\mathbb{R}^{n}$,
-    $\Delta_{j}(u):=2(\gamma_{j}(u)-\theta_{j}).$ (82)
-Remark 51.
-  • 1.
-    Observe that $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{H}=1$ and $\left(\widehat{v}^{\left(j\right)}\right)^{H}\widehat{v}^{\left(j\right)}=1$.
-On the other hand, $\widehat{w}^{(j)}\left(\widehat{w}^{(j)}\right)^{T}$ and $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$ are complex numbers and $W_{j}$ and $V_{j}$ are their modules, respectively.
-  • 2.
-    At point 1 of Remark 49, we have seen that $W_{j}<1$, The same holds for $V_{j}$. In fact, by the Cauchy-Schwarz inequality we have
-    $V_{j}=\left|\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}\right|=\left|\left(\overline{\widehat{v}^{\left(j\right)}}\right)^{H}\widehat{v}^{\left(j\right)}\right|\leq 1.$
-    Moreover, $V_{j}<1$ holds, otherwise $\widehat{v}^{\left(j\right)}$ and $\overline{\widehat{v}^{\left(j\right)}}$ were linearly dependent, which is impossible since $v^{\left(j\right)}$ and $\overline{v^{\left(j\right)}}$ are the columns of the non-singular matrix $V$ corresponding to the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, respectively.
-Therefore, we have $V_{j},W_{j}\in[0,1)$.
-7.5. The norm $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$
-Next proposition provides an expression for the norm $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$.
-Proposition 52.
-Assume that $A$ is a real matrix. Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex. For $u\in\mathbb{R}^{n}$, we have
-$\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\Delta_{j}(u)\right)}{2}},$ (83)
-where $x_{j}(t)$ and $\Delta_{j}(u)$ are defined in (81) and (82).
-Proof.
-Let
-$\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}=V_{j}\mathrm{e}^{\sqrt{-1}\delta_{j}}$
-be the polar form of $\left(\widehat{v}^{\left(j\right)}\right)^{T}\widehat{v}^{\left(j\right)}$ (remind Subsection 7.4). We have (remind (66))
-$\left\|Q_{j0}(t)u\right\|_{2}^{2}=\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u+\overline{\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u}\right\|_{2}^{2}$
-$=2\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right\|_{2}^{2}+2\mathrm{Re}\left(\left(\overline{\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u}\right)^{H}\mathrm{e}^{\sqrt{-1}\omega_{j}}v^{(j)}w^{(j)}u\right)$
-$=2\left\|\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right\|_{2}^{2}+2\mathrm{Re}\left(\left(\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right)^{T}\mathrm{e}^{\sqrt{-1}\omega_{j}t}v^{(j)}w^{(j)}u\right)$
-$=2\left(\left\|v^{(j)}\right\|_{2}^{2}\left|w^{(j)}u\right|^{2}+\mathrm{Re}\left(\mathrm{e}^{2\sqrt{-1}\omega_{j}t}\left(\left(v^{(j)}\right)^{T}v^{(j)}\right)\left(w^{(j)}u\right)^{2}\right)\right)$
-$=2\left(f_{j}^{2}|\widehat{w}^{(j)}u|^{2}+\mathrm{Re}\left(f_{j}^{2}|\widehat{w}^{(j)}u|^{2}\mathrm{e}^{2\sqrt{-1}\omega_{j}t}V_{j}\mathrm{e}^{\sqrt{-1}\delta_{j}}\mathrm{e}^{2\sqrt{-1}\gamma_{j}(u)}\right)\right)$
-$=2f_{j}^{2}\left|\widehat{w}^{(j)}u\right|^{2}\left(1+V_{j}\cos\left(2\omega_{j}t+\delta_{j}+2\gamma_{j}(u)\right)\right)$ (84)
-$=2f_{j}^{2}\left|\widehat{w}^{(j)}u\right|^{2}\left(1+V_{j}\cos\left(x_{j}(t)+\Delta_{j}(u)\right)\right).$
-Now, (83) follows by (75).
-When $V_{j}=0$, we have
-$\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1}{2}}$
-On the other hand, when $V_{j}\neq 0$, $\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}$, considered as $\frac{\pi}{\omega_{j}}$-periodic function of $t$, oscillates monotonically between the values
-$\max\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1+V_{j}}{2}}$ (85)
-and
-$\min\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t,u\right)\right\|_{2}=\sqrt{\frac{1-V_{j}}{2}}.$ (86)
-The maximum value is achieved if and only if $x_{j}(t)+\Delta_{j}(u)$ is an even multiple di $\pi$ and the minimum value is achieved if and only if $x_{j}(t)+\Delta_{j}(u)$ is an odd multiple di $\pi$.
-Remark 53.
-By using $t=0$ in (84), we obtain
-$\left\|Q_{j0}(0)u\right\|_{2}=f_{j}\left|\widehat{w}^{(j)}u\right|\sqrt{2\left(1+V_{j}\cos\left(\delta_{j}+2\gamma_{j}(u)\right)\right)}.$
-and this is norm of the projection of $u\in\mathbb{R}^{n}$ onto the sum of the eigenspaces corresponding to $\lambda_{j}$ and $\overline{\lambda_{j}}$ (remind Remark 44).
-The norm of the projection on each eigenspace is $f_{j}\left|\widehat{w}^{(j)}u\right|$ (see Subsection 7.2).
-7.6. The functions $f_{VW}$, $f_{VW}^{\max}$ and $f_{VW}^{\min}$
-For the analysis of the norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$, as well as for the analysis of the oscillating terms in Section 9, we introduce the functions $f_{VW}$, $f_{VW}^{\max}$ and $f_{VW}^{\min}$, where $V$ and $W$ are parameters in the interval $[0,1)$.
-(In this subsection, $V$ is NOT the matrix $V$, with columns the Jordan basis, and W is NOT the inverse of the matrix $V$).
-We use the symbols $V$ and $W$ because we have in mind the values $V_{j},W_{j}\in[0,1)$.
-The function $f_{VW}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is given by
-$f_{VW}\left(\alpha,x\right):=\frac{1+V\cos\left(x+\alpha\right)}{1-W\cos\alpha},\ \left(\alpha,x\right)\in\mathbb{R}^{2}.$ (87)
-For $x\in\mathbb{R}$, let
-$U_{VW}(x):=V\mathrm{e}^{\sqrt{-1}x}+W.$
-If $U_{VW}(x)=0$, i.e. $V=W=0$ or $V=W\neq 0$ and $x$ is an odd multiple of $\pi$, then
-$f_{VW}\left(\alpha,x\right)=1,\ \alpha\in\mathbb{R}.$
-The next lemma states the form of the $2\pi$-periodic function $f_{VW}(\cdot,x)$, for a given $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$.
-Lemma 54.
-Let $V,W\in[0,1)$ and let $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$.
-The function $f_{VW}(\cdot,x)$ is not constant and it oscillates monotonically between its maximum value $f_{VW}(\alpha^{\max}_{VW}(x),x)$ and minimum value $f_{VW}(\alpha^{\min}_{VW}(x),x)$, where $\alpha^{\max}_{VW}(x)$ and $\alpha^{\min}_{VW}(x)$ are given by
-$\alpha^{\max}_{VW}(x)=\arcsin\frac{VW\sin x}{|U_{VW}(x)|}-\theta_{VW}(x)$
-and
-$\alpha^{\min}_{VW}(x)=\pi-\arcsin\frac{VW\sin x}{|U_{VW}(x)|}-\theta_{VW}(x)$
-with $\theta_{VW}(x)$ the angle in the polar form of $U_{VW}(x)$.
-Proof.
-In this proof, to simplify the notation, we write $U(x)$ for $U_{VW}(x)$ and $\theta(x)$ for $\theta_{VW}(x)$.
-We determine the stationary points of $f_{VW}(\cdot,x)$. By deriving $f_{VW}$ with respect to the argument $\alpha$, we obtain
-$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)-W\sin\alpha+VW\sin x}{\left(1-W\cos\alpha\right)^{2}},\ \alpha\in\mathbb{R}.$
-Thus, $\alpha\in\mathbb{R}$ is a stationary point of $f_{VW}(\cdot,x)$ if and only if
-$V\sin(x+\alpha)+W\sin\alpha=VW\sin x,$ (88)
-which can equivalently rewritten as
-$\mathrm{Im}\left(U(x)\mathrm{e}^{\sqrt{-1}\alpha}\right)=VW\sin x$
-since
-$V\sin(x+\alpha)+W\sin\alpha=\mathrm{Im}\left(U(x)\mathrm{e}^{\sqrt{-1}\alpha}\right)$ (89)
-By writing $U(x)=|U(x)|\mathrm{e}^{\sqrt{-1}\theta(x)}$ in polar form, we obtain
-$\sin(\alpha+\theta(x))=\frac{VW\sin x}{|U(x)|}$
-and then the function $f_{VW}(\cdot,x)$ has the stationary points
-$\alpha=\arcsin\frac{VW\sin x}{|U(x)|}-\theta(x)+2\pi k$ (90)
-and
-$\alpha=\pi-\arcsin\frac{VW\sin x}{|U(x)|}-\theta(x)+2\pi k,$ (91)
-where $k\in\mathbb{Z}$. In Appendix A, we show that
-$\left|\frac{VW\sin x}{|U(x)|}\right|=\frac{|VW\sin x|}{|U(x)|}<1.$ (92)
-The maximum and the minimum points of $f_{VW}(\cdot,x)$ are these stationary points. Since (recall (89))
-$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=|U(x)|\frac{-\sin(\alpha+\theta(x))+\frac{VW\sin x}{|U(x)|}}{\left(1-W\cos\alpha\right)^{2}},\ \alpha\in\mathbb{R},$
-the second derivative of $f_{VW}$ with respect to the argument $\alpha$ is
-$\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}\left(\alpha,x\right)=|U(x)|\frac{-\cos(\alpha+\theta(x))}{\left(1-W\cos\alpha\right)^{2}},$
-when $\alpha$ is stationary point. Thus (90) are maximum points and (91) are minimum points. ∎
-Remark 55.
-Observe that the functions $\alpha_{VW}^{\max}$ and $\alpha_{VW}^{\min}$ are differentiable at every point where they are defined, i.e.
-at every point $x\in\mathbb{R}$ such that $U_{VW}(x)\neq 0$.
-This is a consequence of the fact that the argument of $\mathrm{arcsin}$ in the definitions of $\alpha_{VW}^{\max}$ and $\alpha_{VW}^{\min}$ in Lemma 54 is neither $-1$ nor $1$, as can be seen in (92).
-The functions $f_{VW}^{\max},f_{VW}^{\min}:\mathbb{R}\rightarrow\mathbb{R}$ are given by
-$f_{VW}^{\max}(x)=\max\limits_{\alpha\in\mathbb{R}}f_{VW}\left(\alpha,x\right),\ x\in\mathbb{R},$ (93)
-and
-$f_{VW}^{\min}(x)=\min\limits_{\alpha\in\mathbb{R}}f_{VW}\left(\alpha,x\right),\ x\in\mathbb{R}.$ (94)
-If $V=0$ or $W=0$, then
-$f_{VW}^{\max}(x)=\frac{1+V}{1-W},\ x\in\mathbb{R},$ (95)
-and
-$f_{VW}^{\min}(x)=\frac{1-V}{1+W},\ x\in\mathbb{R},$
-are constant functions. Thus, assume $V\neq 0$ and $W\neq 0$. The forms of the $2\pi$-periodic functions $f_{VW}^{\max}$ and $f_{VW}^{\min}$ are stated in the next Lemma.
-Lemma 56.
-Let $V,W\in[0,1)$ such that $V\neq 0$ and $W\neq 0$.
-The function $f_{VW}^{\max}$ oscillates monotonically between the values
-$\max\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1-W}$
-and
-$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\left\{\begin{array}[]{l}\frac{1-V}{1-W}\text{\ if\ }V\leq W\\ \\ \frac{1+V}{1+W}\text{\ if\ }V\geq W.\end{array}\right.$
-The maximum value is achieved if and only if $x$ is an even multiple of $\pi$ and the minimum value is achieved if and only if $x$ is an odd multiple of $\pi$.
-The function $f_{VW}^{\min}$ oscillates monotonically between the values
-$\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\left\{\begin{array}[]{l}\frac{1+V}{1+W}\text{\ if\ }V\leq W\\ \\ \frac{1-V}{1-W}\text{\ if\ }V\geq W\end{array}\right.$
-and
-$\min\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1+W}.$
-The maximum value is achieved if and only if $x$ is an odd multiple of $\pi$ and the minimum value is achieved if and only if $x$ is an even multiple of $\pi$.
-Proof.
-We determine the stationary points of two-variable function $f_{VW}$. We have, for $(\alpha,x)\in\mathbb{R}^{2}$,
-$\frac{\partial f_{VW}}{\partial\alpha}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)-W\sin\alpha+VW\sin x}{\left(1-W\cos\alpha\right)^{2}}$
-$\frac{\partial f_{VW}}{\partial x}\left(\alpha,x\right)=\frac{-V\sin(x+\alpha)}{1-W\cos\alpha}$ (98)
-Hence, $(\alpha,x)\in\mathbb{R}^{2}$ is a stationary point of $f_{VW}$ if and only if
-$-V\sin(x+\alpha)-W\sin\alpha+VW\sin x=0$
-$\sin(x+\alpha)=0,$
-which is equivalent to $\alpha$ and $x$ being multiples of $\pi$.
-In the following table, we see the values of
-$f_{VW}(\alpha,x)=\frac{1+V\cos(x+\alpha)}{1-W\cos\alpha}$
-at the stationary points $(\alpha,x)$.
-$x$ even multiple of $\pi$ $x$ odd multiple of $\pi$
-$\alpha$ even multiple of $\pi$ $\frac{1+V}{1-W}$ $\frac{1-V}{1-W}$
-$\alpha$ odd multiple of $\pi$ $\frac{1-V}{1+W}$ $\frac{1+V}{1+W}$
-Observe that
-$\max\limits_{x\in\mathbb{R}}f^{\max}_{VW}(x)=\max\limits_{(\alpha,x)\in\mathbb{R}^{2}}f_{VW}(\alpha,x)\text{\ \ and\ \ }\min\limits_{x\in\mathbb{R}}f^{\min}_{VW}(x)=\min\limits_{(\alpha,x)\in\mathbb{R}^{2}}f_{VW}(\alpha,x)$
-are values of stationary points of $f_{VW}$. Since
-$\frac{1+V}{1-W}\text{\ \ and\ \ }\frac{1-V}{1+W}$
-are, respectively, the maximum and the minimum among the values in (99), we conclude that
-$\max\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1-W}\text{\ \ \ and\ \ \ }\min\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1+W}$ (100)
-with both the maximum and minimum achieved for $x$ being an even multiple of $\pi$.
-Observe that
-$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)\text{\ and \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)$ (101)
-also are values of stationary points of $f_{VW}$. In fact, for $x\in\mathbb{R}$, we have (remind Lemma 54 and Remark 55)
-$\left(f_{VW}^{\max}\right)^{\prime}\left(x\right)$ $=$ $\frac{d}{dx}f_{VW}(\alpha^{\max}_{VW}(x),x)$
-$=$ $\frac{\partial f_{VW}}{\partial\alpha}(\alpha^{\max}_{VW}(x),x)(\alpha^{\max}_{VW})^{\prime}(x)+\frac{\partial f_{VW}}{\partial x}(\alpha^{\max}_{VW}(x),x)$
-$=$ $\frac{\partial f_{VW}}{\partial x}(\alpha^{\max}_{VW}(x),x).$
-and, analogously,
-$\left(f_{VW}^{\min}\right)^{\prime}\left(x\right)=\frac{\partial f_{VW}}{\partial x}(\alpha^{\min}_{VW}(x),x).$
-Therefore, the values (101) are those in (99) different from (100), i.e. those in the second column.
-Indeed, when $V=W$, $x$ cannot be an odd multiple of $\pi$ in the previous argument, since $\alpha^{\max}_{VW}(x)$ and $\alpha^{\min}_{VW}(x)$ are not defined in this case.
-However, when $V=W$, we can still conclude that the values in (101) are stationary points of $f_{VW}$.
-In fact, for $x\in\mathbb{R}$ that is not an odd multiple of $\pi$, the previous argument works and then $\left(f_{VW}^{\max}\right)^{\prime}(x)$ and $\left(f_{VW}^{\min}\right)^{\prime}(x)$ are zero only when $x$ is a multiple of $\pi$, indeed an even multiple of $\pi$.
-Thus, the maximum and minimum values of $\left(f_{VW}^{\max}\right)(x)$ and $\left(f_{VW}^{\min}\right)(x)$ as $x$ varies can be achieved only for $x$ being a multiple of $\pi$.
-Now, we immediately conclude that the values in (101) are stationary points of $f_{VW}$.
-When $V=W$, the values in the second column of (100) are both equal to $1$ and then the values (101) are both equal to $1$ and both achieved for $x$ being an odd multiple of $\pi$.
-Suppose $V\neq W$. Since
-$\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}\left(\alpha,x\right)=\frac{-V\cos(x+\alpha)-W\cos\alpha}{(1-W\cos\alpha)^{2}}$ (102)
-at a stationary point $(\alpha,x)$, we obtain the following table of the values of $\frac{\partial^{2}f_{VW}}{\partial\alpha^{2}}(\alpha,x)$ at stationary points.
-$x$ even multiple of $\pi$ $x$ odd multiple of $\pi$
-$\alpha$ even multiple of $\pi$ $-\frac{V+W}{(1-W)^{2}}$ $\frac{V-W}{(1-W)^{2}}$
-$\alpha$ odd multiple of $\pi$ $\frac{V+W}{(1+W)^{2}}$ $\frac{-V+W}{(1+W)^{2}}$
-where a positive value indicates that $\alpha$ is a minimum point of the function $f_{VW}(\cdot,x)$ and then $f_{VW}(\alpha,x)$ is a value of $f^{\min}_{VW}$, while a negative value indicates that $\alpha$ is a maximum point and then $f_{VW}(\alpha,x)$ is a value of $f^{\max}_{VW}$.
-This confirms (100) and shows that if $V<W$, then
-$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1-V}{1-W}\text{\ \ and\ \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1+V}{1+W},$
-and if $V>W$, then
-$\min\limits_{x\in\mathbb{R}}f_{VW}^{\max}\left(x\right)=\frac{1+V}{1+W}\text{\ \ and\ \ }\max\limits_{x\in\mathbb{R}}f_{VW}^{\min}\left(x\right)=\frac{1-V}{1-W}.$
-with both the maximum and minimum achieved for $x$ being an odd multiple of $\pi$. ∎
-Remark 57.
-When $x$ is an even multiple of $\pi$, $f_{VW}^{\max}(x)$ and $f_{VW}^{\min}(x)$ are at their maximum value and minimum value, respectively.
-On the other hand, when $x$ is an odd multiple of $\pi$, $f_{VW}^{\max}(x)$ and $f_{VW}^{\min}(x)$ are at their minimum value and maximum value, respectively.
-7.7. The norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$
-Next proposition provides an expression for the norm $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$.
-Proposition 58.
-Assume that $A$ is a real matrix. Let $j\in\{1,\ldots,q\}$ such that $\Lambda_{j}$ is complex. We have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1-W_{j}^{2}}{4}f_{V_{j}W_{j}}^{\max}\left(x_{j}(t)\right)},$ (103)
-where $f_{V_{j}W_{j}}^{\max}$ is defined in the previous Subsection 7.6 and $x_{j}(t)$ is defined in (81).
-Proof.
-By (71), Proposition 48 and Proposition 52, we have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}},$
-where $\mathcal{W}_{j}$ is the ellipse $\mathcal{E}_{j}$ of Proposition 48 when $n=2$ and $\mathcal{W}_{j}$ is the union of $\mathcal{E}_{j}$ and its interior when $n>2$. We now show that (103) holds.
-Suppose $n>2$. By using, in the complex plane, the cartesian coordinate with axes the axes of the ellipse $\mathcal{E}_{j}$, the set $\mathcal{W}_{j}$ is the set of point $(x,y)$ such that
-$\left(\frac{x}{\sigma_{j}}\right)^{2}+\left(\frac{y}{\mu_{j}}\right)^{2}\leq 1,$
-where $\sigma_{j}$ and $\mu_{j}$ are the lenghts of the axes of the ellipse. Thus
-$\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}$$\Leftrightarrow$ $\left(\frac{\rho\cos(\theta-\theta_{j})}{\sigma_{j}}\right)^{2}+\left(\frac{\rho\sin(\theta-\theta_{j})}{\mu_{j}}\right)^{2}\leq 1$
-$\Leftrightarrow$ $\rho^{2}\leq R_{j}\left(\theta\right)^{2},$
-where
-$R_{j}\left(\theta\right):=\sqrt{\frac{\sigma_{j}^{2}\mu_{j}^{2}}{\mu^{2}_{j}\cos^{2}(\theta-\theta_{j})+\sigma^{2}_{j}\sin^{2}(\theta-\theta_{j})}}.$ (104)
-Thus
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
-$=\max\limits_{\theta\in[0,2\pi]}\max\limits_{\rho\in\left[0,R_{j}\left(\theta\right)\right]}\rho\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
-$=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}.$
-In the case $n=2$, we have
-$\rho\mathrm{e}^{\sqrt{-1}\theta}\in\mathcal{W}_{j}\Leftrightarrow\rho^{2}=R_{j}\left(\theta\right)^{2}$
-and then
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
-in this case as well.
-Recall (76). In (104), we have
-$\sigma_{j}^{2}\mu_{j}^{2}=\frac{1-W^{2}_{j}}{4}$
-and, by introducing $\alpha=2(\theta-\theta_{j})$,
-$\mu^{2}_{j}\cos^{2}(\theta-\theta_{j})+\sigma^{2}_{j}\sin^{2}(\theta-\theta_{j})$
-$=\mu^{2}_{j}\cos^{2}\frac{\alpha}{2}+\sigma^{2}_{j}\sin^{2}\frac{\alpha}{2}$
-$=\mu^{2}_{j}\frac{1+\cos\alpha}{2}+\sigma_{j}^{2}\frac{1-\cos\alpha}{2}$
-$=\frac{\mu^{2}_{j}+\sigma^{2}_{j}+\left(\mu^{2}_{j}-\sigma^{2}_{j}\right)\cos\alpha}{2}$
-$=\frac{1-W_{j}\cos\alpha}{2}.$
-Now, we can write
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\max\limits_{\theta\in[0,2\pi]}R_{j}\left(\theta\right)\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+2(\theta-\theta_{j})\right)}{2}}$
-$=\max\limits_{\alpha\in[-2\theta_{j},2(2\pi-\theta_{j})]}\sqrt{\frac{1-W_{j}^{2}}{2\left(1-W_{j}\cos\alpha\right)}}\cdot\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\alpha\right)}{2}}$
-$=\max\limits_{\alpha\in[0,2\pi]}\sqrt{\frac{1-W_{j}^{2}}{2\left(1-W_{j}\cos\alpha\right)}}\cdot\sqrt{\frac{1+V_{j}\cos\left(x_{j}(t)+\alpha\right)}{2}},$
-i.e. (103) by recalling the definition (93)-(87) of $f^{\max}_{V_{j}W_{j}}$. ∎
-Recall Lemma 56 (and the observation just before it) and the previous Proposition 58. When $V_{j}=0$, we have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1+W_{j}}{4}}.$
-When $W_{j}=0$, we have
-$\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{1+V_{j}}{4}}.$
-On the other hand, when $V_{j}\neq 0$ and $W_{j}\neq 0$, $\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}$, considered as a $\frac{\pi}{\omega_{j}}$-periodic function of $t$, oscillates monotonically between the values
-$\max\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{(1+W_{j})(1+V_{j})}{4}}$ (105)
-and
-$\min\limits_{t\in\mathbb{R}}\left\|\widehat{\Theta}_{j}\left(t\right)\right\|_{2}=\sqrt{\frac{a(V_{j},W_{j})}{4}},$ (106)
-where
-$a(V_{j},W_{j}):=\left\{\begin{array}[]{l}(1+W_{j})(1-V_{j})\text{\ if\ }V_{j}\leq W_{j}\\ \\ (1-W_{j})(1+V_{j})\text{\ if\ }V_{j}\geq W_{j}.\end{array}\right.$
-The maximum value is achieved if and only if $x_{j}(t)$ is an even multiple di $\pi$ and the minimum value is achieved if and only if $x_{j}(t)$ is an odd multiple di $\pi$.
-8. The generic real case. Part III: asymptotic condition numbers
-In this section, we describe the asymptotic condition numbers $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$ for a real matrix $A$ and in a generic situation for $A$, $y_{0}$ and $\widehat{z}_{0}$.
-This description is given in the next theorem.
-Theorem 59.
-Assume that $A$ is a real matrix, $\Lambda_{1}$ consists of either a real simple eigenvalue(i.e.
-$\Lambda_{1}$ is real), or a single pair of complex conjugate smple eigenvalues (i.e.
-$\Lambda_{1}$ is complex), and $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$.
-We have (recall Subsections 6.3 and 6.4 )
-$K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{g_{1}\left(t,\widehat{z}_{0}\right)}{g_{1}\left(t,\widehat{y}_{0}\right)}$ (107)
-and
-$K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{g_{1}\left(t\right)}{g_{1}\left(t,\widehat{y}_{0}\right)}.$ (108)
-Proof.
-By reminding that $v^{(1)}w^{(1)}$ is the projection onto the eigenspace of $\lambda_{1}$, the projections of $y_{0}$ and $\widehat{z}_{0}$ onto the eigenspace of $\lambda_{1}$ are
-$v^{(1)}w^{(1)}y_{0}=\left(w^{(1)}y_{0}\right)v^{(1)}\neq 0$
-and
-$v^{(1)}w^{(1)}\widehat{z}_{0}=\left(w^{(1)}\widehat{z}_{0}\right)v^{(1)}\neq 0,$
-respectively, and then $j^{\ast}=j^{\ast\ast}=1$. By points 2 of Remarks 33 and 35, we have
-$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|Q_{10}(t)\widehat{z}_{0}\right\|}{\left\|Q_{10}(t)\widehat{y}_{0}\right\|}.$
-and
-$K_{\infty}\left(t,y_{0}\right)=\frac{\left\|Q_{10}(t)\right\|}{\left\|Q_{10}(t)\widehat{y}_{0}\right\|}.$
-Now, (107) and (108) follows by (68) and (69). ∎
-Remark 60.
-The previous theorem considers a generic situation for the real matrix $A$, the initial value $y_{0}$ and the direction of perturbation $\widehat{z}_{0}$.
-In fact, as previously noted, it is a generic situation for $A$ real to have, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ consisting of either a real simple eigenvalue, or a single pair of complex conjugate simple eigenvalues.
-Moreover it is a generic situation for $y_{0}$ to satisfy $w^{(1)}y_{0}\neq 0$ and it is a generic situation for $\widehat{z}_{0}$ to satisfy $w^{(1)}\widehat{z}_{0}\neq 0$.
-Now, we describe the result of the previous theorem by separating the two cases of $\Lambda_{1}$real and $\Lambda_{1}$ complex.
-  • If $\Lambda_{1}$ is real, then
-    $K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
-    $K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
-    by Lemma 45.
-  • If $\Lambda_{1}$ is complex, then
-    $K_{\infty}(t,y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}$ (109)
-    $K_{\infty}(t,y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\cdot\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}$ (110)
-    by Lemma 46.
-8.1. The oscillation scale factors and the oscillating terms
-Consider the case where $\Lambda_{1}$ is complex. We can rewrite (109) and (110) as
-$K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z_{0}})\cdot\mathrm{OT}(t,y_{0},\widehat{z_{0}})$
-$K_{\infty}(t,y_{0})=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
-where the constants
-$\mathrm{OSF}(y_{0},\widehat{z_{0}}):=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$ (111)
-$\mathrm{OSF}(y_{0}):=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$ (112)
-are called the oscillation scale factors and the functions
-$t\mapsto\mathrm{OT}(t,y_{0},\widehat{z_{0}}):=\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|},\ t\in\mathbb{R},$ (113)
-$t\mapsto\mathrm{OT}(t,y_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|},\ t\in\mathbb{R},$ (114)
-which are periodic functions of period $\frac{\pi}{\omega_{1}}$, are called the oscillating terms.
-Remark 61.
-The oscillation scale factor $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ and the oscillating term $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ are two completely unrelated quantities, separating moduli and angles in the polar forms of $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$.
-Specifically, $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ depends on $y_{0}$ and $\widehat{z}_{0}$ through the moduli of the complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$.
-On the other hand, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ depends on $y_{0}$ and $\widehat{z}_{0}$ through the angles $\gamma_{1}(\widehat{y}_{0})$ and $\gamma_{1}(\widehat{z}_{0})$ in these polar forms (refer to the definition of $\widehat{\Theta}(t,u)$ in Lemma 46).
-Similarly, $\mathrm{OSF}(y_{0})$ depends on $y_{0}$ through the modulus of the complex number $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\mathrm{OT}\left(t,y_{0}\right)$ depends on $y_{0}$ through the angle $\gamma_{1}(\widehat{y}_{0})$ in this polar form.
-9. The generic real case. Part IV: oscillation scale factors and oscillating terms in case of the euclidean norm
-In this section, we study the oscillating scale factors and the oscillating terms in case of the Euclidean norm as vector norm.
-9.1. The oscillation scale factor $\mathrm{OSF}(y_{0},\widehat{z}_{0})$
-The oscillating scale factor
-$\mathrm{OSF}(y_{0},\widehat{z}_{0})$ in (111) is the ratio of the modules of the non-zero complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$.
-In case of the Euclidean norm as vector norm, such complex numbers can be arbitrary non-zero points in the set $\mathcal{W}_{1}$ of Proposition 48.
-Thus, as $y_{0}$ and $\widehat{z}_{0}$ vary, $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ can be any number in $(0,+\infty)$ when $n>2$ and any number in
-$\left[\frac{\mu_{1}}{\sigma_{1}},\frac{\sigma_{1}}{\mu_{1}}\right]=\left[\sqrt{\frac{1-W_{1}}{1+W_{1}}},\sqrt{\frac{1+W_{1}}{1-W_{1}}}\right],$
-where $\sigma_{1}$ and $\mu_{1}$ are the lengths of the semi-axes of the ellipse $\mathcal{E}_{1}$ of Proposition 48, when $n=2$.
-Remark 62.
-By (80), we can express $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ in terms of the components $c_{1}(u)$ and $d_{1}(u)$ of $u=\widehat{y}_{0},\widehat{z}_{0}$ along the first two right singular vectors of the matrix $R_{1}$ of Proposition 48.
-We have
-$\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{(1+W_{1})c_{1}(\widehat{z}_{0})^{2}+(1-W_{1})d_{1}(\widehat{z}_{0})^{2}}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$ (115)
-9.2. The oscillating term $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$
-In case of the Euclidean norm as vector norm, by reminding Proposition 52, we can write the oscillating term (113) as
-$\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}\left(\alpha,x\right)},$
-where $f_{V_{1}V_{1}}$ is the function defined in (87),
-$\alpha=x_{1}(t)+\Delta_{1}(\widehat{y}_{0})+\pi,$
-with $x_{1}(t)$ and $\Delta_{1}(\widehat{y}_{0})$ defined in (81) and (82), and
-$x=\Delta_{1}(\widehat{z}_{0})-\Delta_{1}(\widehat{y}_{0})-\pi=2\left(\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})\right)-\pi.$
-In the case $V_{1}=0$, we have
-$\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1.$
-Therefore, we assume $V_{1}\neq 0$. We are interested to the maximum and minimum values
-$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}^{\max}\left(x\right)}$ (116)
-and
-$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{f_{V_{1}V_{1}}^{\min}\left(x\right)}$ (117)
-of the oscillating term, where $f_{V_{1}V_{1}}^{\max}$ and $f_{V_{1}V_{1}}^{\min}$ are defined in (93) and (94). Observe that (116) and (117) hold since
-$\alpha=x_{1}(t)+\Delta_{1}(\widehat{y}_{0})+\pi=2(\omega_{1}t+\gamma_{1}(\widehat{y}_{0}))+\delta_{1}+\pi$
-varies in $\mathbb{R}$, as $t$ varies in $\mathbb{R}$.
-Given $y_{0}$, we do not determine (116) and (117) for any $\widehat{z}_{0}$ but we look only at their maximum and minimum values, by varying $\widehat{z}_{0}$.
-The next theorem gives such maximum and minumum values.
-They are independent of $y_{0}$.
-Theorem 63.
-Assume that $A$ is a real matrix, $\Lambda_{1}$ is complex and $w^{(1)}y_{0}\neq 0$. Assume $V_{1}\neq 0$.
-We have
-$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1+V_{1}}{1-V_{1}}}$
-and
-$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1.$
-The maximum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$ and the minimum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$.
-Moreover, we have
-$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$
-and
-$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1-V_{1}}{1+V_{1}}},$
-The maximum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is multiple of $\pi$ and the minimum value is achieved if and and only if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$.
-Proof.
-As $\widehat{z}_{0}$ varies in $\mathbb{R}^{n}$ with $\|\widehat{z}_{0}\|_{2}=1$ and $w^{(1)}\widehat{z}_{0}\neq 0$, $\widehat{w}^{(1)}\widehat{z}_{0}$ can be an arbitrary non-zero point in the set $\mathcal{W}_{1}$ of Proposition 48.
-Then, $\gamma_{1}\left(\widehat{z}_{0}\right)$ varies in $[0,2\pi]$ and then
-$x=2\left(\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})\right)-\pi$
-in (116)-(117) varies in $[-2\gamma_{1}\left(\widehat{y}_{0}\right)-\pi,-2\gamma_{1}\left(\widehat{y}_{0}\right)+3\pi]$, over two periods of the functions $f_{V_{1}V_{1}}^{\max}$ and $f_{V_{1}V_{1}}^{\min}$.
-We conclude that
-$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\max\limits_{x\in\mathbb{R}}f^{\max}_{V_{1}V_{1}}(x)}$
-$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\min\limits_{x\in\mathbb{R}}f^{\max}_{V_{1}V_{1}}(x)}$
-$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\max\limits_{x\in\mathbb{R}}f^{\min}_{V_{1}V_{1}}(x)}$
-$\min\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{R}^{n}\\ \left\|\widehat{z}_{0}\right\|_{2}=1\\ w^{(1)}\widehat{z}_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\min\limits_{x\in\mathbb{R}}f^{\min}_{V_{1}V_{1}}(x)}.$
-Now, the theorem follows by Lemma 56. ∎
-Remark 64.
-  • 1.
-    If $V_{1}$ is not close to $1$, then, for any $y_{0}$ and $\widehat{z}_{0}$, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ does not assume large or small values, as $t$ varies.
-In fact, we have
-    $\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}}.$
-  • 2.
-    When $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$, we have
-    $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1=\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$
-    and then
-    $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1,$
-    as it can be directly verified by the definition of $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ and Proposition 52.
-  • 3.
-    When $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, we have
-    $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1+V_{1}}{1-V_{1}}}$ (118)
-    and
-    $\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1-V_{1}}{1+V_{1}}}$ (119)
-    with $\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ at its maximum value and $\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ at its minimum value, as $\widehat{z}_{0}$ varies.
-  • 4.
-    If $V_{1}$ is close to $1$ and $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies.
-This follows by (118) and (119).
-  • 5.
-    The condition $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$ is equivalent to saying that the vectors $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$ are orthogonal in the complex plane, i.e.
-    $(1+W_{1})c_{1}(\widehat{y}_{0})c_{1}(\widehat{z}_{0})+(1-W_{1})d_{1}(\widehat{y}_{0})d_{1}(\widehat{z}_{0})=0$
-    by reminding (79).
-    The condition $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is a multiple of $\pi$ is equivalent to saying that the vectors $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$ are parallel in the complex plane, i.e.
-    $c_{1}(\widehat{y}_{0})d_{1}(\widehat{z}_{0})-d_{1}(\widehat{y}_{0})c_{1}(\widehat{z}_{0})=0.$
-  • 6.
-    Some specific situations are worth considering.
-    • a)
-      If $y_{0}$ is orthogonal to the first right singular vector of the matrix $R_{1}$ of Proposition 48 and $\widehat{z}_{0}$ is orthogonal to the second right singular vector (i.e.
-$c_{1}(\widehat{y}_{0})=0$ and $d_{1}(\widehat{z}_{0})=0$), then, by previous points 5 and 3, (118)-(119) hold.
-Moreover, by (115) we have
-      $\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{1+W_{1}}{1-W_{1}}}\cdot\frac{|c_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$
-      If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies.
-If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ large.
-Regarding the asymptotic condition number
-      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0}),$
-      it is expected that it assumes large values, as $t$ varies, if $V_{1}$ is close to $1$ or $W_{1}$ is close to $1$: we have
-      $\max\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})=\sqrt{\frac{(1+W_{1})(1+V_{1})}{(1-W_{1})(1-V_{1})}}\cdot\frac{|c_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$ (120)
-    • b)
-      If $y_{0}$ is orthogonal to the second right singular vector of $R_{1}$ and $\widehat{z}_{0}$ is orthogonal to the first right singular vector (i.e.
-$d_{1}(\widehat{y}_{0})=0$ and $c_{1}(\widehat{z}_{0})=0$), then (118)-(119) hold again.
-Moreover, we have
-      $\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{1-W_{1}}{1+W_{1}}}\cdot\frac{|d_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$
-      If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values, as $t$ varies.
-If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ small.
-Regarding the asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$, it is expected that it assumes large values, as $t$ varies, if the ratio $\frac{1-W_{1}}{1-V_{1}}$ is large: we have
-      $\max\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})=\sqrt{\frac{(1-W_{1})(1+V_{1})}{(1+W_{1})(1-V_{1})}}\cdot\frac{|d_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$ (121)
-    • c)
-      If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the first right singular vector of $R_{1}$ (i.e.
-$c_{1}(\widehat{y}_{0})=0$ and $c_{1}(\widehat{z}_{0})=0$), then, by previous point 2 and 5, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$.
-Moreover, we have
-      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})=\frac{|d_{1}(\widehat{z}_{0})|}{|d_{1}(\widehat{y}_{0})|}.$
-    • d)
-      If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the second right singular vector of $R_{1}$ (i.e.
-$d_{1}(\widehat{y}_{0})=0$ and $d_{1}(\widehat{z}_{0})=0$), then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=1$.
-Moreover, we have
-      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})=\frac{|c_{1}(\widehat{z}_{0})|}{|c_{1}(\widehat{y}_{0})|}.$
-9.3. The oscillation scale factor $\mathrm{OSF}(y_{0})$
-The oscillating scale factor $\mathrm{OSF}(y_{0})$ in (112) is the reciprocal of the modulus of the complex number $\widehat{w}^{(1)}\widehat{y}_{0}$.
-In case of the Euclidean norm as vector norm, such complex number can be an arbitrary non-zero point in the set $\mathcal{W}_{1}$ of Proposition 48.
-Thus, by varying $y_{0}$, $\mathrm{OSF}(y_{0})$ can be any number in
-$\left[\frac{1}{\sigma_{1}},+\infty\right)=\left[\sqrt{\frac{2}{1+W_{1}}},+\infty\right)$
-when $n>2$ and any number in
-$\left[\frac{1}{\sigma_{1}},\frac{1}{\mu_{1}}\right]=\left[\sqrt{\frac{2}{1+W_{1}}},\sqrt{\frac{2}{1-W_{1}}}\right].$
-when $n=2$.
-Remark 65.
-By (80), we can express $\mathrm{OSF}(y_{0})$ in terms of the components $c_{1}(\widehat{y}_{0})$ and $d_{1}(\widehat{y}_{0})$ of $\widehat{y}_{0}$ along the first two right singular vectors of the matrix $R_{1}$.
-We have
-$\mathrm{OSF}(y_{0})=\sqrt{\frac{2}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$ (122)
-9.4. The oscillating term $\mathrm{OT}(t,y_{0})$
-In case of the Euclidean norm as vector norm, by reminding Propositions 52 and 58, we can write the oscillating term (114) as
-$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\frac{f_{V_{1}W_{1}}^{\max}(x_{1}(t))}{1+V_{1}\cos(x_{1}(t)+\Delta_{1}(\widehat{y_{0}}))}},$ (123)
-where $x_{1}(t)$ and $\Delta_{1}(\widehat{y}_{0})$ are defined in (81) and (82).
-In the case $V_{1}=0$, we have
-$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1+W_{1}}{2}}$
-by reminding (95). Therefore, we assume $V_{1}\neq 0$. We are interested to the maximum and minimum values
-$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\max\limits_{x\in\mathbb{R}}\frac{f_{V_{1}W_{1}}^{\max}(x)}{1+V_{1}\cos(x+\Delta_{1}(\widehat{y_{0}}))}}$ (124)
-and
-$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\cdot\min\limits_{x\in\mathbb{R}}\frac{f_{V_{1}W_{1}}^{\max}(x)}{1+V_{1}\cos(x+\Delta_{1}(\widehat{y_{0}}))}}.$ (125)
-of the oscillating term. Observe that (124) and (125) hold since $x_{1}(t)$ varies in $\mathbb{R}$, as $t$ varies in $\mathbb{R}$. In the case $V_{1}\neq 0$ and $W_{1}=0$, we have
-$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1+V_{1}}{2(1-V_{1})}}$
-$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1}{2}},$
-by reminding (95). Observe that, for $V_{1}=0$ or $W_{1}=0$, such maximum and minimum are independent of $y_{0}$.
-We proceed by assuming $V_{1}\neq 0$ and $W_{1}\neq 0$.
-We do not determine (124) and (125) for any $y_{0}$, but we look only at their maximum and minimum values as $y_{0}$ varies.
-The next theorem gives such maximum and minumum values.
-Theorem 66.
-Assume that $A$ is a real matrix and $\Lambda_{1}$ is complex. Assume $V_{1}\neq 0$ and $W_{1}\neq 0$.
-We have
-$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\max}(V_{1},W_{1}):=\sqrt{\frac{(1+W_{1})(1+V_{1})}{2(1-V_{1})}}$ (126)
-and
-$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\left\{\begin{array}[]{l}\sqrt{\frac{1+W_{1}}{2}\cdot\frac{1-V_{1}^{2}}{1-Q_{1}V_{1}}}\text{\ if\ }V_{1}\leq\frac{2W_{1}}{1+W_{1}}\\ \\ \sqrt{\frac{(1-W_{1})(1+V_{1})}{2(1-V_{1})}}\text{\ if\ }\frac{2W_{1}}{1+W_{1}}\leq V_{1},\end{array}\right.$ (130)
-where
-$Q_{1}:=\frac{1}{2}\cdot\frac{V_{1}}{W_{1}}(1+W_{1}).$ (131)
-The maximum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$ and the minimum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is a multiple of $\pi$.
-The angle $\theta_{1}$ is defined in Subsection 7.4.
-In the case $V_{1}\leq\frac{2W_{1}}{1+W_{1}}$, we have
-$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)\leq\sqrt{\frac{(1+W_{1})(1+V_{1})}{2}}.$ (132)
-Moreover, we have
-$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1+W_{1}}{2}}$ (133)
-and
-$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\min}(V_{1},W_{1}):=\left\{\begin{array}[]{l}\sqrt{\frac{(1+W_{1})(1-V_{1})}{2(1+V_{1})}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \sqrt{\frac{1-W_{1}}{2}}\text{\ if\ }V_{1}\geq W_{1}.\end{array}\right.$ (137)
-The maximum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is a multiple of $\pi$ and the minimum value is achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$.
-Proof.
-We introduce the function $H:\mathbb{R}^{2}\rightarrow\mathbb{R}$ given by
-$H(x,\beta)=\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)},\ (x,\beta)\in\mathbb{R}^{2}.$ (138)
-Thus, by (123), (124) and (125) we can write
-$\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H(x_{1}(t),\Delta_{1}(\widehat{y}_{0}))}$ (139)
-and
-$\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H^{\max}(\Delta_{1}(\widehat{y}_{0}))}$
-$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,\widehat{y}_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}H^{\min}(\Delta_{1}(\widehat{y}_{0}))},$
-where the functions $H^{\max},H^{\min}:\mathbb{R}\rightarrow\mathbb{R}$ are given by
-$H^{\max}(\beta):=\max\limits_{x\in\mathbb{R}}H(x,\beta)\text{\ \ and\ \ }H^{\min}(\beta):=\min\limits_{x\in\mathbb{R}}H(x,\beta),\ \beta\in\mathbb{R}.$
-Observe that (140) holds since, as $t$ varies, $x_{1}(t)$ in (81) varies in $\mathbb{R}$.
-Therefore, we have
-$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta))}$
-$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta))}$
-$\max\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta))}$
-$\min\limits_{\begin{subarray}{c}{y}_{0}\in\mathbb{R}^{n}\\ w^{(1)}y_{0}\neq 0\end{subarray}}\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=\sqrt{\frac{1-W_{1}^{2}}{2}\min\limits_{\beta\in\mathbb{R}}H^{\min}(\beta))}.$
-Observe that (141) hold since, as $y_{0}\in\mathbb{R}^{n}$ with $w^{(1)}y_{0}\neq 0$ varies, $\widehat{w}^{(1)}\widehat{y}_{0}$ can be an arbitrary non-zero point in $\mathcal{W}_{1}$ and then $\gamma_{1}(\widehat{y}_{0})$ varies in $[0,2\pi]$ and then
-$\Delta_{1}(\widehat{y}_{0})=2(\gamma_{1}(\widehat{y}_{0})-\theta_{1})$
-varies in $[-2\theta_{1},-2\theta_{1}+4\pi]$, over two periods of the functions $H^{\max}$ and $H^{\min}$.
-The theorem will follow once we have determined
-$\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta),\ \min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta),\ \max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)\text{\ and\ }\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)$
-and for which values $\beta$ the maximum and minimum values are achieved.
-By Lemma 56, we immediately see that
-$\max\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)$ $=$ $\max\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}H(x,\beta)=\max\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)}$
-$=$ $\frac{1+V_{1}}{(1-W_{1})(1-V_{1})}$
-with the maximum value achieved if and only if $x$ is an even multiple of $\pi$ (thus $f_{V_{1}W_{1}}^{\max}(x)$ achieves its maximum value
-$\frac{1+V_{1}}{1-W_{1}}$
-as $x$ varies) and $x+\beta$ is an odd multiple of $\pi$ (thus $1+V_{1}\cos(x+\beta)$ achieves its minimum value $1-V_{1}$ as $x$ and $\beta$ vary).
-In other words, the maximum is achieved if and only if $\beta$ is an odd multiple of $\pi$.
-Thus, by (141) and (LABEL:MAXMAX) we have (126) with the maximum achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}=\frac{\Delta_{1}(\widehat{y}_{0})}{2}=\frac{\beta}{2}$ is an odd multiple of $\frac{\pi}{2}$.
-Remind that $\Delta_{1}(\widehat{y}_{0})=\beta$ in (140).
-By Lemma 56, we also immediately see that
-$\min\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)$ $=$ $\min\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}H(x,\beta)=\min\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}\frac{f^{\max}_{V_{1}W_{1}}(x)}{1+V_{1}\cos(x+\beta)}$ (146)
-$=$ $\left\{\begin{array}[]{l}\frac{1-V_{1}}{(1-W_{1})(1+V_{1})}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \frac{1}{1+W_{1}}\text{\ if\ }V_{1}\geq W_{1},\end{array}\right.$
-with the minimum value achieved if and only $x$ is an odd multiple of $\pi$ (thus $f_{V_{1}W_{1}}^{\max}(x)$ achieves its minimum value
-$\left\{\begin{array}[]{l}\frac{1-V_{1}}{1-W_{1}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \frac{1+V_{1}}{1+W_{1}}\text{\ if\ }V_{1}\geq W_{1}\end{array}\right.$
-as $x$ varies) and $x+\beta$ is an even multiple of $\pi$ (thus $1+V_{1}\cos(x+\beta)$ achieves its maximum value $1+V_{1}$ as $x$ and $\beta$ vary).
-In other words, the minimum is achieved if and only if $\beta$ is an odd multiple of $\pi$.
-Thus, by (141) and (LABEL:MINMIN) we have (137) with the minimum achieved if and only if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$.
-It remains to determine
-$\min\limits_{\beta\in\mathbb{R}}H^{\max}(\beta)=\min\limits_{\beta\in\mathbb{R}}\max\limits_{x\in\mathbb{R}}H(x,\beta)\text{\ and\ }\max\limits_{\beta\in\mathbb{R}}H^{\min}(\beta)=\max\limits_{\beta\in\mathbb{R}}\min\limits_{x\in\mathbb{R}}H(x,\beta).$
-By using the property
-$f_{V_{1}W_{1}}^{\max}(-x)=f_{V_{1}W_{1}}^{\max}(x),\ x\in\mathbb{R},$
-(this follows since $\alpha^{\max}_{V_{1}W_{1}}(-x)=-\alpha^{\max}_{V_{1}W_{1}}(x)$, with $\alpha^{\max}_{V_{1}W_{1}}(x)$ defined in Lemma 54), we obtain
-$H^{\max}(\pi-\beta)=H^{\max}(\pi+\beta)\text{\ and\ }H^{\min}(\pi-\beta)=H^{\min}(\pi+\beta),\ \beta\in\mathbb{R}.$
-This fact shows that we can confine our analysis to the case $\beta\in[0,\pi]$ (remind that $H^{\max}$ and $H^{\min}$ have period $2\pi$).
-For the remainder of the proof, we assume $V_{1}\neq W_{1}$.
-This ensures that the function $\alpha_{V_{1}W}^{\max}$ is defined and differentiable at every point $x\in\mathbb{R}$, by reminding Remark 55.
-The annoying case $V_{1}=W_{1}$ is treated as a limiting case of the situation $V_{1}\neq W_{1}$.
-This thanks to the following proposition.
-Proposition 67.
-Let $F:A\times B\rightarrow\mathbb{R}$ be a continuous function, with $A$ and $B$ metric spaces. If
-$\lim\limits_{n\rightarrow\infty}a_{n}=a,\ \lim\limits_{n\rightarrow\infty}b_{n}=b\text{\ \ and\ \ }F(a_{n},b_{n})=\max\limits_{x\in A}F(x,b_{n}),$
-where $a_{n},a\in A$ and $b_{n},b\in B$, then
-$F(a,b)=\max\limits_{x\in A}F(x,b).$
-Proof.
-Let $x\in A$. We have
-$F(a,b)$ $=$ $F(x,b)+F(x,b_{n})-F(x,b)+F(a_{n},b_{n})-F(x,b_{n})+F(a,b)-F(a_{n},b_{n})$
-$\geq$ $F(x,b)+F(x,b_{n})-F(x,b)+F(a,b)-F(a_{n},b_{n})$
-since
-$F(a_{n},b_{n})-F(x,b_{n})\geq 0.$
-By letting $n\rightarrow\infty$, we obtain
-$F(a,b)\geq F(x,b)$
-for the continuity of $F$. ∎
-For a given $\beta\in[0,\pi]$, the values $H^{\max}(\beta)$ and $H^{\min}(\beta)$ are values $H(x,\beta)$, where $x$ is stationary point of the function $H(\cdot,\beta):\mathbb{R}\rightarrow\mathbb{R}$ and satisfies
-$\frac{\partial H}{\partial x}(x,\beta)=0.$ (148)
-Since
-$\frac{\partial H}{\partial x}(x,\beta)=\frac{\partial}{\partial x}\left(\frac{f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)}{1+V_{1}\cos(x+\beta)}\right)$
-$=\frac{1}{(1+V_{1}\cos(x+\beta))^{2}}$
-$\ \ \ \ \cdot\left(\left(\underset{=0}{\underbrace{\frac{\partial f_{V_{1}W_{1}}}{\partial\alpha}(\alpha^{\max}_{V_{1}W_{1}}(x),x)}}(\alpha^{\max}_{V_{1}W_{1}})^{\prime}(x)+\frac{\partial f_{V_{1}W_{1}}}{\partial x}(\alpha^{\max}_{V_{1}W_{1}}(x),x)\right)\right.$
-$\left.\ \ \ \ \ \ \ \ \ \ \ \cdot(1+V_{1}\cos(x+\beta))\right.$
-$\left.\ \ \ \ \ \ \ \ \ \ \ -f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)\left(-V_{1}\sin(x+\beta)\right)\right)$
-$=\frac{1}{(1+V_{1}\cos(x+\beta))^{2}}$
-$\ \ \ \ \cdot\left(\frac{\partial f_{V_{1}W_{1}}}{\partial x}(\alpha^{\max}_{V_{1}W_{1}}(x),x)(1+V_{1}\cos(x+\beta))+f_{V_{1}W_{1}}(\alpha^{\max}_{V_{1}W_{1}}(x),x)V_{1}\sin(x+\beta)\right),$
-we see (remind (87) and (98)) that (148) is equivalent to
-$-\sin(x+\alpha^{\max}_{V_{1}W_{1}}(x))+\sin(x+\beta)-V_{1}\sin(\alpha^{\max}_{V_{1}W_{1}}(x)-\beta)=0.$ (149)
-By solving this equation for $\beta=0$, we obtain (see Appendix B)
-$H^{\max}(0)=\left\{\begin{array}[]{l}\frac{1}{1-W_{1}}\cdot\frac{1-V_{1}^{2}}{1-Q_{1}V_{1}}\text{\ if\ }V_{1}\leq\frac{2W_{1}}{1+W_{1}}\\ \\ \frac{1}{1+W_{1}}\cdot\frac{1+V_{1}}{1-V_{1}}\text{\ if\ }\frac{2W_{1}}{1+W_{1}}\leq V_{1},\end{array}\right.$ (150)
-where $Q_{1}$ is defined in (131), and
-$H^{\min}(0)=\frac{1}{1-W_{1}}.$ (151)
-Moreover, in the case $V_{1}\leq\frac{2W_{1}}{1+W_{1}}$, we have
-$H^{\max}(0)\leq\frac{1}{1-W_{1}}(1+V_{1}).$ (152)
-The stationary points of $H(\cdot,\beta)$ (i.e. the solutions of (149)) form, as function of $\beta\in[0,\pi]$, a family of continuous branches
-$x_{i}:D_{i}\rightarrow\mathbb{R},\ \beta\mapsto x_{i}(\beta),\ i\in I,$
-where $I$ is a finite set of indices and $D_{i}$ a closed interval included in $[0,\pi]$. These branches satisfy the differential equation
-$\frac{d}{d\beta}\frac{\partial H}{\partial x}(x_{i}(\beta),\beta)=\frac{\partial^{2}H}{\partial x^{2}}(x_{i}(\beta),\beta)x_{i}^{\prime}(\beta)+\frac{\partial^{2}H}{\partial\beta\partial x}(x_{i}(\beta),\beta)=0.$ (153)
-The ends of the interval $D_{i}$ are 00, $\pi$ or points $\beta$ where
-$\frac{\partial^{2}H}{\partial x^{2}}(x_{i}(\beta),\beta)=0.$
-Each branch $x_{i}$, $i\in I$, is continuously differentiable in the interior of $D_{i}$.
-The values of $H(\cdot,\beta)$ at the stationary points also form a family of continuous branches
-$h_{i}:D_{i}\rightarrow\mathbb{R},\ \beta\mapsto H(x_{i}(\beta),\beta),\ i\in I,$
-continuously differentiable in the interior of $D_{i}$.
-Since $H^{\max}(\beta)$ and $H^{\min}(\beta)$, $\beta\in[0,\pi]$, are values $H(x_{i}(\beta),\beta)$ for some $i\in I$, we have
-$H^{\max}(\beta)=\max\limits_{\begin{subarray}{c}i\in I\\ \beta\in D_{i}\end{subarray}}h_{i}(\beta)\text{\ \ and\ \ }H^{\min}(\beta)=\min\limits_{\begin{subarray}{c}i\in I\\ \beta\in D_{i}\end{subarray}}h_{i}(\beta),\ \beta\in[0,\pi].$ (154)
-In Figure 8, we see examples of such branches.
-Remark 68.
-Observe that the equation (149) can be factorized as
-$-2\sin\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}\left(\cos\left(x+\frac{\alpha^{\max}_{V_{1}W_{1}}(x)+\beta}{2}\right)+V_{1}\cos\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}\right)=0$
-and then its solutions are the union of the solutions of
-$\alpha^{\max}_{V_{1}W_{1}}(x)-\beta\text{\ is a multiple of $2\pi$}$ (155)
-and
-$\cos\left(x+\frac{\alpha^{\max}_{V_{1}W_{1}}(x)+\beta}{2}\right)+V_{1}\cos\frac{\alpha^{\max}_{V_{1}W_{1}}(x)-\beta}{2}=0.$
-The branches $h_{i}$ corresponding to solutions of (155) are given by
-$h_{i}(\beta)=\frac{1}{1-W_{1}\cos\beta}$
-as can be immediately verified by the definition (138) of $H$. In Figure 9, we see examples of such branches $x_{i}$ and $h_{i}$ corresponding to solutions of (155).
-Figure 8. Left: branches $x_{i}$ for $(\beta,x)\in[0,\pi]\times[-\pi,\pi]$. Right: branches $h_{i}$ for $\beta\in[0,\pi]$.
-Figure 9. Left: branches $x_{i}$ for $(\beta,x)\in[0,\pi]\times[-\pi,\pi]$ corresponding to solutions of (155). Right: branches $h_{i}$ for $\beta\in[0,\pi]$ corresponding to solutions of (155).
-Proposition 69.
-The branches $h_{i}$, $i\in I$, are either monotonically increasing or decreasing.
-Proof.
-Let $i\in I$. We show that there are no interiors points of $D_{i}$ where the derivative $h_{i}^{\prime}$ is zero.
-For $\beta$ interior point of $D_{i}$, we have
-$h_{i}^{\prime}(\beta)$ $=$ $\frac{d}{d\beta}H(x_{i}(\beta),\beta)$
-$=$ $\frac{\partial H}{\partial x}(x_{i}(\beta),\beta)x_{i}^{\prime}(\beta)+\frac{\partial H}{\partial\beta}(x_{i}(\beta),\beta)$
-$=$ $\frac{\partial H}{\partial\beta}(x_{i}(\beta),\beta)$
-$=$ $\frac{V_{1}\sin(x_{i}(\beta)+\beta)}{(1+V_{1}\cos(x_{i}(\beta)+\beta))^{2}}f^{\max}_{V_{1},W_{1}}(x_{i}(\beta)),$
-where $f^{\max}_{V_{1},W_{1}}(x_{i}(\beta))>0$ (see Lemma 56, or the definition (87), by reminding $V_{1},W_{1}\in[0,1)$). Then, for $h^{\prime}_{i}(\beta)=0$, we have
-$-V_{1}\sin(x+\alpha)-W_{1}\sin\alpha+V_{1}W_{1}\sin x=0$
-$-\sin(x+\alpha)+\sin(x+\beta)-V_{1}\sin(\alpha-\beta)=0$
-$\sin(x+\beta)=0,$
-where $x=x_{i}(\beta)$ and $\alpha=\alpha^{\max}_{V_{1}W_{1}}(x_{i}(\beta))$.
-Regarding the equations in (157), observe that:
-  • The first equation is the equation for the stationary points $\alpha$ of $f_{V_{1}W_{1}}(\cdot,x_{i}(\beta))$ (recall (88)): $\alpha=\alpha^{\max}_{V_{1}W_{1}}(x_{i}(\beta))$ is a stationary point of $f_{V_{1}W_{1}}(\cdot,x_{i}(\beta))$.